A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.03

Title word cross-reference

F_{1} [DA00, SA04b]. F_{2} [DA00]. F_{ST} [MN07]. Q_{ST} [MN07]. × [CGAD^{+}02].

-Tagged [BHB^{+}07].

1* [RHR^{+}06]. 16S [AWNT00]. 1930s [WEB00]. 1970s [RWB^{+}08]. 1992 [HK08].

2-Year-Olds [VFT02]. 21st [CW08].

5 [SSSS05]. 5’-Nuclease [SSSS05].

6-Year [DHR07].

Ability [CDT^{+}02, DH04, GMJ^{+}06, MSM^{+}08, SWD04, TAD05]. Abiotic [CW00a, KC05, QH05, SW09b]. Ablation [LFG06]. Abnormalities [HEW^{+}05, SKA^{+}09]. Absence [MOT04, VT06]. Abundance [Ano02c, BL05, BS08, BZW04, DA05, ECW^{+}09a, Gro06, HHDC03, HPB05, IHF08, IH04, JM05, JD07, KM01, LDK^{+}04, MB05a, MS04, NL03, PTG04, RL08, Sch00a, SA04a, SW09b, SWF09, Sto02, Tho03, TAM07, iTE08, TAC08, WEB00, WKCC08, WMH^{+}04, YSC^{+}06]. Acceleration [KGW05]. Access [MMH01]. Accessible [AKPQ08]. Acclimated [BCW^{+}09]. Account [Bre08]. Accounting [BH07]. Accumulation [LDDB08, OW09]. Accuracy [RP07, SS08a]. Acid [HNH^{+}04, SBP^{+}08]. Acidic [KS01, WSJ^{+}05]. Acidification [BL00]. Acidified [BLS07]. Acids [RSBS09]. Acipenser [MSS^{+}00]. Acoustic [CBN^{+}09, Gan08, GA07, HHDC03, LKM04, NAG07, WLBBM09]. Acoustics [FT08, LBBF^{+}08, LPJS08, LMR08]. Across [CW00a, PMP05, RIA^{+}07, BCJ^{+}06, DS04, EWB^{+}06, ECW^{+}09a, dlHFB04, IH04, KM01, MB08b, Sch00a, WBH04]. Across-Species [PMP05]. Actions [PF05]. Activity [BGM06, BH06b, HB00, RE08, SCH04a, SLC06, SHDR03]. Actual [iTE08]. Acute [BH00c, CG04, WCBS06]. Additions [WHC04]. Additive [JD07]. Adfluvial [BV01, BLG01, HB08b]. Adirondack [BLS07, BK03, SCP^{+}05, WJK07]. Adjacent [BSM^{+}08]. Adjoining [LYS06]. Adjustable [KM05]. Adjusted [BKP^{+}04]. Admixture [HSS07, LCNB^{+}06]. Adult [BAS^{+}09, Bet05, BQH04, BKP^{+}04, BGM06, BKHH07, CSPP00, CH05, DSH^{+}09, DGC07, ERS^{+}06, FSMH08, GABC00, GKB^{+}06, GSSB03, GDW03, HBP^{+}09, HB00, HBBG07, HCK06, JJRP08, JCB^{+}05, KJL^{+}07, KPB^{+}04, KD01, KRM03, LS04, LHL06, MKDB03, MOSB02, NAG07, PPR06, QB01, RHM^{+}06, RHB^{+}06, RQ02, SA06, WNH05, WCM^{+}07, WHT^{+}00, WHZR03, YI07, ZBF05]. Adults [CG06]. aeglefinus [BB00]. Aerial [VDG02]. Affect [GDWC04, GJP09, LBVW03, Mic06, NBS07, PF05, TBA^{+}04]. Affected [SHT^{+}00]. Affecting [AHJ^{+}05, BHC03, BBRH06, CDT^{+}02, CPG03, GBMB01, HW04, HB04, LDK^{+}04, PYFS07, QPG^{+}02, SP04, ZDK^{+}04]. Affects [RHR^{+}06, ZG07]. Affinity [PR06]. After [DHR07, AW05, BSN04, BH06b, CRMN04, CLSD00, DC02, FHML08, HM07, HBTT06, LAM04, MKBF00, MSET05, MBC^{+}05, QK03, SKK^{+}08, SBC^{+}05]. Age [AHC^{+}09, BHP^{+}08, BL05, BSEM05, BHMK05, BCV07, BL02, Bro05, BGS^{+}04, BCDF^{+}09, BHV08, CB04, CF07b, CF07a, DFBC^{+}09, DH04, DA05, EH03, FSH07, GWSF00, HPW08, HC05, HJM04, HWM04, HWW^{+}07, Har00, HB09, HRMC07, JN00b, KYWC09, KC09, KSR01, LT01a, LGJ02, LM08b, MC00a, MPVM00, MBS03a, MD05, MB05a, MZ02, MFFI09, NSP03, NTW02, NFP00, NSF04, NBK01, OEO^{+}00, PB03, PLO07, PA01, PAD01, PKM07, PSJ08, SBG01a, SRL^{+}06, Sch05, SGF03, SWHW06, SW08b, SLC06, SWD04, SGG09, SLC00, TRBH08, WMBM02, WKCC08, WW02, YSB^{+}08]. Age-0 [BSEM05, BHMK05, BL02, BHV08, CB04, CF07b, CF07a, DH04, DA05, FSH07, GWSF00, HC05, HJM04, Har00, HRMC07, JN00b, KYWC09, KC09, KSR01, LT01a, LM08b, MPVM00, MBS03a, MD05, MB05a, MZ02, MFFI09, NFP00, NSF04, PA01, PKM07, SBG01a, SGF03, SW08b, SWD04, SGG09, WMBM02]. Age-1 [EH03, TRBH08, WKCC08]. Age-2 [BL05, OEO^{+}00, SLC06]. Age-Dependent [BHP^{+}08]. Age-Specific [BCV07]. Age-Structured [HB09, PAD01]. Agents [CC03]. Ages [PB03]. Aggregation [LBVW03, PP08]. Aging [IC02, SC02]. Agricultural [CGL^{+}09]. Air [COB^{+}03, DS05, vT05]. Alabama [FHP00, DD02, GJP09, HI08, MG06, RBM06]. Alaska [BAWT05, AMB^{+}08, BQH04, BH03, BH04, BBWW02, BZW04, BWFB06, CBMM08, HSS07, HE00, HE01, JTM08, MAW09, MKDB03, MWMT02, PF00, RFOR03, SHT^{+}00, SWHW06, SGS05, SSO09, WSJH01, WRGN07, WORB00]. Alaskan [EM01, GKW^{+}06, RT04]. Albacore [Ess03]. Albemarle [RTBH05, TRBH08]. Alberta [TBA^{+}04]. Alevins [GAH^{+}06]. Alewife [HHB02, KRM03, WSP05, WKCC08]. Alewives [BR09a, HC05, HRMC07, MHD03, MHR^{+}05, OEO^{+}00, OLS04, PW09, PV04, WRK02]. Algorithm [IMM^{+}04, Lin04]. Alien [MMP02]. Alizarin [Bas04]. Alleles [SS08a]. Alleviation [ASK01]. Alligator [AMRM02, HHHK08]. Allocation [SJ08]. Allometric [DW08b]. Allometry [DSSA01]. Allopatric [Heg02]. Allow [MFOA09, OBB^{+}00]. Allows [IHF08]. Allozyme [BTLD08, KMSL02, KGS00, LYS06]. Allozymes [SAT^{+}07]. Alluvial [ML04]. along [BG02, BLGW06, HWM04, MDE^{+}09, MHM08, RA09, SFS04, TFO^{+}09, TTW^{+}09]. Altamaha [PSD^{+}08]. Alteration [WJB06]. Altered [MWN05]. Alternative [CST^{+}05, Ess03, GGCF07, HR09]. Alternatives [Tho03]. Alters [LBS^{+}06]. Alverson [ZM06b]. Always [CP03]. Amberjack [HWW^{+}07]. Ambloplites [KGS00]. America [TFO^{+}09, TTW^{+}09]. American [AO09, BIB04, BH00b, GVHC^{+}09, GSWS04, GA03, HL00a, HM09, HO05, HBO07, IMM^{+}04, Jes00, JSI09, LME^{+}08, MLSD07, MS04, OM02, OLWC06, RH01, SZM01, SFE01, SGG09, TAM07, WA05, WTO08, WMH^{+}04]. Ammonia [BWVN09, MH02]. among [BHHS04, BSS05, CN08, CN01, DW08a, DS05, DFCDR07, DK06, DA00, FWG07a, FGWS00, GW03, GFL00, HSK03, HBJ^{+}08b, KESB05, KGS00, KFB05, MFOA09, MOT04, OR06, OPH09, PT03a, QHR04, RHG^{+}02, SJ06, SLF^{+}05, SMK^{+}07, SBH00, TRK07, VHF05, WC00, WB01, WSJ06, WTSC00, WNG02, ZN07b]. Amphibian [WHLP07]. Amplified [FWG07b]. Anacostia [SJP05]. Anadromous [BCV07, BH06b, DB08, HC05, HTT00, HTBT01, MKP07, OPH09, PW09, RHB^{+}06, SLM07, YMJ00]. Anadromy [Beh02, BBS07, PBR^{+}01, PKR02]. Analog [WHC04]. Analyses [BHL05, DSH^{+}09, LMAF08, LYS06, TA06, WJMM07, WKM^{+}07]. Analysis [AWNT00, AW02, BBP^{+}04, BMM^{+}06, BJS^{+}06, BB00, BRK02, BBC^{+}08, BCT04, BT04a, BPM^{+}04, DKP^{+}06, DWBQ05, Esh09, Ess03, FWG07b, GFL00, GFLF02, HSS07, HMD02, HTBT01, JSH^{+}08, JM05, KAB09, Leg05, LFG06, LCNB^{+}06, MBMO06, MSS^{+}00, ME07, NFP00, PHF07, PLH^{+}08, SSHB02, SLP^{+}01, TR07, WGRW04, WGW06, WDHW09]. Ancestry [PWLL05]. Anchoveta [CH00]. Anchovy [HHS04]. Ancient [BPM^{+}04].

Angler [CKB01, Mir05a, VAR^{+}02, WP08, WS09]. Angler-Caught [WP08, WS09]. Angling [BE00, FCS^{+}03, HGR^{+}08, KSC^{+}06, MSET05, PCC^{+}09, PT03b, SCM01, SCWP09, SMS04, SSL^{+}03, SP04, SKC^{+}04, WJB06]. Anguilla [AWNT00]. Animal [MFC04]. Annual [BGS^{+}04, FRK01, HB07, JN00b, LS07, LHL06, PSD^{+}08, WSLT07, YI02]. Annular [MFC04]. Annuli [BGS^{+}04]. Anode [MK08]. Antenna [BHB^{+}07, HDL07]. Anthropogenic [MLSD07, WAS^{+}00]. Anticipated [RIA^{+}07]. Antithyroid [CC03]. Antitrawling [SJRE00]. Apache [CRA05]. Apex [PKR07]. Appalachian [BIB^{+}00, FRB06, LHM02, PLM05, PG07, URMH07]. Apparent [BPM^{+}04, KBPC07, WM05b, WM07]. Appearance [LAM04]. Application [Ano05, BOMC07, BMH07, BCS^{+}01, FWG07b, GH06, HB07, IHF08, LHOP01, LPN01, McK05, NHFF09, Pap08, PSJ08, SHT^{+}00, SMZB01, WRC^{+}03, vT05]. Applied [DW08b, Kol06, PF05, QH05]. Apportionment [HS08]. Approach [BP01, BL02, BHV08, GGCF07, HB07, HBJ^{+}08a, IDW03, KR07, KHKS00, MWG^{+}08, NN06, PF05, SSCM06, Tho03, TPS^{+}03, VBT08, VBM^{+}09, WLB^{+}00, WHS^{+}07]. Approaches [Ros03a]. Approximate [Mir07]. Aquaculture [TBWD06, VPN09, VGA01]. Aquatic [Bur09, CPB07, FSN^{+}03, KR07, MB05b, MFC04, PHWS09, POH^{+}09, RL08, SS06, Van04]. Aquatic-Vertebrate [PHWS09]. Aragonite [VPN09]. Archaeological [GSWS04]. Archimedes [MLB03]. Architecture [FCC07, VB02]. Arctic [BRK02, DH04, GN00, JT04, LM08b]. Area [BLG01, CMKL01, CMSM01, HB00, HKHC00]. Areas [CPG03, HPB06, HBBG07, Sto02]. argenteus [GK06]. Argentina [BOMC07, Beh02, PBR^{+}01, PKR02, VBM^{+}09]. Arizona [MPK03, Bro05, CRA05, CWB06, VBT08]. Arrangement [MK08]. Array [Hol08]. Arsenic [FSN^{+}03]. Artificial [HK04a, HSH^{+}05, OJ01, OS03, QPG^{+}02, RV07, RQ01, SGS06a, SGS06b, SKP^{+}08, SS05, TF07, VWD^{+}01]. Asian [DGC08, KA07]. Aspects [Spi01]. Assemblage [BJZH05, BWC^{+}07, DW06, GS06, OW01, QK03, SP07, TL06, WC00, WHLP07, WTW03, WTSC00, WS08]. Assemblage-Habitat [BWC^{+}07]. Assemblages [BHHS04, CB06, HW08a, HS08, MM02, MMP02, MB08b, Mea05, MC09, ML04, Mir05b, OW04, OBB^{+}09, PR01c, PHWS09, SJRE00, SK05b, TJNL08, TBA^{+}04, ZWT05]. Assemblagesin [TPS^{+}03]. Assess [AHC^{+}09, BGM06, MWG^{+}08, MOP02]. Assessed [BBC^{+}08, DWBQ05]. Assessing [ASS00, CH05, ESM^{+}03, FT08, HDRB07, HKCL07, HMT^{+}06, JR03, JT04, Kon00, MB08b, MMP03, MC09, PK04, Ros03a, SA01, YMJ00]. Assessment [BOMC07, BWS^{+}05, BBV00, BSW00, BCW^{+}09, GH06, GGO01, HB09, KMV05, LBC^{+}04, LPL01, MBS04, PSB^{+}03, SARC09, SB05, TGW04, WKPS05]. Assessments [ARC^{+}08]. Assignment [MFOA09, SS08a]. Associated [BMN04, BLS^{+}00, BAWT05, BTLD08, CMKL01, DG04, FRK01, HSRD04, HKCL07, MHHS01, MM02, RMRT03, RE08, SJRE00, SRB00, SSHB02, WLBBM09]. Association [CHH^{+}06].

Associations [BAS^{+}09, BPR06, GMCQ08, OR06, TMK03, VHF05, WC00]. Assumptions [HOA04, WA05]. Asymmetric [RTG00]. Atchafalaya [FRK01, RGK01, TRK07]. Atherina [WC03]. Atherinid [WC03]. Atlantic [HWW^{+}07, LT01a, Rot07, RH06a, ASS00, AG02, ACB02, BNB^{+}07, BH06b, CHH^{+}06, CSPP00, CC03, CGM03, CMSM01, CB06, DRMH^{+}02, DCCA01, EKSX08, GMGV00, HMD02, HL00a, HAS06, HSBC02, KBW^{+}00, KHKS00, LLVB04, LS04, LKM04, LMK04, LT01a, LT01b, Leg05, LGJ02, MCS08, MGJ^{+}01, MHK^{+}01, MOT04, MCS06, NKJ03, NFP00, NSF04, PR01a, PLO07, PSD^{+}08, PP02, RSC07, Rot07, RH06b, SP03, SKC^{+}05, SLC06, SKL^{+}03, SFS04, SEP^{+}08, SLP^{+}01, SBH00, VGA01, WNF09, WSLT07, WWR^{+}00, WKM^{+}07]. Attacks [BSO01]. Attempts [YT00]. Attributes [BG02, BWS08]. Auke [BAWT05, MWMT02]. auroguttatus [PACH01]. Australia [WS09, XM00]. Autumn [CB02a]. Availability [CN05b, EM04b, GDWC04, HCK09b, HWHP04, JD07]. Average [SBSO06]. Avian [CRC^{+}02, RSBF03].

Bacillus [JHS02]. Back [HAS06]. Back-Calculation [HAS06]. Backpack [MMP03]. Backscattering [MW06]. Backwater [CG08, GSSB03]. Bacteria [HHB02]. Baffles [WCM^{+}07]. Bahamas [GdMH01]. bairdi [SGS05]. Balancing [POH^{+}09]. Banded [KGS07]. Bank [Sto02, BB00]. Banks [MR00]. Barataria [NTW02]. Barge [CLSD00]. Barging [HWS^{+}09]. Barotrauma [BCW^{+}09, PMRH06]. Barriers [CS06, POH^{+}09, SS06, SML^{+}07]. Based [ARC^{+}08, BE00, BB00, BBRH06, BWS08, CPB07, CKB01, CSBM06, DH04, GMJ^{+}06, HR09, JR03, KAB09, LFG06, LPN01, MBS04, NN06, NAG07, Pap08, PCF06, PP02, PBT04, PSJ08, SWH^{+}01, SBM00, TRGV00, iTE08, TR07, WGW06, WWS03, WWM07, WSLT07, WHC04, WWR^{+}00, ZTP^{+}08]. Baselines [SS08a]. Basin [AZS07, BHW^{+}07, BBWW02, FRK01, GS06, HI08, MM02, MJH^{+}09, PT05, PPDV09, RL08, RGK01, TWF^{+}07, TRK07, WEB00, CCMP06, CMBB01, CST^{+}05, DLES06, GBBP01, KA01, McK05, NPT04, PWLL05, ROC^{+}03, RIA^{+}07, TM06, WC00]. Basin-Scale [PT05]. Basins [SJ06, WMH^{+}04, SS06]. Basinwide [CR07]. Basis [BCM^{+}05]. Bass [ASN^{+}05, ASK01, BH00b, Bet05, BIT00, BC02, COB^{+}03, DF08, DWBE00, DCHS09, DS06, DSD07, FP04, FP06, FGWS00, FCS^{+}03, GWSF00, HC05, HGR^{+}08, Har00, HN05, Hei08, HK04a, HJP01, HSC00, IDW03, JN00a, JN00b, JRM^{+}08, LHOP01, LCNB^{+}06, MG06, MKBF00, MVS07, MSET05, NN06, NNL06, NC05, NAG07, OYB03, OY03, OMM09, PR01a, Pea04, PCC^{+}09, PLD00, PACH01, RCB^{+}04, RA09, RTBH05, RMD08, Sav04, SIH02, SW08a, SBB05, SCM01, SW08b, SW09a, SCWP09, SWD04, SHWS05, SMS04, SGG09, SSL^{+}03, SP04, SKC^{+}04, SN01, TAB00, TRBH08, VB02, VBT08, WJB06, WBI00, WNH05, WJK07, WSLT07, WK05, WA03, Whi00, WHZR03, WP08, WS07a, WSEW09, WNG02, YMJ00, YI02, YI07]. Basses [IMM^{+}04, SGSS01]. Bay [CMSM01, DW08a, DSSA01, GJP09, HSS07, HHS04, MOP02, MCR05, MMCR08, MM08, NTW02, NBK01, RWB^{+}08, RMD08, WLBBM09, WS08, CMMM04, Far01, FSH07, GLHA06, Har00, HLH^{+}07, JM05, OMM09, PSJ08, SC02, SLC00]. Bayesian [HB07, HBJ^{+}08a, WPGH08]. Bays [CM07, EKSX08]. Be [BEV^{+}07, DW08b]. Bear [BQH04, RFOR03, RWL01]. Bearing [BCW^{+}09]. Bears [QB01]. Beaver [MEA^{+}09, SLC06, HK00]. Beavers [WR08]. Bed [MWN05, MSM^{+}08, VB02]. Beds [CM03b]. Beetle [URMH07]. Before [DHR07, WSY^{+}07, CLSD00, DC02, HBTT06]. Before-and-After [DHR07]. Behavior [AD08, AAP03, AVM05, BL05, BVL^{+}05, CW00b, DC02, DCHS09, GDWC04, GHL02, HM07, HL00b, KSCD02, KB06, MSM^{+}08, NBS07, OLWC06, RSC07, SGSS01, SSL^{+}03, SYC04, TRS05, WSSD03a, WSY^{+}07, WC03, YI07]. Behavioral [BK03, BNG08, FT08, GKB^{+}06, HGR^{+}08, HOA04, KGW05, SYC05, TBWD06, WB01]. below [JKWS06]. Benign [Cam04, Cam05, Qui05]. Benthic [BE07b, Gid02, HPM05, HW08b, KBJ08, MG03, Pol07, TAC08, TK01]. Benthopelagic [MBMO06]. Bering [MFFI09, RM04, SK08, TJNL08, vT05]. Bertalanffy [HB07]. Best [PLD00]. Better [CP03]. Between [BKHH07, KJL^{+}07, AVM05, BGGL08, BM09, BBC^{+}08, BWC^{+}07, BL02, BTLD08, CHH^{+}06, CPF06, CG08, DYWA^{+}04, DS06, EKSX08, FLLG00, GMGV00, GVS02, GF07, Gro06, HC05, HAS06, HHHK08, HK04b, HCW01, JN00b, KLA02, KLSA02, KR07, KSB^{+}06, KCY^{+}07, LRF^{+}08, LJE08, MGJ^{+}01, MWN05, MZB^{+}07, MCR05, MPP05, MMBG09, PF01b, PAS09, RFOR03, RH02, RV07, RTG00, RWB^{+}08, RHV03, SBR^{+}06, SGF03, SSRM08, SK07, TMK03, iTE08, WF05, WRC^{+}03, WK05, WA03, WMH^{+}04, WRGN07, ZV07]. Bias [SAT^{+}07]. Biased [KESB05]. Bigger [CP03]. Bighead [CD09, PPR06, SBG01b, SGF03]. Bile [RSBS09]. Billy [BV01]. Bimini [GdMH01]. Bimodality [GSS03]. Bioacoustics [RGP06]. Biochemistry [RSB^{+}08]. Biodiversity [Bur09, MC09, Rid04, WWKM07]. Bioeconomic [GGO01]. Bioelectrical [PLH^{+}08]. Bioenergetic [BBV00, CEW00, NFP00, PDP08, TB08]. Bioenergetics [BH06a, Bre08, BR02, CW04, CW08, CHP08, Ess03, HK08, HC08, HSS00, IDW03, KJM^{+}08, KRM03, KFB^{+}06, LRF^{+}08, MOP^{+}06, MC04, NN06, PP05, RMH^{+}08, SCB08, SWD04, VBT08, VBM^{+}09, WHZR03, WBH06]. Bioenergetics-Based [NN06]. Biointegrity [HHK04]. Biological [CSPP00, DRMH^{+}02, ESM^{+}03, GPUH07, Hit04, HB08b, MAW09, MWG^{+}08, MMH03, RJ09, RP07, Sec08, Van04, WA05]. Biology [Ano03a, HWM04, Nev04, Sch03a]. Biomass [CND03, FSN^{+}03, KAB^{+}05, PT05, SVW02, SJ00, WNF09]. Biomass-Dependent [SVW02]. Biomechanics [Bur09]. Biophysical [BAWT05]. Biosphere [MEA^{+}09]. Biota [WBH^{+}05]. Biotic [ASS00, BDD07, BJZH05, CW00a, GF07, HH07, LPN01, MB08b, MHK^{+}01, MWG^{+}08, MMBG09, Pir04, PHWS09, QH05, RMRT03, Sch00b, SHC01, SA01, WHP01, WHS^{+}07]. Biphenyls [OW09]. Bivalve [Nev04]. Black [BS06, CEB^{+}05, DA05, JHS02, PA01, Sav04, SA04b, SCM09, TAC08, WK05, SK05a]. Blackfoot [PPDV09]. Bladder [CGD05, RB05]. Bladders [BBS^{+}08]. Blasting [FWTS08]. Bloaters [EBC01, TF01]. Blood [WFF^{+}02]. Blue [DYWA^{+}04, DA00, EBS09, HBSA09, HLH^{+}07, JM05, LHL06, PG01, PSJ08]. Blue-Marlin-Like [PG01]. Blueback [NGC^{+}02, SLM07, TAM07, WSP05]. Bluefin [MH05]. Bluefish [CGM03, GLHA06, GA07, SBR^{+}06, SC02, TA06]. Bluegill [AHW03, ASN^{+}05, BE00, HK04a, HAW09, OADC^{+}08, PCHB01, SW03a, TP01]. Bluegills [MB05a, Mic06, PW02, SCH03b, SW08b, WSEW09]. Blueline [HWM04]. Bluff [MLB03]. Boat [BA02, BS07, Mea05, MK08]. Bodies [SARC09]. Body [BSO01, Bre08, HAS06, JRM^{+}08, KJL^{+}07, PSKH04, PLO07, SH00a, SHDR03, SWD04, SBH00, TRGV00]. Bomb [BCDF^{+}09, DFBC^{+}09, HPW08]. Bonavista [CMSM01]. Bonneville [FSR^{+}07, JR03, MB06, SRJ03, WR08]. Book [Ano02a, Ano08, Aps06, Bar01, Boe01, Boz01, BWW01, Bur09, Cas00, Cur01, Hea01, Hil00, New01, NH01, Rah00, Sut01]. Bootstrap [Zho02a]. Border [BHHS04]. Boreal [CN01, DT03, HBTT06, TPS^{+}03]. Bottleneck [CF07b]. Bottlenecks [SM08a]. Boulder [FSN^{+}03]. Boundaries [Ano03h]. Boundary [JJRP08]. Brassy [SFB03]. Brauner [Bur09]. Brazilian [GK06]. Brazos [DW09, WTSC00, ZWT05]. Breeding [KFB05, OBB^{+}00, SKP^{+}08]. Bridging [SLM07]. Brief [COB^{+}03, GN00]. Bristol [HSS07]. British [BSC^{+}04, HWMH02, MHMP05, VGA01, BCS^{+}01, BSM^{+}08, ERS^{+}06, HBJ^{+}08b, TMWG00]. Broadcast [WD08b]. Broadcast-Spawning [WD08b]. Broodstocks [MFOA09]. Brook [AFR00, AFR02, BK03, BLS07, BDD07, BHMK05, BR09b, CM03a, CSvdS02, DW08a, FB08, GMCQ08, GVS02, GF07, GHL02, HMD02, HC08, HSBC02, HB08b, IK00, KLA02, KC06, KM06a, KM06b, LDDB08, ML07, MZB^{+}07, MD05, MBS05, MM08, NL03, PLM05, Rid08, RV07, RC04, RDU02, SJFP08, SH01, SCH04b, SH08, URMH07, VHF05, WSJ^{+}05, ZTP^{+}08, ZN07a, ZN07b]. Brown [ANE06, BIB^{+}00, Car06, CN01, CWT06, DHC04, GMGV00, GSS03, GVS02, GSP01, GF07, HSS00, Heg02, LC03, MPVM00, MB06, MD05, NA04, PK02, RHV03, SJP05, SBP^{+}08, Spi01, WB09, ZN07a, ZN07b]. Brunswick [ACB02, KBPC07, KC07, RCB^{+}04, SLP^{+}01]. Bubble [MWM00, WSSD03b]. Bull [ACB08, Ano05, BV01, BLG01, BCV07, GHL02, HWLC02, HB08a, KA01, KLA02, MZB^{+}07, MK05, MTSS06, RMRT03, RIA^{+}07, RV07, SMZB01, WSRA06]. Bullhead [SJP05]. Bump [Ano03e]. Buoyancy [PMRH06, PAR01, Wei08, ZKE06]. Burbot [HF02, NS04, SMW06, WEB00]. Butterfly [NNL06]. Bycatch [KCK07]. Bypass [FSR^{+}07, JAJ^{+}00]. Bythotrephes [PRME01].

C [Bur09, WNT03]. C-Start [WNT03]. Cadmium [FSN^{+}03]. Cage [BR09a]. Cages [TF07]. Calcein [Bas04, LBR^{+}01]. Calcified [IC02, SC02]. Calculation [HAS06]. California [BD07, BH07, CSBM06, EM04b, GSWS04, HWN02, HBH^{+}08, Log01, Mat01, MMP02, MDE^{+}09, MLB03, MM01, PACH01, QM05, RL08, SBC^{+}09, SWF09, VT07, WF05, WSFJ05, WM07, ZEP09]. californiensis [BS03]. Call [GGG06]. Callinectes [JM05, PSJ08]. Can [BEV^{+}07, HH05, HH06, KM06a, KM06b, WJB06, YSB^{+}08]. Canada [BSC^{+}04, CBN^{+}09, ISH07, MWN05, MHMP05, PB03, PR01a, PRR03, RMD08, SKL^{+}03, TMWG00]. canadensis [MEA^{+}09]. Canadian [JT04, NSP03, PP01]. Canals [CR07, RR08]. Candidate [ASS00]. Cannibalism [STR^{+}07]. Cantabrian [LC03]. Canyon [PP05, PCF06]. Capacity [JT04, WSA00]. Cape [NBK01, HKCP09, MEA^{+}09]. Captive [HDH00, HWW^{+}06]. Captively [BVL^{+}05]. Captivity [HHHK08, MWS08]. Capture [ASK01, BA02, HM07, KYWC09, LDOS06, MHHS01, PAD01]. Capture-Related [LDOS06]. Captured [ZEP09]. Carbon [SBP^{+}08, SDSB08]. Carcass [WHC04, Zho02b]. Carcasses [HNH^{+}04, WHCC03]. Cardiac [SCM01]. Cardiovascular [COB^{+}03]. Cards [MB02]. Care [SSL^{+}03]. Carney [ZM06b]. Carolina [YI02, BIT00, CSPP00, CPRS02, CCP^{+}03, EBS09, GI06, HJP01, RTBH05, RWB^{+}08, SARC09, SS08b, WSP05]. carolinae [KGS07]. Carp [BS06, BGH06, BSW00, BGS^{+}04, CD09, DGC07, PP08, PPR06, PGW08, SBG01b, SGF03, WG05, WSEW09]. Carps [DGC08]. Cascade [MHM08]. Cascades [OR06, RHB^{+}06]. Case [Beh02, MMM07, PBR^{+}01, PKR02]. Cases [DWEL08]. Caspian [ARL^{+}05]. Castor [MEA^{+}09]. Catadromous [AWNT00]. Catastrophic [RDU02, RTH^{+}01, RB05]. Catch [ARC^{+}08, BHP^{+}08, HG07, LBVW03, Mir05a, Qui04, RWB^{+}08, SMS04, SSL^{+}03, iTE08]. Catch-and-Release [LBVW03, SMS04, SSL^{+}03]. Catchability [IHF08, PT03b, iTE08]. Catches [LAM04]. Catfish [BKCL05, DA00, Mic06, PKWR05, SJ04, SJ06, VR05]. Catskill [WEB09]. Caudal [STC^{+}09]. Caught [NS04, WP08, WS09]. Cause [CF07b]. Caused [KJM^{+}08]. Causes [GWM^{+}07]. Caution [DW08b]. Cautionary [MMM07]. Cech [Bur09]. Central [WM07, BL01, CRA05, CH00, CG08, HBH^{+}08, HCK09b, MHMP05, OS03, PLM05, PT03a, PT05, SBC^{+}09, Sch00b, URMH07, VT07, ZEP09]. Centrarchid [SCH04a]. Centrarchidae [KGS00]. Centrarchids [WJMM07]. Century [Ano03h, CW08]. Cephalopods [Bur09]. cerebralis [DCG07, dlHFB04, PPDV09, SBW00]. Chain [AWNT00, MB02]. Challenged [DCG07]. Chamber [MFC04]. Champlain [EM04a, SM07, WGW06, ZM06a]. Chance [OJPN02]. Change [CPF06, FSMH08, IHF08, QPG^{+}02, RQ02]. Changed [SCP^{+}05]. Changes [AFR02, BSLN08, BLS^{+}00, Bre08, CB04, GMJ^{+}06, GS06, LS07, LAM04, MHHS01, OW04, PP05, QK03, SJRE00, SIH02, SHW00, SSHB02, SFE01, TK01, WC05]. Changing [GBC^{+}05, MMP02, PJSL09, PT03b]. Channel [AFR00, BHHS04, DA00, HW04, HH03, LCM07, Mic06, PR01b, PR01c, PJSL09, SJ04, SJ06, VGA01]. Channels [RQ01]. Char [BRK02, GN00]. Characteristics [ACB08, AG02, BH00b, BL02, CDT^{+}02, CJFM05, CSPP00, GA03, HB08b, Lin01, MSE07, MJH^{+}09, PF01a, PR01b, PSD^{+}08, PT05, PCMS05, QHR04, RE08, RMD08, SBG01a, SB05, SW09b, WPS02, ZM05, ZM06a]. Characterization [LJC^{+}07, MBC00, MGM02, MDP^{+}06, MK05]. Characterize [MMP03]. Characterizing [FSN^{+}03, GRR05, Mea05, OM05, WMO03]. Charleston [Ano03e]. Charlotte [LM08a]. Chemical [BT04a, MSC08]. Chemicals [BH00c]. Chemistries [WRC^{+}03]. Chemistry [BCV07, BBS07, WRC^{+}03]. Chemosterilization [YBSL04]. Chesapeake [GLHA06, Har00, HLH^{+}07, JM05, OMM09, PSJ08, SC02, WS08]. Chile [CH00, MEA^{+}09, PAS09]. Chinese [BGH06]. Chinook [Ano06b, BV01, AZS07, ARC^{+}08, BCJ^{+}06, BJS^{+}06, BLS^{+}00, BL05, BGPL08, BM06, BBHD09, BEV^{+}07, BBC^{+}08, BKP^{+}04, BGM06, BCW^{+}09, BKHH07, CN08, CHH02, CN05a, CLSD00, CMBB01, CPG03, CST^{+}05, CG06, DBW09, EM04b, FNL^{+}08, GABC00, GAH^{+}06, GKB^{+}06, GJP^{+}05, HCK^{+}09a, HWS^{+}09, HBJ^{+}08a, HF09, HCK06, HCK09b, JDSS01, JCB^{+}05, KPB^{+}04, KPC08, KSCD02, KSB^{+}06, KSB^{+}08, KFB^{+}06, KB06, LBC^{+}04, LBS^{+}06, Lin01, MBC00, MBS04, MB04, MLB03, MMS00, MJH^{+}09, MMS^{+}06, NPT04, NSC07, NSVT08, OW09, PE01, PF01b, PF05, PAR01, PKM07, QPG^{+}02, QVPG04, ROC^{+}03, RTG00, Sch05, SSD^{+}06, SKP^{+}08, SB06, SSSS05, SSR06, SAT^{+}07, SYC04, SJS09, TMWG00, TBK^{+}09, TRGV00, TKH^{+}09, TFO^{+}09, VMS01, VDG02, VT07, VPQ04, WKCC08, WF05, WC05, WSJ06, WM05b, WM07, Zho02b]. Chionoecetes [FSMH08, SGS05]. Chipola [WA03]. Chloride [ASK01]. Choctawhatchee [FHP00]. Choice [BMH07]. Choose [CWT06]. Chromosome [BT04b]. Chronic [BWVN09, WCBS06]. Chronology [BBHD09, LB00]. Chub [Bro05, DG04, MVD^{+}00, MM05, PCF06, PP05, Sch02, SLGH08, WD08b]. Chum [FWG07a, FWG07b, KMV05, MG03, OSS^{+}04, PMP^{+}02, SS08a, TRS05]. Cichlid [Ano03b]. Cichlids [NNL06]. Circle [KCK07]. Circulatory [FCS^{+}03]. Cirrhinus [BGH06]. Cisco [YSB^{+}08]. Ciscoes [JJRP08, MJSY09, OA07]. Clair [WRC^{+}07]. Clam [Ano03a]. Clarity [DA05]. Clark [WSSD03a, WSSD03b]. Class [HNH^{+}04, JN00b, SAT^{+}07, WKCC08, WEB09, WS07b]. Classes [SCH03b, SW08b]. Classification [BLBH05, BWS08, CTT04, HBBG07, MW06, PJSL09, SZSS08, WFF^{+}02, ZTP^{+}08]. Classifying [WWS03]. Clear [PP08]. Climate [BSC^{+}04, DS06, FRB06, OLS04, RIA^{+}07, RMH^{+}08, RL08, SFE01]. Climates [RHM^{+}06]. Climatic [BSN04]. Clonal [RT04]. Close [SYC05, WSY^{+}07]. Closed [Sto02]. Clupeid [CN05b]. Clupeids [CND03]. Clustering [ZTP^{+}08]. Co [OY03]. Co-occurring [OY03]. Coast [BL01, CEB^{+}05, HH03, LT01b, TTW^{+}09, CGM03, HWM04, HWW^{+}07, NSP03, SBC^{+}09, Sch00a, SFS04]. Coastal [AW05, BBM02, BS03, BHL05, BBHD09, BCT04, BTLD08, CMKL01, CMSM01, CBMM08, DBW09, DHR07, GLHA06, GH07, HWN02, HI08, HBTT06, KM01, LMK04, LJC^{+}07, MPP05, PKWR05, PKR07, RL08, SGS06a, SGS06b, SLK01, SWF09, SSO09, WSLT07, WB01, WRGN07]. Coaster [DW08a, GMCQ08, HB08b]. Coasters [Rid08]. Cod [BNG08, CHH^{+}06, CC03, CMSM01, DCCA01, HL00a, NBK01, Rot07, RH06a, RH06b, SB03a, WKM^{+}07]. Coded [MWMT02, TFO^{+}09]. Coeur [MM02, WRC^{+}03]. Coherence [Rot07]. Coho [AKPQ08, BS03, BCS^{+}01, BMN04, BSLN08, BFC^{+}04, BDR01, BD07, BDLB04, BBDM05, BH07, BBWW02, BAWT05, CEB^{+}05, CBF^{+}06, CBN^{+}09, DBW09, EWB^{+}06, ECW^{+}09a, ECW^{+}09b, HL00b, HNH^{+}04, KSR01, LFA09, MAW09, MN07, MS03, MSM^{+}08, OSS^{+}04, QPG^{+}02, QVPG04, RSB^{+}08, RQ01, RMG^{+}06, SB08, TAD05, TBWD06, VFT02, VPQ04, WB08, Wei08]. Cohort [HO05, JN00a, SBR^{+}06, TA06]. Cohort-Specific [HO05]. Cohorts [CGM03, PLD00, PA01, SW03a, SBR^{+}06]. Cold [CF07b, CF07a, MBS03a]. Coldwater [HHK04, JMS02]. Coleoptera [URMH07]. Collapse [RRN^{+}08]. Collecting [ADP00, BS06]. Collection [LMAF08]. Colonial [CRC^{+}01]. Colonization [AKPQ08]. Colony [CRC^{+}02]. Colorado [GSSB03, MPK03, PP05, WJMM07, BBRH06, BHW^{+}07, CF07b, CF07a, HJM04, HKHC00, KLSA02, PCF06]. Coloration [PR01a]. Columbia [Ano06a, BSC^{+}04, CCK^{+}09, CRC^{+}02, ERS^{+}06, GBBP01, ISH07, MHMP05, MOSB02, PPWvdL08, SSD^{+}06, SS06, TMWG00, WE05, ARL^{+}05, BS03, BCS^{+}01, BSM^{+}08, BL05, BM06, BKP^{+}04, BGM06, CN05a, CRC^{+}01, EK08, ERS^{+}06, FACS06, FSR^{+}07, GKB^{+}06, HBJ^{+}08b, HWMH02, HPB06, JCB^{+}05, KPB^{+}04, KPC08, KMSL02, NPT04, ROC^{+}03, RIA^{+}07, RQ02, RSBF03, TRS05, VDG02, VGA01]. Combahee [BIT00]. Combination [BBL06]. Combined [BH06a]. Combining [PJH04]. Comment [Beh02, Cam05, EBC01, Esh09, GRR05, HE01, OM05, PKR02, Qui05, SHC01, TF01, WDHW09, WSJH01, WHP01]. Comments [HL00a, HHHK08]. Commercial [Chi09, GVHC^{+}09]. Common [BGS^{+}04, HGR^{+}08, LBVW03, PP08, PGW08, WSEW09]. Communities [BL00, LJE08, MG03, Sch00b]. Community [ASN^{+}05, CB04, GMCQ08, LJC^{+}07, MB08a, MMP03, MM03, PKR07, RRN^{+}08, TK01, WWS03, WJK07]. Comparative [BMZ07]. Compared [BGH06, BR09a, RSP09]. Comparison [BK03, Bas04, BWC^{+}07, CWB06, CW00a, CGAD^{+}02, ERS^{+}06, Ess03, FNL^{+}08, dlHFB04, GFL00, HK04a, IC02, JD07, KRM03, KSB^{+}06, KSB^{+}08, KSR01, LBS^{+}06, Lin01, LFG06, NSP03, RWB^{+}08, SC02, SAT^{+}07, TAM07, VAR^{+}02, WNT03, ZM06a, Zho02a]. Comparisons [Gid02, GM04, LRF^{+}08, MN07, PF01a, PMP05]. Compatibility [NNL06]. Compensation [IRJ^{+}09, JT04, PAR01]. Compensatory [WBH06]. Competition [BGGL08, BM09, Cam04, Cam05, FLLG00, ML07, MZB^{+}07, Qui05, RV07, RWL01, WF05]. Competitive [SGF03, SCWP09, TAD05]. Complementary [FHB^{+}07]. Complementation [WR08]. Complex [BCS^{+}01, BLC^{+}04, BCM^{+}05, KLSA02, KGS07, RA07, RJ09]. Complexity [BNB^{+}07, HKHC00, SP07]. Components [KAB09, SCWP09, SBSO06].

Composite [JD07]. Composition [BL00, BCM^{+}05, BCJ^{+}06, Bre08, BZW04, CRC^{+}02, GLHA06, HNH^{+}04, HDH00, HPB05, JRM^{+}08, KJL^{+}07, LFG06, SHDR03, TBK^{+}09, TAC08, WTJ00]. Concentration [CD09, CDF^{+}08, EH03, WCC08]. Concentrations [CG04, DWEL08, FSN^{+}03, HBP^{+}09, MH02]. Concepts [Wil07]. Concordance [DS06]. Condition [BH06a, BWS^{+}05, BH04, BG04, BWFB06, ESM^{+}03, FSMH08, HBJ^{+}08a, HK04b, KFT^{+}09, KJL^{+}07, MHD03, MWG^{+}08, PSKH04, SLC06, SHDR03, SB03a, SGG09, SBH00, TPS^{+}03]. Conditional [XM00]. Conditions [Ano06b, FNL^{+}08, GJP^{+}05, HW08b, JN00b, KTNK01, LFA09, MJS^{+}03, MPP05, NA04, SH00a, VLD^{+}06, WSFJ05, WH07]. Conductivity [BHW03, Kol06]. Confinement [SZM01]. Connecticut [SP03]. Connecting [LCM07]. Connectivity [BDD07, BBW03, Mir05b, NDR^{+}09, Sch02, SS06, ZWT05]. Consequences [CGD05, DSD07, MSET05, OBW04, TBWD06, VT06]. Conservation [Bur09, FB08, HBTT06, MFOA09, MVD^{+}00, PK04, PWLL05, Rid08, Sav04, Sch02, TCG09]. Conserving [GGG06]. Consistent [HH07]. Constant [JR03]. Constituents [DS05]. Constraining [QH05]. Constraints [CW08, HGW04]. Constructed [RMG^{+}06]. Construction [KJM^{+}08]. Consumption [BBV00, BBHD09, BW02, CEW00, EBS02, GW03, GBC^{+}05, HK04a, LRF^{+}08, RWL01, SKK^{+}08, SH08, TGW04, TW05]. Consumptive [DHS06]. Contact [HMD02, WSY^{+}07]. Contemporary [GFLF02, PBT04]. Content [CPB07, CH05]. Contents [SH00a]. Continental [DGK^{+}08, TL06, TFO^{+}09]. Contingent [WS07a]. Contrasting [MGP06, NA04, WHC04]. Contribute [GVS02, KMP03]. Contributing [KSC^{+}06, KM01]. Contribution [CGM03, FLLA09, SH08, YMJ00]. Contributions [Ano03d]. Control [BH00c, GFL00, IDW03, POH^{+}09, SMW06, WWR^{+}00]. Controlled [VGA01]. Convenient [BSW00]. Conventional [FNL^{+}08]. Cook [SHT^{+}00]. Coolwater [HPB06]. Cooper [CCP^{+}03]. Copepoda [BS03]. Copper [FSN^{+}03, HWLC02, WRGN07]. Coregonine [BBS07]. Coregonus [EBS02]. Cormorant [VAR^{+}02]. Correlation [PPDV09]. Correlations [JM09]. Correspondence [DS05]. Corridors [RA07]. Cortisol [FHML08]. Cost [BHB^{+}07, Pol07]. Costocking [IK00]. Costs [AVM05, SHBFE07]. Cottus [KGS07, Nat07]. Count [WPGH08]. Countervailing [QPG^{+}02]. Counts [MTSS06]. Coupled [LFG06, RMH^{+}08]. Coupling [MBMO06]. Course [OADC^{+}08, PLO07]. Courtship [SGSS01]. Covariation [PMP^{+}02]. Cover [GP05, KB06, RA07, WLK03]. Cover-Seeking [KB06]. Crab [EBS09, FSMH08, GJP09, HB04, HBSA09, HLH^{+}07, JM05, KESB05, SGS05, vT05]. Crabs [DYWA^{+}04, LHL06, PSJ08, SGS06a, SGS06b, SGS05]. Crappie [BMH07, BSS05, DA05, DD02, PA01]. Crappies [MBS03a, SBG01a, SA04b, TAC08]. Crayfish [BE07b, Sch03a]. Crayfishes [AW05]. Creek [BD07, BDD07, BAWT05, Car06, HK00, MS03, MWMT02, SLGH08, WW02, ZM05]. Crevice [BJ01, Sut07]. Crevice-Spawning [BJ01, Sut07]. Crisis [WWKM07]. Criteria [MB04, Pea04, SRJ03, WA05].

Critical [AAP03, CM01, HF02, LPJS08, Pea04]. Croakers [LT01a, LT01b, RGP06]. Crude [WHT^{+}00]. Crustaceans [BJ03, MMCR08]. Cryopreservation [CDFW06]. Cryptic [DGK^{+}08]. Crystal [VPN09]. Cue [DGFS04]. Cui [SRB00, SR07]. Cui-ui [SRB00, SR07]. Cull [HBSA09]. Culture [Nev04]. Cultured [SBB05]. Culvert [MSM^{+}08, NHFF09, Pea04]. Culverts [LDDB08]. Cumberland [TH00]. Cunner [NBK01]. Current [HKCL07, TGW04]. Cutthroat [BBM02, BBG02, BS03, BHL05, BMZ07, BPM^{+}04, CDP02, CCMP06, CF07b, CF07a, DHR07, DCG07, DB08, GH07, HF04, HK00, HK04b, HKHC00, JR03, KLSA02, KCY^{+}07, ML07, MB06, MSEL03, MSL^{+}06, MMTF05, MMBG09, NS02, NQ02, OR06, PK04, PJC07, QH05, SRJ03, WPS02, WRC^{+}03, WB01, WR08, WRGN07, ZBF05]. Cycling [JR03]. Cyprinid [BHSR09, DW09, ME07, WD08a, WD08b]. Cyprinids [GGG06, QHR04]. Cytometric [BSW00]. Cytonuclear [OR06].

D.C. [SJP05]. Dace [ZG07]. Dactyl [FSMH08]. Daily [BBHD09, Har00, HHS04, PKM07, TMK03, WD08a]. Dakota [BW01, FS06, JKWS06]. d’Alene [MM02, WRC^{+}03]. Dam [BIB04, BH00b, BGM06, BH06b, EBFC04, JKWS06, KPB^{+}04, FSR^{+}07, JAJ^{+}00, MJH^{+}09]. Damaged [CTT04]. Dammed [GK06]. Dams [Ano06a, BKP^{+}04, FACS06, MOSB02, TGWE04, WE05, ZDK^{+}04]. Danios [CB05]. Darter [LAM04]. Darwin [Tur05]. Data [BBL06, FSMH08, GH06, HB09, HH05, HH06, HMT^{+}06, IMM^{+}04, JTPW09, KAB^{+}05, NBK01, OLWC06, PG01, STC^{+}09, SA04a, TAM07, WRC^{+}07, WPGH08, XM00]. Data-Limited [SA04a]. Dates [HRMC07]. Day [NSVT08]. Daytime [BMH06]. Dead [CTT04, RBR^{+}00]. Dean [HWMH02]. Debris [RDU02]. Decade [CRMN04]. Decades [AFR02, RQ02]. Decapod [BJ03, MMCR08]. Deciphering [BR09a]. Decline [MPK03, PBT04, SLSF07]. Declining [TCG09]. Decommissioning [MSE07]. Decompression [BCW^{+}09, MSET05, NS04, RB05]. Decreases [SCP^{+}05]. Deep [DGK^{+}08]. Deep-Sea [DGK^{+}08]. Deepwater [DGK^{+}08, RRN^{+}08]. Defined [DCT02]. Defines [HSS07]. Defining [RHB^{+}06]. Degradation [OSS^{+}04, Pir04]. Delaware [EKSX08]. Delay [SA06]. Delayed [FACS06, IRJ^{+}09, MMS^{+}06, SSD^{+}06]. Delineation [MB01b]. Delivery [CC03]. Delta [CBR^{+}08, HBBG07, WSY^{+}07, WRGN07, Zho02a]. Demand [BBV00, BSM^{+}04, DHS06, RMD07]. Demersal [RRN^{+}08]. Demographic [ACB08, DSD07, GA03, JSH^{+}08, PACH01]. Demographics [MLSD07, MSEL03]. Densities [SBG01a]. Density [Bre08, BNB^{+}07, BR09a, BL02, CPB07, DGC07, DA05, GF07, HKCP09, HPB05, HKHC00, IRJ^{+}09, KWH03, KYWC09, LC03, MGMM01, MD05, OBW04, PT03a, PT03b, PT05, PLH^{+}08, RSP09, SGS06a, SBG01b, SBSO06, SGG09, WNF09, WJH06, ZM05, SGS06b]. Density-Dependent [IRJ^{+}09, MGMM01]. Dentaries [SRL^{+}06]. Dependence [HS08]. Dependent [BHP^{+}08, BH07, IRJ^{+}09, MGMM01, PG07, RH02, SBC^{+}09, SVW02, SN01, Zho02b]. Depensatory [IRJ^{+}09]. Depletion [GdMH01].

Deployment [SJRE00]. Deposition [CJFM05, OA07, SCP^{+}05]. Depth [BCW^{+}09, GSP01, HGW04, MB08a, OEO^{+}00, PV04, ZKE06]. Depth-Acclimated [BCW^{+}09]. Depths [BM06, JCB^{+}05]. Derivative [MGJ^{+}01]. Derived [BCT04, HMT^{+}06, SRJ03, TFO^{+}09, Zho02a]. Deschutes [ZR02]. Describe [WRC^{+}03]. Describing [HPW08]. Description [AMRM02]. Design [GJP09, HB04]. Designing [AAG03]. Despite [NSVT08]. Detect [Sch00b, WPS02, WNG02]. Detectable [Bas04]. Detected [BEV^{+}07, DFCDR07]. Detecting [SGS05]. Detection [CDP02, DRG09, GA07, HI08, KBPC07, LBR^{+}01, LMA09, SKK^{+}08]. Determination [BGS^{+}04, Pea04, PSJ08, SWHW06, SRH07, TAB00]. Determine [BSM^{+}08, BCV07, CBN^{+}09, LBBF^{+}08, SSR06]. Determined [BMM^{+}06, BJS^{+}06, BVMM06, TA06]. Determining [Ano05, CW00a, SMZB01, SA01, WB09]. Determinism [ML04]. Deterministic [HPW08]. Detroit [MC00a, MC00b, LCM07]. Developing [PDP08, WHS^{+}07]. Development [BOMC07, BMZ07, BHC03, BJZH05, CW04, CHP08, ESM^{+}03, Ess03, GAH^{+}06, HSS00, HDH00, HGM03, LBS^{+}06, LB00, LPN01, MHK^{+}01, MBS05, PP05, PJSL09, Pir04, RP03, RT04, SHWS05, TW05, WHZR03, vT05]. Device [ADP00]. Did [NBS07]. Diel [BBG02, BBHD09, GI06, LM08a, NQ02, OS03, PPWvdL08, PW02, SCH03b, SCH04a, TRS05, YI07]. Diet [BGH06, CPA^{+}06, CRC^{+}02, DRQ09, GLHA06, HHS04, HPB05, KFB^{+}06, MBMO06, ME07, OY03, OMM09, PKWR05, PV04, RTBH05, SVW02, SHS04, SGG09, TWF^{+}07, TAC08, VT06, WA03, ZV07]. Diets [BH03, DBW09, LPL01, SH00b, WG05]. Differ [BHL05, DS06, GMS^{+}06]. Difference [SJ06]. Differences [BTLD08, BKHH07, CDF^{+}08, DYWA^{+}04, GW03, HK04b, ZKE06]. Different [BL01, CRC^{+}01, GW03, GMP^{+}05, LT01b, MPVM00, MKDB03, OM02, PR01a, PAR01, Ros03a, SCM01, SJ04, SCWP09, WH07]. Differential [BBC^{+}08, PA01]. Differentiated [BFM08]. Differentiates [NPT04]. Differentiation [CN01, HBJ^{+}08b, KESB05, MGJ^{+}01, MOT04]. Differing [OBB^{+}09]. Difficult [HB00]. Digestion [BGH06, CB04, SH00b, WK05]. Dimensions [BHMK05]. Diploid [CGAD^{+}02, WSA00]. Diporeia [HPB05]. Direct [AHW03, HLH^{+}07, MMM07]. Discarded [HMT^{+}06]. Discharge [BMH06, DC02, DW06, KPC08, RHM^{+}06, Spi01, WSJ^{+}05]. Discovery [SSCM06]. Discrete [BMH07, KYWC09]. Discriminant [MW06]. Discriminate [VT07]. Discriminates [ROC^{+}03]. Discrimination [AWNT00, PLO07]. Disease [MMS00, VMS01, WSSD03b]. Disentangling [IRJ^{+}09]. Dispersal [BKCL05, CMBB01, KESB05, RBR^{+}00]. Displacement [ML07]. Disproportionate [URMH07]. Dissolved [BM06, CG04, CD09, GAH^{+}06, JCB^{+}05, KA07, MH02, OW01, PP02, SBP^{+}08, WSSD03a]. Dissolved-Solid [CD09]. Distance [AAG03, HH07, KVD^{+}09, MOT04, Mea05, SH01]. Distinct [GMP^{+}05, GMS^{+}06]. Distinction [WK05, WWKM07]. Distinctions [MFOA09]. Distinctive [NPT04]. Distinguish [BBM02]. Distribution [AFR02, AHJ^{+}05, AKPQ08, BFM08, BBG02, BAS^{+}09, BH00b, BL05, BH06b, CPF06, CPG03, DG04, FRB06, FRK01, dlHFB04, HJM04, HPM05, Hol08, JM05, JTM08, KBJ08, MB08a, MFFI09, OW09, PCF06, PP01, PP08, RM04, RSB^{+}08, RB07, Ros03b, SFS04, SJ00, TFO^{+}09, WSRA06, WAS^{+}00, YW07]. Distributions [AAG03, BR09a, CJMB06, KC05, LHM02, Mir07, OEO^{+}00, OJPN02, SZSS08, SCM09, YA01, ZSW02]. Disturbance [KSC^{+}06, PF00, VT06]. Disturbance-Prone [VT06]. Disturbances [DT03]. Disturbed [ML04]. Divergence [HSS07, KGW05]. Diverse [IH04]. Diversion [SYC04]. Diversity [BRK02, DWBE00, DK06, HBTT06, McK05, MTWS09, MM05, NDR^{+}09, PSBC04, PJC07, SHT^{+}00, SW09b, TM06, WM07, ZKE06]. Division [GSSB03]. DNA [BVMM06, DFCDR07, DWBQ05, Esh09, FHB^{+}07, FWG07a, GFL00, GFLF02, GBN05, HMD02, HG00, HM08, KA01, KGS07, LYS06, MCS08, Mil03, PWDG07, SWH^{+}01, SB03a, SBM00, WGRW04, WGW06, WDHW09, WM07, WWR^{+}00, WKM^{+}07]. DNA-Restriction [DFCDR07]. Do [CWT06, Cou07, EBS09, GMS^{+}06, HJM04, Mic06, PF05, VO01, WG03]. Documentation [GKW^{+}06]. Documented [Beh02, PBR^{+}01, PKR02]. Does [LBVW03, TBA^{+}04]. Dolly [BBWW02, DRQ09, OPH09]. Domain [AWNT00]. Domestic [CM01]. Domesticated [TBWD06]. Donor [MGJ^{+}01]. Dorsal [PR01a]. Downstream [LDDB08, WE05, Ano06a]. Drainage [AFR02, BBWW02, BBS07, BH06b, CN01, GA03, HSRD04, HBP^{+}09, KMSL02, KCY^{+}07, MTWS09, MK05, OR06, PK04, QHR04, RR08, SML^{+}07]. Drainages [DRMH^{+}02, WBH^{+}05]. Draining [CGL^{+}09]. Dreissenids [OEO^{+}00]. Drift [HWN02, HSS00, ZG07]. Drift-Feeding [HSS00, ZG07]. Drifting [ADP00, CMKL01, SH00a]. Drought [AW05, HW08a, WR08]. Drum [BG04, DFBC^{+}09, FWL^{+}07, GGCF07, GMS^{+}08, Hol08, LBBF^{+}08, RBM06, Sch00a, SS08b]. Drums [RGP06]. Dry [HNW06]. Dry-Season [HNW06]. Due [vT05]. Dummy [HCK^{+}09a, LKM04, MH05, PAR01]. duorarum [DSSA01]. During [CBMM08, RH06a, WSY^{+}07, AKPQ08, Ano06b, BFM08, CH00, CPG03, DC02, GJP^{+}05, Gro06, HGM03, IRJ^{+}09, KSC^{+}06, Log01, MA09, PD05, PW09, RH06b, SBP^{+}08, SZM01, SH06, SMS04, SEP^{+}08, SKC^{+}04, VR05, Wei08]. Dwelling [DC02, HR09, LGJ02, MHMP05, NA04, PTG04, TGWE04, TRK07]. Dynamic [Bre08, FSH07]. Dynamics [BSM^{+}04, BLS^{+}00, BC07, BBW03, DW09, HOA04, HSC00, JN00a, Kan01, LC03, MB08a, MVS07, PBI05, PLM05, RB07, Rot07, SARC09, SBR^{+}06, SLSF07, Spi01, TA06, VBT08, YSB^{+}08, ZN07a, ZN07b].

Early [Ano03d, BMN04, BSM^{+}04, BWVN09, CH00, CBMM08, DD02, GBBP01, IRJ^{+}09, LBS^{+}06, MBS05, PLO07, PLD00, RCP00, SBG01a, VPQ04, WSP05, WEB00, WB08, WB09]. East [NSP03, CRA05, CG08].

East-Central [CRA05, CG08]. Eastern [CCR^{+}07, CPF06, CM07, DF08, GFLF02, Kan01, MC09, MFFI09, PRME01, PP01, SKC^{+}05, SK08, TL06, ZTP^{+}08]. Eating [SMK^{+}07]. Eco [DSD07]. Eco-Genetic [DSD07]. Ecological [Bro04, CGD05, HSS07, NNL06, RE08, SMK^{+}07]. Ecology [AS07, Ano02d, Ano03f, BSLN08, BR09b, Bur09, CSL09, FB08, KC07, Nev04, NTB^{+}06, Sav04, SW03b, Tur05, WM05a]. Economic [Bro04]. Ecoregion [BWS^{+}05]. Ecosystem [BSN04, LJE08, SVW02, Van04]. Ecosystems [Lon04, Mil04, Rid04, SYC05, WHCC03]. Ecotone [MS03]. Ecotype [MFOA09]. Edge [AHJ^{+}05]. edited [Bur09]. edwardsii [IHF08]. Eel [GA03, Jes00, MLSD07, OM02, WMH^{+}04]. Eels [AWNT00, JSI09, MS04]. Effect [ACB02, BGPL08, BMH06, BR09a, CS06, CD09, DCCA01, DWBE00, EH03, GSP01, GHL02, HB09, HN00, HPB06, HCK06, JM09, KFT^{+}09, KAB09, KZ06, OLS04, SMZB01, SW08b, SA04b, SSL^{+}03, WCC08, WS09, YBSL04, Ano05, PGW08]. Effective [DWBE00, GMS^{+}08, VFT02]. Effectiveness [JT04, PLO07, RBR^{+}00]. Effects [AHW03, AAG03, ASK01, AW02, AVM05, BCT^{+}05, BE00, BH00b, BVL^{+}05, BHSR09, BBDM05, BW02, BBW03, BH07, BHMK05, BNB^{+}07, BSS05, BS07, BJ01, CM03a, Car04, CN01, CEW00, CW00a, CCK^{+}09, DHR07, DCG07, FWTS08, FCC07, FWL^{+}07, FRK01, FLLA09, FP06, GP05, GJP^{+}05, HW08a, HCK^{+}09a, HWS^{+}09, HWLC02, HC05, HDRB07, Hei08, HNH^{+}04, HB00, HKHC00, HAW09, JHS02, Kan01, KPC08, KSCD02, KBW^{+}00, KA07, KYWC09, KC09, KB06, LS07, LKM04, LGJ02, LC03, LM08b, MAW09, MD07, MC04, MSE07, MB06, MM03, MBS05, MM01, MG03, MJSY09, NHFF09, PCHB01, PKR07, PG01, PMP05, QH05, RHM^{+}06, RIA^{+}07, RSC07, RV07, RQ01, RMH^{+}08, REK06, RB05, RGK01, SA06, SW03a, SBR^{+}06, SVW02, SBW00, Sch00b, SR07, SB06, SCH03b, SW09a, SHW00, SHDR03]. Effects [SWD04, SBC^{+}05, SGG09, Sut07, SJS09, TRS09, TGWE04, TPS^{+}03, TRBH08, TAC08, TR07, VB02, VMS01, VPQ04, WSB00, WBI00, WMBM02, WCM^{+}07, WJF03, WBH^{+}05, WBH04, WS08, WHC04, WSEW09, ZBF05, ZJPB08, Ano06b]. Efficacy [MMP03]. Efficiencies [HDL07]. Efficiency [AJF09, BH00a, BA02, Chi09, MOSB02, SJ08]. Effort [RWB^{+}08, SJ08, iTE08]. Efforts [SJFP08]. Egg [CD09, CHH02, CJFM05, JM09, KTNK01, QVPG04, WSJ^{+}05]. Eggs [ADP00, CTT04, FWTS08, RTH^{+}01, WC03]. Electric [BNG08, HKCL07]. Electrical [Kol06]. Electrofisher [BA02]. Electrofishing [BS06, Car04, DM03, HKCL07, HH07, KVD^{+}09, Kol06, KYWC09, MMP03, MM03, Mea05, MD03, MK08, PTG04]. Electromyogram [BEV^{+}07, BGM06, GABC00]. Electroshock [BHSR09, HGJO04, VMS01]. Electroshock-Induced [HGJO04]. Electroshocking [CHH02, HGM03]. Electroshocking-Induced [HGM03]. Element [BBS07, SW08a, WTJ00]. Elemental [BCT04, HBBG07, PCWC08]. Elements [AHC^{+}09]. Elevated [ANC^{+}06, FSN^{+}03, MH02]. Elevation [AHJ^{+}05, BLGW06, CF07a, MZB^{+}07]. Elvers [Jes00]. Embryo [CBM^{+}00].

Embryonic [RT04, WHT^{+}00]. Embryos [BHSR09, DKP^{+}06, GAH^{+}06, WNT03]. Emerged [CB02b]. Emergence [BGPL08]. Emigration [HDL07]. Emphasis [AMRM02, GSWS04, SCM09, SLC00, WTW^{+}00]. Empirical [BOMC07, CPB07, LPL01]. Enclosures [RHV03, WF05]. Encounter [KGW05]. Encounters [SYC05, WSY^{+}07]. Encyclopedia [Tur05]. Endangered [BHSR09, HKCP09, MD07, MVD^{+}00, PBI05, PF05, PDP08, STC^{+}09, Sch02, SLSF07, SYC05, WG03, WS07b]. Endemic [DW09, RWL01]. Endocrine [BFC^{+}04]. Endoscopy [SRH07]. Energetic [GMJ^{+}06, MFFI09, PF02]. Energetics [Pap08, RHM^{+}06]. Energy [Bre08, CPB07, CH05, FWL^{+}07, GABC00, HPB05, HSC00, Log01, PSKH04, PLH^{+}08, RSP09, SHBFE07, SLM07, WLB^{+}00, WJH06]. Engineered [SB08]. Engraulis [CH00]. Enhanced [BDLB04, TAD05]. Enhancement [DYWA^{+}04, GMS^{+}08, Sch00a, WBH04]. Enhancing [Bre08, JT04, POH^{+}09]. Enriched [TRS09]. Enrichment [EH03]. Entrainment [CR07, GDW03, HDRB07]. Entry [AHC^{+}09, DGFS04, JTPW09]. Enumeration [EBFC04]. Environment [FSH07, FLLA09, JM09, KPC08, NDČ^{+}04, TRS09, TBA^{+}04]. Environmental [ANE06, Ano06b, BL02, FRK01, dlHFB04, GJP^{+}05, HPM05, HAW09, HGW04, LJE08, MAW09, MM02, MISS07, MJS^{+}03, PCMS05, PMP05, QPG^{+}02, RCP00, RQ02, TL06, WC00, WBH^{+}05, WTSC00, WB09, ZWT05]. Environments [BC02, Heg02, HK04a, Wei08]. Episodes [KS01]. Episodically [BLS07]. Erie [WRC^{+}07, MKBF00, PRME01, RTH^{+}01, SMW06, TK01]. Errata [Ano01, Ano02b, Ano03c, Ano04]. Erratum [Ano05, Ano06a, Ano06b, HH06, HSH^{+}05, KM06a, RH06a, SGS06b, SMZB05]. Error [JTPW09, MTSS06]. Errors [HB09]. Escape [KB06, WG03]. Escaped [LS04]. Escapement [BKP^{+}04, SA04a]. Esocid [SCB08]. Established [DK06, MB05b, WG05]. Establishment [OEO^{+}00, SSHB02, SGG09]. Estimate [BP01, CBM^{+}00, GABC00, HJP01, HH07, KAB^{+}05, Mir07, NHFF09, PHF07, PJH04, YSC^{+}06]. Estimated [AZS07, EKS^{+}05, GdMH01, KYWC09, Log01, PYFS07]. Estimates [BKP^{+}04, BR09a, CCK^{+}09, FSN^{+}03, HHDC03, Har00, HLH^{+}07, Jes00, KBPC07, LRF^{+}08, Mea05, MMCR08, NDČ^{+}04, PB03, SSD^{+}06, VFT02, WA05, WWM07, WSLT07]. Estimating [AHC^{+}09, Ano02c, BHP^{+}08, BHW03, BH07, CGM03, DSH^{+}09, EBS02, GH06, GM02, GDW03, HG07, HDL07, KVD^{+}09, MCS08, OLWC06, PTG04, PJSL09, PLH^{+}08, RHG^{+}02, SA04a, SM07, Tho03, WJH06, ZM06b, Zho02a, vT05]. Estimation [BCM^{+}05, BCJ^{+}06, CJMB06, CND03, GGCF07, JRM^{+}08, JJRP08, LHL06, MSC08, PG01, SRL^{+}06, SBSO06, Sto02, XM00]. Estimator [IHF08]. Estradiol [HBP^{+}09]. Estuaries [Sch00a, LT01a, LT01b]. Estuarine [BHP^{+}08, DGFS04, GA07, HBH^{+}08, KAB^{+}05, KHKS00, LH03, MMP02, RCP00, SBR^{+}06, SS08b]. Estuarine-Dependent [BHP^{+}08].

Estuary [Ano06b, BJ03, GJP^{+}05, HPM05, HBO07, MS03, QM05, SSD^{+}06, CRC^{+}01, HDRB07, MS04, NAG07, RB07, RSBF03]. Ethanol [LMAF08]. Ethanol-Free [LMAF08]. Ethylene [HBC^{+}02, SBC^{+}05]. Eurasian [BL02, FLLG00, SHBFE07]. European [NC05]. Evacuation [BBHD09, BBL06, HN00, PKM07, SCH04b, WK05]. Evaluate [CB02a, PE01]. Evaluating [BWC^{+}07, GWSF00, IDW03, KM05, OSS^{+}04, PR01b, Pir04, REK06, SRJ03, SWHW06, EM04b]. Evaluation [BBP^{+}04, BS06, BCS^{+}01, BFC^{+}04, BBL06, BJZH05, Bro04, Chi09, CW04, CHP08, Gro06, HL00a, JAJ^{+}00, JTPW09, KJM^{+}08, KM06a, KM06b, LMA09, MOP^{+}06, MB04, MWG^{+}08, MCS06, NNL06, NS04, PSB^{+}05, Pea04, PP05, PTG04, PJSL09, Ros03a, RP07, SBB05, SM08a, SP07, WGW06, WHZR03, YSC^{+}06]. Events [DC02, SCWP09]. Evidence [BMN04, BSS05, DGFS04, FACS06, GMGV00, HWMH02, KLA02, MBMO06, MKP07, MKS01, MBC^{+}05, PAS09, PR06, RFOR03, RBLF06, SBR^{+}06, SSD^{+}06, SLSF07, SMW06, WB01]. Evolution [Ano03b, FWG07a, FB08, MTWS09, Tur05]. Evolutionary [CSL09, DSD07, OPH09]. Examination [NSP03]. Examining [HAW09]. Example [SRJ03]. Examples [WHS^{+}07]. Excavation [MG03]. Exchange [BHW03, CG08]. Excretion [Gid02, PW09]. Exercise [BVL^{+}05, BC02, COB^{+}03, MBS03b, SCM01, WMBM02]. Exhaustive [COB^{+}03, MBS03b]. Existing [GSSB03]. Exotic [HCW01, MB05b, QHR04, VB02, WB09]. Expansion [OO07]. Expenditure [GABC00]. Experience [KMP03]. Experiment [Ano03b, CB02a, DHR07, KD01]. Experimental [LH03, MGG^{+}06, MB06, MSM^{+}08, YT00]. Experiments [BBRH06, FLLG00, GdMH01]. Explain [NN06, WJB06]. Explaining [OW01]. Explanatory [OJ01]. Explicit [BWS^{+}05, NFP00]. Exploitation [GVHC^{+}09, OLWC06, RMD08, TH00]. Exploited [KBJ08, SB05, SYC05]. Exploration [DD02]. Exploring [HR09, RA09, SK07]. Exposed [GAH^{+}06, HBC^{+}02]. Exposure [COB^{+}03, DS05, HWLC02, HKCL07, OADC^{+}08, SBC^{+}05, WWM07, WHT^{+}00, WC03, vT05]. Expression [BBC^{+}08]. Expulsion [MH05]. Extent [CPF06]. External [SB03b]. Extinction [OBB^{+}00]. Extirpation [KBW^{+}00, MMM07]. Extrapolation [KRM03]. Exxon [CRMN04, SHT^{+}00, WHT^{+}00].

Factor [CPG03]. Factors [ANE06, AHJ^{+}05, BHC03, BBRH06, BAWT05, CDT^{+}02, CW00a, CCK^{+}09, DG04, DD02, GBMB01, HW04, HB04, HI08, HEW^{+}05, KSC^{+}06, KC06, LDK^{+}04, MD05, MSEL03, MMBG09, PK02, PYFS07, QPG^{+}02, RP07, SP04, SBH00, TAB00, WB09, ZDK^{+}04]. Fall [BBC^{+}08, CN05a, CPG03, CST^{+}05, CG06, GABC00, GAH^{+}06, GKB^{+}06, MBC00, SH00a, SHW00, TRGV00, TKH^{+}09, VDG02, VR05, VT07, WEB09, WM07, WS08]. Fall- [BBC^{+}08, WEB09]. Fall-Run [CN05a, MBC00, VT07, WM07]. Fall-to-Winter [SHW00]. Fallback [BKP^{+}04]. Falls [RHB^{+}06]. Families [CHH^{+}06]. Family [GWM^{+}07, RGP06, URMH07]. Farfantepenaeus [DSSA01]. Farmed [LS04, Rei01, SLP^{+}01]. Farrell [Bur09]. Fasting [SHDR03]. Fat [SBH00]. Fathead [DT03, MD07]. Fatty [HNH^{+}04]. Fauna [CMKL01]. Fear [HKCP09]. Feather [EM04b]. Features [BWC^{+}07, KBJ08, RMRT03, SJ00]. Fecundity [LS07, MC00b, SRB00, SM07, WG05, Wyd01]. Fed [WK05]. Feed [FWL^{+}07, SF07]. Feeding [AMB^{+}08, ACB02, AVM05, BSM^{+}04, BBHD09, EK08, GHL02, HM09, HSS00, Lin04, Lin01, NBS07, NTB^{+}06, OS03, OA07, RM04, Rei01, SLM07, SB03a, WH07, ZG07]. Female [BEV^{+}07, CBF^{+}06, KSB^{+}08, LHL06, SKP^{+}08, WM07]. Fertilization [HWW^{+}06]. Fertilizing [CDT^{+}02]. Fidelity [BDR01, MA09, MKS01, NAG07, SWF09, YI02]. Field [BHL05, BBRH06, BBL06, BWC^{+}07, BR09a, BSW00, DHR07, JJRP08, LRF^{+}08, PR06, SRJ03, WWM07]. Field-Based [WWM07]. Field-Measured [BWC^{+}07]. Fin [AHC^{+}09, PB03]. Final [DCCA01]. Fine [KGW05, LDDB08, SML^{+}07, WSRA06]. Fine-Scale [KGW05, SML^{+}07, WSRA06]. Fine-Sediment [LDDB08]. Finfishes [WLB^{+}00]. Fingerling [SBW00]. Fingerprinting [HBBG07]. Fingerprints [BCT04]. Fins [STC^{+}09]. Fire [BH00c, TPS^{+}03, TBA^{+}04]. Fire-Control [BH00c]. First [Beh02, BG04, BKHH07, GP08, GKW^{+}06, JN00a, KSB^{+}06, KSB^{+}08, Lin01, MBC^{+}05, NL03, PBR^{+}01, PKR02, PLD00, RBLF06, SW03a, SKP^{+}08, SS08b, SN01]. First-Generation [BKHH07, KSB^{+}06, KSB^{+}08, SKP^{+}08]. First-Order [NL03]. First-Summer [PLD00]. First-Winter [SW03a]. First-Year [BG04, JN00a, SS08b, SN01]. Fish [ADP00, Ano02c, AW02, BHP^{+}08, BOMC07, BH06a, BBP^{+}04, BL00, BHHS04, BP01, BSB09, BHB^{+}07, BJZH05, Bre08, BWC^{+}07, BWS08, BCDF^{+}09, Bur09, Car04, CR07, CWB06, CS06, CMKL01, CW00a, CPF06, CGL^{+}09, CW00b, CG08, CB06, DT03, DC02, DM03, DCT02, DW06, DW09, FHB^{+}07, FCC07, GP08, GM04, GMCQ08, GDWC04, GJP09, GF07, GS06, GDW03, HPW08, HL00a, HDRB07, Hei08, HGM03, HGJO04, HS08, HMT^{+}06, HOA04, JHS02, JD07, KD01, KAB^{+}05, KAB09, Koe04, KWH03, KM05, KYWC09, KMV05, LMAF08, LBS^{+}06, LHM02, LBR^{+}01, LJC^{+}07, LPL01, Lin04, LM08a, LJE08, LFG06, LPN01, MM02, MRC^{+}09, MB08b, MW06, McK05, MLB03, MMP03, MM03, Mea05, MC09, MOP02, ML04, Mir05b, ME07, MEA^{+}09, MB02, MFMB06, NFG00, NDČ^{+}04, OJ01, OJPN02, OW01, OW04]. Fish [OBB^{+}09, Pap08, PR01c, PJSL09, PKR07, Pir04, PLH^{+}08, POH^{+}09, QB01, QK03, RRN^{+}08, Ros03b, RWB^{+}08, RP07, RBR^{+}00, SARC09, SJRE00, STC^{+}09, SGS06a, SGS06b, Sch00b, SCB08, SM08b, SP07, SW09b, SCP^{+}05, SK05b, SSRM08, SZSS08, SFE01, SH00b, SYC05, TJNL08, Tho03, TPS^{+}03, TBA^{+}04, TRBH08, VO01, WC00, WMBM02, WPGH08, WWS03, WJK07, WSSD03a, WSSD03b, WSY^{+}07, WBH06, WHLP07, WS09, WTW03, WC03, WTSC00, WS08, WNG02, WJH06, YSB^{+}08, ZWT05, Bur09, Bur09]. Fish-Based [LPN01]. Fish-Habitat [DCT02]. Fished [EBS02]. Fisheries [ARC^{+}08, Ano02d, Ano03d, Ano03e, Ano03f, Ano03g, Bro04, Chi09, DW08b, GGCF07, Gan08, GSWS04, JTPW09, LMR08, PJH04, Qui04, Sec08, SK07, WRC^{+}07, Wil07, Bur09]. Fishery [Bur09, Cop02, HFHR00, SGS05, vT05]. Fishes [AW05, AAG03, ARJ08, Ano03b, BJ03, BHSR09, BBDM05, BW02, BGH06, BG02, BPR06, BE07b, BBS07, BH06b, Gid02, HWN02, HI08, HPB05, JGMC05, KBJ08, LCM07, MMP02, MM01, NHFF09, OW01, PAS09, Pir04, Pol07, PG04, RMD03, RCP00, RGP06, RE08, RA07, RR08, Ron02, Ros03a, RSP09, Rul03, RGK01, RWL01, STC^{+}09, SKK^{+}08, SMZB01, SCH04a, SHS04, SMK^{+}07, TF07, TGWE04, TRK07, Tur05, VT06, ZSW02, Ano05]. Fishing [BNG08, CKB01, HJP01, KBJ08, LBVW03, PJH04, SSCM06, WNH05, WSLT07, XM00]. Fishway [BKP^{+}04]. Five [KBPC07, LHM02, MG06, OW01, SM08b, WMH^{+}04]. Flannelmouth [WMBM02]. Flathead [BKCL05, MSE07, PKWR05, VR05, KA01, SHS04]. Flexible [Heg02]. Flies [JHS02]. Flood [BDR01, FRK01, RDU02, RGK01, SBP^{+}08]. Floodplain [BBW03, HW04, LJE08, ML04, MMH01, RMG^{+}06, Sch02, SW09b, SCO^{+}02, TBK^{+}09]. Floods [ACB02, CM03a, WEB09]. Florida [BCT^{+}05, DWEL08, DA05, DA08, FHP00, HPM05, HMW07, HM09, HBC^{+}02, LM08a, LCNB^{+}06, PA01, RA09, SBC^{+}05, TAM07, TAC08, WLBBM09, WA03]. Flounder [BR02, FLLA09, MJS^{+}03, MCR05, SLC00]. Flow [AVM05, BBW03, BPR06, BSW00, FPD^{+}09, GBC^{+}05, HCK06, KGW05, KYWC09, LFA09, MM01, NSVT08, PG04, RHB^{+}06, RDU02, SA06, SJS09, WB01, WJF03, WS08]. Flow-Regulated [BPR06]. Flowing [ERS^{+}06, TKH^{+}09]. Flows [DA08, MPVM00, SYC04]. Fluctuating [SBSO06, WCBS06]. Fluctuation [RCP00, WWM07]. Fluvial [ACB08, KTNK01, Nat07, PPDV09]. fluviatilis [BL02, SHBFE07]. following [MVS07, OEO^{+}00, WHT^{+}00]. Food [BBV00, BSM^{+}04, BBHD09, CEW00, DCCA01, EBS02, EM04b, FGWS00, MFFI09, PSKH04, SVW02, TAD05, WHC04]. Forage [OBB^{+}09]. Foraging [ASN^{+}05, BR02, CGAD^{+}02, GDWC04, GMJ^{+}06, GHL02, HSS00, OBW04, Pol07, SCH03b, SH01, TAD05, VB02, ZG07]. Forced [SCM01]. Forest [CM03a, KR07, MDE^{+}09, TPS^{+}03]. Forests [BWS^{+}05, DHR07]. Fork [KHT08, KCY^{+}07, MSE07, MDE^{+}09, RSB^{+}08, WSSD03a, WSSD03b, MZ02]. Forms [GMP^{+}05, HBJ^{+}08b, OPH09]. Formulation [VO01]. Fort [JKWS06]. Four [HGM03, OM02, RSP09]. Fragment [AWNT00, DFCDR07, FWG07b, GFL00]. Fragmentation [POH^{+}09]. Fragmented [MMM07, PJC07]. Framework [SA01]. Frameworks [ASS00]. Francisco [FSH07, RB07]. Fraser [BLC^{+}04, BSC^{+}04, EKS^{+}05, PRR03, RHM^{+}06]. Frazil [SHW00]. Frecklebelly [PBT04]. Free [LMAF08, TKH^{+}09]. Free-Flowing [TKH^{+}09]. Frequency [BNG08, EKSX08, GSS03, HB09, TA06]. Freshwater [AS07, BD07, BG04, CHP08, DFBC^{+}09, HW08a, HNH^{+}04, HTT00, HTBT01, KD01, KB06, LS04, MM01, QVPG04, Ros03b, Rul03, RBM06, Sch03a, SZM01, VPQ04, WRK02, Wei08, WTB06, WRC^{+}03, WHCC03]. Fringe [BZW04]. Fry [BSM^{+}04, GAH^{+}06, TRS09, TAD05]. Full [CHH^{+}06].

Full-Sibling [CHH^{+}06]. Function [HB07, LDOS06, Pap08, TF07, VMS01]. Fundamental [Bet05]. Fundy [RMD08]. Future [PBI05, RHM^{+}06]. Fyke [BS06].

Gag [RBLF06]. Gags [LH03]. Gaining [OJ01]. Galaxiids [McD03, VBM^{+}09]. Galveston [MMCR08]. Gap [SLM07]. Gar [AMRM02, HHHK08, OBW04]. Gars [HHHK08]. Gas [BM06, GGCF07, JCB^{+}05, MWM00, WSSD03a, WSSD03b]. Gaston [HJP01]. Gastric [BBHD09, HCK^{+}09a, HN00, PKM07, SCH04b, WK05]. Gear [MM03]. Gene [BRK02, BBC^{+}08, NSVT08, WB01]. General [Lin04]. Generality [RA09]. Generalized [JD07]. Generated [NDČ^{+}04]. Generating [GM02]. Generation [BKHH07, KSB^{+}06, KSB^{+}08, SKP^{+}08]. Generic [MB04]. Genetic [BHL05, BSB09, BPM^{+}04, CDP02, CN01, CCMP06, DW08a, DLES06, DKP^{+}06, DFCDR07, DWBE00, DK06, DSD07, GMS^{+}08, HSS07, HBJ^{+}08b, HBTT06, HG00, HM08, HWMH02, IMM^{+}04, KA01, KLSA02, KMSL02, KFB05, LMAF08, LMA09, MBC00, MGJ^{+}01, MFOA09, MISS07, MTWS09, MOT04, MSS^{+}00, MKS01, Mil03, MMY09, NBMP06, NSC07, NSVT08, NDR^{+}09, NS02, PSB^{+}03, PSBC04, PSB^{+}05, PSSV^{+}09, PSD^{+}08, PJC07, RFOR03, ROC^{+}03, RC04, SMG03, SM08a, SHT^{+}00, SLGH08, SJFP08, SLF^{+}05, SBM00, TMWG00, TBK^{+}09, TM06, VFT02, WB01, Whi00, WKPS05, WSRA06, WRGN07, WM05b, WWR^{+}00]. Genetically [BFM08, Cam04, Cam05, GMP^{+}05, GMS^{+}06, Qui05]. Genetics [PWLL05]. Genotype [CHH^{+}06, RHR^{+}06, TBWD06]. Genotypes [PWDG07]. Genotypic [WORB00]. Genotyping [SSSS05]. Genus [AWNT00, SGSS01]. Geochemical [MMTF05, WTO08]. Geographic [BH03, BWFB06, HB06, WTJ00]. Geologies [BL01]. Geomorphological [CN01]. Georges [BB00, Sto02]. Georgia [CBN^{+}09, CPRS02, GI06, BSLN08, LBBF^{+}08, PSD^{+}08, RJ00, Sch00b, SLF^{+}05, YI02]. Geostatistical [JM05]. Geostatistics [Ano02c]. Giant [Ron02]. Gill [ARC^{+}08, BHS05, HFHR00, MHHS01, PT03b]. Gill-Net [ARC^{+}08, HFHR00, MHHS01, PT03b]. Gizzard [AHW03, ASN^{+}05, IDW03, MB05a, SVW02]. Glacial [YW07]. Glacially [YW07]. Global [Bur09]. Glycol [HBC^{+}02, SBC^{+}05]. Gobies [LAM04]. Goby [BM09, MC00a, MC00b, Mat01, RSP09, SMS04]. Golden [CSBM06, PACH01]. Gonadal [WFF^{+}02]. Goosefish [GH06]. Gradient [BG02, BZW04, DS04, HH03, KC05, RA09, SEP^{+}08]. Gradient-Related [SEP^{+}08]. Gradients [BLGW06, MDE^{+}09, MHM08]. Grand [Ano03b, PCF06, PP05]. Grande [DCG07, MJH^{+}09, NBMP06]. Granite [JAJ^{+}00, MJH^{+}09]. Grass [BSW00]. Gravel [HI08, Kon00]. Gravitation [DGFS04]. Gray [DS04]. Grayling [DH04, LM08b, MPVM00]. Grazing [SF07]. Great [BJZH05, DLES06, RE08, RTG00, GFL00, PSB^{+}03, PSBC04, PSB^{+}05, SFB03, WD08a, ZKE06]. Greater [HWW^{+}07, SZM01]. Green [ANC^{+}06, AHC^{+}09, LME^{+}08, MC04, MMH01, MBK05, VWD^{+}01, VLD^{+}06, BHW^{+}07]. Greenback [KLSA02, ML07]. Grilse [LLVB04]. Groundfishes [TL06]. Grounds [CM03b, DSH^{+}09, HM09]. Groundwater [BHW03, WSJ^{+}05]. Groups [SHBFE07]. Grow [TF07]. Grow-Out [TF07]. Growth [AZS07, AHW03, ASN^{+}05, AMRM02, ANC^{+}06, AKPQ08, Ano05, ACB02, BH06a, BMN04, BFC^{+}04, BDLB04, BG04, Bro05, BS07, CBF^{+}06, CHH^{+}06, CEW00, CN05a, CW00a, CB02b, CB02a, CB04, CCR^{+}07, CMBB01, CBMM08, DHR07, DS04, DRQ09, DSSA01, DH04, DS06, EWB^{+}06, EM01, FLLA09, FGWS00, FLLG00, GW03, GWSF00, GAH^{+}06, GSS03, GDWC04, GBMB01, GHL02, HWLC02, HWM04, HWW^{+}07, HNW06, HSS00, HBH^{+}08, HB07, HAS06, HN00, HK04b, HO05, HRMC07, HB06, HWHP04, HAW09, IRJ^{+}09, IK00, KFT^{+}09, KA07, KWH03, KC09, KSR01, LBS^{+}06, LB00, LM08b, MC00a, MCS08, MHD03, MGMM01, MN07, MJS^{+}03, Mic06, MBC^{+}05, MCS06, NSP03, NA04, NTW02, NSF04, NBK01, OM02, OY03, PB03, PCHB01, PT03a, PA01, QVPG04, RM04, RCP00, RMH^{+}08, REK06, RBM06, SW03a, Sch00a, SBR^{+}06]. Growth [SVW02, SR07, SMZB01, SKC^{+}05, SWHW06, SJ06, SG07, SLC06, SH06, SB08, SWF09, SSHB02, SS08b, SEP^{+}08, SGG09, SLC00, SN01, SB03b, SH08, TRS09, TMK03, TP01, TCG09, TTW^{+}09, TAC08, TAD05, TBWD06, VPN09, VT06, VPQ04, WNF09, WBH06, WG05, WHCC03, WSEW09, WW02, YMJ00, ZBF05, ZV07]. Growth-Enhanced [BDLB04, TAD05]. Growth-Related [MN07]. Guarding [HGR^{+}08]. Guide [Bur09]. Guilds [VO01]. Gulf [BH03, BWFB06, HL00a, HSH^{+}05, JPLSH09, LJC^{+}07, OS03, PWSC01, PCWC08, PACH01, RBLF06, RCB^{+}04, SMG03, SS05, WCC08, AMB^{+}08, CBMM08, FPD^{+}09, FHP00, GGCF07, GGO01, HPM05, HSRD04, HBP^{+}09, HG00, PAD01, XM00]. Gulls [QB01]. Guppies [Bas04]. Gut [BBL06].

Habitat [AKPQ08, BK03, BL00, BAS^{+}09, BH00b, BLBH05, BDD07, BSEM05, BIT00, BBW03, BW01, BCT04, BMH06, BWC^{+}07, BPR06, BZW04, BHV08, CRA05, CRMN04, CMMM04, CDF^{+}08, CJFM05, CSPP00, CPRS02, DF08, DGC08, DCT02, DA08, EM01, FHP00, GMCQ08, GI06, GA07, GHL02, HW04, HF04, HWN02, Heg02, HKCP09, HBO07, HF02, HRMC07, HKHC00, HSH^{+}04, JM09, JT04, JKWS06, JD07, KR07, KVZD02, KYWC09, KC09, LDDB08, LMK04, LC03, MGP06, MD07, MSE07, MDE^{+}09, MBS04, MB04, MOP02, MCR05, MZ02, MM08, MMBG09, MR00, NDR^{+}09, NAG07, NFP00, OJ01, OSS^{+}04, OYB03, PSSV^{+}09, PW02, PF05, PP08, PRR03, PF00, PR06, PP02, Pir04, POH^{+}09, QM05, QHR04, RBLF06, RMRT03, RR08, RSB^{+}08, RDU02, Ron02, Ros03a, RJ00, RMD08, SIH02, SB06, SP07, SS06]. Habitat [SG07, SCH03b, SHW00, SK05a, SK05b, SB08, SFE01, SFS04, SJ00, TRS09, TRK07, VO01, VHF05, WJB06, WLBBM09, WCC08, WA03, WMH^{+}04, WBH04, YI02, YI07, ZKE06, SMZB05]. Habitat-Specific [CDF^{+}08, HRMC07, SG07]. Habitats [BHHS04, BDD07, BBWW02, EWB^{+}06, GSSB03, JTM08, KHKS00, LH03, LPJS08, RIA^{+}07, RMG^{+}06, RBM06, SP03, TJNL08, WR08, WBH04, YW07].

Habits [MFFI09]. Haddock [BB00]. Hake [EK08, GVHC^{+}09]. Halibut [DS05]. Hampshire [NL03]. Handling [BC02, SZM01, vT05]. Hankin [Tho03]. Haplotype [WM07]. Haplotypes [GBN05]. Harbor [HDRB07, LM08a, YT00]. Hard [Ano03a, MM02, WRC^{+}03]. Hard-Part [WRC^{+}03]. Hard-Rock [MM02]. Hardness [CD09]. Harmful [Car04]. Harvest [BS08, DWBE00, HAW09, VAR^{+}02]. Harvested [LS07]. Harvesting [TPS^{+}03]. Hatch [HRMC07]. Hatcheries [Cam04, Cam05, Qui05, WNG02]. Hatchery [AMB^{+}08, BSLN08, BSB09, BH04, BKHH07, CLSD00, CG06, DSH^{+}09, DYWA^{+}04, DK06, FNL^{+}08, FP04, GMS^{+}08, GSSB03, HBTT06, HE00, HE01, HCK06, JKWS06, KHT08, KSB^{+}06, KSB^{+}08, KMP03, KZ06, KSR01, LBC^{+}04, LBS^{+}06, MBK05, MJH^{+}09, PSBC04, PSB^{+}05, PE01, PBI05, RHR^{+}06, SKP^{+}08, SLK01, SH06, SML^{+}07, SB08, TRS09, VFT02, WF05, Wei08, WC05, WSJH01, WSJ06]. Hatchery-Produced [DSH^{+}09]. Hatchery-Raised [DYWA^{+}04, GMS^{+}08]. Hatchery-Reared [GSSB03, HCK06, JKWS06, LBS^{+}06, MBK05, WF05]. Hatching [CD09]. Hawaiian [BLGW06]. Head [BIB04, BH00b, BH06b]. Headwater [BZW04, DHR07, Nat07, NDR^{+}09, SLGH08, VHF05, WBH^{+}05]. Healing [BBS^{+}08, PCHB01, WSB00]. Health [FSN^{+}03, KMV05]. Hearing [MWSP09]. Heat [FHML08, WSFJ05, WC03]. Heavily [EBS02]. Held [KSR01]. Henry [MZ02]. Hepatic [FHML08]. Heritability [DWBQ05]. Hermaphroditism [HWW^{+}06]. Herring [BSM^{+}08, HEW^{+}05, KM01, MOT04, NGC^{+}02, PSKH04, RMH^{+}08, SLM07, SLF^{+}05, TAM07, WSP05, YMJ00, YSC^{+}06]. Heterogeneity [dlHFB04]. Hickory [HMW07, TAM07]. Hidden [ME07]. Hidrostal [MLB03]. Hierarchical [ECW^{+}09a, HB07, HBJ^{+}08a, KHKS00, WB08]. High [AHJ^{+}05, AMB^{+}08, BHP^{+}08, BZW04, CF07a, CBMM08, DC02, GKB^{+}06, GN00, HGW04, KM06a, KM06b, LBC^{+}04, SBP^{+}08, SGS05, WJF03]. High-Discharge [DC02]. High-Elevation [AHJ^{+}05, CF07a]. High-Gradient [BZW04]. High-Throughput [SGS05]. High-Velocity [WJF03]. Highland [ASS00, KGS07]. Highlands [CB06, MHK^{+}01]. Highly [PHF07]. Hill [AW05]. Histocompatibility [BCS^{+}01, BLC^{+}04, BCM^{+}05]. Historical [GFLF02, HMD02, MB02, MVS07, PBT04, SCM09, TAM07]. Histories [Ess03, MTWS09, OM02, Rid08]. History [AHC^{+}09, BE00, BGPL08, BG02, CBF^{+}06, CN08, CST^{+}05, DWBQ05, DD02, FSH07, GBBP01, HMW07, HBJ^{+}08b, JSI09, KSB^{+}06, LGJ02, MGJ^{+}01, MMM07, MISS07, MKDB03, NSC07, NL03, OW09, OPH09, PR01a, PDP08, PPDV09, PCMS05, SBG01a, SBC^{+}09, TMWG00, VHF05, WSP05, WD08b, Wyd01, ZM05, ZM06a, ZEP09]. Hoh [BCV07]. Home [MA09, SJP05, VR05]. Homing [WS07a]. Hood [MFOA09]. Hook [HM07, WS09]. Hook-and-Line [HM07]. Hooks [KCK07]. Horizontal [NQ02]. Hormone [CN05a]. Horn [MEA^{+}09]. Horseshoe [HB04, KESB05]. Hosts [SMK^{+}07]. Hudson [ARJ08, HHS04, Hei08, HSC00, MLSD07, MS04, SLM07, WS07a, WS07b]. Humpback [MVD^{+}00, PCF06, PP05]. Huron [BT04a, GFLF02, HB07, HBJ^{+}08a, RRN^{+}08]. Hybrid [GPUH07, SA04b, WBI00, WSJ06, WSA00]. Hybridization [BPM^{+}04, CDP02, CSBM06, HHHK08, KLA02, KLSA02, KCY^{+}07, MMBG09, OR06, RTG00, WPS02, WRGN07]. Hybridized [PK04]. Hybrids [BBM02, BHL05, DA00, OR06]. Hydraulic [BHW03, CBM^{+}00, ZDK^{+}04]. Hydroacoustic [BR09a, CJMB06]. Hydrodynamics [Pap08]. Hydrodynamics-Based [Pap08]. Hydroelectric [Ano06a, EBFC04, GM02, WE05]. Hydrologic [NSF04, ZWT05, ZDK^{+}04]. Hydrology [SBG01a, ZSW02]. Hydropeaking [KC09]. Hydrophone [Hol08]. Hydropower [CW00b, MOSB02, MMS^{+}06]. Hydrosystem [ERS^{+}06, KPB^{+}04]. Hypotheses [RA09]. Hypothesis [KBW^{+}00, MZB^{+}07, MTWS09, QH05]. Hypoxia [FRK01, FCS^{+}03].

I* [CHH^{+}06]. Iberian [DFCDR07]. IBI [MMH03, HHK04, HH07]. Ice [SHW00]. Icelandic [AG02, DGK^{+}08]. ICES [Ano03h]. Ichthyofaunal [SJ08]. Ichthyology [Tur05]. Ictalurid [WTW^{+}00]. Idaho [HK00, AZS07, AFR02, BDD07, CB04, GBC^{+}05, HCK09b, KCY^{+}07, MM02, MSEL03, MSL^{+}06, MZ02, RWL01, WRC^{+}03]. Identification [BCS^{+}01, BLC^{+}04, BCM^{+}05, BCJ^{+}06, BB00, CN08, CB06, EBFC04, GKW^{+}06, HTBT01, MK05, PWDG07, ROC^{+}03, SA04b, SLP^{+}01, TRGV00, WKM^{+}07, ZR02]. Identifications [BHL05]. Identified [Hol08]. Identify [WLBBM09]. Identifying [BT04a, CSBM06, KLSA02, LPJS08]. Identity [OBB^{+}00, SJFP08]. II [WSY^{+}07]. Illegally [SGS05]. Illinois [PPR06, SW09b]. Ilyophagous [BGH06]. Immobilization [DM03]. Immune [LDOS06, VMS01]. Immunoglobulin [EKSX08]. Immunology [CHH02]. Impact [CMMM04, OSS^{+}04, WA05]. Impacts [GGCF07, MLSD07, McD03, SAT^{+}07, WLK03, WNF09]. Impair [LDOS06]. Impedance [PLH^{+}08]. Impede [YSB^{+}08]. Impediments [SLSF07]. Imperial [GSSB03]. Imperiled [DW09]. Impingement [HDRB07]. Implantation [WBI00]. Implanted [LKM04, MH05, PCHB01, PAR01, ZBF05]. Implicate [WM05b]. Implications [ASN^{+}05, BMZ07, BHSR09, BKCL05, CDP02, CB00, CB04, CRC^{+}02, DRG09, DYWA^{+}04, DA08, EK08, FPD^{+}09, FCS^{+}03, HF02, HRMC07, KRM03, LT01a, LT01b, MC09, MVD^{+}00, MH05, MB05a, Mir05a, MEA^{+}09, MTSS06, NKJ03, OLWC06, Sch00a, SVW02, Sch02, SW03b, SH00b, TCG09, TTW^{+}09, VB02]. Importance [Gid02, HBSA09, KWH03, LPL01, MRC^{+}09, MBS03a, STR^{+}07, URMH07, Wil07, WJH06, YW07]. Important [DBW09, SP03]. Impounded [SLGH08, TKH^{+}09]. Impoundment [NGC^{+}02, QK03]. Impoundments [BMH07, DD02, GS06, IDW03, Mic06]. Improve [CJMB06]. Improved [Ano05, BE07b, FSR^{+}07, SF07, SMZB01]. Improvement [BSN04, SCB08, WBH06]. Improvements [MBS04]. Improves [AJF09]. Improving [MSC08]. In-Lake [CM03b]. In-River [CM03b, OLWC06]. Inbreeding [Kan01]. Incidence [KCY^{+}07, LBS^{+}06, MSET05]. Inconnu [HTT00, HTBT01]. Incorporating [WWM07]. Increase [SR07, WHCC03]. Increased [HCK06, Sut07, WNF09]. Increases [SF07]. Increment [OA07]. Incubation [CPG03, Spi01]. Independent [HEW^{+}05]. Index [BJZH05, DRMH^{+}02, ESM^{+}03, GF07, HHK04, HH07, IHF08, LPN01, MHK^{+}01, MWG^{+}08, MOP02, Pir04, PHWS09, Sch00b, SHC01, SHWS05, WHP01, WHLP07, MMH03]. Index-Removal [IHF08]. Indexes [Pir04]. Indian [DWEL08]. Indiana [GS06, LAM04]. Indicates [Esh09, WGRW04, WDHW09]. Indicator [BBS07, PR01a, SJP05, VPN09]. Indicators [AMRM02, Van04, WFF^{+}02]. Indices [BFC^{+}04, CLSD00, HBJ^{+}08a, JD07, KJL^{+}07, LPL01, PF01b, SA04a, SHDR03, WHS^{+}07]. Indigenous [Esh09, WGRW04, WDHW09]. Indirect [AHW03, HLH^{+}07]. Individual [BCM^{+}05, BCJ^{+}06, BE00, BFC^{+}04, BBRH06, FT08, HR09, MWS08, MB06, OBB^{+}00, SCWP09, TR07, TB08]. Individual- [MB06]. Individual-Based [BE00, BBRH06, HR09, TR07]. Individually [HAS06, Sch05]. Individuals [BSS05, SRH07]. Induced [HGM03, HGJO04, MWMT02]. Inducing [Bas04]. Inductively [LFG06]. Infected [SBW00]. Infection [KA07, MMS00, PPDV09]. Infections [KMV05]. Inference [MTSS06]. Inferences [HF09, PAD01, SAT^{+}07, WB08]. Inferred [FSMH08, HM08, KFB05, WRC^{+}07]. Infestation [BS03, CEB^{+}05]. Influence [BSC^{+}04, CEB^{+}05, DW06, HJM04, HI08, HPB05, HOA04, IK00, KS01, KVD^{+}09, KBJ08, KC06, MSEL03, OBB^{+}09, OA07, PGW08, PF02, RP07, Sch05, Sec08, SP07, SLC06, SJ00, SB03b, SH01, WKCC08, WEB09, WTW03, ZM05]. Influenced [WR08, YW07]. Influences [CHH02, CJFM05, DT03, DCT02, MISS07, NDR^{+}09, NL03, QVPG04, RSB^{+}08, SLSF07, WAS^{+}00, ZN07a]. Influencing [DD02, KC05, MD05, MMBG09, PK02, SN01]. Information [FHB^{+}07, KR07, ME07, PF05]. Information-Theoretic [KR07, PF05]. Inhabiting [KGS00, OPH09, PLM05, RC04]. Inheritance [MGM02, WM05b]. Inhibition [MB02]. Initial [AHC^{+}09, VT06]. Inland [Ano02d, BTLD08, CHP08, EKSX08, PWLL05]. Inlet [SHT^{+}00]. Inner [RMD08]. Inopsetta [GPUH07]. Input [KJM^{+}08]. Inputs [SF07]. Insertions [HCK^{+}09a]. Insight [OJ01]. Insights [BE00, CW08, GMJ^{+}06, SBC^{+}09]. Installing [BHW03]. Instantaneous [XM00, ZM06b]. Institutionalize [Cam04, Cam05, Qui05]. Instream [BHB^{+}07, HKCP09, Jes00, JT04, MB01a, RDU02]. Integrated [BHB^{+}07, BH07, CRC^{+}01, FHML08, HDL07, NBS07, REK06, RSBF03]. Integrating [SYC05]. Integrity [ASS00, BJZH05, DRMH^{+}02, GF07, HH07, LPN01, MB08b, MHK^{+}01, MWG^{+}08, MMH03, Pir04, PHWS09, Sch00b, SHC01, SA01, WHP01, WHS^{+}07]. Intensity [BHMK05]. Interact [GWSF00]. Interaction [BR09b, MMS00]. Interactions [BMZ07, BV01, BR09b, CB00, FT08, FCC07, FGWS00, GVS02, GHL02, HC05, HCW01, JMS02, MBMO06, MB05a, RH02, RHV03, SBR^{+}06, SGF03, SMK^{+}07, ZV07]. Interactive [PAS09, WBH^{+}05]. Interannual [AMB^{+}08, PV04, SW09b]. Intercept [WS07a]. Interconnected [RHG^{+}02]. Interior [BPM^{+}04, NPT04, RIA^{+}07]. Intermittent [CGL^{+}09, WTW03]. Interplay [IRJ^{+}09, PSSV^{+}09]. Interpopulation [HAW09]. Interrelationships [PT03a]. Interslat [SJ04]. Interspecies [TM06]. Interspecific [BM09, Gid02, RV07, RHV03]. Interstitial [CJFM05]. Intertidal [RMD08, SCM09]. Interval [HDL07]. Interval-Specific [HDL07]. Intervention [PBI05]. Intra [TM06]. Intra- [TM06]. Intragravel [KTNK01]. Intraguild [BR09b]. Introduced [Beh02, BBDM05, BKCL05, Cou07, DS06, KZ06, McD03, MM01, MMBG09, OR06, PBR^{+}01, PKR02, PAS09, PKWR05, PKR07, RA07, RSP09, VBM^{+}09, WG03]. Introduction [MD07]. Introductions [MEA^{+}09]. Introgression [CDP02, HW09, RTG00, TBWD06]. Introgressive [CSBM06, KLA02, WPS02]. Invader [Mat01]. Invalid [WA05]. Invasion [BDD07, BLGW06, MKBF00, MVS07, WB09]. Invasions [IMM^{+}04, VB02]. Invasive [BM09, DGC07, DGC08, POH^{+}09]. Inventories [SJ08]. Invertebrate [CB00, HPM05, JD07, KR07, SF07, TK01]. Invertebrates [SH00a, SH08]. Investigation [WRGN07]. Investment [CBF^{+}06]. Involving [SRJ03]. Iowa [PP08]. Irminger [DGK^{+}08]. Irrigation [CR07, RR08]. ischyra [GPUH07]. Island [CPA^{+}06, HDRB07, HE00, HE01, LS07, VGA01, WSJH01, Ano03e, MOP02, MCR05]. Isle [GMCQ08]. Isolated [GH07, OW04, SM08a]. Isolation [HHB02, MOT04, RFOR03, Sch02]. Isotope [AW02, JGMC05, STC^{+}09, WLB^{+}00, WJMM07]. Isotopes [HBO07, JMS02, MBMO06]. israelensis [JHS02]. Items [SHS04].

J [Bur09]. J. [YI02]. Jack [EK08]. Jackson [HF04]. James [AO09]. Japan [IMM^{+}04]. Japanese [Nat07]. Jasus [IHF08]. Jersey [HDRB07, NAG07]. John [NSVT08]. Johnny [LAM04]. Johns [HMW07, HM09, TAM07]. Jr [Bur09]. Juan [PG04]. Jump [KM06a, KM06b]. Jumping [BHMK05, KM05, KM06a, KM06b]. June [BC07]. Just [MR00]. Juvenile [AAP03, ASN^{+}05, ANC^{+}06, AKPQ08, ARL^{+}05, AMB^{+}08, ACB02, BCT^{+}05, Bas04, BHL05, BSN04, BLS^{+}00, BLBH05, BM06, BDR01, BD07, BBHD09, BH03, BBDM05, BH07, BBWW02, BC02, BCW^{+}09, BWFB06, CEB^{+}05, CG04, CBN^{+}09, CHH02, CPA^{+}06, CCK^{+}09, CCR^{+}07, CPRS02, CRC^{+}01, CRC^{+}02, CLSD00, CST^{+}05, CMSM01, DBW09, DB08, EWB^{+}06, ECW^{+}09a, EM01, EK08, FACS06, FSR^{+}07, FCC07, FWL^{+}07, GBC^{+}05, GGG06, GdMH01, HCK^{+}09a, HWLC02, HC05, HL00b, Hei08, HNH^{+}04, HSBC02, HO05, HCK06, HCK09b, HB08a, HBC^{+}02, JN00a, JKWS06, KFT^{+}09, KSCD02, KFB^{+}06, KWH03, KB06, KM01, KMV05, LKM04, LH03, LFA09, MRC^{+}09, MLB03, MJS^{+}03, MCR05, MMS00, MWM00, MS03, MMH01, MSM^{+}08, NDČ^{+}04, NC05, NKJ03, NBS07, NTB^{+}06, OYB03, PSKH04, PCWC08, PF05, PAR01, QM05, RH02, RBLF06, RSC07, RSB^{+}08, RV07]. Juvenile [RQ01, RSBF03, SMG03, SW08a, Sch00a, Sch05, SCB08, SSD^{+}06, SB06, SG07, SZM01, SH00a, SHW00, SHDR03, SB03a, SHWS05, SK05a, SBO05, SWF09, SBC^{+}05, SS08b, SCM09, SSO09, SB03b, SYC04, TMK03, TRGV00, TAB00, TAM07, TFO^{+}09, TTW^{+}09, TRBH08, VMS01, VT07, WNF09, WF05, WTJ00, WCC08, WSFJ05, WSJ06, YT00]. Juveniles [BS03, BH04, LS04, Rei01, SLP^{+}01, WD08a].

Kamchatka [OPH09]. Kansas [SGG09]. Kaskaskia [SW09b]. Kelts [Ano06a, EBFC04, WE05]. Key [LDK^{+}04]. kHz [MW06]. Kill [RBR^{+}00]. King [Chi09, XM00]. Klamath [BAS^{+}09, BHV08, JSH^{+}08, MD07, MRC^{+}09, TMK03, AHC^{+}09, TM06, VWD^{+}01, VLD^{+}06]. Know [Cou07]. Known [BCDF^{+}09]. Known-Age [BCDF^{+}09]. Knows [Ano03h]. Kodiak [HE00, HE01, WSJH01]. Kokanee [CB02b, CB02a, CB04, Gro06, HJM04, SW03b]. Kootenai [KMSL02, PB03, PBI05, PD05]. Korea [CMKL01]. Kronotsky [OPH09]. Kuskokwim [MTWS09].

Laboratory [BBL06, CF07b, FLLG00, HW08b, KM06a, KM06b, NDČ^{+}04, PR06, RH02, SBB05, SRJ03, WHZR03]. Laboratory-Derived [SRJ03]. Laboratory-Generated [NDČ^{+}04]. Lack [CN05a, DS05]. Lacustrine [BLG01]. Lacustrine-Adfluvial [BLG01]. Lagoon [DWEL08]. Lagoons [CM07]. Lags [MPP05]. Lahontan [NS02]. Lake [BAS^{+}09, BSEM05, BSO01, BFHO05, BQH04, BM07, BS08, BL02, BCDF^{+}09, BHV08, CMMM04, CDF^{+}08, CDFW06, CJFM05, CB02b, CB02a, CB04, CCR^{+}07, CM03b, CJMB06, CND03, CN05b, DW08a, DLES06, DKP^{+}06, EM04a, EH03, FB08, GFLF02, GMCQ08, HW09, HSK03, HB07, HBJ^{+}08a, HPB05, HCW01, JSH^{+}08, JM09, KVZD02, KFB^{+}06, LRF^{+}08, LB00, LJE08, MGP06, MOP^{+}06, MD07, MRC^{+}09, MISS07, MGM02, MKS01, Mil04, MB01b, MVS07, MM08, NTB^{+}06, OEO^{+}00, OPH09, PSB^{+}03, PSBC04, PSKH04, PRME01, PSS09, PW02, PP08, PWLL05, PYFS07, RFOR03, RTH^{+}01, RWL01, SW09b, SH06, SK05a, SSHB02, SHS04, SDSB08, SB03b, TMK03, TP01, TK01, VAR^{+}02, VBM^{+}09, WEB00, WSJ^{+}05, WJK07, WORB00, WNT03, YSC^{+}06, ZKE06, ARC^{+}08, BV01, BSM^{+}04, BCT04, BM07, BS08, BT04a, DKP^{+}06, EM04a]. Lake [Esh09, HSK03, HBJ^{+}08a, HL00b, HJP01, HRMC07, HB08b, KBW^{+}00, KFB^{+}06, LAM04, MKBF00, MHD03, MHR^{+}05, MB08a, MPK03, MMM07, MBMO06, McK05, Mil03, MMY09, MB01b, MJSY09, NQ02, OLS04, PSB^{+}05, PT03b, PA01, PYFS07, PV04, PKM07, RRN^{+}08, RSP09, SW08a, SCH04a, SH06, SJFP08, SSR06, SM07, SMW06, SDSB08, TWF^{+}07, VBT08, WGRW04, WGW06, WDHW09, WEB00, WKCC08, WWKM07, WHC02, YSC^{+}06, YSB^{+}08, ZM06a, ZKE06]. Lake-Migratory [FB08]. Lake-Rearing [KFB^{+}06]. Lake-Spawning [WSJ^{+}05]. Lake-Specific [SW09b]. Lakes [BOMC07, BWS^{+}05, BHC03, BBDM05, CB00, DT03, DLES06, DA05, EM01, JJRP08, MGP06, ML04, OJ01, OBB^{+}09, PSB^{+}05, PT03a, PT05, PCMS05, RHG^{+}02, RA09, SW09b, SFE01, TPS^{+}03, TBA^{+}04, TAC08, VT06, GFL00, Mir05b, PSB^{+}03, PSBC04, RE08, RTG00, WRC^{+}07, WTSC00]. Lamprey [BSO01, CDT^{+}02, GBMB01, PSS09, POH^{+}09, RSBS09, SMW06, ZM05, ZM06a]. Lampreys [BT04a, Esh09, MWS08, MBS05, MBS03b, MOSB02, RSBS09, SM07, WGRW04, WGW06, WDHW09, YBSL04]. Lampricide [GBMB01, ZM06a]. Lampricide-Treated [GBMB01, ZM06a]. Lances [JTM08]. Land [PF01b, Sch00b, WLK03]. Landlocked [BRK02, Bet05, NKJ03, OPH09, SM07]. Lands [CGL^{+}09]. Landscape [IH04, KC06, LM08b, MB08b, SK05b, SZSS08]. Landscape-Scale [KC06]. Landscapes [Rid08]. Large [BLBH05, CG08, DGC07, DGC08, DHC04, DG04, ERS^{+}06, Gid02, GM04, JHS02, KYWC09, KC09, LCM07, LJC^{+}07, LJE08, LPN01, MKS01, MFMB06, RC04, SP07, WTB06]. Large-Scale [DG04, LJC^{+}07, WTB06]. Largemouth [ASN^{+}05, COB^{+}03, FGWS00, FCS^{+}03, GWSF00, HK04a, IDW03, JN00a, JN00b, LCNB^{+}06, MG06, NN06, NNL06, OYB03, OY03, PCC^{+}09, PLD00, RA09, SBB05, SW08b, SW09a, SCWP09, SWD04, SHWS05, SGG09, SP04, SKC^{+}04, TAB00, VB02, WJB06, WNH05, WA03, WP08, WSEW09, WNG02, YMJ00]. Largest [HEW^{+}05]. Larvae [AMRM02, CM03b, GP05, HEW^{+}05, KD01, WC03]. Larval [ANC^{+}06, BBL06, CW00a, CBR^{+}08, CG08, CGD05, DGC07, FRK01, FLLA09, GDWC04, GKW^{+}06, GBMB01, HWN02, HBBG07, HWHP04, JPLSH09, LFG06, MD07, MRC^{+}09, MWS08, MM01, OA07, RJ00, SKK^{+}08, SBG01b, SA04b, SH00b, TAB00, VWD^{+}01, WG03, WJF03, WC03, ZM05, ZM06a]. Laser [LFG06]. Late [ECW^{+}09a, EKS^{+}05]. Late-Run [EKS^{+}05]. Late-Summer [ECW^{+}09a]. Latent [HF09]. Lateral [CG08]. latisulcatus [XM00]. Latitudes [GW03]. Latitudinal [BG02, DS04, FLLA09, HGR^{+}08, RA09]. Laurentian [RE08, RTG00]. lavaretus [EBS02]. Law [Pap08]. Lawrence [Far01, NTB^{+}06, RCB^{+}04, HL00a]. Lead [FSN^{+}03]. Lean [HSK03, SH06]. Leaping [MSM^{+}08]. Learn [HH05, HH06]. Learned [YSB^{+}08]. Least [MM05]. Led [SMW06]. Legacy [MMY09]. Lemon [GdMH01]. Length [AWNT00, BH06a, DW08b, DFCDR07, FWG07b, GH06, GSS03, GF07, GBMB01, HN05, HBJ^{+}08a, HB09, KAB^{+}05, MVD^{+}00, MSEL03, Mir07, TA06, WMBM02]. Length-Frequency [TA06]. Lentic [PF01a, RBM06, SB06]. Lernaeopodidae [BS03]. Lessons [YSB^{+}08]. Lethal [JJRP08]. Lethality [BSO01, PSS09]. Level [CEW00, GP05, GBC^{+}05, MB06, SJ08]. Levels [CMMM04, EKSX08, FHML08, SCH04a]. Lewis [ZM05]. Lice [KMV05]. License [AJF09]. Life [AHC^{+}09, Ano03d, BE00, BGPL08, BG02, BWVN09, CN08, CH00, CST^{+}05, DWBQ05, DD02, Ess03, FSH07, GBBP01, HMW07, HBJ^{+}08b, IRJ^{+}09, JSI09, KSB^{+}06, LGJ02, MGJ^{+}01, MISS07, MTWS09, MBS05, MKDB03, NSC07, OW09, OPH09, PR01a, PDP08, PPDV09, PCMS05, Rid08, SBG01a, SBC^{+}09, TMWG00, WSP05, WD08b, Wyd01, ZM05, ZM06a]. Lifetime [HSS00]. Lifts [MLB03]. Light [BHMK05, GP05, GBC^{+}05, LDK^{+}04]. Like [PG01]. Likelihood [XM00]. Limit [WSRA06]. Limitations [Tho03]. Limited [SA04a]. Limiting [KC06, TAB00]. Limits [CF07a, HF02, WWM07]. Limnetic [KD01]. Limulus [HB04, KESB05].

Line [HM07, RT04]. Lineages [HMD02]. Linear [BBL06]. Lingcod [BE07a, LYS06, MDP^{+}06, PR06]. Linkage [NSF04]. Linkages [MWN05]. Linked [BR02]. Linking [DGC07, EM01, dlHFB04, MDE^{+}09]. Links [SK07]. Lip [WJF03]. Lipid [HNH^{+}04, WC05]. Lipids [PLH^{+}08]. Lipofuscin [PSJ08]. Lipofuscin-Based [PSJ08]. Liquefied [GGCF07]. Littoral [BL02, TRK07, WJK07]. Live [CTT04, FCS^{+}03, KSC^{+}06, MSET05, SKC^{+}04]. Live-Release [FCS^{+}03, KSC^{+}06, MSET05, SKC^{+}04]. Livestock [WR08]. Living [FSH07]. Loach [WCBS06]. Load [RSB^{+}08]. Lobster [HL00a, IHF08, WA05]. Lobsters [RH01]. Local [BDD07, DT03, GWSF00, MMBG09, RMRT03, SK05b]. Local- [GWSF00]. Local-Habitat [RMRT03]. Localized [NSVT08, OSS^{+}04]. Locally [YA01]. Location [BSO01, CBN^{+}09, DWEL08, GJP09, HBSA09, HEW^{+}05, MRC^{+}09, PD05]. Locations [CWT06, SW08a]. Loci [HG00, HM08, KMSL02, LCNB^{+}06, MGM02, Mil03]. Lock [BIB04]. Locus [NPT04, SAT^{+}07, SSCM06]. Locus-Specific [SAT^{+}07]. Logging [DHR07, NL03, VHF05]. Logistic [GM02, HMT^{+}06]. Logperch [BM09]. Long [BBC^{+}08, CMBB01, LS07, LMA09, MH05, QK03, SRL^{+}06, SLSF07, SS05, CPA^{+}06, HDRB07, HSH^{+}05, LS07]. Long-Range [CMBB01]. Long-Term [LS07, LMA09, MH05, QK03, SRL^{+}06, SLSF07, SS05, HSH^{+}05]. Longevity [FSMH08, SRB00]. Longfin [RB07]. Longitudinal [MHM08]. Longliners [KCK07]. Longnose [HHHK08]. Look [HH05, HH06, McD03]. Loss [RH01, WBI00]. Losses [CRC^{+}02, Log01, SBB05]. Lost [SS06, BAS^{+}09, BHV08, CM03b, CBR^{+}08, JSH^{+}08, MH02, TMK03, WW02]. Lotic [Heg02, PF01a, RBM06]. Louisiana [BJ03, FRK01, NTW02, RGK01, TRK07, WBH^{+}05]. Low [AMB^{+}08, BIB04, BH00b, BHB^{+}07, BH06b, CG04, CBMM08, HGW04, KMP03, LT01a, MH02, PE01, RJ00]. Low-Cost [BHB^{+}07]. Low-Head [BIB04, BH00b, BH06b]. Low-Temperature [LT01a]. Low-Velocity [RJ00]. Lower [Ano06a, BMN04, GSSB03, MS03, SHBFE07, ZSW02, AZS07, BSEM05, BG04, CCK^{+}09, CRC^{+}02, DG04, GBBP01, GLHA06, GI06, ISH07, JAJ^{+}00, JCB^{+}05, MJH^{+}09, MOSB02, PPWvdL08, SB05, SBG01b, SSD^{+}06, SS06, SW09b, TBK^{+}09, TRGV00, TH00, WSP05, WSSD03a, WSSD03b, WE05]. Lowers [HNW06]. Lowhead [TGWE04]. Lowlands [MB08b]. Lumpfish [HH05, HH06]. Lunar [DGFS04].

M [Bur09]. Maccullochella [BNG08]. Machines [BNG08]. Mackenzie [HTT00, HTBT01]. Mackerel [Chi09, EK08]. Macroinvertebrates [JHS02, TGWE04, VHF05, WTW03]. Macrophyte [VB02]. Macrophytes [MB05a]. Madtom [PBT04, WTW^{+}00, WAS^{+}00]. Main [BHHS04, BBWW02, WEB00, WTW^{+}00]. Main-Channel-Border [BHHS04]. Main-Stem [BBWW02, WTW^{+}00]. Maine [SKC^{+}05, BRK02, Leg05, OM02, SKL^{+}03]. Mainstream [DHC04, SN01]. Major [BCS^{+}01, BLC^{+}04, BCM^{+}05, HMD02]. Male [FSMH08, GMGV00, LBC^{+}04, LBS^{+}06, SSL^{+}03, SP04, VPQ04, WB08, YBSL04]. Managed [CDP02, Lon04]. Management [Ano02d, CDP02, GGO01, HB09, MVD^{+}00, PSBC04, PKR07, SF07, Sav04, SHT^{+}00, SK07, SW03b, Wil07, WM05a]. Manistee [TR07]. Manual [BFHO05]. Marine [Ano03f, Ano03h, AJF09, BMN04, BAWT05, CBMM08, DBW09, GN00, HNH^{+}04, LME^{+}08, MAW09, OW09, Qui04, QVPG04, SGS06a, SGS06b, SKC^{+}05, SFS04, SSO09, VPN09, VPQ04, WLB^{+}00, Wei08, WTB06, WHCC03]. Mark [AAG03, Bas04, FSMH08, GdMH01, LBR^{+}01, MKS01, NSP03, PYFS07, XM00]. Mark-Recapture [FSMH08]. Marked [MWMT02, Sch05]. Marker [BT04b, ROC^{+}03, SAT^{+}07]. Markers [BBM02, CSBM06, GKW^{+}06, KA01, KLSA02, RP03]. Marking [REK06]. Marks [LMA09, MWMT02]. Marlin [KCK07, PG01]. Marsh [MM01, MMP02, WLB^{+}00]. Mary [MK05]. Maryland [EKSX08, Pir04, WMH^{+}04, WS08]. Mass [HBJ^{+}08a, LFG06]. Mass-Length [HBJ^{+}08a]. Massachusetts [KHKS00, NBK01]. Masu [KTNK01, Rei01]. Matagorda [SLC00]. Maternal [BSS05, ZEP09]. Mating [MKP07]. Maturation [HAW09, LBC^{+}04, LBS^{+}06, SR07]. Mature [GMGV00, SRH07, WB08]. Maturity [Gro06, MSEL03, Sch05, SH06, SK08, SRH07, VPQ04, WA05, WFF^{+}02]. Maxima [CM01]. May [GVS02]. Meadows [BS07, SJRE00]. Mean [GH06, GF07]. Means [BT04a, WNT03, YA01]. Measure [PHF07, SB03a]. Measured [BWC^{+}07]. Measurements [LRF^{+}08, iTE08]. Measuring [BHW03]. Mechanically [CTT04]. Mechanism [WM05b]. Mechanisms [ML07, OYB03, SN01]. Mechanistic [NSF04]. Mediterranean [RL08]. Melania [MB05b]. Melanogrammus [BB00]. Melanoides [MB05b]. melanostomus [MC00a, MC00b]. Members [KGS07]. Menhaden [EKSX08]. Menu [EM04b]. Mercury [BBP^{+}04, CDF^{+}08, EH03, SARC09, SHS04, WCC08]. Mesocosm [KD01]. Mesocosms [LH03]. Metabolic [BC02, Log01]. Metabolism [GW03]. Metamorphosis [GBMB01, MWS08]. Methanol [BCT^{+}05, SBC^{+}05]. Method [Ano05, BSW00, CH05, HG07, LMAF08, LBR^{+}01, MSC08, SMZB01, WBI00]. Methodology [LMA09]. Methods [Ano03g, BBP^{+}04, Chi09, HAS06, HJP01, JDSS01, JTPW09, PDP08, VO01, YSC^{+}06, Zho02a]. Metric [PG01]. Metrics [ASS00, MG06]. Mexico [GGCF07, GGO01, HPM05, HG00, HSH^{+}05, JPLSH09, LJC^{+}07, OS03, PWSC01, PCWC08, PAD01, RBLF06, SMG03, SS05, WCC08]. Mg [VPN09]. Michigan [ARC^{+}08, BWS08, BS08, DKP^{+}06, DHC04, HPB05, HRMC07, LAM04, MHD03, MHR^{+}05, MB08a, MD05, Mil03, PV04, RSP09, SH06, SK05a, SZSS08, TR07, VHF05, WEB00, WKCC08, WBH04, WHC02, ZSW02, ZN07a, ZN07b]. Microchemical [WJMM07]. Microchemistry [FHB^{+}07, ZR02].

Microelemental [LFG06]. Microhabitat [AVM05, BJ03, CMSM01, HL00b, HCK09b, LCM07, MPVM00]. Microsatellite [BCS^{+}01, BCM^{+}05, BCJ^{+}06, BMM^{+}06, BJS^{+}06, BVMM06, BRK02, FHB^{+}07, HG00, HM08, KMSL02, LYS06, LCNB^{+}06, MSS^{+}00, MGM02, Mil03, NPT04, NFG00, RP03, WKM^{+}07]. Microsatellites [BLC^{+}04, BSM^{+}08, KA01, OBB^{+}00]. Microstructure [SSR06, TA06, VT07]. Mid [LT01a, ASS00, ARL^{+}05, CB06, DRMH^{+}02, ERS^{+}06, HMD02, MHK^{+}01]. Mid-Atlantic [LT01a, ASS00, CB06, DRMH^{+}02, HMD02, MHK^{+}01]. Mid-Columbia [ARL^{+}05, ERS^{+}06]. Middle [SMZB05, HSH^{+}04, KCY^{+}07, MMH01, MBK05, TCG09, WG05]. Midsummer [KSR01]. Midwestern [TGWE04, WTW^{+}00, WAS^{+}00]. Migrating [BGM06, CLSD00, CH05, RHM^{+}06, TRGV00, WC05, YT00]. Migration [AZS07, AO09, BSM^{+}08, BL05, BM06, BQH04, BLG01, CBN^{+}09, CM03b, EKS^{+}05, EBC01, FHP00, GK06, GKB^{+}06, GN00, HPB06, HB00, HB08a, HTT00, JCB^{+}05, KPC08, LMK04, LME^{+}08, MJH^{+}09, PW09, Rul03, SA06, SZM01, SJS09, TF01, TKH^{+}09, TFO^{+}09, WS07a]. Migrations [ERS^{+}06, HB08b, TTW^{+}09]. Migratory [BCV07, FHB^{+}07, FS06, FB08, HBJ^{+}08b, HWMH02, MKDB03, MK05, OLWC06, RQ02, ZEP09]. Military [WBH^{+}05]. Mimicry [PG04]. Minijack [BL05]. Minimum [DA08]. Mining [HI08, MM02]. Minipiezometers [BHW03]. Minnesota [MMY09, JJRP08, PT03a, PT03b, PT05, TP01, WLK03]. Minnow [BJ01, MD07, Sch02, SFB03, SLGH08, Sut07, WCBS06]. Minnows [DT03]. Miramichi [DCHS09]. Misapplication [WA05]. Mislabeling [GVHC^{+}09]. Mississippi [BHHS04, HSH^{+}04, KVZD02, Koe04, PWLL05, SMZB05, TCG09, WG05, ZDK^{+}04, HSRD04, HBP^{+}09, ML04, Mir05b, SJ06]. Missouri [BG04, SBG01b, BMH07, BG02, BW01, DG04, DK06, FS06, JKWS06, QHR04, VR05]. Misspecified [Sec08]. Mitigation [BSB09]. Mitochondrial [AWNT00, DFCDR07, Esh09, FWG07a, GFL00, GFLF02, GKW^{+}06, GBN05, HMD02, KA01, KGS07, SWH^{+}01, SBM00, WGRW04, WGW06, WDHW09, WM07, WWR^{+}00]. Mixed [FWG07b, PG01, WSLT07]. Mixed-Metric [PG01]. Mixed-Stock [FWG07b]. Mobile [HK04b, NAG07, GJP09, HI08]. mochon [WC03]. Model [BHP^{+}08, BH06a, BE00, BHC03, BBL06, BWC^{+}07, BR02, CEW00, CW04, CHP08, DW08b, DH04, DSD07, EM04b, GM02, GSSB03, HPW08, HC08, HR09, HSS00, HDL07, HMT^{+}06, KJM^{+}08, KRM03, Lin04, MOP^{+}06, MBS04, MD03, NHFF09, PP05, PHWS09, PG01, SBC^{+}09, SCB08, SWD04, SCO^{+}02, TW05, TR07, WPS02, WHZR03, WBH06, WD08b, WP08, WHC02, ZM06b, ZTP^{+}08]. Model-Based [MBS04, ZTP^{+}08]. Model-Predicted [BWC^{+}07]. Modeling [BWS^{+}05, BV01, BBRH06, CS06, CW08, CJMB06, ECW^{+}09a, ECW^{+}09b, FRB06, HK08, HB07, HBJ^{+}08a, HS08, PKR07, RMD03, RJ00, TGW04, TW05, VBT08, VBM^{+}09, WPGH08]. Modelling [Ano03g]. Models [BOMC07, BMH07, Bre08, Ess03, HB09, HS08, HF09, HOA04, JD07, MC09, OJPN02, PDP08, PAD01, PJH04, RMH^{+}08, RP07, SZSS08, TGW04, VO01, WB08, vT05]. Moderately [HCK06]. Modification [FSR^{+}07, WC03]. Modified [FPD^{+}09]. Modifying [CC03]. Modular [MGG^{+}06]. Modulation [LBS^{+}06]. Mohave [MPK03]. Mohawk [SLM07]. Molecular [BBM02, CSBM06, KLSA02, LMAF08, MM05]. molitorella [BGH06]. Molluscs [Nev04]. Monitor [VDG02]. Monitoring [DRG09, Mir05a, SJFP08]. Monongahela [CM03a]. Montana [BW01, FSN^{+}03, FS06, KMSL02, MSE07, MK05, MTSS06, PPDV09, RMRT03, SHS04]. Montane [IH04]. Moran [PGW08]. Morbidity [SBW00]. Moroko [AVM05]. Morone [SGSS01]. Morphological [AMRM02, BM07, BKHH07, DWBQ05, KGS00, SHWS05]. Morphology [HFHR00, PSSV^{+}09, RSB^{+}08, TRGV00, WSJ06, ZKE06]. Morphometric [SM07]. Morphometrics [KAB09, SKC^{+}05]. Morphometry [TP01]. Morphotypes [MB01b]. Mortalities [XM00]. Mortality [BHP^{+}08, BLS07, BG04, BH07, CJFM05, CBM^{+}00, DS05, DSD07, FSN^{+}03, FWTS08, FACS06, GH06, GM02, GDW03, HGM03, HGJO04, HLH^{+}07, HJP01, HF09, HH05, HH06, HO05, HRMC07, HMT^{+}06, IRJ^{+}09, JN00a, Jes00, MBS03a, MMS00, MBC^{+}05, MMS^{+}06, PT03a, PJH04, PF02, QB01, RMD03, RHM^{+}06, RWB^{+}08, REK06, SW03a, Sch00a, SBW00, SSD^{+}06, TH00, TAD05, WBI00, WNH05, WSLT07, ZM06b, vT05]. Mosquitofish [CW04, Kan01]. Motility [CDT^{+}02]. Mottled [LAM04, PG07, REK06]. Mountain [CND03, CN05b, iTE08, Wyd01, WW02, HKHC00, WEB09]. Mountains [BLS07, PP01]. Mounting [MB02]. Mouth [EK08]. Movement [AAG03, BIB04, BBG02, BSEM05, BIT00, BIB^{+}00, CG06, DGC08, DRQ09, DHC04, GBC^{+}05, GH07, HSRD04, HK00, HSBC02, HOA04, KVZD02, MHMP05, MH05, MZ02, MM08, Nat07, NGC^{+}02, NAG07, NHFF09, PWSC01, PCF06, RHG^{+}02, RA07, RMD08, SJP05, SK05a, SSRM08, SEP^{+}08, WRC^{+}07, WHC02]. Movements [AFR00, ARC^{+}08, AHC^{+}09, BW01, CPRS02, CSvdS02, FS06, GI06, JKWS06, LFA09, MA09, MS03, NQ02, PPWvdL08, PPR06, RCB^{+}04, RQ01, SP03, SHW00, WTB06, WRC^{+}03]. Moxostoma [RJ00]. Much [ME07]. Mud [BGH06]. Multilocus [PWDG07]. Multimetric [BWS^{+}05, ESM^{+}03]. Multipass [PTG04]. Multiple [BH06a, DWEL08, HWMH02, HG07, KJM^{+}08, MCR05, MPP05, SBW00]. Multiple-Regression [BH06a]. Multiple-Scale [MCR05]. Multiplexing [NFG00]. Multiscale [CCMP06, SFB03]. Multistate [NHFF09]. Multistream [DHR07]. Multiyear [LHOP01]. Murray [BNG08]. Muscle [AW02]. Muskellunge [BMH06, CEW00, Far01, MMY09]. Mussel [HW08a]. Mussels [HKCL07]. mykiss [BPM^{+}04, GBN05, RL08]. Mysis [CB00, CB02b, CB02a, SSHB02]. Mystery [MMS^{+}06]. Myxobolus [DCG07, dlHFB04, PPDV09, SBW00].

N [Bur09]. Narragansett [MOP02, MCR05]. Natal [MSC08, MKS01, SSR06, WTO08, WHT^{+}00]. Natal-Site [MKS01]. National [CM03a, OR06, SKA^{+}09]. Native [BM09, CDP02, CWB06, CSBM06, DS06, FP06, GGG06, GSWS04, HWN02, HCW01, IMM^{+}04, MMP02, McD03, MB06, MM01, MEA^{+}09, MMBG09, OR06, PP01, PAS09, PT05, PWLL05, PG04, RSP09, TBWD06, VBM^{+}09, WGW06, ZV07]. Natural [AHJ^{+}05, BSB09, BLG01, CBM^{+}00, DT03, FS06, GGCF07, GMGV00, HLH^{+}07, HJP01, Lon04, MJH^{+}09, PW02, PJH04, PG04, RQ01, RMG^{+}06, VFT02, WNH05, WAS^{+}00, WRGN07, XM00, ZM06b]. Naturalized [KHT08]. Naturally [BLS^{+}00, ERS^{+}06, GH07, JMS02, KMP03, KCY^{+}07, MBC00, MM05, PKM07, TRS09]. Nature [HHHK08, Ano03b]. Navarro [WSFJ05]. near [AHJ^{+}05, BIB04, HF04, SYC04]. Nearshore [BHV08, GMCQ08, KC09, PMRH06, TJNL08]. Nebraska [JKWS06, SBM00]. Nechako [MWN05]. Need [Cam04, Cam05, Qui05]. Needed [HH07]. Negative [Gro06]. Negatively [NBS07]. Neighboring [HWN02, OPH09, OBB^{+}09]. Nekton [BS07]. Neogobius [MC00a, MC00b]. Neosho [WTW^{+}00, WAS^{+}00]. nerka [CB00]. Nest [BL01, HGR^{+}08, WJB06]. Nest-Guarding [HGR^{+}08]. Nesting [SP04]. Nests [KFB05, OADC^{+}08, SMS04]. Net [ARC^{+}08, BH00a, BR09a, CB02a, HFHR00, MHHS01, PT03b]. Net-Cage [BR09a]. Net-Pen [CB02a]. Netting [BS06]. Network [EWB^{+}06, ECW^{+}09a, ECW^{+}09b, GH07]. Network-Scale [ECW^{+}09b]. Networks [McK05, OJ01]. Neural [McK05, OJ01]. Nevada [MPK03, PK04]. Neversink [BL00]. Newfoundland [CMSM01, HSBC02, HH05, HH06, JSI09]. Newly [AKPQ08, CB02b, WG05]. Niche [Bet05, HS08, JJRP08]. Nine [HM07]. Ninety [RP03]. Ninety-Seven [RP03]. Nipigon [DW08a, MM08]. Nitrogenous [BH00c]. No [Ano03h]. Nocturnal [EK08]. Nomenclature [GBN05]. Nominal [KGS00]. Nonannual [SRB00]. Nonequilibrium [GH06]. Nonindigenous [HWN02]. Noninflation [CGD05]. Nonlethal [BBP^{+}04, CH05, KMV05, STC^{+}09, SB03a]. Nonlethally [Bas04, MCS08]. Nonlinear [BL02, IH04]. Nonmigratory [MK05]. Nonmixing [LHOP01]. Nonnative [AFR00, AFR02, BGGL08, BDD07, CWB06, FP06, MB06, PP01, PG04, WJMM07, ZV07]. Nonspawning [BFM08]. Nonstructured [HK04a]. Nontarget [JHS02]. Normalization [BWS^{+}05]. Normandy [SBG01a]. North [FS06, HJP01, OR06, RTBH05, RWB^{+}08, TFO^{+}09, TTW^{+}09, WSP05, BW01, EBS09, GVHC^{+}09, IMM^{+}04, LME^{+}08, MHMP05, OS03, PT03a, PT05, SARC09, SFE01, SS08b]. North-Central [MHMP05, OS03, PT03a, PT05]. Northeast [HSBC02, MPP05, PMP^{+}02, PMP05]. Northeastern [HSH^{+}05, HB08a, PWDG07, SFS04, SS05]. Northern [BWS^{+}05, CN01, Far01, FLLA09, GBBP01, HCW01, LCNB^{+}06, MBS03a, MKS01, OO05, OO07, PT03a, PT03b, PT05, VT06, BHC03, BH03, BH07, DRMH^{+}02, dlHFB04, GN00, JPLSH09, LJC^{+}07, MMY09, PWSC01, PCWC08, PAS09, PACH01, RHG^{+}02, SMG03, TBA^{+}04, VHF05, VBM^{+}09, WCC08]. Northwest [BBDM05, BPM^{+}04, MMH03, MG03, Rot07]. Northwestern [BJZH05]. Norwegian [LLVB04]. Note [OJPN02]. Notes [KHKS00]. Nova [Jes00]. Novel [CN08]. Nuclear [PWDG07]. Nuclease [SSSS05]. Nucleotide [BTLD08, CN08, SSSS05, SAT^{+}07, SS08a, SSCM06, WKM^{+}07]. Number [QVPG04, SS08a]. Numbers [MTSS06]. Nursery [CM07, HBBG07, HBO07, RBLF06, WS08]. Nutrient [CB00, EH03, Gid02, PW09, RMH^{+}08]. Nutrient-Phytoplankton-Zooplankton [RMH^{+}08]. Nutrients [BBF^{+}03, Rid04]. Nutritional [SHWS05].

Observational [MKP07]. Observations [LHM02]. Observatory [GA07]. Observer [CTT04, MTSS06]. Occupancy [BHV08, SB08]. Occurrence [AHJ^{+}05, RMRT03]. Occurring [JMS02, MM05, OY03]. Ocean [Ano06b, GLHA06, GJP^{+}05, KPC08, MBC^{+}05, MPP05, NPT04, ROC^{+}03, Rot07]. Ocean-Type [NPT04, ROC^{+}03]. Oceanic [DGK^{+}08]. Oceanography [Ano03e]. Oconee [RJ00]. Off [AVM05, CH00, HW04, EK08, HWW^{+}07]. Off-Channel [HW04]. Offs [HDH00]. Offshore [GGCF07, GVHC^{+}09]. Offspring [SMS04]. Ohio [DHS06, ESM^{+}03, SCH04a, WKPS05]. Oil [CRMN04, SHT^{+}00, WHT^{+}00]. Oklahoma [MMM07, SW08a, DF08]. Old [CW08, McD03]. Olds [VFT02]. Olfactory [RSBS09]. Oligotrophic [CB00, CB02b, CB02a]. Oligotrophication [CMMM04]. Olympic [MKP07]. Omnivorous [SVW02]. Once [HEW^{+}05, PHF07]. Once-Largest [HEW^{+}05]. Oncorhynchus [BPM^{+}04, CB00, GBN05, RL08]. One [CPG03, MMP02]. Oneida [LRF^{+}08, VAR^{+}02]. Onslow [RWB^{+}08]. Ontario [McK05, CMMM04, CCR^{+}07, Esh09, HW09, KBW^{+}00, OEO^{+}00, OLS04, PKM07, SSR06, SJ00, WGRW04, WDHW09]. Ontogenetic [BE07a, DBW09, GMJ^{+}06, HF02, MBMO06, RCP00, TWF^{+}07]. Ontogenic [LCM07]. Ontogeny [SB06, WJH06]. Operation [HBTT06]. opilio [FSMH08]. Optimize [SJ08]. Optimum [FNL^{+}08]. Order [NL03]. Oregon [NBMP06, ACB08, BAS^{+}09, BL01, BV01, BBHD09, BHV08, CEB^{+}05, CGL^{+}09, CBR^{+}08, HH03, HB08a, HHK04, HH07, JSH^{+}08, MD07, MRC^{+}09, MFOA09, MS03, MJH^{+}09, RSB^{+}08, Ron02, Sch02, SLK01, TBK^{+}09, TMK03, WC00, ZR02]. Oreille [CB04]. Organic [SBP^{+}08]. Organisms [CPB07, VBM^{+}09]. Origin [BL05, CLSD00, FLLA09, FGWS00, MMTF05, RCB^{+}04, RT04, SLK01, SSR06, TMWG00, WC05, WSJ06, ZEP09]. Origins [BT04a, WTO08]. Osmolality [ASK01]. Osmotic [CC03]. Other [BGH06, CCK^{+}09, HL00a, HI08, RR08, SM07, VBM^{+}09, MLB03]. Otolith [BB00, BCV07, BGS^{+}04, BBS07, Chi09, FHB^{+}07, HTBT01, LFG06, MWMT02, NSP03, OA07, PCWC08, SW08a, SSR06, TA06, VT07, WRC^{+}03, ZR02]. Otoliths [BCT04, LMA09, MSC08, SSHB02, VPN09, WTO08, WJMM07]. Out-Migrating [YT00]. Out-Migration [CM03b]. Output [LBVW03]. Overexpansion [RB05]. Overfishing [SB05]. Overlap [PP02]. Overview [PDP08, Ros03a]. Overwinter [BH07, ECW^{+}09b, FGWS00, HRMC07, JN00a, MGMM01, MBS03a, PSKH04, PF02, SW08b]. Overwintering [CP03, HL00b]. Oxbow [WTSC00, ZWT05, Mir05b]. Oxidative [WC05]. Oxygen [CG04, GBC^{+}05, GAH^{+}06, KA07, MH02, OW01, PP02, TGW04, TW05]. Oxytetracycline [LMA09]. Oxythermal [JJRP08]. Oyster [GJP09, TF07]. Ozark [KGS07, PK02, PR01b, PR01c, QK03].

Pacific [BCM^{+}05, BCJ^{+}06, BMM^{+}06, BJS^{+}06, BSM^{+}08, BSN04, BBDM05, BPM^{+}04, BT04b, CH05, DS05, DGFS04, DB08, EK08, FACS06, HM07, HEW^{+}05, JTM08, KCK07, KGW05, Log01, MWS08, MMH03, MBS05, MBS03b, MG03, MOSB02, MPP05, PHF07, PWDG07, PMP^{+}02, PMP05, RSBS09, Ron02, RMH^{+}08, SA04a, SLF^{+}05, SSO09, TL06, TGW04, TW05]. Paddlefish [FS06, HHDC03, HM08, PF01a, RVJ01, SRL^{+}06, SB05, SGF03, SBM00, TH00, ZDK^{+}04]. Pallid [AAP03, BW01, GGG06, HSH^{+}04, JKWS06, SMZB05]. Pamunkey [HO05]. Pan [CHH^{+}06]. Panacea [KCK07]. Panther [BDD07]. Paralabrax [PACH01]. Parameters [HPM05, KJM^{+}08, KRM03, PACH01, Zho02a]. Parametric [Zho02a]. Parasite [RSB^{+}08]. Parasitic [KMV05]. Parasitism [PSS09]. Parentage [DWBQ05]. Parental [SSL^{+}03]. Park [OR06]. Parks [SKA^{+}09]. Parr [AZS07, GMGV00, PF01b, SLC06, SEP^{+}08, SBH00, TAD05]. Parr-to-Smolt [AZS07, PF01b]. Part [KCK07, WRC^{+}03]. Partial [WCM^{+}07]. Particle [GRR05, OM05, WMO03]. Partitioning [WA03]. Party [Mir05a]. parva [AVM05]. Pascagoula [HSRD04, HBP^{+}09]. Pass [MMP03, Mea05]. Passage [Ano06a, CS06, CW00b, FCC07, HB00, KPB^{+}04, MLB03, MOSB02, POH^{+}09, RHB^{+}06, WE05, ZDK^{+}04]. Passing [FACS06]. Passive [BHB^{+}07, BH07, CRC^{+}01, FHML08, FT08, Gan08, LBBF^{+}08, LPJS08, LMR08, NBS07, REK06, RSBF03, WLBBM09, HDL07]. Past [BGM06, SLSF07]. Patagonia [VBM^{+}09, Beh02, PBR^{+}01, PKR02]. Patagonian [BOMC07, CPB07, PAS09]. Patch [BHV08]. Paternal [FLLA09]. Pattern [DSH^{+}09, HB00, Mil04, NA04]. Patterns [AMB^{+}08, BJ03, BE07a, BSEM05, BLGW06, BCV07, DF08, DHC04, DB08, GKW^{+}06, GA07, HBH^{+}08, HSRD04, HS08, HK00, HB08a, HTT00, JM05, KM01, KFB05, MB01a, MB04, MKDB03, MHMP05, MM05, MM08, Nat07, NGC^{+}02, NDR^{+}09, NQ02, OY03, OADC^{+}08, OW01, OS03, OMM09, PR01a, PV04, RB07, RL08, SJP05, Sch00a, SWHW06, SWF09, TM06, TRK07, VFT02, VR05, WSB00, WWS03, WS07a, ZWT05]. Peacock [NNL06]. Pearl [PBT04]. Pectoral [AHC^{+}09]. peelii [BNG08, BNG08]. Pelagic [MFMB06]. Pen [CB02a]. Penaeus [XM00]. Pend [CB04]. Peninsula [MKP07, DFCDR07, ZSW02]. Pennsylvania [Car06, KC05, KC06]. PEPA [RHR^{+}06]. PEPA-1* [RHR^{+}06]. Peppered [WD08b]. Perca [BL02, SHBFE07]. Percent [PLH^{+}08]. Perch [BCT04, BL02, BR09b, CGD05, EH03, FLLG00, GDWC04, GMJ^{+}06, HN05, HCW01, HGW04, IRJ^{+}09, MKBF00, MA09, Mil03, PSSV^{+}09, PCMS05, SHBFE07, SM08a, TK01, VAR^{+}02]. Performance [BCT^{+}05, BDLB04, BHMK05, BNB^{+}07, CGAD^{+}02, DCG07, FWL^{+}07, HBC^{+}02, KM05, KM06b, LT01b, MDE^{+}09, MB06, MBS03b, NC05, NBS07, RJ00, SM08b, SBC^{+}05, SYC04, WMBM02, WCM^{+}07, WG03, KM06a]. Period [GN00, VR05]. Periodicity [LM08a]. Periods [PD05]. Peroxidation [WC05]. Persistence [HR09, SFB03, SS06]. Persistent [BLS07]. Persisting [VT06]. Persists [NSVT08]. Perspective [CR07, FB08, QK03]. Perspectives [GSWS04]. Peshtigo [BSEM05]. pH [KC05, MH02, NL03]. Phase [MS04]. Phenotype [JM09]. Phenotypes [DGK^{+}08]. Phenotypic [HWMH02, RFOR03, RT04, WPS02, WORB00, ZKE06]. Phenotypically [GMS^{+}06]. Photographic [GRR05, OM05, WMO03]. Photography [VDG02]. Phylogenetic [GFL00, RE08]. Phylogenetics [AS07, SWH^{+}01]. Physical [BHHS04, BWS08, KR07, MMS00, PR01b, SB06, TL06, WMH^{+}04]. Physicochemical [RGK01]. Physiological [BLS^{+}00, HGR^{+}08, JDSS01, KSC^{+}06, LT01b, LBS^{+}06, MMS00, MBS03b, MDP^{+}06, MSET05, PSKH04, RB05, SHDR03, SKC^{+}04, SYC05]. Physiology [Bur09, CHH02, FSN^{+}03, HCK06, KSCD02]. Phytoplankton [RMH^{+}08]. Pigmentation [GKW^{+}06]. Pike [Far01, MKS01, PT03a, PT03b, PT05, VT06]. Pikeminnow [BBRH06, BHW^{+}07, GBBP01, RH02]. Pinfish [SG07]. Pink [AMB^{+}08, BSC^{+}04, BH03, BH04, CTT04, CRMN04, CBM^{+}00, CBMM08, DSSA01, DWBQ05, GWM^{+}07, HE00, HE01, KMV05, MWMT02, MBC^{+}05, OBB^{+}00, RTG00, WHT^{+}00, WSJH01]. Piscine [WK05]. Piscivores [CN05b]. Piscivorous [CRC^{+}02, GGG06, ME07, RMD03, SW09a]. Piscivory [KWH03, MVS07]. PIT [HDL07, BHB^{+}07, KBPC07, NBS07]. PIT-Tagged [KBPC07]. Placement [HBSA09, RQ01]. Plain [AW05, HI08]. Plains [TPS^{+}03, BJZH05, SFB03, WD08a]. Planktivores [CB00]. Planktivorous [HPB05]. Plasma [DS05, EKSX08, FHML08, HBP^{+}09, LFG06, WFF^{+}02]. Pleuronectidae [GPUH07]. Pleuronectiformes [GPUH07]. Ploidy [BSW00]. Poeciliid [LBR^{+}01]. Points [RJ09, Sec08]. Poisoning [DWEL08]. Policies [GGO01]. Pollock [BWFB06, MWSP09, MFFI09, NSP03, SK08]. pollux [Nat07]. Polychlorinated [OW09]. Polymerase [AWNT00, MB02]. Polymorphic [RP03]. Polymorphism [AWNT00, SSSS05, SSCM06, WKM^{+}07, SS08a]. Polymorphisms [BTLD08, CN08, DFCDR07, FWG07b, SAT^{+}07]. polyphemus [HB04, KESB05]. Pompano [BCT^{+}05, HBC^{+}02, SBC^{+}05]. Ponds [BS06, HW04]. Pool [MGG^{+}06]. Pools [GSP01, KSR01, OW04, SCM09]. Poor [WG03]. Population [AHW03, ACB08, BIB04, BMM^{+}06, BJS^{+}06, BVMM06, BSM^{+}08, BSLN08, Beh02, BHW^{+}07, BC07, CDP02, Car06, CJMB06, CSvdS02, DW08a, DGK^{+}08, DLES06, DWBE00, DWBQ05, DW09, FWG07b, GH07, GN00, HSS07, HEW^{+}05, HKCP09, HOA04, ISH07, Jes00, KA01, Kan01, KTNK01, KZ06, LS07, Leg05, MG06, MB06, MTWS09, MMCR08, NKJ03, NS02, OO07, OBB^{+}00, OSS^{+}04, OPH09, PBI05, PBR^{+}01, PKR02, PHF07, PAD01, PYFS07, RT04, RC04, RDU02, RB07, RMD08, RVJ01, SB05, SR07, SLSF07, SH06, SML^{+}07, SAT^{+}07, SKL^{+}03, Spi01, SMW06, SBM00, TMWG00, WNF09, YSB^{+}08, ZN07a, ZN07b]. Population-Level [MB06]. Population-Scale [GH07]. Populations [BFM08, BSB09, BKCL05, BSS05, CM03a, CN01, CSBM06, DW08a, DF08, DRG09, DHR07, DLES06, DFCDR07, DK06, DS06, FWG07a, FPD^{+}09, FB08, GMS^{+}08, HSS07, HW09, HMD02, HR09, HWMH02, HH03, JSH^{+}08, KR07, KESB05, KMSL02, KC06, Lin01, MGJ^{+}01, MN07, MB04, MOT04, MM05, MK05, NSVT08, OY03, PSBC04, PF01a, PHF07, PK04, PLM05, PT05, PJC07, RIA^{+}07, SM08a, SHT^{+}00, SKC^{+}05, SCP^{+}05, SLF^{+}05, SM07, SKL^{+}03, TR07, VFT02, WB01, WTW^{+}00, WWR^{+}00, WORB00]. Position [SK05b]. Positive [HHB02]. Possible [GKW^{+}06, HHHK08]. Post [MMS^{+}06]. Post-Hydropower [MMS^{+}06]. Posthatching [HGM03]. Postrelease [BNB^{+}07, MFMB06]. Postsmolt [BFC^{+}04]. Postsmolts [LMK04]. Postspawning [DCHS09, RHG^{+}02]. Poststocking [JKWS06, SBB05]. Potential [BGGL08, DYWA^{+}04, FPD^{+}09, FLLG00, GGCF07, Gid02, HHDC03, HL00a, HDRB07, Hei08, IDW03, MBS04, MMH01, OSS^{+}04, SM08a, SJFP08, WFF^{+}02, WRC^{+}03]. Potomac [GA03, OO05, OO07]. Powell [VBT08]. Power [MD03, PE01, Pap08]. Power-Law [Pap08]. Practices [HGR^{+}08, PE01, TGW04]. Prairie [BW02, BPR06, DW06, OW01, OW04, BD07]. Prawn [XM00]. Pre [BNB^{+}07]. Pre- [BNB^{+}07]. Precocious [LBC^{+}04]. Precursor [SMK^{+}07]. Predacious [LPL01]. Predation [ARL^{+}05, BE07a, BBRH06, CRC^{+}01, CRC^{+}02, DB08, EK08, FP04, FP06, GP05, HL00a, Hei08, KD01, ML07, OLS04, Pol07, PF02, RMD03, RSBF03, SW03a, SCH03b, SW08b, SMS04, SSO09, TRBH08, VBM^{+}09, WTW03, YT00]. Predator [BR09b, CDF^{+}08, FP06, PP02, PKR07, RMD07, VBT08]. Predator-Prey [BR09b, VBT08]. Predators [CJFM05, HL00a, HJM04]. Predatory [IDW03, MJSY09, SBB05]. Predict [BMH07, DH04, Mat01]. Predictability [STR^{+}07]. Predicted [BWC^{+}07]. Predicting [IMM^{+}04, MFMB06, PCMS05, SZSS08, WP08]. Prediction [CB06, McK05]. Predictions [CEW00, CPB07, DSD07, OJPN02, WBH06]. Predictive [GSSB03, MC09, OJ01, OJPN02, PHWS09, SWD04, SM07]. Predictor [SS08a]. Predictors [GP08]. Preference [GMJ^{+}06, MFC04]. Preferences [BSEM05, CKB01, HF02, SB06]. Preferendum [DCCA01]. Prehatch [WNT03]. Presence [BP01, JD07, MB05a, RP07, WTW^{+}00]. Preservation [AW02, BSW00]. Prespawn [PD05]. Prespawning [CG06, DCHS09]. Pressure [KBJ08]. Prevalence [dlHFB04]. Previously [BD07]. Prey [ASN^{+}05, BW02, BR09b, CB04, CN05b, DHS06, FP06, GDWC04, GMJ^{+}06, HPM05, HW08b, HWHP04, KWH03, LPL01, OBW04, PP02, PV04, RMD07, RTBH05, RL08, RWL01, SBR^{+}06, SW09a, SH00b, SH08, TAB00, TAC08, URMH07, VBT08, VT06, WNF09, WKCC08, WK05]. Prickly [GP05, TWF^{+}07]. Primarily [HEW^{+}05]. Primary [CN05b]. Primer [MSS^{+}00]. Prince [BH03, BH04, CBMM08, HE00, HE01, WSJH01].

Principal [KAB09]. Prismatoidal [SBO05]. Pristine [FB08]. Probabilities [HI08, KYWC09]. Probability [BP01, WB08]. Problem [McD03]. Procedure [NS04]. Processes [BE00, GWSF00, HS08, KM01, MAW09, QH05, SFB03]. Prochilodus [GK06]. Produced [DSH^{+}09, PKM07, SBO05]. Production [AD08, ANE06, BSN04, FNL^{+}08, KMP03, LDK^{+}04, LC03, LM08a, McK05, MMCR08, PT03a, PG01, Rid04, RMG^{+}06, RH06a, RH06b, YBSL04]. Productive [JT04]. Productivity [KR07, KZ06, WHC04]. Progeny [BGPL08, KSB^{+}08, ZR02]. Program [LBC^{+}04, PSB^{+}05, SRL^{+}06]. Programs [HE00, HE01, WSJH01]. Progress [HK08]. Progression [MMS00, MWM00, VMS01]. Project [FRB06]. Projected [SFE01]. Prone [VT06]. Proofing [JTPW09]. Propeller [BS07]. Propellers [GDW03]. Proper [OJPN02]. Proportion [LLVB04, Lin04, VFT02]. Proportions [HAS06]. Prospects [PBI05]. Prospectus [Gan08]. Protect [EBS09, SYC05]. Protection [BMZ07]. Protein [FHML08]. Proteins [WSFJ05]. Protocol [CBM^{+}00]. Protocols [Cam04, Cam05, Qui05]. Prototype [JAJ^{+}00]. Provides [FHB^{+}07, HSS07, HWMH02, SB03a]. Proximate [HDH00, JRM^{+}08, KJL^{+}07]. Proxy [VPN09]. Pseudogene [CN05a]. Pseudorasbora [AVM05]. Puerto [NNL06]. Puffer [DWEL08]. Puffers [DWEL08]. Puget [Bis04, DB08, LYS06, MB08b, OW09, SLF^{+}05]. Pulse [BNG08, FRK01, RGK01, SBP^{+}08]. Pulsed [WJF03]. Pump [CC03, MLB03].

Quality [BDD07, Kon00, LC03, MDE^{+}09, MOP02, NFP00, SIH02, TMK03, TP01]. Quantification [LDDB08]. Quantify [Pol07]. Quantifying [SH00b, VBM^{+}09]. Quantiles [DCT02]. Quantitative [Ano03g]. Quinn [PK04]. Quinte [CMMM04].

Raceway [HCK06]. Rack [FCC07]. Radiated [BBL06]. Radiation [HGW04, OADC^{+}08]. Radio [BAS^{+}09, BE07b, GK06, HSBC02, JDSS01, MBS03b, PPR06]. Radio-Tagged [BAS^{+}09, GK06, HSBC02, MBS03b, PPR06]. Radio-Tagging [JDSS01]. Radio-Tracking [BE07b]. Radiocarbon [BCDF^{+}09, DFBC^{+}09, HPW08]. Radiotelemetry [BCV07, BCW^{+}09, EKS^{+}05]. Raftable [HH07]. Rainbow [BK03, BBM02, BMZ07, Beh02, BBS^{+}08, BWVN09, BTLD08, BHS05, BH00c, DK06, DCG07, FWTS08, FHML08, HNW06, HBJ^{+}08b, HCW01, IK00, KLSA02, KYWC09, KC09, KCY^{+}07, MKP07, MKDB03, MHMP05, MZ02, MMBG09, MJSY09, NBS07, OEO^{+}00, OR06, PRME01, PBR^{+}01, PKR02, PP02, PPDV09, RP03, RT04, SBW00, SH00a, SHW00, SHDR03, SML^{+}07, SBO05, SBSO06, SSCM06, STR^{+}07, TB08, VPN09, VBM^{+}09, WSB00, WPS02, WG03, WRGN07, WNT03, ZR02, ZEP09]. Raised [DYWA^{+}04, GMS^{+}08]. Raker [BHS05]. Ramp [WCM^{+}07]. Ranched [Rei01]. Randall [JKWS06]. Random [Zho02a]. Range [BL01, CMBB01, COB^{+}03, IMM^{+}04, MA09, VR05, CEB^{+}05, HH03]. Rangeland [SF07]. Ranges [SJP05]. Rangewide [HM08]. Rank [Lin04]. Rapid [BCW^{+}09, CM03b, CH05, LMAF08, RT04, SGS05]. Rappahannock [LHOP01]. Rare [KVD^{+}09, PDP08]. Rate [BBS^{+}08, GN00, HN00, HG07, Jes00, Log01, RHR^{+}06, TKH^{+}09, ZM06b]. Rates [AZS07, ARC^{+}08, ACB08, Ano06b, BHP^{+}08, CW00a, CB02a, GJP^{+}05, GDW03, HK04a, HPB06, HG07, HCK06, HDL07, HAW09, KPB^{+}04, LHL06, LBC^{+}04, MCS08, MJS^{+}03, Mir05a, MPP05, PJH04, PYFS07, PW09, PMP^{+}02, PMP05, SHBFE07, SCH04b, TGW04, TW05, TB08, VPQ04, WD08a, WS09, WHCC03, XM00]. Ratio [DSH^{+}09, HH05, HH06, SB03a]. Ration [CEW00, DCCA01, Har00, HHS04, HN00, PSKH04, PKM07]. Ratios [MCS08]. Rays [AHC^{+}09, PB03]. Razorback [GSSB03, MPK03, MMH01, MBK05, WG03]. Reaches [KVD^{+}09, SP07]. Reaction [AWNT00, MB02, SSSS05]. Reactive [SH01]. Real [HR09, MFOA09]. Real-Time [MFOA09]. Real-World [HR09]. Really [PLD00]. Reared [BLS^{+}00, BVL^{+}05, CF07b, FNL^{+}08, GSSB03, HCK06, JKWS06, LBS^{+}06, MWS08, MBK05, TRS09, WF05]. Rearing [BNB^{+}07, CRC^{+}01, CM03b, HBH^{+}08, JM09, KFB^{+}06, OA07, PE01, SCO^{+}02, TRS09, VWD^{+}01]. Reascension [BKP^{+}04]. Reassessment [MJSY09]. Recapture [AAG03, FSMH08, MKS01, NSP03, NBK01, PAD01, PYFS07, WS09, XM00]. Recaptured [HAS06]. Recaptures [PCF06]. Receiving [BSB09]. Recolonization [AW05]. Recompression [HM07]. Reconnected [GSSB03]. Reconstruction [MVS07]. Record [MMTF05, WTO08]. Recovered [DKP^{+}06]. Recoveries [ARC^{+}08, TFO^{+}09]. Recovering [CRMN04, PYFS07]. Recovery [BC02, OLWC06, SLSF07, SH06, SMW06, WS07b, Zho02b]. Recreational [ARC^{+}08, AJF09, Bro04, HGR^{+}08, Qui04]. Recruit [HF09]. Recruitment [BSC^{+}04, BHC03, BBRH06, CF07b, CF07a, HRMC07, LT01a, MKBF00, MHR^{+}05, MRC^{+}09, MB01a, MWN05, OLS04, PSB^{+}03, PGW08, RA09, SW03a, Sch00a, SS08b, TCG09]. Recurrent [MTWS09]. Red [Bas04, CB05, FWL^{+}07, GGCF07, GGO01, GMS^{+}08, HG00, Hol08, HSH^{+}05, JPLSH09, LBBF^{+}08, MGMM01, MMM07, MB05b, OS03, PWSC01, PCWC08, RB05, SMG03, Sch00a, SWHW06, SS08b, SS05, WCC08, MLB03]. Red-Rim [MB05b]. Redband [CSL09, KMSL02]. Redd [HW04, MG03, MTSS06]. Redds [KTNK01, WSJ^{+}05]. Redfish [DGK^{+}08]. Redhorse [RJ00, WJF03]. Redhorses [GI06]. Redside [HB06]. Reduce [AAG03, NS04, YT00]. Reduced [BMN04, CWT06, HNW06, MB05a, WSA00]. Reduction [MKBF00]. Reef [GJP09, KBJ08, RWB^{+}08, SJRE00]. Reefs [HSH^{+}05, OS03, RTH^{+}01, SGS06a, SS05, TF07, SGS06b]. Reeves [Tho03]. Reference [RJ09, Sec08, SKL^{+}03, TPS^{+}03]. Refinement [HC08]. Refining [Pir04]. Reflect [WSFJ05]. Refuge [BC07, MZB^{+}07, YI07]. Refugial [MM05]. Regeneration [CB00]. Regime [BSN04, Mil04, NSF04, OA07, PG04, RMH^{+}08, WWS03]. Regimes [ANC^{+}06, CF07b, FPD^{+}09, GAH^{+}06, HR09, JR03, WCBS06]. Region [GFL00, MHK^{+}01, SC02, WWR^{+}00]. Regional [ASS00, BWS^{+}05, BSN04, BHC03, DT03, GWSF00, KESB05, KM01, Pir04, SA01, WWS03, ZN07b]. Regional-Scale [GWSF00]. Regions [BSM^{+}08]. Registry [AJF09]. Regression [BH06a, DCT02, GM02, YA01]. Regulate [GWSF00, SFB03]. Regulated [BPR06, FS06, KC09, SP07]. Regulation [BBW03, Car06, PMRH06, Wei08]. Rehabilitation [SJFP08]. Reintroductions [SM08a]. Related [CPA^{+}06, LDOS06, MN07, SEP^{+}08]. Relatedness [KFB05, SMG03]. Relating [RCP00]. Relation [AFR00, BL00, BSO01, CW00b, DA05, HPM05, JCB^{+}05, KTNK01, LMK04, MG06, PSKH04, PR01c, PT05, SIH02, SW09b, SML^{+}07, SK05b, TL06, WTW^{+}00, WTSC00, WHC02, ZV07]. Relations [HBJ^{+}08a, QHR04, TBA^{+}04, WMH^{+}04]. Relationship [BHHS04, EKSX08, Gro06, HN05, HAS06, PF01b, RT04, TP01, ZSW02]. Relationships [BWC^{+}07, BL02, DW08b, DFCDR07, DCT02, GFL00, GF07, HWN02, HSK03, JN00b, KJL^{+}07, KR07, KAB^{+}05, LJE08, MCR05, MVD^{+}00, MPP05, OJ01, OPH09, PSSV^{+}09, Pir04, PF02, SM07, SBH00, iTE08, WRC^{+}03]. Relative [BM06, BS08, BHS05, CGM03, DM03, DCG07, GMJ^{+}06, HPB05, HMT^{+}06, KWH03, Mir05a, Mir05b, MK08, PLO07, QH05, RP07, RSBF03, SP07, SAT^{+}07, SS08a, URMH07, WEB00, WNF09, WSSD03a, WG03]. Release [BHP^{+}08, FCS^{+}03, HM07, KSC^{+}06, LBVW03, MSET05, Qui04, SMS04, SSL^{+}03, SKC^{+}04]. Released [GMS^{+}08, HMT^{+}06, WP08]. Relevance [PSBC04]. Reliability [PLH^{+}08]. relicta [CB00, CB02b, CB02a, SSHB02]. Remains [SKK^{+}08]. Remnant [DLES06]. Removal [BH06b, IHF08, RTH^{+}01, WJK07, WS09]. Renibacterium [MMS00]. Repeat [SS08a]. Repeats [SAT^{+}07]. Replacement [GVS02, RJ09]. Reproduction [DGC07, DA00, EM04a, GMGV00, HWW^{+}07, KS01, KBW^{+}00, NTW02, VGA01, WSEW09, WW02]. Reproductive [BVL^{+}05, BDLB04, BJ01, CBF^{+}06, DW06, Far01, HWM04, HDH00, KC07, KSB^{+}08, KMP03, LBVW03, MC00b, MSEL03, PK02, RFOR03, SKA^{+}09, SRB00, Sec08, SCWP09, Sut07, VLD^{+}06, WSA00]. Required [Mir07]. Requirements [BMZ07, Ros03a]. Research [Wil07]. Reserve [MEA^{+}09]. Reservoir [BM06, CCP^{+}03, Gid02, HJM04, JN00b, KPB^{+}04, MMM07, MHM08, PF01a, RMD03, RMD07, SBG01a, SIH02, SSRM08, SN01, WNH05, WTW^{+}00, WW02, YI02, BBV00]. Reservoir-Stocked [SSRM08]. Reservoirs [CW00a, DHS06, JMS02, MG06, NN06, NNL06, SGG09, WBH04]. Residence [GN00, HSH^{+}05, MS03, SS05]. Residency [Ano06b, BD07, GJP^{+}05]. Resident [GJP09, HBJ^{+}08b, HK00, HK04b, MKP07, WLB^{+}00, WSSD03a, WSSD03b, WS07a, WHCC03, WHC04, ZR02]. Resistance [BDD07, SH00b]. Resolving [GFL00]. Resource [BM09, FLLG00, PSSV^{+}09]. Resources [HPM05, Bur09]. Respiration [KRM03, SHBFE07]. Respiratory [FCS^{+}03]. Response [BDR01, BFHO05, Cam05, CBR^{+}08, DRQ09, GBC^{+}05, GKB^{+}06, HE01, Hit04, HWHP04, IH04, JDSS01, KB06, MBS04, MDP^{+}06, NKJ03, OM05, PKR02, RDU02, SBG01a, SCM01, SHC01, SZM01, SCP^{+}05, SBO05, SFE01, TF01, TK01, WDHW09, WJK07, ZWT05]. Responses [ASN^{+}05, BNG08, BC02, CCR^{+}07, COB^{+}03, DS05, FLLG00, FCS^{+}03, GJP09, HGR^{+}08, KGW05, MPVM00, MMS00, MBS03b, PJSL09, PG04, RA07, SGS06a, SGS06b, SHDR03, WNT03]. Restoration [Bis04, CBR^{+}08, MPVM00, NKJ03, PSBC04, TR07]. Restored [CG08, GJP09]. Restoring [Bur09, Lon04, WHC04]. Restraint [MHHS01]. Restricted [OW04, WB01]. Restriction [AWNT00, DFCDR07]. Result [HEW^{+}05]. Resulting [BR09b]. Results [NSP03, SYC05]. Retention [LMA09, MRC^{+}09, PRME01, SB03b]. Reticulate [BL01]. Retrospective [SSHB02]. Return [Ano06b, BHP^{+}08, GMJ^{+}06, GJP^{+}05, GN00, HCK06, LS04, LHL06, RHR^{+}06, SCH03b, WRC^{+}07]. Returning [HSS07]. Reveal [Mil03, OBB^{+}00]. Revealed [KMSL02, NDR^{+}09, WJMM07, ZR02]. Reveals [BPM^{+}04, SM08a]. Reversal [WM05b]. Review [Ano02a, Aps06, Bar01, Boe01, Boz01, BWW01, Bur09, Cas00, CW00b, Cur01, Gan08, Hil00, HE00, HE01, New01, NH01, Rah00, Sut01, WSJH01]. Reviewing [CW08]. Reviews [Ano08, Hea01]. Revised [ZM06b]. Revisiting [CW08]. RFLP [GFL00]. Rheotactic [NKJ03]. Rhode [MOP02, MCR05]. Ribosomal [AWNT00]. Richness [BP01, GP08, KVD^{+}09, Koe04, Mea05]. Rico [NNL06]. Riffle [MGG^{+}06, TGWE04]. Riffle-Dwelling [TGWE04]. Riffles [KSR01]. Rim [BCM^{+}05, BCJ^{+}06, BMM^{+}06, BJS^{+}06, MB05b]. Ring [HBSA09]. ringens [CH00]. Rings [LB00]. Rio [DCG07]. Riparian [BBF^{+}03]. Risk [OBB^{+}00, PF02, SCH03b]. Risks [HKCL07]. River [AZS07, Ano06b, AO09, BK03, BL00, BSC^{+}04, BLBH05, BSEM05, BHW^{+}07, BIT00, BMH06, BWS08, BCV07, BPR06, Bro05, BKHH07, CN01, CCMP06, CPRS02, CRC^{+}01, CGL^{+}09, CMBB01, CST^{+}05, CM03b, CBR^{+}08, CG08, DGC07, DGC08, DHC04, DW09, EKS^{+}05, EM04b, Far01, FS06, FRK01, FHP00, FP04, GBBP01, GBC^{+}05, GI06, GJP^{+}05, GS06, GDW03, HW04, HPM05, HMW07, HM09, HHS04, HI08, HBTT06, HSRD04, HBP^{+}09, HWMH02, HO05, HBO07, JHS02, Jes00, JAJ^{+}00, KA01, KVD^{+}09, KHT08, KPC08, KYWC09, KC09, KM01, LHOP01, Lon04, LJE08, MM02, MPK03, MFOA09, MGMM01, MMM07, MWN05, MSE07, MA09, MD05, Mea05, MZ02, MMH01, MK05, MJH^{+}09, MS04, MOSB02, NPT04, NBMP06, OLWC06, OR06, PB03, PCF06, PF05, PRR03, PSD^{+}08, PPDV09, PWLL05, QK03, QM05, QHR04]. River [RHM^{+}06, ROC^{+}03, RIA^{+}07, RSB^{+}08, RC04, RJ00, RGK01, RSBF03, SJP05, Sch02, SFB03, SMZB05, SS06, SW09b, SLGH08, TBK^{+}09, TGWE04, TM06, TAM07, TRK07, TR07, WSP05, WF05, WRC^{+}03, WSFJ05, WA03, WMH^{+}04, WRGN07, WTSC00, WS07b, ZR02, AZS07, AHC^{+}09, ARJ08, ARL^{+}05, BAS^{+}09, BHHS04, BS03, BLC^{+}04, BLS^{+}00, BL05, BGPL08, BM06, BKP^{+}04, BG02, BG04, BGM06, BBS07, BHV08, BIB^{+}00, CN05a, CCK^{+}09, CF07b, CRC^{+}02, CPG03, CG06, CM03b, CBR^{+}08, DWEL08, DCHS09, ESM^{+}03, EK08, EKS^{+}05, EBFC04, FSN^{+}03, FACS06, FSR^{+}07, FWG07a, FWG07b, GK06, GKB^{+}06, GA03, GSSB03, HF04, HWN02, Hei08, HPB06, HF09, HKHC00, HTT00, HTBT01, HSH^{+}04, HSC00, ISH07, JSH^{+}08, JKWS06, KLSA02, KPC08, KVZD02, KMSL02, KSB^{+}06, KSB^{+}08, Koe04, KCY^{+}07, LCM07, MC00a, MC00b, MLSD07]. River [MBC00, MDE^{+}09, MBS04, MTWS09, MH02, Mir05b, MBK05, MJH^{+}09, MMS^{+}06, NSC07, NSVT08, NTB^{+}06, OO05, OO07, OSS^{+}04, PBI05, PD05, PPWvdL08, PF01b, PK04, PPR06, PP05, PBT04, PG04, QHR04, RHM^{+}06, RQ02, RVJ01, SB05, SBG01b, SSD^{+}06, SLSF07, SML^{+}07, TMK03, TRGV00, TRS05, TKH^{+}09, TCG09, VWD^{+}01, VLD^{+}06, VDG02, WSSD03a, WSSD03b, WKPS05, WJMM07, WG05, WS07a, ZWT05, ZDK^{+}04, ZN07a]. River-Floodplain [LJE08]. River-Greenback [KLSA02]. Riverine [CSvdS02, PKR07, SSRM08]. Rivers [AHJ^{+}05, AG02, Bis04, FS06, GM04, LPN01, MWG^{+}08, MMH03, OM02, PAS09, PK02, PKWR05, SLK01, SLM07, SSRM08, SLP^{+}01, WHLP07, WHS^{+}07, WM05a, ZEP09, Ano06a, BW01, CSPP00, DG04, DA08, ERS^{+}06, HH07, JCB^{+}05, KPB^{+}04, KHKS00, LLVB04, TH00, WE05, ZN07b]. RNA [AWNT00, FLLG00, MCS08, SB03a]. Road [MSE07]. Roadmap [Rid08]. Roanoke [WSP05]. Robust [GI06, HPW08, RJ00, WJF03]. robustum [RJ00]. Rock [HH03, IHF08, MM02, PACH01]. Rockfish [GMP^{+}05, GMS^{+}06, HM07, PMRH06, PWDG07, SCM09]. Rockfishes [BE07a, GKW^{+}06]. Rocky [PP01, SCM09]. Role [ANE06, FP06, OW01, WLB^{+}00, WB09]. Roles [BDD07, HGW04]. Ronde [MJH^{+}09, NBMP06]. Rosyside [ZG07]. Rougheye [GMP^{+}05, GMS^{+}06]. Round [BM09, MC00a, MC00b, RSP09, SMS04, LAM04]. Roundtail [Bro05]. Royale [GMCQ08]. Ruffe [FLLG00, HN00, SHBFE07]. Run [BBC^{+}08, CN05a, EKS^{+}05, HF09, KS01, KHT08, MBC00, PSD^{+}08, RQ02, VT07, WM07]. Runoff [KS01]. Runs [RFOR03]. Russia [BVMM06].

S [Bas04]. Sable [DHC04, MD05]. Sablefish [LDOS06, SSO09]. Sacramento [RH02, SM08a, WF05]. Sakakawea [SRL^{+}06]. Salamanders [Ron02]. Salinity [BKCL05, Kan01, Mat01, OW01, ZJPB08]. Salmincola [BS03]. Salmon [AZS07, ARC^{+}08, AKPQ08, Ano06b, AG02, AMB^{+}08, ACB02, BS03, BCS^{+}01, BLC^{+}04, BCM^{+}05, BCJ^{+}06, BMM^{+}06, BJS^{+}06, BVMM06, BMN04, BSC^{+}04, BSN04, BSLN08, BSM^{+}04, BLS^{+}00, BFC^{+}04, BL05, BGPL08, BM06, BDR01, BD07, BBHD09, BEV^{+}07, BBC^{+}08, BDLB04, BBF^{+}03, BQH04, BKP^{+}04, BH03, BH04, BBDM05, BH07, BBWW02, BAWT05, BNB^{+}07, BGM06, BCW^{+}09, BT04b, BKHH07, CEB^{+}05, CBF^{+}06, CN08, Cam04, Cam05, CTT04, CRMN04, CBN^{+}09, CHH02, CN05a, CBM^{+}00, CLSD00, CMBB01, CPG03, CST^{+}05, CG06, CBMM08, CH05, DBW09, DGFS04, DRQ09, DWBQ05, DB08, EWB^{+}06, ECW^{+}09a, ECW^{+}09b, EM01, EK08, EKS^{+}05, EM04b, FNL^{+}08, FACS06, FSR^{+}07, FWG07a, FWG07b, GWM^{+}07, GABC00, GAH^{+}06, GMGV00, GSWS04, GKB^{+}06, GJP^{+}05, HSS07, HW04, HCK^{+}09a, HWS^{+}09, HFHR00, HBJ^{+}08a, HL00b, HAS06, HNH^{+}04, HDH00]. Salmon [HE00, HE01, HB00, HF09, HSBC02, HCK06, HCK09b, JDSS01, JCB^{+}05, KPB^{+}04, KPC08, KSCD02, KBW^{+}00, KSB^{+}06, KSB^{+}08, KFB^{+}06, KTNK01, KB06, KCY^{+}07, KMV05, KSR01, LLVB04, LS04, LKM04, LMK04, LBC^{+}04, LBS^{+}06, Leg05, LGJ02, Lin01, LFA09, MCS08, MAW09, MBC00, MGJ^{+}01, MN07, MBS04, MB04, MLB03, MTWS09, MMS00, MS03, MG03, MJH^{+}09, MWMT02, MBC^{+}05, MSM^{+}08, MPP05, MMS^{+}06, MCS06, NPT04, NSC07, NSVT08, NKJ03, NFP00, NSF04, OW09, OBB^{+}00, OSS^{+}04, PE01, PF01b, PF05, PHF07, PLO07, PAR01, PF00, PP02, PKM07, PMP^{+}02, PMP05, QB01, QPG^{+}02, QVPG04, Qui05, RM04, RFOR03, RHM^{+}06, ROC^{+}03, Rei01, RSC07, RSB^{+}08, RQ01, RMG^{+}06, RTG00, SA06, Sch05, SSD^{+}06, SKP^{+}08, SHT^{+}00, SLSF07, SB06, SKC^{+}05, SS06, SA04a, SLC06, SSSS05, SSR06]. Salmon [SAT^{+}07, SB08, SS08a, SKL^{+}03, SW03b, SEP^{+}08, SLP^{+}01, SSO09, SBH00, SYC04, SJS09, TMWG00, TBK^{+}09, TRGV00, TRS05, TKH^{+}09, TGW04, TW05, TFO^{+}09, TTW^{+}09, TAD05, TBWD06, VFT02, VMS01, VDG02, VT07, VPQ04, VGA01, WNF09, WKCC08, WB08, WF05, Wei08, WC05, WHT^{+}00, WSJH01, WSJ06, WM05b, WM07, WHCC03, WHC04, WORB00, YW07, Zho02b, AZS07]. Salmon-Based [WHC04]. Salmonid [CRC^{+}01, FSN^{+}03, FP06, HH03, JAJ^{+}00, KGW05, Kon00, McK05, Rid04]. Salmonids [ARL^{+}05, BBV00, BLBH05, BZW04, CCK^{+}09, CRC^{+}01, CRC^{+}02, FP04, HW08b, McD03, MWM00, PP01, PTG04, QM05, RHB^{+}06, RSBF03, SRH07, iTE08, WEB09, WHCC03, WHC04, YT00]. salmoninarum [MMS00]. Salmonine [RWL01]. Salmonines [CCR^{+}07, SJ00]. Salt [WLB^{+}00]. Sample [KAB09, Mir07]. Sampled [MCS08, SMG03]. Samples [GFLF02]. Sampling [AJF09, Chi09, CBM^{+}00, JRM^{+}08, MM03, Mea05, STC^{+}09, SJ08]. San [FSH07, PG04, RB07]. Sanctuaries [EBS09]. Sand [JTM08]. Sandbar [CM07]. Santee [CCP^{+}03]. sapidus [JM05, PSJ08]. Sardine [Log01]. Sauger [CGAD^{+}02]. Saugers [AHJ^{+}05, BGGL08]. Saugeyes [CGAD^{+}02, DHS06]. Savannah [GI06, CPRS02]. Saxitoxin [DWEL08]. Scale [ACB02, BE07b, DSH^{+}09, DG04, ECW^{+}09b, GWSF00, GH07, HAS06, HS08, KGW05, KC06, LJC^{+}07, MB01a, MCR05, ME07, MB02, NBMP06, PT05, SML^{+}07, SLP^{+}01, WTB06, WRC^{+}03, WSRA06]. Scales [MWG^{+}08, ME07, MMTF05, PMP05, SB03a, WTJ00]. Scaling [GA07]. Scallop [Sto02]. Scaphirhynchus [MSS^{+}00, SWH^{+}01]. Scarabaeidae [URMH07]. Scarring [BS07]. Scenarios [PKR07]. Schedules [Sec08]. Schistosomes [SMK^{+}07]. Schooling [WC03]. Sciaenid [LPJS08]. Sciaenidae [RGP06]. Science [Ano03d, Ano03h, Gan08, LMR08]. Scores [HH07]. Scotia [Jes00]. Screen [SYC05, WSY^{+}07]. Screened [SYC04]. Sculpin [BL01, BM09, KS01, KBPC07, KC07, KGS07, LAM04, MB08a, Nat07, PG07, RHV03, TWF^{+}07, ZV07]. Sculpins [AS07, GP05, REK06].

Sea [Ano03h, BSO01, CDT^{+}02, DGK^{+}08, Esh09, GBMB01, KMV05, PSS09, POH^{+}09, Rei01, SWHW06, SM07, SMW06, Sto02, WGRW04, WGW06, WDHW09, YBSL04, ZM05, ZM06a, DGK^{+}08, MFFI09, NC05, RM04, SK08, TJNL08, vT05]. Sea-Ranched [Rei01]. Seabass [AD08]. Seagrass [BS07, SJRE00]. Seals [YT00]. Season [BFM08, HNW06, HWMH02, LGJ02, MC00b, OADC^{+}08]. Seasonal [ARC^{+}08, BBG02, BV01, BLS^{+}00, BSEM05, BIT00, BH03, BBWW02, BWFB06, BIB^{+}00, DHS06, DB08, EWB^{+}06, GI06, HSC00, IHF08, LCM07, MPVM00, MHMP05, MS03, MZ02, NQ02, PPWvdL08, PW02, PP08, PLM05, PV04, RFOR03, RSC07, RL08, SJP05, SIH02, SWF09, TBK^{+}09, TAM07, TTW^{+}09]. Seasons [CH00]. Seatrout [IC02, NTW02, WLBBM09]. Seawater [AHC^{+}09, HCK^{+}09a, PE01, SZM01]. Seaweed [CMKL01]. Second [DHR07]. Second-Growth [DHR07]. Sediment [BJ01, KB06, LDDB08, RSC07, Sut07]. Seeking [KB06]. Segment [BWS08]. Segments [FS06]. Segregation [PAS09]. Seine [BH00a]. Selection [AVM05, BMH07, BL01, BMH06, BPR06, BHS05, CRA05, CB04, DGC08, DA08, FPD^{+}09, GDWC04, HW04, HFHR00, Heg02, HF02, HCK09b, LCM07, LH03, MB04, PW02, PP08, PCC^{+}09, PP02, PV04, QPG^{+}02, Rei01, RSB^{+}08, SJ04, SW09a, VHF05, WKCC08]. Selective [BG04, DSD07, MBC^{+}05, POH^{+}09, QB01]. Selectivity [BHP^{+}08, Lin04, PKWR05, RTBH05]. Self [HWW^{+}06]. Self-Fertilization [HWW^{+}06]. Semen [ASK01, CDFW06]. Semiannual [LHL06]. Semibuoyant [ADP00]. Seminatural [FNL^{+}08]. senegalensis [DFCDR07]. Senegalese [DFCDR07]. Sensitivity [CG04, Ess03, MWSP09, RSBS09]. Sensory [HWS^{+}09]. Sequences [MSS^{+}00, SWH^{+}01, WWR^{+}00]. Sequencing [GFL00, WGW06]. Serial [BBC^{+}08]. series [Bur09]. Serological [MHHS01]. Settlement [DS04]. Seven [PT03a, RP03]. Severe [AW05, HW08a]. Severity [FGWS00, MWM00, MCS06, PPDV09]. Sex [CN05a, HH05, HH06, KESB05, QB01, RSB^{+}08, SRH07, WFF^{+}02, WM05b, YBSL04]. Sex-Biased [KESB05]. Sex-Selective [QB01]. Sexual [MSEL03]. Sexually [OM02]. Shad [AHW03, ASN^{+}05, AO09, BIB04, BH00b, FRK01, HMW07, HM09, HO05, HBO07, IDW03, MB05a, OLWC06, SVW02, SZM01, TAM07, VBT08, WTO08]. Shallow [BBDM05, BBW03, CDF^{+}08, JTM08, LCM07, PW02, TJNL08]. Shallow-Water [BBW03, JTM08]. Shanny [BBL06]. Shape [BB00, PLO07, SB03b]. Shapes [SBO05]. Shark [CM07, KCK07]. Sharks [GdMH01, MHHS01]. Shear [NDČ^{+}04]. Shelf [TFO^{+}09]. Shell [FSMH08]. Shellfish [GJP09]. Shift [BSN04, HF02]. Shifts [BR09b, DBW09, LCM07, Mil04, OEO^{+}00, SVW02, TWF^{+}07]. Shimofuri [Mat01]. Shiner [BJ01, DW09, HKCP09, HB06, KA07, MMM07]. Shiners [MGMM01]. Shoal [WA03]. Shock [FHML08, WSFJ05]. Shore [CM07, HB08b, SJFP08]. Short [GN00, NBS07, SAT^{+}07, SS08a]. Short-Term [NBS07]. Shortnose [BAS^{+}09, BHV08, CG04, CPRS02, CCP^{+}03, CM03b, CBR^{+}08, HWW^{+}06, JSH^{+}08, KHKS00, TMK03, WS07b, ZJPB08]. Should [DW08b]. Shovelnose [AAP03, BW01, KFT^{+}09, TCG09]. Show [JM09]. Showed [PE01]. Shown [KA01]. Shrimp [DSSA01, SMG03, YA01]. Sibling [CHH^{+}06, GMS^{+}06]. Sicklefin [DG04]. Signal [HPW08]. Signatures [Hit04, MSC08, MMTF05, PCWC08, SW08a, WTO08]. Significance [SKC^{+}04]. Significant [SM08a]. Silver [WG05]. Silverside [CHP08]. Simple [HG07, LHOP01, NS04, WP08]. Simulated [FWTS08, KSC^{+}06, RBR^{+}00, SFE01, SMS04, SYC04, WBI00]. Simulating [BH06a, NGC^{+}02]. Simulation [FCC07, PG01, SCM01]. Simulations [LRF^{+}08]. since [HK08]. Single [BTLD08, CN08, MMP03, Mea05, PR01a, SSSS05, SAT^{+}07, SSCM06, WKM^{+}07, SS08a]. Single-Nucleotide [BTLD08, CN08, SAT^{+}07]. Single-Nucleotide-Polymorphism [SS08a]. Single-Pass [MMP03, Mea05]. Sink [RR08]. Sire [BGPL08]. Siscowet [BM07, BS08, HSK03]. Site [BL01, FPD^{+}09, HW04, MA09, MB04, MKS01, MBK05, NAG07, SWF09, YI02]. Site-Specific [MB04]. Sites [dlHFB04, Hol08, KYWC09, WLBBM09]. Situ [WRK02, WSJ^{+}05]. Situations [GH06, SA04a]. Six [CRA05, MGM02, RQ02, WD08a]. Size [AHW03, BIB04, BE00, BGPL08, BH04, BG04, BH07, BHS05, CD09, CPA^{+}06, CRC^{+}02, CJMB06, DWBE00, DA05, DM03, ECW^{+}09b, FP06, FGWS00, GWM^{+}07, GMS^{+}08, GDWC04, Gro06, HC05, HAW09, JM09, Jes00, KBJ08, KAB09, KYWC09, LS07, Mic06, Mir05a, MBC^{+}05, NN06, NDR^{+}09, OLS04, PSKH04, PCHB01, PLO07, PAR01, PSD^{+}08, PG07, QB01, QVPG04, Sch05, SJ04, SCH03b, SW08b, SWD04, SK08, SN01, SB03b, TAB00, TAM07, WA05, Zho02b, ZSW02]. Size- [QB01]. Size-at-Maturity [WA05]. Size-Classes [SCH03b, SW08b]. Size-Dependent [BH07, PG07, SN01, Zho02b]. Size-Related [CPA^{+}06]. Size-Selective [BG04, MBC^{+}05]. Sizes [GRR05, Mir07, OM05, WMO03]. Skagit [Ano06b, GJP^{+}05]. Skin [EKSX08]. Skipjack [Ess03]. Skipped [Sec08]. Skykomish [KHT08]. Slat [SJ04]. Slave [ZKE06]. Slimy [BM09, KS01, KBPC07, KC07, RHV03, ZV07]. Slope [AFR00, BH06b, DRMH^{+}02, TL06, DGK^{+}08]. Slopes [PP01]. Slough [MS03]. Slowed [GKB^{+}06]. Small [AHJ^{+}05, BS03, BS06, BHB^{+}07, BBWW02, BE07b, GP08, GM04, HNW06, HBH^{+}08, IDW03, KBPC07, KC07, KM01, LHM02, MB08b, MA09, Mic06, NBMP06, NHFF09, PF00, PT03b, Ron02, SCH04a, SWF09, SRH07, Tho03, TAB00, ZBF05]. Small-Scale [BE07b, NBMP06]. Small-Stream [MB08b, PF00]. Smalleye [DW09]. Smallmouth [DF08, DS06, DSD07, FP04, FP06, FCS^{+}03, HC05, HGR^{+}08, MVS07, MSET05, OYB03, OY03, Pea04, SCM01, SCWP09, SMS04, SSL^{+}03, SP04, WJK07, WHZR03]. Smelt [HBBG07, HCW01, MJSY09, OEO^{+}00, PRME01, PP02, RB07, STR^{+}07, WSY^{+}07]. Smith [RSB^{+}08, CND03, CN05b, QM05]. Smolt [AZS07, AG02, ECW^{+}09b, KMP03, PF01b, PLO07, RMG^{+}06]. Smolting [BLS^{+}00]. Smolts [DGFS04, HWS^{+}09, JDSS01, JAJ^{+}00, KGW05, MCS08, MJH^{+}09, PE01, SJS09, WC05]. Snail [MB05b, SMK^{+}07]. Snail-Eating [SMK^{+}07]. Snake [AZS07, Ano06a, CCMP06, TRGV00, WE05, AZS07, BKP^{+}04, CMBB01, CPG03, CST^{+}05, CG06, EBFC04, GBC^{+}05, HF04, HF09, JAJ^{+}00, JCB^{+}05, KPB^{+}04, MBC00, MBS04, MZ02, MJH^{+}09, MMS^{+}06, NSC07, PF01b, PF05, SLSF07, TKH^{+}09]. Snakehead [OO07]. Snakeheads [OO05]. Snapper [DS04, GGCF07, GGO01, HG00, HSH^{+}05, JPLSH09, OS03, PWSC01, PCWC08, RB05, SMG03, SS05, WCC08]. Snook [LBVW03]. Snow [FSMH08, SGS05, vT05]. SNP [SS08a]. SNPs [SSCM06]. Social [Bro04]. Sockeye [BLC^{+}04, BCM^{+}05, BMM^{+}06, BVMM06, BSC^{+}04, BSM^{+}04, BQH04, DRQ09, EM01, EKS^{+}05, HSS07, HW04, HFHR00, HDH00, HB00, MTWS09, PF00, QB01, RM04, RFOR03, RHM^{+}06, SHT^{+}00, SLSF07, TTW^{+}09, WORB00, YW07]. Sodium [ASK01]. Solar [HGW04]. Sole [DFCDR07, GPUH07]. Solea [DFCDR07]. Solid [CD09]. Solution [KCK07, LFG06]. Solution-Based [LFG06]. Somatic [CH05, DS06, HB06]. Some [SLF^{+}05]. Soniferous [ARJ08]. Sound [AD08, BH04, LYS06, LM08a, OW09, RH06b, RTBH05, Bis04, BH03, CPA^{+}06, CBMM08, DB08, HDRB07, HE00, HE01, LS07, MB08b, RH06a, SLF^{+}05, TRBH08, WSJH01]. Source [BDD07, DW08a, RMD03, URMH07, WLB^{+}00]. Sources [Lon04, NSVT08, PSB^{+}03, SDSB08, WJMM07]. South [BIT00, CSPP00, CPRS02, CCP^{+}03, GI06, HB08b, JKWS06, MS03, SJFP08, YI02, KHT08, MSE07, MDE^{+}09]. Southeast [BZW04, GKW^{+}06, HPB05, MAW09, SWHW06, SSO09]. Southeastern [SP07, BBWW02, HWM04, HWW^{+}07, JTM08, KBJ08, RMD03, RMD07, WNG02]. Southern [BR02, IHF08, NGC^{+}02, SIH02, SLC00, WSRA06, BIB^{+}00, FRB06, HW09, HL00a, HB07, NAG07, OEO^{+}00, PG07, RBLF06, RCB^{+}04, SLF^{+}05, SJ00]. Southwest [MKDB03]. Southwestern [MGMM01, BLS07, DFCDR07, PF01a, SLP^{+}01]. Space [SJ04, VR05]. Spacing [BHS05]. Spatial [ANE06, AMB^{+}08, ACB02, BJ03, BSM^{+}04, BE07a, CSvdS02, DF08, DCT02, FRB06, HB08a, Koe04, LLVB04, LDDB08, MKP07, MM03, MWG^{+}08, MBK05, MFFI09, OMM09, PCWC08, PP01, PLM05, PMP^{+}02, PMP05, RP07, SW08a, SWHW06, SLF^{+}05, SK08, SCM09, WPGH08]. Spatial-Temporal [ACB02]. Spatially [BWS^{+}05, NFP00]. Spatiotemporal [LC03, MVD^{+}00, PGW08, SBG01b]. Spawned [CGM03, MBC00, SBR^{+}06, TA06]. Spawner [HF09, YSC^{+}06]. Spawning [AD08, AO09, BH00b, BEV^{+}07, BBF^{+}03, BLG01, BPR06, BH06b, BJ01, Cam04, Cam05, CRMN04, CH00, CJFM05, CPG03, CM03b, DSH^{+}09, DS04, DCHS09, EBS09, FPD^{+}09, FHP00, GK06, GVS02, Gro06, HFHR00, HM09, HSRD04, HEW^{+}05, Hol08, HGW04, KWH03, Kon00, KTNK01, KMP03, LBVW03, LBBF^{+}08, LPJS08, MB04, MKS01, MBK05, OADC^{+}08, PD05, PRR03, PF00, PPDV09, PLD00, PW09, QPG^{+}02, Qui05, RH06b, SGSS01, SW08a, SKP^{+}08, SRB00, Sec08, SCO^{+}02, Sut07, TRS05, VWD^{+}01, VDG02, WLBBM09, WSJ^{+}05, WEB09, WD08b, YW07, RH06a]. Spawning-Site [MKS01]. Specie [AHJ^{+}05]. Species [BP01, BMZ07, BM09, BZW04, CWB06, CW00a, CRC^{+}01, Cou07, DGK^{+}08, DWEL08, GVHC^{+}09, GPUH07, GMP^{+}05, GMS^{+}06, GP08, GM04, GVS02, GJP09, HM07, HGM03, HGJO04, KVD^{+}09, Koe04, KWH03, KGS00, LHM02, MHHS01, MD07, MLB03, Mea05, ME07, OJPN02, OBW04, PMRH06, PDP08, POH^{+}09, PMP05, QHR04, RP07, SJP05, SM08b, SJ00, SCM09, SYC05, WD08a, WS09, Wil07, WBH04, WWKM07]. Specific [BCV07, BT04b, CDF^{+}08, GSS03, HO05, HRMC07, HDL07, JRM^{+}08, MB04, PSB^{+}03, SG07, SW09b, SAT^{+}07, TTW^{+}09]. Specificity [CN05a]. Spectrometry [LFG06]. Speed [AAP03, GBC^{+}05, HB00, KRM03, Pea04, SA06]. Speeds [CS06]. Sperm [Cam04, Cam05, Qui05]. Spermatozoa [CDT^{+}02]. Spill [CRMN04, SHT^{+}00]. Spines [PRME01]. Splittail [BFM08, FSH07, SCO^{+}02]. Spokane [SML^{+}07]. Spoonhead [BM09]. Spot [CEB^{+}05]. Spotted [DA08, IC02, NTW02, OBW04, PACH01, WLBBM09]. Spread [MMBG09]. Spring [AZS07, BLS^{+}00, BGPL08, BBC^{+}08, BGM06, CMBB01, CGM03, FNL^{+}08, FS06, GN00, HF09, JAJ^{+}00, JCB^{+}05, KPC08, KSB^{+}06, KSB^{+}08, LBC^{+}04, LBS^{+}06, MBS04, MB04, MJH^{+}09, PF01b, SBR^{+}06, SSD^{+}06, SKP^{+}08, SBP^{+}08, TRGV00, VMS01, WJB06, WEB09, WC05, WM07, BKHH07]. Spring- [CGM03, SBR^{+}06]. Spring-Run [BBC^{+}08, WM07]. Spring-Spawning [WEB09]. Sprint [NC05]. Spruce [Car06]. Sr [VPN09]. St [Far01, HL00a, RCB^{+}04, WRC^{+}07, XM00]. St. [HMW07, HM09, MK05, NTB^{+}06, TAM07]. Stability [BSB09, KAB09, MOT04, SP07]. Stable [AW02, HBO07, JGMC05, JMS02, MBMO06, STC^{+}09, WLB^{+}00, WJMM07]. Stage [BWVN09, MBS05, WFF^{+}02]. Stages [Ano03d, IRJ^{+}09]. Staging [HPB06]. Stamina [ZBF05]. Standard [DW08b, VO01]. Standardization [RCP00]. Standardized [MD03, PE01]. Standards [BMZ07]. Start [WNT03]. Starvation [Log01]. State [HEW^{+}05, SBC^{+}09]. State-Dependent [SBC^{+}09]. States [HWM04, KBJ08, MB05b, RMD07, SFS04, ZTP^{+}08]. Statewide [SARC09]. Stations [GM02]. Statistical [BWS^{+}05, PE01, PF01b, RP07]. Statistics [RVJ01]. Statolith [BT04a]. Status [BLS^{+}00, BHW^{+}07, CC03, ISH07, MSL^{+}06, MFFI09, OLS04, PBI05, SHWS05, WGW06]. Staudinger [Bur09]. Steelhead [Ano06a, BBM02, BS03, BHL05, BM06, BVL^{+}05, BSB09, BKP^{+}04, BBWW02, CLSD00, CP03, DSH^{+}09, ERS^{+}06, EBFC04, GSWS04, HBH^{+}08, HBTT06, HWMH02, HPB06, KHT08, KPB^{+}04, KSCD02, KMP03, KZ06, KFB05, MFOA09, MDE^{+}09, MKP07, NBMP06, RH02, RHR^{+}06, RQ02, SBC^{+}09, SLK01, SWF09, TRS09, TGW04, TW05, TR07, WSFJ05, WE05, WHC02, ZR02, ZEP09]. Stehekin [OR06]. Stem [BBWW02, WTW^{+}00]. Steroid [YBSL04]. Stimulus [WNT03]. Stock [BCS^{+}01, BLC^{+}04, BCM^{+}05, BCJ^{+}06, BSC^{+}04, BHC03, BB00, BQH04, CGM03, DYWA^{+}04, DA05, EBS02, EBS09, FWG07b, GFLF02, GSS03, GMS^{+}08, HB09, KZ06, LT01b, OLS04, PG01, Rot07, SRL^{+}06, Sch00a, SLSF07, TBK^{+}09, TTW^{+}09, WSLT07, WKM^{+}07, YSB^{+}08]. Stock-Recruitment [BHC03]. Stock-Specific [GSS03, TTW^{+}09]. Stocked [CCR^{+}07, DHS06, PWLL05, SSRM08, SN01]. Stocking [HW09, MMY09, SJFP08, SML^{+}07]. Stockings [Mic06]. Stocks [CRC^{+}01, DK06, HDRB07, SM08a, VT07]. Stomach [SH00a]. Stomachs [PRME01]. Stone [AVM05, KS01]. Stores [PSKH04]. Storm [RTH^{+}01]. STR [SS08a]. Strain [PSB^{+}03, TBWD06]. Strain-Specific [PSB^{+}03]. Strains [CM01, NKJ03]. Strait [SLF^{+}05, BSLN08, CBN^{+}09]. Strategies [Lon04]. Stratified [NGC^{+}02]. Strawberry [BBV00]. Straying [MWMT02, SLK01, WHT^{+}00]. Stream [AFR00, AAG03, Ano03e, Ano06b, BWS^{+}05, BL00, BHW03, BW02, BH07, BWC^{+}07, BT04a, CEB^{+}05, DF08, DC02, DW06, EWB^{+}06, ECW^{+}09a, ECW^{+}09b, GP08, GSP01, GJP^{+}05, GH07, HNW06, HW08b, HR09, HS08, HK00, HCK09b, IH04, JGMC05, KVD^{+}09, KC05, LHM02, LGJ02, LC03, LFA09, MGG^{+}06, MB08b, MSE07, MHMP05, MG03, MMTF05, MMS^{+}06, NPT04, Nat07, NA04, NL03, OW01, OW04, PF00, PR01b, PR01c, PTG04, PJSL09, PG07, Pir04, ROC^{+}03, RA07, RV07, RQ01, Ros03a, RHV03, RP07, RBR^{+}00, Sch00b, SKP^{+}08, SM08b, SP07, SJ06, SCP^{+}05, SK05b, SB08, SWF09, SZSS08, SEP^{+}08, SJ00, iTE08, VO01, VGA01, WWS03, WHT^{+}00, WH07, WTW^{+}00, WAS^{+}00, WTW03, WBH^{+}05, WHCC03, WHC04, ZSW02]. Stream-Dwelling [DC02, HR09, LGJ02, MHMP05, NA04, PTG04]. Stream-Resident [HK00, WHCC03, WHC04]. Stream-Type [MMS^{+}06, NPT04, ROC^{+}03]. Streambed [GRR05, MPVM00, OW04, OM05, WMO03]. Streamer [RH01]. Streamflow [HNW06, PJSL09, WH07]. Streams [AW05, ASS00, ACB02, BLS07, BL01, BHB^{+}07, BJZH05, BLGW06, BWS08, BE07b, BZW04, CRA05, CPF06, CF07b, CF07a, CB06, FSN^{+}03, GM04, GF07, GBMB01, HKHC00, HHK04, KBPC07, KC07, KR07, KC05, KC06, LM08b, McD03, MWG^{+}08, MC09, MTSS06, NDR^{+}09, NL03, PP01, PHWS09, RH02, RMRT03, Ron02, SF07, SBP^{+}08, SJFP08, SA01, TRS09, Tho03, VHF05, VR05, WC00, WLK03, WEB09, WPGH08, WHLP07, WHS^{+}07, Wyd01]. Strength [BR09a, CJMB06, HN05, JN00b, WRK02, WKCC08, WEB09, WS07b]. Stress [ASK01, BC02, CLSD00, Kan01, MMS00, MDP^{+}06, SZM01, Van04, WC05]. Stressors [LDOS06, SBW00]. Striped [BH00b, Bet05, BIT00, BC02, DWBE00, DCHS09, Har00, HN05, Hei08, HJP01, HSC00, JRM^{+}08, LHOP01, NAG07, OMM09, PR01a, RCB^{+}04, RTBH05, RMD08, SIH02, SW08a, SN01, TRBH08, VBT08, WBI00, WSLT07, WS07a, YI02, YI07]. Strom [YI02]. Strontium [HTBT01]. Structural [BNB^{+}07, PR06]. Structure [AHW03, ACB08, BMM^{+}06, BJS^{+}06, BVMM06, BSM^{+}08, BE00, BQH04, CCMP06, DW08a, DGK^{+}08, DLES06, FWG07b, GFLF02, GS06, HSS07, HC05, HG00, HM08, HH03, HAW09, KA01, KMSL02, KFB05, LT01b, MKP07, MTWS09, MMP03, Mic06, Mil03, MTSS06, NBMP06, NSC07, NSVT08, NS02, OSS^{+}04, OBB^{+}09, RC04, RJ09, SP07, SLGH08, SLF^{+}05, SML^{+}07, SKL^{+}03, SBM00, TMWG00, TL06, VPN09, WC00, WSRA06, WTW03, WTSC00, WWR^{+}00]. Structured [HB09, HK04a, PAD01, WHS^{+}07]. Structures [CN01, IC02, JT04, SC02, WBH04]. Studies [AAG03, BBRH06, DW08b, LHOP01, MH05, MFC04, NSP03, WM05b]. Study [PG01, YA01, ZTP^{+}08]. Sturgeon [AAP03, ANC^{+}06, AHC^{+}09, BSEM05, BW01, BCDF^{+}09, CG04, CDFW06, CSPP00, CPRS02, CCP^{+}03, DLES06, FPD^{+}09, FHP00, GP05, GBC^{+}05, GGG06, HPM05, HSRD04, HBP^{+}09, HWW^{+}06, HSH^{+}04, ISH07, JKWS06, KFT^{+}09, KVZD02, KHKS00, LB00, LME^{+}08, MC04, MWN05, MSS^{+}00, MGM02, NTB^{+}06, PB03, PBI05, PD05, PPWvdL08, PSS09, PRR03, PSD^{+}08, PAD01, SP03, SK05a, SFS04, SDSB08, SB03b, TCG09, VWD^{+}01, VLD^{+}06, WFF^{+}02, WCM^{+}07, WTB06, WWR^{+}00, WS07b, ZJPB08, SMZB05]. Subadult [DBW09, HB08a, SP03, WHZR03]. Subarctic [SGS06a, SGS06b, TBA^{+}04]. Subchronic [MH02]. Subjected [NDČ^{+}04, SCWP09]. Sublethal [BCT^{+}05, SBC^{+}05]. Subsidies [DRQ09, HNH^{+}04, WHCC03]. Substantially [BHL05]. Substrate [GSP01]. Substrates [MWN05]. Subtropical [BS07, DSSA01]. Subyearling [CMBB01, TBK^{+}09, TKH^{+}09]. Success [BVL^{+}05, BJ01, CD09, DW06, Far01, HB00, HRMC07, KMP03, Mat01, Mir05a, PK02, SKP^{+}08, SCWP09, Sut07, SH01, WJB06, WH07, WB09, ZG07]. Successful [GMGV00]. Sucker [BC07, BHV08, CM03b, JSH^{+}08, MD07, MPK03, WMBM02]. Suckers [BAS^{+}09, CBR^{+}08, GSSB03, MH02, MMH01, MBK05, TMK03, TM06, WG03, WW02]. Suggest [PB03]. Suggests [GVHC^{+}09]. Suisun [MM01]. Suitability [Bas04, HKCP09, JD07, LB00, LBR^{+}01, MB04, MMH01, VO01]. Sulfate [SCP^{+}05]. Summer [AZS07, AKPQ08, ACB02, BK03, BSB09, CEB^{+}05, CF07b, CF07a, CMBB01, CGM03, CM07, ECW^{+}09a, Heg02, HPB06, HF09, HCK09b, IH04, JCB^{+}05, KHT08, KZ06, LJC^{+}07, LFA09, MA09, MBS04, MBC^{+}05, NSF04, PF01b, PLD00, RQ02, SBR^{+}06, SSD^{+}06, SLC06, TA06, VR05, WS08, YI07]. Summer-Fall [WS08]. Summer-Run [HF09, KHT08, RQ02]. Summer-Spawned [CGM03, SBR^{+}06, TA06]. Sunfish [DA08, FRK01, WSA00]. Sunshine [HK04a]. Superior [BCT04, BM07, BS08, DW08a, GMCQ08, HSK03, HL00b, HB08b, MB01b, MM08, MJSY09, PYFS07, SH06, SJFP08, YSC^{+}06, YSB^{+}08]. Supersaturation [BM06, JCB^{+}05, WSSD03a]. Supplemental [HW09]. Supplementation [LBC^{+}04]. Supply [BBV00, BSM^{+}04, RMD07]. Support [MTWS09]. Supportive [OBB^{+}00]. Suppressed [WNF09]. Surface [BSO01, JAJ^{+}00, Rei01]. Surfactants [BH00c]. Surgery [PCHB01]. Surgical [HCK^{+}09a, WSB00]. Surgically [MH05, ZBF05]. Surplus [CND03]. Survey [AJF09, HB04, KAB^{+}05, LJC^{+}07, RBR^{+}00, WLBBM09, WNG02]. Surveys [BR09a, PBT04, WPGH08, Zho02b]. Survival [AZS07, Ano05, AMB^{+}08, BMN04, BDR01, BH07, BAWT05, CH00, CHH02, CB02b, CCK^{+}09, CCR^{+}07, CBMM08, DBW09, EWB^{+}06, ECW^{+}09b, EKS^{+}05, FNL^{+}08, FSR^{+}07, FGWS00, GAH^{+}06, GDWC04, GdMH01, HWLC02, HG07, HCK06, HRMC07, HWHP04, HMT^{+}06, KFT^{+}09, KBPC07, KA07, KWH03, KTNK01, KSR01, LHL06, LGJ02, MAW09, MD07, MGMM01, MBS05, MZ02, MWMT02, MFMB06, MPP05, NDČ^{+}04, PSKH04, PCHB01, PF01b, PF05, PLD00, PA01, PYFS07, PMP^{+}02, PMP05, SMZB01, SBP^{+}08, SW08b, SS08b, SN01, TRS09, WSJ^{+}05, WD08a, WP08, WSEW09, WB09, ZJPB08]. Susceptibility [DCG07, HGJO04]. Suspended [BJ01, KB06, RSC07, Sut07]. Sustainable [Cop02, Sec08]. Sustaining [Rid04]. Suture [WSB00]. Suwannee [HPM05]. Svalbard [GN00]. Swallowed [TAB00]. Swim [BBS^{+}08, CS06, CGD05, HB00, RB05, SA06]. Swimming [AAP03, AVM05, BCT^{+}05, BGM06, GBC^{+}05, HBC^{+}02, KRM03, MBS03b, NC05, NBS07, Pap08, Pea04, RJ00, SM08b, SHDR03, SBC^{+}05, SYC04, WMBM02, WCM^{+}07, ZBF05]. Sympatric [Far01, GVS02, HBJ^{+}08b, KCY^{+}07, MKP07, MB01b, OPH09]. Sympatry [BGGL08]. Symposium [HK08]. Synchronous [HWW^{+}06]. Synchrony [PGW08, ZN07b]. Synergistic [SBC^{+}05]. Synthesis [BBRH06]. System [BHB^{+}07, CCP^{+}03, FSR^{+}07, FHP00, HTT00, HTBT01, KVZD02, Koe04, LMK04, LDOS06, MGG^{+}06, MMM07, MKP07, MMS^{+}06, NTB^{+}06, OO05, OO07, SLGH08, ZM06a]. Systematic [HB09, YSB^{+}08]. Systematics [SK07]. Systems [CHP08, Cop02, HWS^{+}09, KM01, Lon04].

Tactile [WNT03]. Tag [ARC^{+}08, BHP^{+}08, HDL07, LHL06, NBK01, OLWC06, RH01, SB03b, TFO^{+}09, WRC^{+}07, WSLT07]. Tag-Based [WSLT07]. Tagged [BAS^{+}09, BHB^{+}07, CRC^{+}01, GK06, HAS06, HSBC02, LHOP01, MBS03b, PB03, PWSC01, PPR06, RSBF03, KBPC07]. Tagging [BH07, FHML08, HMT^{+}06, JDSS01, LHOP01, NBS07, PJH04]. Tags [CBN^{+}09, MWMT02, REK06]. Tailwater [PK02, SH00a, SHW00, YI07]. Tailwaters [SSRM08]. Taken [SGS05]. Tampa [WLBBM09]. Tandem [SAT^{+}07, SS08a]. Tank [SGSS01]. Tanner [SGS05]. Tapeworm [KA07]. Target [BR09a, CJMB06, HN05, WRK02]. Targeted [SSCM06]. Targets [HHDC03]. Tautog [LS07]. Tautogs [CPA^{+}06]. Taxonomy [SMK^{+}07]. Technique [BE07b, SGS05]. Techniques [BHW03, EKS^{+}05, Gan08, GRR05, OM05, Pol07, WMO03]. Telemetered [NQ02]. Telemetry [BBG02, BEV^{+}07, BGM06, GABC00, GA07, HJP01, NAG07, PJH04, WS07a]. Teleostei [GPUH07]. Temperate [HCW01, JGMC05, SGSS01, TF07]. Temperature [ANC^{+}06, Ano05, BMH07, CEW00, CMMM04, CF07b, CF07a, CPG03, CB05, DH04, EM01, FWL^{+}07, GBC^{+}05, HN00, HF02, HGW04, JR03, KFT^{+}09, KPC08, KA07, LT01a, LDK^{+}04, Mat01, MC04, MBS03a, MZB^{+}07, MBS05, MB05b, OW01, OA07, PP05, PP02, RHM^{+}06, RH02, RV07, SA06, SMZB01, SA04b, SJS09, WBI00, WMBM02, WWM07, WCBS06, WS08, WHC02, ZJPB08]. Temperature-Based [DH04].

Temperature-Dependent [RH02]. Temperatures [CEB^{+}05, COB^{+}03, GKB^{+}06, IH04, SCM01]. Temporal [ACB02, BBV00, BJ03, BSM^{+}04, BE07a, BSB09, CSvdS02, DF08, DRG09, DCT02, HB08a, LLVB04, LDDB08, MKP07, MOT04, MM03, MBK05, OMM09, PCWC08, SW08a, SLF^{+}05, SK08, VFT02, WORB00]. Temporary [HPB06]. Ten [HGJO04]. Tennessee [SB05, TH00, SBG01a]. Term [LS07, LMA09, MH05, NBS07, QK03, SRL^{+}06, SLSF07, SS05, HSH^{+}05]. Terminals [GGCF07]. Tern [ARL^{+}05]. Terrestrial [RL08, SF07, SH08, URMH07]. Territoriality [PG07]. Territory [MGP06]. Test [HSS00, HDH00, LHOP01, MZB^{+}07, MD03, MSM^{+}08, QH05, SHC01, WHP01, WNT03]. Tested [GP08]. Testing [BSS05, CB04, DH04, HC08]. Tetrodotoxin [DWEL08]. Texas [MMM07, DW09, MMCR08, SW08a, Sch00a, SLC00]. Texoma [MMM07, SW08a]. Their [ASK01, BBM02, BHL05, BLBH05, DA00, HOA04, KBW^{+}00, OR06, OW01, PW09, SHS04, WSRA06, Wil07, HB09, WHT^{+}00, ZSW02]. Theoretic [KR07, PF05]. Theory [Pol07, RJ09]. theOzarksRegion [KGS00]. There [HK04b]. Thermal [Ano05, BMZ07, Bet05, CM01, CWB06, DCCA01, GAH^{+}06, HF02, JR03, MGP06, NA04, SRJ03, SMZB01, WWS03, WWM07, WSFJ05, WCBS06, vT05]. Thermally [IH04, MWMT02]. Thermoregulation [BK03, GKB^{+}06]. Thiaminase [HHB02, RE08]. Thiaminase-Positive [HHB02]. Thiamine [KBW^{+}00]. Threadfin [VBT08]. Threatened [BHSR09, STC^{+}09, WTW^{+}00, WAS^{+}00]. Three [BM09, BBDM05, Chi09, DWEL08, DA05, DD02, DA08, KBJ08, KGS00, MHHS01, SKC^{+}05, SJ00, TAC08, WHS^{+}07, WBH^{+}05]. Thresholds [DM03, SBP^{+}08]. Throughput [SGS05]. thuringiensis [JHS02]. Thurmond [YI02]. Thyroid [CC03]. Thyroidal [CC03]. Tidal [MA09, OO05, WLB^{+}00]. Tide [SCM09]. Tiger [CEW00]. Tilefish [HWM04]. Time [MFOA09, MS03, PLO07, WWM07]. Timing [AZS07, AG02, BGPL08, BQH04, CBN^{+}09, DGFS04, EKS^{+}05, KPC08, KWH03, QPG^{+}02, RQ02, SJS09]. Tissue [BBP^{+}04, BSW00, JRM^{+}08, LMAF08, RSB^{+}08]. Tissue-Specific [JRM^{+}08]. Tissues [JGMC05, MCS08]. Toadfish [FT08]. Tolerance [Ano05, BKCL05, JR03, LT01a, MB05b, Pir04, SMZB01, WWM07, WHLP07, WCBS06]. Tolerances [CWB06, CB05, Mat01, OW01]. Tongyeong [CMKL01]. Tool [HSS07, LMR08]. Top [CDF^{+}08]. Topeka [KA07]. Total [BM06, EKSX08, RWB^{+}08, WSSD03a]. Tournaments [FCS^{+}03, KSC^{+}06, MSET05, SKC^{+}04]. Towboats [GDW03]. Towed [Hol08]. Toxicity [BCT^{+}05, BWVN09, BH00c, MH02]. Trace [AHC^{+}09, BBS07, HBBG07, SW08a, WTJ00]. Trace-Element [SW08a]. Tracking [BFHO05, BHB^{+}07, BE07b, HBO07, JMS02]. Trade [AVM05, HDH00]. Trade-Off [AVM05]. Trade-Offs [HDH00]. Tradeoff [QVPG04]. Training [WBH^{+}05]. Traits [DWBQ05, FSH07, GM04, KSB^{+}06, KSB^{+}08, MN07, MISS07, OW09, PCMS05, WORB00]. Transfer [BBF^{+}03, MD03]. Transformer [ZM05]. Transgenic [BDLB04, CB05, TAD05]. Transient [GJP09, WLB^{+}00]. Transintestinal [MH05]. Transition [Wei08]. Transmitter [PCHB01, SB03b, ZBF05]. Transmitters [BCW^{+}09, HCK^{+}09a, LKM04, MH05, PAR01, WBI00]. Transponder [BHB^{+}07, BH07, FHML08, HDL07, NBS07, REK06]. Transponders [CRC^{+}01, RSBF03]. Transport [JPLSH09, MRC^{+}09, MB01a]. Transportation [ASK01, CCK^{+}09, CLSD00]. Transported [MMS^{+}06]. Trap [NS04]. Trap-Caught [NS04]. Traps [HBSA09, SJ04]. Trash [FCC07]. Trauma [MWM00, NS04]. Trawl [HB04, LAM04, LJC^{+}07]. Trawling [WCC08]. Trawls [SMG03]. Treated [GBMB01, ZM06a]. Tree [SZSS08]. Trend [DRG09]. Trends [TAM07]. Trevally [MH05]. Tributaries [CCR^{+}07, FSH07, HWN02, MLSD07, MMM07, MA09, WHC02]. Tributary [BS03, BBWW02, HL00b, HB08b, RSP09, ZM06a]. Tricolor [BJ01]. Trinity [MDE^{+}09]. Triploid [CGAD^{+}02, WSA00]. Trophic [BV01, BR09b, CB00, HC05, HSK03, JMS02, MDE^{+}09, MBMO06, OLS04, OBB^{+}09]. Tropical [AMRM02, GP08, NN06, WNH05]. Trout [AFR00, AFR02, ACB08, ANE06, Ano05, BK03, BBM02, BLS07, BBG02, BS03, BHL05, BMZ07, BV01, Beh02, BBS^{+}08, BDD07, BSO01, BFHO05, BHMK05, BLG01, BCV07, BWVN09, BM07, BS08, BPM^{+}04, BR09b, BTLD08, BHS05, BH00c, BIB^{+}00, CDP02, CRA05, CM01, CM03a, Car06, CN01, CCMP06, CJFM05, CF07b, CF07a, CSBM06, CWT06, CJMB06, CSL09, CSvdS02, DW08a, DRG09, DHR07, DKP^{+}06, DHC04, DK06, DCG07, DB08, EM04a, FWTS08, FHML08, FRB06, FB08, GMGV00, GSS03, GMCQ08, GVS02, GSP01, GH07, GF07, GHL02, HW09, HMD02, HWLC02, HF04, HC08, HSK03, HNW06, HR09, HSS00, HB07, HBJ^{+}08a, HBJ^{+}08b, Heg02, HK00, HK04b, HB08a, HKHC00, HB08b, IH04, IK00, JM09, JR03, KA01, KLA02, KLSA02, KMSL02, KC05, KC06, KM06a, KM06b]. Trout [KYWC09, KC09, KCY^{+}07, LDDB08, LC03, MGP06, MPVM00, MISS07, ML07, MB06, MZB^{+}07, MKP07, MD05, MKDB03, MHMP05, MSEL03, MSL^{+}06, MZ02, MK05, MB01b, MEA^{+}09, MVS07, MM08, MMTF05, MTSS06, MMBG09, MR00, NDR^{+}09, NBS07, NA04, NS02, NL03, NQ02, OEO^{+}00, OR06, PSB^{+}03, PSBC04, PSB^{+}05, PBR^{+}01, PKR02, PK04, PAS09, PK02, PLM05, PPDV09, PWLL05, PYFS07, PJC07, QH05, RP03, RMRT03, Rid08, RIA^{+}07, RR08, RT04, RV07, RC04, RDU02, RQ01, RHV03, SF07, SBW00, SRJ03, SKA^{+}09, SMZB01, SBP^{+}08, SH00a, SHW00, SHDR03, SH06, SJFP08, SML^{+}07, SBO05, SBSO06, Spi01, SSCM06, SSHB02, SH01, SCH04b, SH08, TB08, URMH07, VHF05, VPN09, VBM^{+}09, WSB00, WLK03, WEB00, WSJ^{+}05, WWM07, WPS02, WRC^{+}03, WB01, WG03, WH07, WR08, WSRA06, WRGN07, WB09]. Trout [WNT03, ZBF05, ZTP^{+}08, ZR02, ZKE06, ZV07, ZEP09, ZN07a, ZN07b]. Trout/Steelhead [BBM02]. True [MTSS06, PB03]. Truss [KAB09]. Truss-Based [KAB09]. tuberculatus [MB05b]. Tuna [Ess03]. Tundra [LM08b]. Turbid [HW08b, QHR04]. Turbid-River [QHR04]. Turbidity [BW02, GP05, HR09, SW09a, SH01, WH07, ZG07]. Turbine [GM02]. Turbines [CW00b, FACS06]. Turbulence [CWT06, SBO05, TKH^{+}09]. Tustumena [WORB00]. Twenty [MMP02]. Twenty-One [MMP02]. Twice [PHF07]. Two [AFR02, ACB02, BMH07, BL01, BD07, CSPP00, CST^{+}05, GVHC^{+}09, GMP^{+}05, HAS06, JDSS01, KM05, KHKS00, Lin01, MGP06, MD07, MN07, MA09, MB04, Mea05, NKJ03, PMRH06, PAR01, SCH03b, SW08b, SYC04, VR05, WORB00, WHC02]. Two-Pass [Mea05]. Two-Vector [SYC04]. Two-Year [BD07]. Type [BGPL08, CB05, HH03, MM03, MMS^{+}06, NPT04, ROC^{+}03, WSB00]. Types [BLBH05, CN08, CRC^{+}01, CST^{+}05, MKDB03, NSC07, TMWG00]. Typology [CKB01].

U.S [HWW^{+}07]. U.S. [LT01a, MWG^{+}08, MC09, PF01a, PHWS09, RMD03, SKA^{+}09, TL06, WKM^{+}07]. ui [SRB00, SR07]. Uinta [HKHC00]. Ulceration [EKSX08]. Ultraoligotrophic [VBM^{+}09]. Ultrasonic [BBG02, HCK^{+}09a]. Ultraviolet [HGW04, OADC^{+}08]. Uncertainty [KJM^{+}08, Lon04]. Undercut [MR00]. Underestimate [PB03]. Underestimation [YSB^{+}08]. Undergoing [WBH06]. Underlying [OYB03]. Underreported [GVHC^{+}09]. Understanding [BHC03, Bur09, SMK^{+}07, YSB^{+}08]. Undocumented [BD07]. Unexploited [Car06]. Unimpounded [BHHS04, CCK^{+}09]. Unionid [HKCL07]. Unique [Ano03d]. Unit [RWB^{+}08, iTE08, WPGH08]. Unit-Count [WPGH08]. United [KBJ08, RMD07, ZTP^{+}08, HWM04, MB05b, SFS04]. Units [PR01b, PR01c]. Unraveling [MMS^{+}06]. Unregulated [PRR03, SP07]. Untagged [MBS03b]. Untreated [ZM06a]. Upland [BE07b, McD03, Sut07]. Upper [AW05, CWB06, CPF06, WCBS06, BAS^{+}09, BHHS04, Bro05, BHV08, BKHH07, CCMP06, CGL^{+}09, DLES06, Far01, GS06, JSH^{+}08, KVZD02, KSB^{+}06, KSB^{+}08, Koe04, MC00a, MC00b, MD07, MRC^{+}09, PWLL05, TMK03, WJMM07, ZDK^{+}04]. Upstream [BL05, GS06, HBH^{+}08, RHB^{+}06, WCM^{+}07]. Urban [MB08b, OBB^{+}09, WLK03]. Urbanization [BLGW06]. Urchins [SWHW06]. USA [PB03, CG08, CB06, KVZD02, Leg05, MOSB02, NBMP06, SKC^{+}05, Sto02, WHLP07, WHS^{+}07, WNG02]. Use [BK03, BJ03, BCJ^{+}06, BSM^{+}08, BLBH05, BIT00, BW01, BBWW02, BMH06, BCV07, BGM06, BZW04, BHV08, CBN^{+}09, CPRS02, CGL^{+}09, CMSM01, DHS06, FT08, GABC00, GI06, GSP01, GA07, GSSB03, HPW08, HF04, HW08b, HL00b, HJP01, HBBG07, HBO07, HCK09b, HSH^{+}04, JTM08, JKWS06, KC09, LMK04, LBBF^{+}08, LPJS08, MGP06, MPVM00, MOP02, ME07, MZ02, MMH01, MBK05, MM08, NAG07, OYB03, PSSV^{+}09, PF01b, Pea04, Pol07, QM05, Ron02, RMD08, RL08, SIH02, Sch00b, SM08a, SMZB05, SCH03b, SHW00, SK05a, SSSS05, SBSO06, SFS04, TRS09, TRK07, VDG02, VR05, VT07]. Used [BH00b, KAB^{+}05, KHKS00, MWG^{+}08]. Uses [Sch00b]. Using [AHC^{+}09, AWNT00, BWS^{+}05, BLC^{+}04, BCM^{+}05, BBL06, EKS^{+}05, FSN^{+}03, GKW^{+}06, HAS06, HF09, HBO07, HMT^{+}06, IMM^{+}04, JMS02, LMA09, MCS08, Mat01, MB08b, NBK01, OJ01, PWDG07, RMH^{+}08, SSR06, SSHB02, TR07, VPN09, WLBBM09, WFF^{+}02, WPS02, WKM^{+}07, YI07, BCDF^{+}09, DSH^{+}09, DFBC^{+}09, HB07, LCNB^{+}06]. Utah [RWL01, VBT08, BHW^{+}07, dlHFB04, HK00, MMH01, Wyd01, WW02]. Utilization [CSPP00, YI02].

V [WJF03]. V-Lip [WJF03]. Valdez [CRMN04, SHT^{+}00, WHT^{+}00]. Validating [BGS^{+}04]. Validation [BCDF^{+}09, DFBC^{+}09, HPW08, HAS06, KJM^{+}08, LPN01, OJPN02, PSJ08, SRL^{+}06, SRH07]. Valley [BWS08, CGL^{+}09, ML04, VT07, ZEP09, WM07]. Valuable [STC^{+}09]. Value [PK04]. Values [WHLP07]. Vancouver [VGA01]. Varden [BBWW02, DRQ09, OPH09]. Variability [AG02, FSH07, GWM^{+}07, Har00, HB08a, JGMC05, KAB^{+}05, KM01, Lon04, MHR^{+}05, MM03, PCWC08, SW08a, SAT^{+}07, SK08, WBH^{+}05, ZWT05]. Variable [GAH^{+}06, HW09, HCK06, LFA09, MAW09, MKDB03, PHF07, SBR^{+}06, SCM01, WBH04]. Variables [HAW09, LJE08, MM02, SK05b, SZSS08, TL06, WC00, Zho02a]. Variation [ANE06, BCS^{+}01, BLC^{+}04, BCM^{+}05, BCJ^{+}06, BVMM06, BSB09, BH03, BM07, BWFB06, CN08, CPA^{+}06, DRG09, DS04, DCT02, ECW^{+}09b, FWG07a, FWG07b, GFL00, GFLF02, HGR^{+}08, HB07, HBJ^{+}08a, HWMH02, HB06, HAW09, JM09, KGS07, Koe04, KGS00, LLVB04, LGJ02, MSS^{+}00, MCR05, MVD^{+}00, Mil04, NBMP06, NSC07, OSS^{+}04, PT03b, SMG03, SBG01b, SW09b, SBM00, TB08, VFT02, WB08, WWS03, WTJ00, WSJ06, Whi00, WTSC00, WORB00]. Variations [JN00b, PD05]. Various [BC02]. Varying [CS06, MWG^{+}08, MJS^{+}03, SH00b]. Vastly [ML04]. Vaterite [VPN09]. Vector [SYC04]. Vegetated [SCH03b]. Vegetation [BBF^{+}03, OBW04, TRK07]. Vegetation-Dwelling [TRK07]. Velocity [AVM05, CS06, FCC07, Pea04, RJ00, SBSO06, TKH^{+}09, WJF03]. Verde [Bro05]. Verification [NBK01, SRL^{+}06]. Vermont [ZM05]. versus [LFG06, MB04, Mea05, TRGV00]. Vertebrate [CB00, KR07, PHWS09]. Vertical [EBC01, NQ02, PD05, TF01]. via [ASK01]. Viability [FPD^{+}09, Leg05, PHF07, PAD01]. Victoria [WWKM07]. Vincent [XM00]. Virginia [AO09, BMH06, CM03a, CM07, CND03, CN05b, GLHA06, GA03, HO05, HBO07, LHOP01, MA09, OLWC06, SA01, SN01]. Virus [MG06, WNG02]. Vital [ACB08]. Vitamin [WC05]. Vitellogenin [HBP^{+}09]. Volitional [SBSO06]. Volume [Bur09, MD07]. Vulnerability [CRC^{+}01, PCC^{+}09, RSBF03, SP04].

Wabash [GS06]. Walleye [BHC03, BWFB06, CMMM04, CGAD^{+}02, LDK^{+}04, LMA09, MWSP09, MFFI09, RTH^{+}01, SK08, VAR^{+}02]. Walleyes [BGGL08, GW03, GFL00, GFLF02, HWHP04, KJL^{+}07, KSC^{+}06, KWH03, LRF^{+}08, PF02, RHG^{+}02, WRC^{+}07, WKPS05]. Warmer [FRB06]. Warming [RIA^{+}07, SFE01]. Warmwater [AW05, LPN01, SM08b, SP07, WTW^{+}00, WAS^{+}00]. Was [GKW^{+}06].

Washington [NQ02, SJP05, TWF^{+}07, Ano06b, BSM^{+}04, BBF^{+}03, BCV07, CPF06, FP04, GJP^{+}05, HEW^{+}05, HHK04, JAJ^{+}00, KHT08, KFB^{+}06, LYS06, MB08b, MBMO06, MKP07, NQ02, OW09, Ron02, TWF^{+}07]. Water [BHW03, BBW03, CD09, CMMM04, CPB07, COB^{+}03, DA05, FCC07, GBC^{+}05, GKB^{+}06, JTM08, Pea04, PP05, QB01, RV07, SARC09, SA06, SIH02, SGG09, SBH00, SYC04, TMK03, TKH^{+}09, TP01, WRC^{+}03, WJF03, WHC02]. Waterbird [CRC^{+}01]. Waterbirds [CRC^{+}02]. Watercourses [CGL^{+}09]. Waterfall [BHMK05]. Waterfalls [KM05]. Waters [BBHD09, BS08, DBW09, GKW^{+}06, LCM07, LAM04, LYS06, LBBF^{+}08, SP03, SH06, SLC00, SSO09, WKM^{+}07]. Watershed [FSN^{+}03, dlHFB04, HBH^{+}08, MB01a, MDE^{+}09, MKDB03, MMBG09, PR01a, PLM05, RMRT03, RC04, SJ08, WSFJ05, YW07]. Watershed-Level [SJ08]. Watersheds [SS06, WR08]. Weakfish [WTJ00]. Wear [FSMH08]. Weathered [WHT^{+}00]. Web [BBV00, SVW02]. Webs [WHC04]. Weekly [PA01]. Weigh [SKC^{+}04]. Weigh-In [SKC^{+}04]. Weight [DW08b, ECW^{+}09a, KAB^{+}05, MVD^{+}00, SBH00, ZJPB08]. Weighted [YA01]. Weighting [AAG03]. Weights [ZBF05]. West [CM03a, RSB^{+}08, Sch00b, TTW^{+}09]. West-Central [Sch00b]. Western [AHJ^{+}05, CW04, MBS05, TFO^{+}09, XM00, BBF^{+}03, BWFB06, HHK04, JSI09, MB08b, MWG^{+}08, NBK01, PHWS09, RM04, Ron02, RTH^{+}01, RTBH05, SKA^{+}09, TRBH08, TK01, WHLP07, WHS^{+}07, WBH^{+}05]. Westslope [BMZ07, BPM^{+}04, KCY^{+}07, MMTF05, MMBG09, OR06, WPS02, WRC^{+}03, ZBF05]. Wetland [CBR^{+}08, MMH01, SCO^{+}02]. Wetlands [BCT04, MMCR08, TBK^{+}09]. White [AD08, ASK01, BMH07, BSS05, GP05, GBC^{+}05, HN05, ISH07, MKBF00, MWN05, MBS03a, MA09, PB03, PBI05, PD05, PPWvdL08, PRR03, SBG01a, SA04b, WFF^{+}02, WCM^{+}07, WTB06]. Whitefish [MOP^{+}06, Wyd01]. Whole [EH03, HSS00, JRM^{+}08, PLH^{+}08]. Whole-Body [JRM^{+}08]. Whole-Fish [PLH^{+}08]. Whole-Lake [EH03]. Whole-Lifetime [HSS00]. Width [BNG08]. Widths [SJ04]. Wild [AZS07, AMB^{+}08, BS03, BSLN08, BH04, BBDM05, BKHH07, CM01, CLSD00, CMBB01, CG06, CB05, DSH^{+}09, DYWA^{+}04, DKP^{+}06, DWBQ05, FP04, HFHR00, Heg02, HAS06, HWMH02, KSB^{+}06, KSB^{+}08, KZ06, LBS^{+}06, MBK05, NSVT08, PSBC04, Sch05, SKP^{+}08, SH06, SBH00, WFF^{+}02, WF05, WPS02, Wei08, WSJ06]. Wild-Type [CB05]. Wilderness [HCK09b]. Wildfire [NDR^{+}09]. Will [WB08]. Willamette [CGL^{+}09, SS06, TBK^{+}09, Sch02, WC00]. William [BH03, BH04, CBMM08, HE00, HE01, WSJH01]. Williamson [CBR^{+}08]. Willow [SGG09]. Winchester [MS03]. Windows [RHB^{+}06]. Winnebago [SDSB08]. Winter [BG04, CH00, FLLA09, FGWS00, HF04, HSBC02, JM05, KZ06, MKP07, MJS^{+}03, MCR05, MCS06, SW03a, SLK01, SH00a, SHW00, SW03b, SEP^{+}08, SN01, WH07, WS08]. Winterkill [DT03]. Winters [MBS03a]. Wire [MWMT02, TFO^{+}09]. Wisconsin [BHC03, BSEM05, GF07, LPN01, RHG^{+}02, RVJ01, SDSB08, WLK03]. within [BSB09, CHH^{+}06, CM07, FWG07a, GS06, ISH07, OR06, PR01a, PJC07, SJP05, ZTP^{+}08]. without [PBI05]. Wood [MB01a, RQ01, WM05a]. Work [VO01]. World [HR09, WM05a]. Wound [PCHB01, WSB00]. Wyoming [HF04, QHR04, RR08, SF07, SH00a, SHW00].

XY [WM07].

Y-Chromosome-Specific [BT04b]. Yakima [BKHH07, KSB^{+}06, KSB^{+}08, BLS^{+}00, BGPL08, FP04]. Year [BD07, BG04, CP03, DHR07, EM04b, GLHA06, HCW01, HSC00, IK00, JN00a, JN00b, KRM03, LT01b, MCS06, PR06, PF02, RCB^{+}04, SBR^{+}06, SS08b, SN01, TB08, VFT02, WKCC08, WEB09, WS07b, ZJPB08]. Year-Class [JN00b, WKCC08, WEB09, WS07b]. Yearling [SK05a]. Yearly [SH08]. Years [AMB^{+}08, CBMM08, HG07, MMP02]. Yellow [BCT04, BR09b, CGD05, EH03, FLLG00, GDWC04, GMJ^{+}06, HCW01, HGW04, IRJ^{+}09, Mil03, MS04, PSSV^{+}09, PCMS05, TK01, VAR^{+}02]. Yellow-Phase [MS04]. Yellowstone [DG04, BW01, CDP02, CCMP06, FS06, KLSA02, MSEL03, MSL^{+}06, SRL^{+}06]. Yield [BOMC07]. Yields [STC^{+}09]. York [HDRB07, BK03, BL00, BLS07, HHS04, HBO07, LRF^{+}08, MLSD07, MA09, OLWC06, SLM07, VAR^{+}02, WS07b]. Young [BBRH06, CP03, EM04b, GLHA06, HBO07, HCW01, HSC00, IK00, KRM03, LT01b, MCS06, PSSV^{+}09, PR06, PF02, RCB^{+}04, SBR^{+}06, TB08, ZJPB08]. Young-of-the-Year [EM04b, GLHA06, HCW01, HSC00, KRM03, LT01b, MCS06, PR06, RCB^{+}04, SBR^{+}06]. Young-of-Year [CP03, IK00, PF02, TB08, ZJPB08]. Yukon [BBS07, FWG07a, FWG07b, MGP06, OSS^{+}04].

Zander [KJM^{+}08]. Zealand [McD03]. Zebra [CB05]. Zinc [FSN^{+}03]. Zone [BL02, HMD02]. Zooplanktivorous [EBS02]. Zooplanktivory [CB00]. Zooplankton [BHS05, CB04, DA05, MW06, RMH^{+}08, SBG01a, SH00b]. Zulega [GK06].

References

Albanese:2003:DMR

[AAG03] Brett Albanese, Paul L. Angermeier, and Charles Gowan. Designing mark–recapture studies to reduce effects of distance weighting on movement distance distributions of stream fishes. Transactions of the American Fisheries Society, 132(5):925–939, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:2003:CSS

[AAP03] S. Reid Adams, Ginny L. Adams, and Glenn R. Parsons. Critical swimming speed and behavior of juvenile shovelnose sturgeon and pallid sturgeon. Transactions of the American Fisheries Society, 132(2):392–397, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arndt:2002:ESF

[ACB02] Steven K. A. Arndt, Richard A. Cunjak, and Tillmann J. Benfey. Effect of summer floods and spatial-temporal scale on growth and feeding of juvenile Atlantic salmon in two New Brunswick streams. Transactions of the American Fisheries Society, 131(4):607–622, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Chokhachy:2008:DCP

[ACB08] Robert Al-Chokhachy and Phaedra Budy. Demographic characteristics, population structure, and vital rates of a fluvial population of bull trout in Oregon. Transactions of the American Fisheries Society, 137(6):1709–1722, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aalbers:2008:WSS

[AD08] Scott A. Aalbers and Mark A. Drawbridge. White seabass spawning behavior and sound production. Transactions of the American Fisheries Society, 137(2):542–550, February 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Altenbach:2000:NDC

[ADP00] Christopher S. Altenbach, Robert K. Dudley, and Steven P. Platania. A new device for collecting drifting semibuoyant fish eggs. Transactions of the American Fisheries Society, 129(1):296–300, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:2000:MNB

[AFR00] Susan B. Adams, Christopher A. Frissell, and Bruce E. Rieman. Movements of nonnative brook trout in relation to stream channel slope. Transactions of the American Fisheries Society, 129(3): 623–638, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:2002:CDN

[AFR02] Susan B. Adams, Christopher A. Frissell, and Bruce E. Rieman. Changes in distribution of nonnative brook trout in an Idaho drainage over two decades. Transactions of the American Fisheries Society, 131(4):561–568, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Antonsson:2002:VTC

[AG02] Thorolfur Antonsson and Sigurdur Gudjonsson. Variability in timing and characteristics of Atlantic salmon smolt in Icelandic rivers. Transactions of the American Fisheries Society, 131(4): 643–655, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:2009:UTE

[AHC^{+}09] Peter J. Allen, James A. Hobbs, Joseph J. Cech, Jr., Joel P. Van Eenennaam, and Serge I. Doroshov. Using trace elements in pectoral fin rays to assess life history movements in sturgeon: Estimating age at initial seawater entry in Klamath River green sturgeon. Transactions of the American Fisheries Society, 138(2): 240–250, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amadio:2005:FAO

[AHJ^{+}05] Craig J. Amadio, Wayne A. Hubert, Kevin Johnson, Dennis Oberlie, and David Dufek. Factors affecting the occurrence of saugers in small, high-elevation rivers near the western edge of the species’ natural distribution. Transactions of the American Fisheries Society, 134(1):160–171, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aday:2003:DIE

[AHW03] D. Derek Aday, R. John H. Hoxmeier, and David H. Wahl. Direct and indirect effects of gizzard shad on bluegill growth and population size structure. Transactions of the American Fisheries Society, 132(1):47–56, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashford:2009:LRI

[AJF09] Julian R. Ashford, Cynthia M. Jones, and Lynn Fegley. A license registry improves sampling efficiency for a marine recreational survey. Transactions of the American Fisheries Society, 138(5):

984–989, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:2008:SDG

[AKPQ08] Joseph H. Anderson, Peter M. Kiffney, George R. Pess, and Thomas P. Quinn. Summer distribution and growth of juvenile Coho salmon during colonization of newly accessible habitat. Transactions of the American Fisheries Society, 137(3):772–781, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:2008:ISF

[AMB^{+}08] Janet L. Armstrong, Katherine W. Myers, David A. Beauchamp, Nancy D. Davis, Robert V. Walker, Jennifer L. Boldt, John J. Piccolo, Lewis J. Haldorson, and Jamal H. Moss. Interannual and spatial feeding patterns of hatchery and wild juvenile pink salmon in the Gulf of Alaska in years of low and high survival. Transactions of the American Fisheries Society, 137(5): 1299–1316, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aguilera:2002:MDA

[AMRM02] Carlos Aguilera, Roberto Mendoza, Gabino Rodríguez, and Gabriel Márquez. Morphological description of alligator gar and tropical gar larvae, with an emphasis on growth indicators. Transactions of the American Fisheries Society, 131(5):899–909, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:2006:GLJ

[ANC^{+}06] Peter J. Allen, Mary Nicholl, Stephanie Cole, Amy Vlazny, and Joseph J. Cech, Jr. Growth of larval to juvenile green sturgeon in elevated temperature regimes. Transactions of the American Fisheries Society, 135(1):89–96, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Almodovar:2006:SVB

[ANE06] Ana Almodóvar, Graciela G. Nicola, and Benigno Elvira. Spatial variation in brown trout production: The role of environmental factors. Transactions of the American Fisheries Society, 135(5): 1348–1360, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2001:E

[Ano01] Anonymous. Errata. Transactions of the American Fisheries Society, 130(6):1240, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2002:BR

[Ano02a] Anonymous. Book review. Transactions of the American Fisheries Society, 131(4):589, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2002:E

[Ano02b] Anonymous. Errata. Transactions of the American Fisheries Society, 131(4):796, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2002:GEF

[Ano02c] Anonymous. Geostatistics for estimating fish abundance. Transactions of the American Fisheries Society, 131(5):1004–1005, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2002:IFE

[Ano02d] Anonymous. Inland fisheries ecology and management. Transactions of the American Fisheries Society, 131(5):1005–1006, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:BHC

[Ano03a] Anonymous. Biology of the hard clam. Transactions of the American Fisheries Society, 132(1):174–175, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:CFN

[Ano03b] Anonymous. The cichlid fishes: Nature’s grand experiment in evolution. Transactions of the American Fisheries Society, 132 (1):173–174, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:E

[Ano03c] Anonymous. Errata. Transactions of the American Fisheries Society, 132(4):824, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:FSU

[Ano03d] Anonymous. Fisheries science: The unique contributions of early life stages. Transactions of the American Fisheries Society, 132 (5):1033, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:ISO

[Ano03e] Anonymous. Island in the stream: Oceanography and fisheries of the Charleston bump. Transactions of the American Fisheries Society, 132(1):172–173, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:MFE

[Ano03f] Anonymous. Marine fisheries ecology. Transactions of the American Fisheries Society, 132(3):619, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:MQM

[Ano03g] Anonymous. Modelling and quantitative methods in fisheries. Transactions of the American Fisheries Society, 132(5):1033– 1034, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2003:SKN

[Ano03h] Anonymous. The sea knows no boundaries: a century of marine science under ICES. Transactions of the American Fisheries Society, 132(3):620, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2004:E

[Ano04] Anonymous. Errata. Transactions of the American Fisheries Society, 133(1):235, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2005:EET

[Ano05] Anonymous. Erratum: Effect of Temperature on Growth and Survival of Bull Trout, with Application of an Improved Method for Determining Thermal Tolerance in Fishes. Transactions of the American Fisheries Society, 134(2):548, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [SMZB01].

Anonymous:2006:EDP

[Ano06a] Anonymous. Erratum: Downstream Passage of Steelhead Kelts through Hydroelectric Dams on the Lower Snake and Columbia Rivers. Transactions of the American Fisheries Society, 135(1): 268, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [WE05].

Anonymous:2006:EEE

[Ano06b] Anonymous. Erratum: Effects of Environmental Conditions during Stream, Estuary, and Ocean Residency on Chinook Salmon Return Rates in the Skagit River, Washington. Transactions of the American Fisheries Society, 135(2):456, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [GJP^{+}05].

Anonymous:2008:BR

[Ano08] Anonymous. Book reviews. Transactions of the American Fisheries Society, 137(5):1567–1570, September 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aunins:2009:MSA

[AO09] Aaron Aunins and John E. Olney. Migration and spawning of American shad in the James River, Virginia. Transactions of the American Fisheries Society, 138(6):1392–1404, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Apse:2006:BR

[Aps06] Colin D. Apse. Book review. Transactions of the American Fisheries Society, 135(6):1578–1579, November 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adlerstein:2008:SMC

[ARC^{+}08] Sara A. Adlerstein, Edward S. Rutherford, Randall M. Claramunt, David F. Clapp, and John A. Clevenger. Seasonal movements of Chinook salmon in Lake Michigan based on tag recoveries from recreational fisheries and catch rates in gill-net assessments. Transactions of the American Fisheries Society, 137(3): 736–750, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:2008:SFH

[ARJ08] Katie A. Anderson, Rodney A. Rountree, and Francis Juanes. Soniferous fishes in the Hudson River. Transactions of the American Fisheries Society, 137(2):616–626, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Antolos:2005:CTP

[ARL^{+}05] Michelle Antolos, Daniel D. Roby, Donald E. Lyons, Ken Collis, Allen F. Evans, Mike Hawbecker, and Brad A. Ryan. Caspian tern predation on juvenile salmonids in the Mid-Columbia River. Transactions of the American Fisheries Society, 134(2):466–480, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:2007:FSP

[AS07] Susan B. Adams and David A. Schmetterling. Freshwater sculpins: Phylogenetics to ecology. Transactions of the American Fisheries Society, 136(6):1736–1741, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allyn:2001:ECT

[ASK01] M. Lizabeth Allyn, Robert J. Sheehan, and Christopher C. Kohler. The effects of capture and transportation stress on white bass semen osmolality and their alleviation via sodium chloride. Transactions of the American Fisheries Society, 130(4):706–711, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aday:2005:PCR

[ASN^{+}05] D. Derek Aday, Daniel E. Shoup, Justin A. Neviackas, Jeffrey L. Kline, and David H. Wahl. Prey community responses to bluegill and gizzard shad foraging: Implications for growth of juvenile largemouth bass. Transactions of the American Fisheries Society, 134(5):1091–1102, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Angermeier:2000:RFC

[ASS00] Paul L. Angermeier, Roy A. Smogor, and Jay R. Stauffer. Regional frameworks and candidate metrics for assessing biotic integrity in Mid-Atlantic highland streams. Transactions of the American Fisheries Society, 129(4):962–981, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Asaeda:2005:EFV

[AVM05] Takashi Asaeda, Trung Kien Vu, and Jagath Manatunge. Effects of flow velocity on feeding behavior and microhabitat selection of the stone moroko Pseudorasbora parva: a trade-off between feeding and swimming costs. Transactions of the American Fisheries Society, 134(2):537–547, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arrington:2002:PES

[AW02] D. Albrey Arrington and Kirk O. Winemiller. Preservation effects on stable isotope analysis of fish muscle. Transactions of the American Fisheries Society, 131(2):337–342, March 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:2005:RWF

[AW05] Susan B. Adams and Melvin L. Warren, Jr. Recolonization by warmwater fishes and crayfishes after severe drought in upper coastal plain hill streams. Transactions of the American Fisheries Society, 134(5):1173–1192, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aoyama:2000:DCE

[AWNT00] J. Aoyama, S. Watanabe, M. Nishida, and K. Tsukamoto. Discrimination of catadromous eels of genus Anguilla using polymerase chain reaction–restriction fragment length polymorphism analysis of the mitochondrial 16S ribosomal RNA domain. Transactions of the American Fisheries Society, 129(3):873–878, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Achord:2007:MTG

[AZS07] Stephen Achord, Richard W. Zabel, and Benjamin P. Sandford. Migration timing, growth, and estimated parr-to-smolt survival rates of wild Snake River spring–summer Chinook salmon from the Salmon River Basin, Idaho, to the Lower Snake River. Transactions of the American Fisheries Society, 136(1):142–154, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:2002:CEB

[BA02] Peter B. Bayley and Douglas J. Austen. Capture efficiency of a boat electrofisher. Transactions of the American Fisheries Society, 131(4):435–451, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartholow:2001:BR

[Bar01] J. M. Bartholow. Book review. Transactions of the American Fisheries Society, 130(2):350, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bashey:2004:CSA

[Bas04] Farrah Bashey. A comparison of the suitability of alizarin red S and calcein for inducing a nonlethally detectable mark in juvenile guppies. Transactions of the American Fisheries Society, 133(6): 1516–1523, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Banish:2009:DHA

[BAS^{+}09] Nolan P. Banish, Barbara J. Adams, Rip S. Shively, Michael M. Mazur, David A. Beauchamp, and Tamara M. Wood. Distribution and habitat associations of radio-tagged adult Lost River suckers and shortnose suckers in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 138(1):153–168, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Briscoe:2005:BFA

[BAWT05] Ryan J. Briscoe, Milo D. Adkison, Alex Wertheimer, and S. G. Taylor. Biophysical factors associated with the marine survival of Auke Creek, Alaska, Coho salmon. Transactions of the American Fisheries Society, 134(4):817–828, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Begg:2000:SIH

[BB00] Gavin A. Begg and Russell W. Brown. Stock identification of haddock Melanogrammus aeglefinus on Georges Bank based on otolith shape analysis. Transactions of the American Fisheries Society, 129(4):935–945, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bernier:2008:DGE

[BBC^{+}08] Jeremiah C. Bernier, Shanda R. Birkeland, Michael J. Cipriano, Andrew G. McArthur, and Michael A. Banks. Differential gene expression between fall- and spring-run Chinook salmon assessed by long serial analysis of gene expression. Transactions of the American Fisheries Society, 137(5):1378–1388, September 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonar:2005:EIF

[BBDM05] Scott A. Bonar, Bruce D. Bolding, Marc Divens, and William Meyer. Effects of introduced fishes on wild juvenile Coho salmon in three shallow Pacific Northwest lakes. Transactions of the American Fisheries Society, 134(3):641–652, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bilby:2003:TNS

[BBF^{+}03] Robert E. Bilby, Eric W. Beach, Brian R. Fransen, Jason K. Walter, and Peter A. Bisson. Transfer of nutrients from spawning salmon to riparian vegetation in Western Washington. Transactions of the American Fisheries Society, 132(4):733–745, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:2002:SDD

[BBG02] Casey M. Baldwin, David A. Beauchamp, and Chad P. Gubala. Seasonal and diel distribution and movement of cutthroat trout from ultrasonic telemetry. Transactions of the American Fisheries Society, 131(1):143–158, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benkwitt:2009:DFC

[BBHD09] Cassandra E. Benkwitt, Richard D. Brodeur, Thomas P. Hurst, and Elizabeth A. Daly. Diel feeding chronology, gastric evacuation, and daily food consumption of juvenile Chinook salmon in Oregon coastal waters. Transactions of the American Fisheries Society, 138(1):111–120, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bochdansky:2006:MEL

[BBL06] Alexander B. Bochdansky, Naomi D. Brunemeyer, and William C. Leggett. Model evaluation of linear gut evacuation in the larval radiated shanny using a combination of laboratory and field data. Transactions of the American Fisheries Society, 135(2):390–398, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:2002:MMD

[BBM02] Jon Baker, Paul Bentzen, and Paul Moran. Molecular markers distinguish coastal cutthroat trout from coastal rainbow trout/ steelhead and their hybrids. Transactions of the American Fisheries Society, 131(4):404–417, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:2004:ENM

[BBP^{+}04] R. F. Baker, P. J. Blanchfield, M. J. Paterson, R. J. Flett, and L. Wesson. Evaluation of nonlethal methods for the analysis of mercury in fish tissue. Transactions of the American Fisheries Society, 133(3):568–576, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:2006:FAR

[BBRH06] K. R. Bestgen, D. W. Beyers, J. A. Rice, and G. B. Haines. Factors affecting recruitment of young Colorado pikeminnow: Synthesis of predation experiments, field studies, and individual-based modeling. Transactions of the American Fisheries Society, 135(6): 1722–1742, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2007:OTE

[BBS07] Randy J. Brown, Nate Bickford, and Ken Severin. Otolith trace element chemistry as an indicator of anadromy in Yukon River drainage coregonine fishes. Transactions of the American Fisheries Society, 136(3):678–690, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bellgraph:2008:HRS

[BBS^{+}08] Brian J. Bellgraph, Richard S. Brown, John R. Stephenson, Abigail E. Welch, Katherine A. Deters, and Thomas J. Carlson. Healing rate of swim bladders in rainbow trout. Transactions of the American Fisheries Society, 137(6):1791–1794, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:2000:BAT

[BBV00] Casey M. Baldwin, David A. Beauchamp, and Jason J. Van Tassell. Bioenergetic assessment of temporal food supply and consumption demand by salmonids in the Strawberry Reservoir food Web. Transactions of the American Fisheries Society, 129(2):

429–450, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowen:2003:EFR

[BBW03] Zachary H. Bowen, Ken D. Bovee, and Terry J. Waddle. Effects of flow regulation on shallow-water habitat dynamics and floodplain connectivity. Transactions of the American Fisheries Society, 132(4):809–823, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bramblett:2002:SUS

[BBWW02] Robert G. Bramblett, Mason D. Bryant, Brenda E. Wright, and Robert G. White. Seasonal use of small tributary and main-stem habitats by juvenile steelhead, Coho salmon, and Dolly Varden in a Southeastern Alaska drainage basin. Transactions of the American Fisheries Society, 131(4):498–506, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brick:2002:MRJ

[BC02] Marianne E. Brick and Joseph J. Cech, Jr. Metabolic responses of juvenile striped bass to exercise and handling stress with various recovery environments. Transactions of the American Fisheries Society, 131(5):855–864, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Billman:2007:PDJ

[BC07] Eric J. Billman and Todd A. Crowl. Population dynamics of a June sucker refuge population. Transactions of the American Fisheries Society, 136(4):959–965, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bruch:2009:LSA

[BCDF^{+}09] Ronald M. Bruch, Steven E. Campana, Shannon L. Davis-Foust, Michael J. Hansen, and John Janssen. Lake sturgeon age validation using bomb radiocarbon and known-age fish. Transactions of the American Fisheries Society, 138(2):361–372, March 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2006:ESC

[BCJ^{+}06] Terry D. Beacham, John R. Candy, Kimberly L. Jonsen, Janine Supernault, Michael Wetklo, Langtuo Deng, Kristina M. Miller, Ruth E. Withler, and Natalia Varnavskaya. Estimation of stock composition and individual identification of Chinook salmon across the Pacific rim by use of microsatellite variation. Transactions of the American Fisheries Society, 135(4):861–888, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2005:ESC

[BCM^{+}05] Terry D. Beacham, John R. Candy, Brenda McIntosh, Cathy MacConnachie, Amy Tabata, Karia Kaukinen, Langtuo Deng, Kristina M. Miller, Ruth E. Withler, and Natalia Varnavskaya. Estimation of stock composition and individual identification of sockeye salmon on a Pacific rim basis using microsatellite and major histocompatibility complex variation. Transactions of the American Fisheries Society, 134(5):1124–1146, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2001:EAM

[BCS^{+}01] Terry D. Beacham, John R. Candy, K. Janine Supernault, Tobi Ming, Bruce Deagle, Angela Schulze, Debra Tuck, Karia H. Kaukinen, James R. Irvine, Kristina M. Miller, and Ruth E. Withler. Evaluation and application of microsatellite and major histocompatibility complex variation for stock identification of Coho salmon in British Columbia. Transactions of the American Fisheries Society, 130(6):1116–1149, November 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brazner:2004:HFL

[BCT04] John C. Brazner, Steven E. Campana, and Danny K. Tanner. Habitat fingerprints for Lake Superior coastal wetlands derived from elemental analysis of yellow perch otoliths. Transactions of the American Fisheries Society, 133(3):692–704, May 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baltz:2005:TSE

[BCT^{+}05] D. M. Baltz, E. J. Chesney, M. A. Tarr, A. S. Kolok, and M. J. Bradley. Toxicity and sublethal effects of methanol on swimming performance of juvenile Florida pompano. Transactions of the American Fisheries Society, 134(3):730–740, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenkman:2007:UOC

[BCV07] Samuel J. Brenkman, Stephen C. Corbett, and Eric C. Volk. Use of otolith chemistry and radiotelemetry to determine age-specific migratory patterns of anadromous bull trout in the Hoh River, Washington. Transactions of the American Fisheries Society, 136(1):1–11, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2009:ABR

[BCW^{+}09] Richard S. Brown, Thomas J. Carlson, Abigail E. Welch, John R. Stephenson, C. Scott Abernethy, Blaine D. Ebberts, Mike J. Langeslay, Martin L. Ahmann, Dan H. Feil, John R. Skalski, and Richard L. Townsend. Assessment of barotrauma from rapid decompression of depth-acclimated juvenile Chinook salmon bearing radiotelemetry transmitters. Transactions of the American Fisheries Society, 138(6):1285–1301, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:2007:PUT

[BD07] Ethan Bell and Walter G. Duffy. Previously undocumented twoyear freshwater residency of juvenile Coho salmon in Prairie Creek, California. Transactions of the American Fisheries Society, 136(4):966–970, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benjamin:2007:INB

[BDD07] Joseph R. Benjamin, Jason B. Dunham, and Matthew R. Dare. Invasion by nonnative brook trout in Panther Creek, Idaho: Roles of local habitat quality, biotic resistance, and connectivity to source habitats. Transactions of the American Fisheries Society, 136(4):875–888, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bessey:2004:RPG

[BDLB04] Cindy Bessey, Robert H. Devlin, N. Robin Liley, and Carlo A. Biagi. Reproductive performance of growth-enhanced transgenic Coho salmon. Transactions of the American Fisheries Society, 133 (5):1205–1220, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:2001:FSJ

[BDR01] Ethan Bell, Walt G. Duffy, and Terry D. Roelofs. Fidelity and survival of juvenile Coho salmon in response to a flood. Transactions of the American Fisheries Society, 130(3):450–458, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beard:2000:EAL

[BE00] T. Douglas Beard, Jr. and Timothy E. Essington. Effects of angling and life history processes on bluegill size structure: Insights from an individual-based model. Transactions of the American Fisheries Society, 129(2):561–568, March 2000. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beaudreau:2007:STO

[BE07a] Anne H. Beaudreau and Timothy E. Essington. Spatial, temporal, and ontogenetic patterns of predation on rockfishes by lingcod. Transactions of the American Fisheries Society, 136(5):1438– 1452, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Broadhurst:2007:ITS

[BE07b] Ben Broadhurst and Brendan Ebner. An improved technique for small-scale radio-tracking of crayfish and benthic fishes in upland streams. Transactions of the American Fisheries Society, 136(2): 423–427, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Behnke:2002:CFD

[Beh02] R. J. Behnke. Comment: First documented case of anadromy in a population of introduced rainbow trout in Patagonia, Argentina. Transactions of the American Fisheries Society, 131(4):582–585, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [PKR02].

Bettoli:2005:FTN

[Bet05] Phillip W. Bettoli. The fundamental thermal niche of adult landlocked striped bass. Transactions of the American Fisheries Society, 134(2):305–314, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berejikian:2007:SFC

[BEV^{+}07] Barry A. Berejikian, Robert C. Endicott, Donald M. Van Doornik, Richard S. Brown, Christopher P. Tatara, and Jeffery Atkins. Spawning by female Chinook salmon can be detected by electromyogram telemetry. Transactions of the American Fisheries Society, 136(3):593–605, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:2004:EEI

[BFC^{+}04] Brian R. Beckman, William Fairgrieve, Kathleen A. Cooper, Conrad V. W. Mahnken, and Richard J. Beamish. Evaluation of endocrine indices of growth in individual postsmolt Coho salmon. Transactions of the American Fisheries Society, 133(5):1057– 1067, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blanchfield:2005:RLT

[BFHO05] Paul J. Blanchfield, Lori S. Flavelle, Tristan F. Hodge, and Diane M. Orihel. The response of lake trout to manual tracking. Transactions of the American Fisheries Society, 134(2):346–355, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baerwald:2008:DGD

[BFM08] Melinda R. Baerwald, Frederick Feyrer, and Bernie May. Distribution of genetically differentiated splittail populations during the nonspawning season. Transactions of the American Fisheries Society, 137(5):1335–1345, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Braaten:2002:LHA

[BG02] Patrick J. Braaten and Christopher S. Guy. Life history attributes of fishes along the latitudinal gradient of the Missouri River. Transactions of the American Fisheries Society, 131(5): 931–945, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Braaten:2004:FYG

[BG04] Patrick J. Braaten and Christopher S. Guy. First-year growth, condition, and size-selective winter mortality of freshwater drum in the Lower Missouri River. Transactions of the American Fisheries Society, 133(2):385–398, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bellgraph:2008:CPB

[BGGL08] Brian J. Bellgraph, Christopher S. Guy, William M. Gardner, and Steven A. Leathe. Competition potential between saugers and walleyes in nonnative sympatry. Transactions of the American Fisheries Society, 137(3):790–800, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowen:2006:DDC

[BGH06] S. H. Bowen, B. Gu, and Zhanghan Huang. Diet and digestion in Chinese mud carp Cirrhinus molitorella compared with other ilyophagous fishes. Transactions of the American Fisheries Society, 135(5):1383–1388, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2006:UET

[BGM06] Richard S. Brown, David R. Geist, and Matthew G. Mesa. Use of electromyogram telemetry to assess swimming activity of adult spring Chinook salmon migrating past a Columbia River dam. Transactions of the American Fisheries Society, 135(2):281–287, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:2008:EYR

[BGPL08] Brian R. Beckman, Brad Gadberry, Paul Parkins, and Donald A. Larsen. The effect of Yakima River spring Chinook salmon sire life history type on emergence timing and size of progeny. Transactions of the American Fisheries Society, 137(5):1285– 1291, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2004:VOA

[BGS^{+}04] Paul Brown, Corey Green, K. P. Sivakumaran, Daniel Stoessel, and A. Giles. Validating otolith annuli for annual age determination of common carp. Transactions of the American Fisheries Society, 133(1):190–196, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:2000:ESN

[BH00a] Peter B. Bayley and Robert A. Herendeen. The efficiency of a seine net. Transactions of the American Fisheries Society, 129(4):

901–923, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beasley:2000:ELH

[BH00b] Chris A. Beasley and Joseph E. Hightower. Effects of a lowhead dam on the distribution and characteristics of spawning habitat used by striped bass and American shad. Transactions of the American Fisheries Society, 129(6):1316–1330, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buhl:2000:ATF

[BH00c] Kevin J. Buhl and Steven J. Hamilton. Acute toxicity of firecontrol chemicals, nitrogenous chemicals, and surfactants to rainbow trout. Transactions of the American Fisheries Society, 129 (2):408–418, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boldt:2003:SGV

[BH03] Jennifer L. Boldt and Lewis J. Haldorson. Seasonal and geographic variation in juvenile pink salmon diets in the Northern Gulf of Alaska and Prince William Sound. Transactions of the American Fisheries Society, 132(6):1035–1052, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boldt:2004:SCW

[BH04] Jennifer L. Boldt and Lewis J. Haldorson. Size and condition of wild and hatchery pink salmon juveniles in Prince William Sound, Alaska. Transactions of the American Fisheries Society, 133(1):173–184, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bajer:2006:CMR

[BH06a] Przemyslaw G. Bajer and Robert S. Hayward. A combined multiple-regression and bioenergetics model for simulating fish growth in length and condition. Transactions of the American Fisheries Society, 135(3):695–710, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burdick:2006:DSA

[BH06b] Summer M. Burdick and Joseph E. Hightower. Distribution of spawning activity by anadromous fishes in an Atlantic slope drainage after removal of a low-head dam. Transactions of the American Fisheries Society, 135(5):1290–1300, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brakensiek:2007:EOS

[BH07] Kyle E. Brakensiek and David G. Hankin. Estimating overwinter survival of juvenile Coho salmon in a Northern California stream: Accounting for effects of passive integrated transponder tagging mortality and size-dependent survival. Transactions of the American Fisheries Society, 136(5):1423–1437, September 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bond:2007:NLC

[BHB^{+}07] Morgan H. Bond, Chad V. Hanson, Robert Baertsch, Sean A. Hayes, and R. Bruce MacFarlane. A new low-cost instream antenna system for tracking passive integrated transponder (PIT)- tagged fish in small streams. Transactions of the American Fisheries Society, 136(3):562–566, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beard:2003:DRS

[BHC03] T. Douglas Beard, Jr., Michael J. Hansen, and Stephen R. Carpenter. Development of a regional stock-recruitment model for understanding factors affecting walleye recruitment in Northern Wisconsin lakes. Transactions of the American Fisheries Society, 132(2):382–391, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barko:2004:RAF

[BHHS04] Valerie A. Barko, David P. Herzog, Robert A. Hrabik, and John S. Scheibe. Relationship among fish assemblages and main-channel-border physical habitats in the unimpounded Upper Mississippi River. Transactions of the American Fisheries Society, 133(2): 371–384, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baumsteiger:2005:GAJ

[BHL05] Jason Baumsteiger, David Hankin, and Eric J. Loudenslager. Genetic analyses of juvenile steelhead, coastal cutthroat trout, and their hybrids differ substantially from field identifications. Transactions of the American Fisheries Society, 134(4):829–840, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:2005:EWD

[BHMK05] Mandi M. Brandt, J. Preston Holloway, Christopher A. Myrick, and Matthew C. Kondratieff. Effects of waterfall dimensions and light intensity on age-0 brook trout jumping performance. Transactions of the American Fisheries Society, 134(2):496–502, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bacheler:2008:ADT

[BHP^{+}08] Nathan M. Bacheler, Joseph E. Hightower, Lee M. Paramore, Jeffrey A. Buckel, and Kenneth H. Pollock. An age-dependent tag return model for estimating mortality and selectivity of an estuarine-dependent fish with high rates of catch and release. Transactions of the American Fisheries Society, 137(5):1422– 1432, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budy:2005:ZSS

[BHS05] Phaedra Budy, Tyler Haddix, and Roger Schneidervin. Zooplankton size selection relative to gill raker spacing in rainbow trout. Transactions of the American Fisheries Society, 134(5): 1228–1235, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bohl:2009:EEC

[BHSR09] Russell J. Bohl, Theodore B. Henry, Richard J. Strange, and Patrick L. Rakes. Effects of electroshock on cyprinid embryos: Implications for threatened and endangered fishes. Transactions of the American Fisheries Society, 138(4):768–776, July 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burdick:2008:ALR

[BHV08] Summer M. Burdick, Heather A. Hendrixson, and Scott P. VanderKooi. Age-0 Lost River sucker and shortnose sucker nearshore habitat use in Upper Klamath Lake, Oregon: a patch occupancy approach. Transactions of the American Fisheries Society, 137 (2):417–430, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baxter:2003:MGS

[BHW03] Colden Baxter, F. Richard Hauer, and William W. Woessner. Measuring groundwater–stream water exchange: New techniques for installing minipiezometers and estimating hydraulic conductivity. Transactions of the American Fisheries Society, 132(3): 493–502, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:2007:PSC

[BHW^{+}07] Kevin R. Bestgen, John A. Hawkins, Gary C. White, Kevin D. Christopherson, J. Michael Hudson, Mark H. Fuller, D. Chris Kitcheyan, Ronald Brunson, Paul Badame, G. Bruce Haines, Julie A. Jackson, Cameron D. Walford, and Tasha A. Sorensen. Population status of Colorado pikeminnow in the Green River Basin, Utah and Colorado. Transactions of the American Fisheries Society, 136(5):1356–1380, September 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burrell:2000:SMB

[BIB^{+}00] Kyle H. Burrell, J. Jeffery Isely, David B. Bunnell, Jr., David H. Van Lear, and C. Andrew Dolloff. Seasonal movement of brown trout in a Southern Appalachian River. Transactions of the American Fisheries Society, 129(6):1373–1379, November 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:2004:MPS

[BIB04] Michael M. Bailey, J. Jeffery Isely, and William C. Bridges, Jr. Movement and population size of American shad near a low-head lock and dam. Transactions of the American Fisheries Society, 133(2):300–308, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bisson:2004:RPS

[Bis04] Peter A. Bisson. Restoration of Puget Sound rivers. Transactions of the American Fisheries Society, 133(5):1290–1292, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjorgo:2000:SMH

[BIT00] Kimberly A. Bjorgo, J. Jeffery Isely, and Christopher S. Thomason. Seasonal movement and habitat use by striped bass in the Combahee River, South Carolina. Transactions of the American Fisheries Society, 129(6):1281–1287, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burkhead:2001:ESS

[BJ01] Noel M. Burkhead and Howard L. Jelks. Effects of suspended sediment on the reproductive success of the tricolor shiner, a crevicespawning minnow. Transactions of the American Fisheries Society, 130(5):959–968, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baltz:2003:TSP

[BJ03] Donald M. Baltz and Robert F. Jones. Temporal and spatial patterns of microhabitat use by fishes and decapod crustaceans in a Louisiana estuary. Transactions of the American Fisheries Society, 132(4):662–678, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2006:PRPb

[BJS^{+}06] Terry D. Beacham, Kimberly L. Jonsen, Janine Supernault, Michael Wetklo, Langtuo Deng, and Natalia Varnavskaya. Pacific rim population structure of Chinook salmon as determined from microsatellite analysis. Transactions of the American Fisheries Society, 135(6):1604–1621, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bramblett:2005:DEF

[BJZH05] Robert G. Bramblett, Thomas R. Johnson, Alexander V. Zale, and Daniel G. Heggem. Development and evaluation of a fish assemblage index of biotic integrity for Northwestern Great Plains streams. Transactions of the American Fisheries Society, 134(3): 624–640, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baird:2003:BTB

[BK03] Owen E. Baird and Charles C. Krueger. Behavioral thermoregulation of brook and rainbow trout: Comparison of summer habitat use in an Adirondack river, New York. Transactions of the American Fisheries Society, 132(6):1194–1206, November 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bringolf:2005:STF

[BKCL05] Robert B. Bringolf, Thomas J. Kwak, W. Gregory Cope, and Michael S. Larimore. Salinity tolerance of flathead catfish: Implications for dispersal of introduced populations. Transactions of the American Fisheries Society, 134(4):927–936, July 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Busack:2007:MDB

[BKHH07] Craig Busack, Curtis M. Knudsen, Germaine Hart, and Paul Huffman. Morphological differences between adult wild and firstgeneration hatchery Upper Yakima River Spring Chinook salmon. Transactions of the American Fisheries Society, 136(4):1076– 1087, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boggs:2004:FRA

[BKP^{+}04] C. T. Boggs, M. L. Keefer, C. A. Peery, T. C. Bjornn, and L. C. Stuehrenberg. Fallback, reascension, and adjusted fishway escapement estimates for adult Chinook salmon and steelhead at Columbia and Snake River dams. Transactions of the American Fisheries Society, 133(4):932–949, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldigo:2000:CFC

[BL00] B. P. Baldigo and G. B. Lawrence. Composition of fish communities in relation to stream acidification and habitat in the Neversink River, New York. Transactions of the American Fisheries Society, 129(1):60–76, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bateman:2001:NSS

[BL01] Douglas S. Bateman and Hiram W. Li. Nest site selection by reticulate sculpin in two streams of different geologies in the Central Coast Range of Oregon. Transactions of the American Fisheries Society, 130(5):823–832, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brosse:2002:RBE

[BL02] Sebastien Brosse and Sovan Lek. Relationships between environmental characteristics and the density of age-0 Eurasian perch Perca fluviatilis in the littoral zone of a lake: a nonlinear approach. Transactions of the American Fisheries Society, 131(6): 1033–1043, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:2005:UMM

[BL05] Brian R. Beckman and Donald A. Larsen. Upstream migration of minijack (age-2) Chinook salmon in the Columbia River: Behavior, abundance, distribution, and origin. Transactions of the American Fisheries Society, 134(6):1520–1541, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beechie:2005:CHT

[BLBH05] T. J. Beechie, M. Liermann, E. M. Beamer, and R. Henderson. A classification of habitat types in a large river and their use by juvenile salmonids. Transactions of the American Fisheries Society, 134(3):717–729, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2004:SIF

[BLC^{+}04] Terry D. Beacham, Michael Lapointe, John R. Candy, Brenda McIntosh, Cathy MacConnachie, Amy Tabata, Karia Kaukinen, Langtuo Deng, Kristina M. Miller, and Ruth E. Withler. Stock identification of Fraser River sockeye salmon using microsatellites and major histocompatibility complex variation. Transactions of the American Fisheries Society, 133(5):1117–1137, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenkman:2001:SML

[BLG01] Samuel J. Brenkman, Gary L. Larson, and Robert E. Gresswell. Spawning migration of lacustrine-adfluvial bull trout in a natural area. Transactions of the American Fisheries Society, 130(5):981– 987, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brasher:2006:IPA

[BLGW06] Anne M. D. Brasher, Corene D. Luton, Steven L. Goodbred, and Reuben H. Wolff. Invasion patterns along elevation and urbanization gradients in Hawaiian streams. Transactions of the American Fisheries Society, 135(4):1109–1129, July 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:2000:PSN

[BLS^{+}00] Brian R. Beckman, Donald A. Larsen, Cameron Sharpe, Beeda Lee-Pawlak, Carl B. Schreck, and Walton W. Dickhoff. Physiological status of naturally reared juvenile spring Chinook salmon in the Yakima River: Seasonal dynamics and changes associated with smolting. Transactions of the American Fisheries Society, 129(3):727–753, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldigo:2007:PMB

[BLS07] Barry P. Baldigo, Gregory Lawrence, and Howard Simonin. Persistent mortality of brook trout in episodically acidified streams of the Southwestern Adirondack Mountains, New York. Transactions of the American Fisheries Society, 136(1):121–134, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beeman:2006:MDJ

[BM06] John W. Beeman and Alec G. Maule. Migration depths of juvenile Chinook salmon and steelhead relative to total dissolved gas supersaturation in a Columbia River reservoir. Transactions of the American Fisheries Society, 135(3):584–594, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bronte:2007:MVS

[BM07] Charles R. Bronte and Seth A. Moore. Morphological variation of Siscowet lake trout in Lake Superior. Transactions of the American Fisheries Society, 136(2):509–517, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bergstrom:2009:IRC

[BM09] Margot A. Bergstrom and Allen F. Mensinger. Interspecific resource competition between the invasive round goby and three native species: Logperch, slimy sculpin, and spoonhead sculpin. Transactions of the American Fisheries Society, 138(5):1009– 1017, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenden:2006:EDD

[BMH06] Travis O. Brenden, Brian R. Murphy, and Eric M. Hallerman. Effect of discharge on daytime habitat use and selection by muskellunge in the New River, Virginia. Transactions of the American Fisheries Society, 135(6):1546–1558, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bajer:2007:ADC

[BMH07] Przemyslaw G. Bajer, Joshua J. Millspaugh, and Robert S. Hayward. Application of discrete choice models to predict white crappie temperature selection in two Missouri impoundments. Transactions of the American Fisheries Society, 136(4):889–901, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2006:PRPa

[BMM^{+}06] Terry D. Beacham, Brenda McIntosh, Cathy MacConnachie, Kristina M. Miller, Ruth E. Withler, and Natalia Varnavskaya. Pacific rim population structure of sockeye salmon as determined from microsatellite analysis. Transactions of the American Fisheries Society, 135(1):174–187, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:2004:ERE

[BMN04] R. J. Beamish, C. Mahnken, and C. M. Neville. Evidence that reduced early marine growth is associated with lower marine survival of Coho salmon. Transactions of the American Fisheries Society, 133(1):26–33, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bear:2007:CTR

[BMZ07] Elizabeth A. Bear, Thomas E. McMahon, and Alexander V. Zale. Comparative thermal requirements of westslope cutthroat trout and rainbow trout: Implications for species interactions and development of thermal protection standards. Transactions of the American Fisheries Society, 136(4):1113–1121, July 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brockmark:2007:ERD

[BNB^{+}07] Sofia Brockmark, Lena Neregård, Torgny Bohlin, Björn Thrandur Björnsson, and Jörgen I. Johnsson. Effects of rearing density and structural complexity on the pre- and postrelease performance of Atlantic salmon. Transactions of the American Fisheries Society, 136(5):1453–1462, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bearlin:2008:BRM

[BNG08] Andrew R. Bearlin, Simon J. Nicol, and Terry Glenane. Behavioral responses of Murray cod Maccullochella peelii peelii to pulse frequency and pulse width from electric fishing machines. Transactions of the American Fisheries Society, 137(1):107–113, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boehlert:2001:BR

[Boe01] George W. Boehlert. Book review. Transactions of the American Fisheries Society, 130(2):349, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baigun:2007:AFY

[BOMC07] Claudio Baigún, Norberto Oldani, Adrián Madirolas, and Gustavo Alvarez Colombo. Assessment of fish yield in Patagonian lakes (Argentina): Development and application of empirical models. Transactions of the American Fisheries Society, 136(3): 846–857, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bozek:2001:BR

[Boz01] Michael A. Bozek. Book review. Transactions of the American Fisheries Society, 130(1):175–176, January 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:2001:AEP

[BP01] Peter B. Bayley and James T. Peterson. An approach to estimate probability of presence and richness of fish species. Transactions of the American Fisheries Society, 130(4):620–633, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2004:GAI

[BPM^{+}04] Kim H. Brown, Steve J. Patton, Kyle E. Martin, Krista M. Nichols, Robyn Armstrong, and Gary H. Thorgaard. Genetic analysis of interior Pacific Northwest Oncorhynchus mykiss reveals apparent ancient hybridization with westslope cutthroat trout. Transactions of the American Fisheries Society, 133(5): 1078–1088, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brewer:2006:SHA

[BPR06] Shannon K. Brewer, Diana M. Papoulias, and Charles F. Rabeni. Spawning habitat associations and selection by fishes in a flowregulated prairie river. Transactions of the American Fisheries Society, 135(3):763–778, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boatright:2004:TAM

[BQH04] Chris Boatright, Thomas Quinn, and Ray Hilborn. Timing of adult migration and stock structure for sockeye salmon in Bear Lake, Alaska. Transactions of the American Fisheries Society, 133(4):911–921, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burke:2002:LFB

[BR02] Brian J. Burke and James A. Rice. A linked foraging and bioenergetics model for southern flounder. Transactions of the American Fisheries Society, 131(1):120–131, January 2002. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brooking:2009:HTS

[BR09a] Thomas E. Brooking and Lars G. Rudstam. Hydroacoustic target strength distributions of alewives in a net-cage compared with field surveys: Deciphering target strength distributions and effect on density estimates. Transactions of the American Fisheries Society, 138(3):471–486, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Browne:2009:STE

[BR09b] David R. Browne and Joseph B. Rasmussen. Shifts in the trophic ecology of brook trout resulting from interactions with yellow perch: an intraguild predator-prey interaction. Transactions of the American Fisheries Society, 138(5):1109–1122, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breck:2008:EBM

[Bre08] James E. Breck. Enhancing bioenergetics models to account for dynamic changes in fish body composition and energy density. Transactions of the American Fisheries Society, 137(1):340–356, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bernatchez:2002:MGD

[BRK02] Louis Bernatchez, James G. Rhydderch, and Frederick W. Kircheis. Microsatellite gene diversity analysis in landlocked Arctic char from Maine. Transactions of the American Fisheries Society, 131(6):1106–1118, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2004:RFE

[Bro04] Tommy L. Brown. Recreational fisheries: Ecological, economic, and social evaluation. Transactions of the American Fisheries Society, 133(1):233–234, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brouder:2005:AGR

[Bro05] Mark J. Brouder. Age and growth of roundtail chub in the Upper Verde River, Arizona. Transactions of the American Fisheries Society, 134(4):866–871, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barndt:2003:ISC

[BS03] Scott Barndt and Jen Stone. Infestation of Salmincola californiensis (Copepoda: Lernaeopodidae) in wild Coho salmon, steelhead, and coastal cutthroat trout juveniles in a small Columbia River tributary. Transactions of the American Fisheries Society, 132(5): 1027–1032, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Basler:2006:EEF

[BS06] Matthew C. Basler and Harold L. Schramm, Jr. Evaluation of electrofishing and fyke netting for collecting black carp in small ponds. Transactions of the American Fisheries Society, 135(2): 277–280, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burfeind:2007:EBP

[BS07] Dana D. Burfeind and Gregory W. Stunz. The effects of boat propeller scarring on nekton growth in subtropical seagrass meadows. Transactions of the American Fisheries Society, 136(6): 1546–1551, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bronte:2008:HRA

[BS08] Charles R. Bronte and Shawn P. Sitar. Harvest and relative abundance of Siscowet lake trout in Michigan waters of Lake Superior, 1929–1961. Transactions of the American Fisheries Society, 137(3):916–926, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blankenship:2009:TSG

[BSB09] Scott M. Blankenship, Maureen P. Small, and Joseph D. Bumgarner. Temporal stability of genetic variation within natural populations of summer steelhead receiving mitigation hatchery fish. Transactions of the American Fisheries Society, 138(5):1052– 1064, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:2004:ICS

[BSC^{+}04] R. J. Beamish, J. T. Schnute, A. J. Cass, C. M. Neville, and R. M. Sweeting. The influence of climate on the stock and recruitment of pink and sockeye salmon from the Fraser River, British Columbia, Canada. Transactions of the American Fisheries Society, 133(6): 1396–1412, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benson:2005:SMP

[BSEM05] Angela C. Benson, Trent M. Sutton, Robert F. Elliott, and Thomas G. Meronek. Seasonal movement patterns and habitat preferences of age-0 lake sturgeon in the Lower Peshtigo River, Wisconsin. Transactions of the American Fisheries Society, 134 (5):1400–1409, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:2008:CPE

[BSLN08] R. J. Beamish, R. M. Sweeting, K. L. Lange, and C. M. Neville. Changes in the population ecology of hatchery and wild Coho salmon in the Strait of Georgia. Transactions of the American Fisheries Society, 137(2):503–520, February 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beauchamp:2004:STD

[BSM^{+}04] David A. Beauchamp, Christopher J. Sergeant, Michael M. Mazur, Jennifer M. Scheuerell, Daniel E. Schindler, Mark D. Scheuerell, Kurt L. Fresh, David E. Seiler, and Thomas P. Quinn. Spatial–temporal dynamics of early feeding demand and food supply for sockeye salmon fry in Lake Washington. Transactions of the American Fisheries Society, 133(4):1014–1032, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2008:UMD

[BSM^{+}08] T. D. Beacham, J. F. Schweigert, C. MacConnachie, K. D. Le, and L. Flostrand. Use of microsatellites to determine population structure and migration of Pacific herring in British Columbia and adjacent regions. Transactions of the American Fisheries Society, 137(6):1795–1811, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:2004:IJP

[BSN04] R. J. Beamish, R. M. Sweeting, and C. M. Neville. Improvement of juvenile Pacific salmon production in a regional ecosystem after the 1998 climatic regime shift. Transactions of the American Fisheries Society, 133(5):1163–1175, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bergstedt:2001:LSL

[BSO01] Roger A. Bergstedt, Clifford P. Schneider, and Robert O’Gorman. Lethality of sea lamprey attacks on lake trout in relation to location on the body surface. Transactions of the American Fisheries Society, 130(2):336–340, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bunnell:2005:TEM

[BSS05] David B. Bunnell, Matthew A. Scantland, and Roy A. Stein. Testing for evidence of maternal effects among individuals and populations of white crappie. Transactions of the American Fisheries Society, 134(3):607–619, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:2000:CFM

[BSW00] Bonnie L. Brown, Summer L. W. Schultz, and Frances K. H. White. A convenient field method of tissue preservation for flow cytometric ploidy assessment of grass carp. Transactions of the American Fisheries Society, 129(6):1354–1359, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brothers:2004:SCA

[BT04a] Edward Brothers and Ronald Thresher. Statolith chemical analysis as a means of identifying stream origins of lampreys in Lake Huron. Transactions of the American Fisheries Society, 133(5):

1107–1116, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brunelli:2004:NCS

[BT04b] Joseph P. Brunelli and Gary H. Thorgaard. A new Ychromosome-specific marker for Pacific salmon. Transactions of the American Fisheries Society, 133(5):1247–1253, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brunelli:2008:SNP

[BTLD08] Joseph P. Brunelli, Gary H. Thorgaard, Robb F. Leary, and James L. Dunnigan. Single-nucleotide polymorphisms associated with allozyme differences between inland and coastal rainbow trout. Transactions of the American Fisheries Society, 137(5): 1292–1298, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burkhead:2009:BRF

[Bur09] Noel M. Burkhead. Book review: Fish Conservation: a Guide to Understanding and Restoring Global Aquatic Biodiversity and Fishery Resources (N. M. Burkhead); Fish Biomechanics (Fish Physiology, Volume 23, series edited by A. Farrell and C. Brauner) (J. J. Cech, Jr.); Cephalopods: Ecology and Fisheries (M Staudinger). Transactions of the American Fisheries Society, 138(5):1198–1205, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beauchamp:2001:MST

[BV01] David A. Beauchamp and Jason J. Van Tassell. Modeling seasonal trophic interactions of adfluvial bull trout in Lake Billy Chinook, Oregon. Transactions of the American Fisheries Society, 130(2): 204–216, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berejikian:2005:EEB

[BVL^{+}05] Barry Berejikian, Donald Van Doornik, Anita LaRae, Skip Tezak, and Joy Lee. The effects of exercise on behavior and reproductive success of captively reared steelhead. Transactions of the American Fisheries Society, 134(5):1236–1252, September 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:2006:PSS

[BVMM06] T. D. Beacham, N. V. Varnavskaya, B. McIntosh, and C. Mac-Connachie. Population structure of sockeye salmon from Russia determined with microsatellite DNA variation. Transactions of the American Fisheries Society, 135(1):97–109, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bramblett:2001:HUM

[BW01] Robert G. Bramblett and Robert G. White. Habitat use and movements of pallid and shovelnose sturgeon in the Yellowstone and Missouri Rivers in Montana and North Dakota. Transactions of the American Fisheries Society, 130(6):1006–1025, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonner:2002:ETP

[BW02] Timothy H. Bonner and Gene R. Wilde. Effects of turbidity on prey consumption by prairie stream fishes. Transactions of the American Fisheries Society, 131(6):1203–1208, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenden:2007:CBM

[BWC^{+}07] Travis O. Brenden, Lizhu Wang, Richard D. Clark, Jr., Paul W. Seelbach, and John Lyons. Comparison between model-predicted and field-measured stream habitat features for evaluating fish assemblage-habitat relationships. Transactions of the American Fisheries Society, 136(3):580–592, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buchheister:2006:SGV

[BWFB06] Andre Buchheister, Matthew T. Wilson, Robert J. Foy, and David A. Beauchamp. Seasonal and geographic variation in condition of juvenile walleye pollock in the Western Gulf of Alaska. Transactions of the American Fisheries Society, 135(4):897–907, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:2005:MAS

[BWS^{+}05] Edward A. Baker, Kevin E. Wehrly, Paul W. Seelbach, Li Wang, Michael J. Wiley, and Thomas Simon. A multimetric assessment of stream condition in the northern lakes and forests ecoregion using spatially explicit statistical modeling and regional normalization. Transactions of the American Fisheries Society, 134(3): 697–710, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenden:2008:RVS

[BWS08] Travis O. Brenden, Lizhu Wang, and Paul W. Seelbach. A river valley segment classification of Michigan streams based on fish and physical attributes. Transactions of the American Fisheries Society, 137(6):1621–1636, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brinkman:2009:CTA

[BWVN09] Stephen F. Brinkman, John D. Woodling, Alan M. Vajda, and David O. Norris. Chronic toxicity of ammonia to early life stage rainbow trout. Transactions of the American Fisheries Society, 138(2):433–440, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bunkley-Williams:2001:BR

[BWW01] Lucy Bunkley-Williams and Ernest H. Williams, Jr. Book review. Transactions of the American Fisheries Society, 130(1):179–180, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryant:2004:SFA

[BZW04] M. D. Bryant, N. D. Zymonas, and B. E. Wright. Salmonids on the fringe: Abundance, species composition, and habitat use of salmonids in high-gradient headwater streams, Southeast Alaska. Transactions of the American Fisheries Society, 133(6):1529– 1538, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campton:2004:SCS

[Cam04] Donald E. Campton. Sperm competition in salmon hatcheries: The need to institutionalize genetically benign spawning protocols. Transactions of the American Fisheries Society, 133(5): 1277–1289, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campton:2005:SCS

[Cam05] Donald E. Campton. Sperm competition in salmon hatcheries — the need to institutionalize genetically benign spawning protocols: Response to comment. Transactions of the American Fisheries Society, 134(6):1495–1498, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Qui05].

Carline:2004:EHE

[Car04] Robert F. Carline. Electrofishing and its harmful effects on fish. Transactions of the American Fisheries Society, 133(6):1539, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carline:2006:RUB

[Car06] Robert F. Carline. Regulation of an unexploited brown trout population in Spruce Creek, Pennsylvania. Transactions of the American Fisheries Society, 135(4):943–954, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Castro:2000:BR

[Cas00] JoséI. Castro. Book review. Transactions of the American Fisheries Society, 129(3):888, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chipps:2000:ZNR

[CB00] Steven R. Chipps and David H. Bennett. Zooplanktivory and nutrient regeneration by invertebrate (Mysis relicta) and vertebrate (Oncorhynchus nerka) planktivores: Implications for trophic interactions in oligotrophic lakes. Transactions of the American Fisheries Society, 129(2):569–583, March 2000. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clarke:2002:NPE

[CB02a] Lance R. Clarke and David H. Bennett. A net-pen experiment to evaluate Kokanee growth rates in autumn in an oligotrophic lake with Mysis relicta. Transactions of the American Fisheries Society, 131(6):1061–1069, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clarke:2002:NEK

[CB02b] Lance R. Clarke and David H. Bennett. Newly emerged Kokanee growth and survival in an oligotrophic lake with Mysis relicta.

Transactions of the American Fisheries Society, 131(1):176–185, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clarke:2004:ZCC

[CB04] Lance R. Clarke and David H. Bennett. Zooplankton community changes at Pend Oreille Lake, Idaho: Testing implications for age-0 Kokanee prey selection, digestion, and growth. Transactions of the American Fisheries Society, 133(5):1221–1234, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cortemeglia:2005:TTW

[CB05] Cheryl Cortemeglia and Thomas L. Beitinger. Temperature tolerances of wild-type and red transgenic zebra danios. Transactions of the American Fisheries Society, 134(6):1431–1437, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cyterski:2006:IPF

[CB06] Michael Cyterski and Craig Barber. Identification and prediction of fish assemblages in streams of the Mid-Atlantic Highlands, USA. Transactions of the American Fisheries Society, 135(1):40– 48, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2006:RIG

[CBF^{+}06] B. Campbell, B. R. Beckman, W. T. Fairgrieve, J. T. Dickey, and P. Swanson. Reproductive investment and growth history in female Coho salmon. Transactions of the American Fisheries Society, 135(1):164–173, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:2000:HSP

[CBM^{+}00] Keya M. Collins, Ernest L. Brannon, Lawrence L. Moulton, Matthew A. Cronin, and Keith R. Parker. Hydraulic sampling protocol to estimate natural embryo mortality of pink salmon. Transactions of the American Fisheries Society, 129(3):827–834, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cross:2008:EMG

[CBMM08] Alison D. Cross, David A. Beauchamp, Katherine W. Myers, and Jamal H. Moss. Early marine growth of pink salmon in Prince William Sound and the coastal Gulf of Alaska during years of low and high survival. Transactions of the American Fisheries Society, 137(3):927–939, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chittenden:2009:UAT

[CBN^{+}09] C. M. Chittenden, R. J. Beamish, C. M. Neville, R. M. Sweeting, and R. S. McKinley. The use of acoustic tags to determine the timing and location of the juvenile Coho salmon migration out of the Strait of Georgia, Canada. Transactions of the American Fisheries Society, 138(6):1220–1225, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crandall:2008:RLL

[CBR^{+}08] John D. Crandall, Leslie B. Bach, Nathan Rudd, Mark Stern, and Matt Barry. Response of larval Lost River and shortnose suckers to wetland restoration at the Williamson River Delta, Oregon. Transactions of the American Fisheries Society, 137(2):402–416, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Comeau:2003:MTS

[CC03] L. A. Comeau and S. E. Campana. Modifying thyroidal status in Atlantic cod by osmotic pump delivery of thyroid and antithyroid agents. Transactions of the American Fisheries Society, 132(5): 1021–1026, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:2009:ETO

[CCK^{+}09] Benjamin J. Clemens, Shaun P. Clements, Mark D. Karnowski, David B. Jepsen, Alix I. Gitelman, and Carl B. Schreck. Effects of transportation and other factors on survival estimates of juvenile salmonids in the unimpounded Lower Columbia River. Transactions of the American Fisheries Society, 138(1):169–188, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cegelski:2006:MGS

[CCMP06] Christine C. Cegelski, Matthew R. Campbell, Kevin A. Meyer, and Madison S. Powell. Multiscale genetic structure of Yellowstone cutthroat trout in the Upper Snake River Basin. Transactions of the American Fisheries Society, 135(3):711–726, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:2003:SSS

[CCP^{+}03] Mark R. Collins, Doug Cooke, Bill Post, John Crane, James Bulak, Theodore J. I. Smith, Thomas W. Greig, and Joseph M. Quattro. Shortnose sturgeon in the Santee–Cooper Reservoir System, South Carolina. Transactions of the American Fisheries Society, 132(6):1244–1250, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coghlan:2007:SGR

[CCR^{+}07] Stephen M. Coghlan, Jr., Michael J. Connerton, Neil H. Ringler, Donald J. Stewart, and Jerry V. Mead. Survival and growth responses of juvenile salmonines stocked in Eastern Lake Ontario tributaries. Transactions of the American Fisheries Society, 136 (1):56–71, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:2009:EWH

[CD09] Duane C. Chapman and Joseph E. Deters. Effect of water hardness and dissolved-solid concentration on hatching success and egg size in bighead carp. Transactions of the American Fisheries Society, 138(6):1226–1231, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chumchal:2008:HSD

[CDF^{+}08] Matthew M. Chumchal, Ray W. Drenner, Brian Fry, K. David Hambright, and Leo W. Newland. Habitat-specific differences in mercury concentration in a top predator from a shallow lake. Transactions of the American Fisheries Society, 137(1):195–208, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ciereszko:2006:CSL

[CDFW06] Andrzej Ciereszko, Konrad Dabrowski, Julia Froschauer, and Tobie D. Wolfe. Cryopreservation of semen from lake sturgeon.

Transactions of the American Fisheries Society, 135(1):232–240, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2002:HIM

[CDP02] Matthew R. Campbell, Jeff Dillon, and Madison S. Powell. Hybridization and introgression in a managed, native population of Yellowstone cutthroat trout: Genetic detection and management implications. Transactions of the American Fisheries Society, 131(4):364–375, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ciereszko:2002:FAM

[CDT^{+}02] Andrzej Ciereszko, Konrad Dabrowski, Gregory P. Toth, Susan A. Christ, and Jan Glogowski. Factors affecting motility characteristics and fertilizing ability of sea lamprey spermatozoa. Transactions of the American Fisheries Society, 131(2):193–202, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cairns:2005:ISS

[CEB^{+}05] Michael A. Cairns, Joseph L. Ebersole, Joan P. Baker, Parker J. Wigington, Jr., Henry R. Lavigne, and Sheila M. Davis. Influence of summer stream temperatures on black spot infestation of juvenile Coho salmon in the Oregon Coast Range. Transactions of the American Fisheries Society, 134(6):1471–1479, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chipps:2000:GFC

[CEW00] Steven R. Chipps, Lisa M. Einfalt, and David H. Wahl. Growth and food consumption by tiger muskellunge: Effects of temperature and ration level on bioenergetic model predictions. Transactions of the American Fisheries Society, 129(1):186–193, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coleman:2007:CSTb

[CF07a] Mark A. Coleman and Kurt D. Fausch. Cold summer temperature limits recruitment of age-0 cutthroat trout in high-elevation Colorado streams. Transactions of the American Fisheries Society, 136(5):1231–1244, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coleman:2007:CSTa

[CF07b] Mark A. Coleman and Kurt D. Fausch. Cold summer temperature regimes cause a recruitment bottleneck in age-0 Colorado River cutthroat trout reared in laboratory streams. Transactions of the American Fisheries Society, 136(3):639–654, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2004:ASJ

[CG04] Jed G. Campbell and Larry R. Goodman. Acute sensitivity of juvenile shortnose sturgeon to low dissolved oxygen concentrations. Transactions of the American Fisheries Society, 133(3):772–776, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connor:2006:PMW

[CG06] William P. Connor and Aaron P. Garcia. Prespawning movement of wild and hatchery fall Chinook salmon adults in the Snake River. Transactions of the American Fisheries Society, 135(2): 297–305, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Csoboth:2008:LEL

[CG08] Laura A. Csoboth and James E. Garvey. Lateral exchange of larval fish between a restored backwater and a large river in the East-Central USA. Transactions of the American Fisheries Society, 137(1):33–44, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Czesny:2002:CFP

[CGAD^{+}02] Sergiusz Czesny, Mary Ann Garcia-Abiado, Konrad Dabrowski, Przemyslaw Bajer, and Maciej Zalewski. Comparison of foraging performance of diploid and triploid saugeyes (Sauger × walleye). Transactions of the American Fisheries Society, 131(5):980–985, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Czesny:2005:ECS

[CGD05] Sergiusz J. Czesny, Brian D. S. Graeb, and John M. Dettmers. Ecological consequences of swim bladder noninflation for larval yellow perch. Transactions of the American Fisheries Society, 134(4):1011–1020, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colvin:2009:FUI

[CGL^{+}09] Randall Colvin, Guillermo R. Giannico, Judith Li, Kathryn L. Boyer, and William J. Gerth. Fish use of intermittent watercourses draining agricultural lands in the Upper Willamette River Valley, Oregon. Transactions of the American Fisheries Society, 138(6):1302–1313, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conover:2003:ERC

[CGM03] David O. Conover, Todd Gilmore, and Stephan B. Munch. Estimating the relative contribution of spring- and summer-spawned cohorts to the Atlantic Coast bluefish stock. Transactions of the American Fisheries Society, 132(6):1117–1124, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Castro:2000:ELS

[CH00] Leonardo R. Castro and Eduardo H. Hernandez. Early life survival of the anchoveta Engraulis ringens off Central Chile during the 1995 and 1996 winter spawning seasons. Transactions of the American Fisheries Society, 129(5):1107–1117, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crossin:2005:NRM

[CH05] Glenn T. Crossin and Scott G. Hinch. A nonlethal, rapid method for assessing the somatic energy content of migrating adult Pacific salmon. Transactions of the American Fisheries Society, 134 (1):184–191, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cho:2002:EIC

[CHH02] Grace K. Cho, John W. Heath, and Daniel D. Heath. Electroshocking influences Chinook salmon egg survival and juvenile physiology and immunology. Transactions of the American Fisheries Society, 131(2):224–233, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Case:2006:ABG

[CHH^{+}06] Richard A. J. Case, William F. Hutchinson, Lorenz Hauser, Vivian Buehler, Catriona Clemmesen, Geir Dahle, Olav S. Kjesbu, Erlend Moksness, Håkon Otterå, Helge Paulsen, Terje Svåsand, Anders Thorsen, and Gary R. Carvalho. Association between growth and Pan I* genotype within Atlantic cod full-sibling families. Transactions of the American Fisheries Society, 135(1):241– 250, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chih:2009:ESE

[Chi09] Ching-Ping Chih. Evaluation of the sampling efficiency of three otolith sampling methods for commercial king mackerel fisheries. Transactions of the American Fisheries Society, 138(5):990–999, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chizinski:2008:DEB

[CHP08] Christopher J. Chizinski, Caleb G. Huber, and Kevin L. Pope. Development and evaluation of a bioenergetics model for the inland silverside in freshwater systems. Transactions of the American Fisheries Society, 137(1):254–261, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Claramunt:2005:ISH

[CJFM05] Randall M. Claramunt, Jory L. Jonas, John D. Fitzsimons, and J. Ellen Marsden. Influences of spawning habitat characteristics and interstitial predators on lake trout egg deposition and mortality. Transactions of the American Fisheries Society, 134(4): 1048–1057, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crockett:2006:MTS

[CJMB06] Harry J. Crockett, Brett M. Johnson, Patrick J. Martinez, and Dan Brauch. Modeling target strength distributions to improve hydroacoustic estimation of lake trout population size. Transactions of the American Fisheries Society, 135(4):1095–1108, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connelly:2001:ATB

[CKB01] Nancy A. Connelly, Barbara A. Knuth, and Tommy L. Brown. An angler typology based on angler fishing preferences. Transactions of the American Fisheries Society, 130(1):130–137, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Congleton:2000:SIM

[CLSD00] James L. Congleton, William J. LaVoie, Carl B. Schreck, and Lawrence E. Davis. Stress indices in migrating juvenile Chinook salmon and steelhead of wild and hatchery origin before and after barge transportation. Transactions of the American Fisheries Society, 129(4):946–961, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carline:2001:CTM

[CM01] Robert F. Carline and James F. Machung. Critical thermal maxima of wild and domestic strains of trout. Transactions of the American Fisheries Society, 130(6):1211–1216, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carline:2003:EFB

[CM03a] Robert F. Carline and Brian J. McCullough. Effects of floods on brook trout populations in the Monongahela National Forest, West Virginia. Transactions of the American Fisheries Society, 132(5):1014–1020, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooperman:2003:RML

[CM03b] Michael Cooperman and Douglas F. Markle. Rapid out-migration of Lost River and shortnose sucker larvae from in-river spawning beds to in-lake rearing grounds. Transactions of the American Fisheries Society, 132(6):1138–1153, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conrath:2007:SSS

[CM07] Christina L. Conrath and John A. Musick. The sandbar shark summer nursery within bays and lagoons of the Eastern shore of Virginia. Transactions of the American Fisheries Society, 136(4): 999–1007, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connor:2001:GLR

[CMBB01] William P. Connor, Anne R. Marshall, Theodore C. Bjornn, and Howard L. Burge. Growth and long-range dispersal by wild subyearling spring and summer Chinook salmon in the Snake River Basin. Transactions of the American Fisheries Society, 130(6):

1070–1076, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cho:2001:FFA

[CMKL01] Sun-Hyung Cho, Jung-Goo Myoung, Jong-Man Kim, and Jin Hwan Lee. Fish fauna associated with drifting seaweed in the coastal area of tongyeong, Korea. Transactions of the American Fisheries Society, 130(6):1190–1202, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chu:2004:IOT

[CMMM04] C. Chu, C. K. Minns, J. E. Moore, and E. S. Millard. Impact of oligotrophication, temperature, and water levels on walleye habitat in the Bay of Quinte, Lake Ontario. Transactions of the American Fisheries Society, 133(4):868–879, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cote:2001:MUJ

[CMSM01] D. Cote, S. Moulton, D. A. Scruton, and R. S. McKinley. Microhabitat use of juvenile Atlantic cod in a coastal area of Bonavista Bay, Newfoundland. Transactions of the American Fisheries Society, 130(6):1217–1223, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlsson:2001:EGS

[CN01] Jens Carlsson and Jan Nilsson. Effects of geomorphological structures on genetic differentiation among brown trout populations in a northern boreal river drainage. Transactions of the American Fisheries Society, 130(1):36–45, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chowen:2005:LSS

[CN05a] Trevor R. Chowen and James J. Nagler. Lack of sex specificity for growth hormone pseudogene in fall-run Chinook salmon from the Columbia River. Transactions of the American Fisheries Society, 134(1):279–282, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cyterski:2005:ACP

[CN05b] Michael J. Cyterski and John J. Ney. Availability of clupeid prey to primary piscivores in Smith Mountain Lake, Virginia.

Transactions of the American Fisheries Society, 134(5):1410– 1421, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:2008:INS

[CN08] Nathan R. Campbell and Shawn R. Narum. Identification of novel single-nucleotide polymorphisms in Chinook salmon and variation among life history types. Transactions of the American Fisheries Society, 137(1):96–106, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cyterski:2003:ESB

[CND03] Michael Cyterski, John Ney, and Michael Duval. Estimation of surplus biomass of clupeids in Smith Mountain Lake, Virginia. Transactions of the American Fisheries Society, 132(2):361–370, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooke:2003:CRL

[COB^{+}03] Steven J. Cooke, Kenneth G. Ostrand, Christopher M. Bunt, Jason F. Schreer, David H. Wahl, and David P. Philipp. Cardiovascular responses of largemouth bass to exhaustive exercise and brief air exposure over a range of water temperatures. Transactions of the American Fisheries Society, 132(6):1154–1165, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Copes:2002:SFS

[Cop02] Parzival Copes. Sustainable fishery systems. Transactions of the American Fisheries Society, 131(5):1004, September 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Courtenay:2007:ISW

[Cou07] Walter R. Courtenay, Jr. Introduced species: What species do you have and how do you know? Transactions of the American Fisheries Society, 136(4):1160–1164, July 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connolly:2003:BAB

[CP03] Patrick J. Connolly and James H. Petersen. Bigger is not always better for overwintering young-of-year steelhead. Transactions of the American Fisheries Society, 132(2):262–274, March 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:2006:SRV

[CPA^{+}06] Paul E. Clark, Jose J. Pereira, Linda A. Auker, Christopher J. Parkins, and Lauren M. Vinokur. Size-related variation in the diet of juvenile tautogs from Long Island Sound. Transactions of the American Fisheries Society, 135(5):1361–1370, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ciancio:2007:EDP

[CPB07] Javier E. Ciancio, Miguel A. Pascual, and David A. Beauchamp. Energy density of Patagonian aquatic organisms and empirical predictions based on water content. Transactions of the American Fisheries Society, 136(5):1415–1422, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cole:2006:CUE

[CPF06] Michael B. Cole, David M. Price, and Brian R. Fransen. Change in the upper extent of fish distribution in Eastern Washington streams between 2001 and 2002. Transactions of the American Fisheries Society, 135(3):634–642, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connor:2003:TDI

[CPG03] William P. Connor, Charles E. Piston, and Aaron P. Garcia. Temperature during incubation as one factor affecting the distribution of Snake River fall Chinook salmon spawning areas. Transactions of the American Fisheries Society, 132(6):1236–1243, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:2002:HUM

[CPRS02] Mark R. Collins, William C. Post, Daniel C. Russ, and Theodore I. J. Smith. Habitat use and movements of juvenile shortnose sturgeon in the Savannah River, Georgia–South Carolina. Transactions of the American Fisheries Society, 131(5):975–979, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:2007:BPE

[CR07] Andrew J. Carlson and Frank J. Rahel. A basinwide perspective on entrainment of fish in irrigation canals. Transactions of the American Fisheries Society, 136(5):1335–1343, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cantrell:2005:HSA

[CRA05] Christopher J. Cantrell, Anthony T. Robinson, and Lorraine D. Avenetti. Habitat selection by Apache trout in six East-Central Arizona streams. Transactions of the American Fisheries Society, 134(5):1382–1388, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collis:2001:CWP

[CRC^{+}01] Ken Collis, Daniel D. Roby, David P. Craig, Brad A. Ryan, and Richard D. Ledgerwood. Colonial waterbird predation on juvenile salmonids tagged with passive integrated transponders in the Columbia River Estuary: Vulnerability of different salmonid species, stocks, and rearing types. Transactions of the American Fisheries Society, 130(3):385–396, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collis:2002:CSD

[CRC^{+}02] Ken Collis, Daniel D. Roby, David P. Craig, Stephanie Adamany, Jessica Y. Adkins, and Donald E. Lyons. Colony size and diet composition of piscivorous waterbirds on the Lower Columbia River: Implications for losses of juvenile salmonids to avian predation. Transactions of the American Fisheries Society, 131(4): 537–550, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carls:2004:PSS

[CRMN04] Mark G. Carls, Stanley D. Rice, Gary D. Marty, and Diane K. Naydan. Pink salmon spawning habitat is recovering a decade after the Exxon Valdez oil spill. Transactions of the American Fisheries Society, 133(4):834–844, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Castro-Santos:2006:MEV

[CS06] Theodore Castro-Santos. Modeling the effect of varying swim speeds on fish passage through velocity barriers. Transactions of the American Fisheries Society, 135(5):1230–1237, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cordes:2006:IIH

[CSBM06] Jan F. Cordes, Molly R. Stephens, Marc A. Blumberg, and Bernie May. Identifying introgressive hybridization in native populations of California golden trout based on molecular markers. Transactions of the American Fisheries Society, 135(1):110–128, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Currens:2009:EER

[CSL09] Kenneth P. Currens, Carl B. Schreck, and Hiram W. Li. Evolutionary ecology of redband trout. Transactions of the American Fisheries Society, 138(4):797–817, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:2000:HUB

[CSPP00] Mark R. Collins, Theodore I. J. Smith, William C. Post, and Oleg Pashuk. Habitat utilization and biological characteristics of adult Atlantic sturgeon in two south Carolina Rivers. Transactions of the American Fisheries Society, 129(4):982–988, July 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connor:2005:TAJ

[CST^{+}05] William P. Connor, John G. Sneva, Kenneth F. Tiffan, R. Kirk Steinhorst, and Doug Ross. Two alternative juvenile life history types for fall Chinook salmon in the Snake River Basin. Transactions of the American Fisheries Society, 134(2):291–304, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Curry:2002:STM

[CSvdS02] R. Allen Curry, David Sparks, and Jacob van de Sande. Spatial and temporal movements of a riverine brook trout population. Transactions of the American Fisheries Society, 131(4):551–560, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carls:2004:OCL

[CTT04] M. G. Carls, J. F. Thedinga, and R. E. Thomas. Observer classification of live, mechanically damaged, and dead pink salmon eggs. Transactions of the American Fisheries Society, 133(2):245–251, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Currie:2001:BR

[Cur01] Rebecca J. Currie. Book review. Transactions of the American Fisheries Society, 130(1):176–178, January 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Claramunt:2000:EAB

[CW00a] Randall M. Claramunt and David H. Wahl. The effects of abiotic and biotic factors in determining larval fish growth rates: a comparison across species and reservoirs. Transactions of the American Fisheries Society, 129(3):835–851, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:2000:FBR

[CW00b] Charles C. Coutant and Richard R. Whitney. Fish behavior in relation to passage through hydropower turbines: a review. Transactions of the American Fisheries Society, 129(2):351–380, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chipps:2004:DEW

[CW04] Steven R. Chipps and David H. Wahl. Development and evaluation of a western mosquitofish bioenergetics model. Transactions of the American Fisheries Society, 133(5):1150–1162, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chipps:2008:BMC

[CW08] Steven R. Chipps and David H. Wahl. Bioenergetics modeling in the 21st century: Reviewing new insights and revisiting old constraints. Transactions of the American Fisheries Society, 137 (1):298–313, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carveth:2006:CUT

[CWB06] Corissa J. Carveth, Ann M. Widmer, and Scott A. Bonar. Comparison of upper thermal tolerances of native and nonnative fish species in Arizona. Transactions of the American Fisheries Society, 135(6):1433–1440, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cotel:2006:DBT

[CWT06] Aline J. Cotel, Paul W. Webb, and Hans Tritico. Do brown trout choose locations with reduced turbulence? Transactions of the American Fisheries Society, 135(3):610–619, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:2000:RAC

[DA00] Rex A. Dunham and Brad J. Argue. Reproduction among channel catfish, blue catfish, and their F_{1} and F_{2} hybrids. Transactions of the American Fisheries Society, 129(1):222–231, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dockendorf:2005:ABC

[DA05] Kevin J. Dockendorf and Micheal S. Allen. Age-0 black crappie abundance and size in relation to zooplankton density, stock abundance, and water clarity in three Florida lakes. Transactions of the American Fisheries Society, 134(1):172–183, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dutterer:2008:SSH

[DA08] Andrew C. Dutterer and Micheal S. Allen. Spotted sunfish habitat selection at three Florida Rivers and implications for minimum flows. Transactions of the American Fisheries Society, 137 (2):454–466, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duffy:2008:SPP

[DB08] Elisabeth J. Duffy and David A. Beauchamp. Seasonal patterns of predation on juvenile Pacific salmon by anadromous cutthroat trout in Puget Sound. Transactions of the American Fisheries Society, 137(1):165–181, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daly:2009:OSD

[DBW09] Elizabeth A. Daly, Richard D. Brodeur, and Laurie A. Weitkamp. Ontogenetic shifts in diets of juvenile and subadult Coho and Chinook salmon in coastal marine waters: Important for marine survival? Transactions of the American Fisheries Society, 138(6): 1420–1438, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

David:2002:BSD

[DC02] Bruno O. David and Gerard P. Closs. Behavior of a streamdwelling fish before, during, and after high-discharge events. Transactions of the American Fisheries Society, 131(4):762–771, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Despatie:2001:FTP

[DCCA01] S-P. Despatie, M. Castonguay, D. Chabot, and C. Audet. Final thermal preferendum of Atlantic cod: Effect of food ration. Transactions of the American Fisheries Society, 130(2):263–275, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DuBey:2007:RSE

[DCG07] Robert J. DuBey, Colleen A. Caldwell, and William R. Gould. Relative susceptibility and effects on performance of Rio Grande cutthroat trout and rainbow trout challenged with Myxobolus cerebralis. Transactions of the American Fisheries Society, 136 (5):1406–1414, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Douglas:2009:PSP

[DCHS09] Scott G. Douglas, Gerald Chaput, John Hayward, and Joseph Sheasgreen. Prespawning, spawning, and postspawning behavior of striped bass in the Miramichi River. Transactions of the American Fisheries Society, 138(1):121–134, January 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:2002:IST

[DCT02] Jason B. Dunham, Brian S. Cade, and James W. Terrell. Influences of spatial and temporal variation on fish-habitat relationships defined by regression quantiles. Transactions of the American Fisheries Society, 131(1):86–98, January 2002. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dubuc:2002:EFI

[DD02] Russell A. Dubuc and Dennis R. DeVries. An exploration of factors influencing crappie early life history in three Alabama impoundments. Transactions of the American Fisheries Society, 131(4):476–491, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauwalter:2008:STP

[DF08] Daniel C. Dauwalter and William L. Fisher. Spatial and temporal patterns in stream habitat and smallmouth bass populations in Eastern Oklahoma. Transactions of the American Fisheries Society, 137(4):1072–1088, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis-Foust:2009:AVF

[DFBC^{+}09] Shannon L. Davis-Foust, Ronald M. Bruch, Steven E. Campana, Robert P. Olynyk, and John Janssen. Age validation of freshwater drum using bomb radiocarbon. Transactions of the American Fisheries Society, 138(2):385–396, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diaz-Ferguson:2007:GRA

[DFCDR07] Edgardo Díaz-Ferguson, Ismael Cross, Maria Mar Del Barrios, and Laureana Rebordinos. Genetic relationships among populations of the Senegalese sole Solea senegalensis in the Southwestern Iberian Peninsula detected by mitochondrial DNA-restriction fragment length polymorphisms. Transactions of the American Fisheries Society, 136(2):484–491, March 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dieterman:2004:LSF

[DG04] Douglas J. Dieterman and David L. Galat. Large-scale factors associated with sicklefin chub distribution in the Missouri and Lower Yellowstone Rivers. Transactions of the American Fisheries Society, 133(3):577–587, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeGrandchamp:2007:LAR

[DGC07] Kelly L. DeGrandchamp, James E. Garvey, and Laura A. Csoboth. Linking adult reproduction and larval density of invasive carp in a large river. Transactions of the American Fisheries Society, 136(5):1327–1334, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeGrandchamp:2008:MHS

[DGC08] Kelly L. DeGrandchamp, James E. Garvey, and Robert E. Colombo. Movement and habitat selection by invasive Asian carps in a large river. Transactions of the American Fisheries Society, 137(1):45–56, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeVries:2004:ELG

[DGFS04] Paul DeVries, Fred Goetz, Kurt Fresh, and David Seiler. Evidence of a lunar gravitation cue on timing of estuarine entry by Pacific salmon smolts. Transactions of the American Fisheries Society, 133(6):1379–1395, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Danielsdottir:2008:PSD

[DGK^{+}08] Anna Kristín Daníelsdóttir, Davíd Gíslason, Kristján Kristinsson, Magnúsörn Stefánsson, Torild Johansen, and Christophe Pampoulie. Population structure of deep-sea and oceanic phenotypes of deepwater redfish in the Irminger Sea and Icelandic Continental Slope: Are they cryptic species? Transactions of the American Fisheries Society, 137(6):1723–1740, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dion:2004:TAT

[DH04] Cheryl A. Dion and Nicholas F. Hughes. Testing the ability of a temperature-based model to predict the growth of age-0 Arctic grayling. Transactions of the American Fisheries Society, 133(4): 1047–1050, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diana:2004:MPL

[DHC04] James S. Diana, John P. Hudson, and Richard D. Clark, Jr. Movement patterns of large brown trout in the mainstream Au Sable River, Michigan. Transactions of the American Fisheries Society, 133(1):34–44, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeGroot:2007:ELS

[DHR07] Jennifer D. De Groot, Scott G. Hinch, and John S. Richardson. Effects of logging second-growth forests on headwater populations of coastal cutthroat trout: a 6-year, multistream, before-and-after field experiment. Transactions of the American Fisheries Society, 136(1):211–226, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denlinger:2006:SCD

[DHS06] Jonathan C. Sieber Denlinger, R. Scott Hale, and Roy A. Stein. Seasonal consumptive demand and prey use by stocked saugeyes in Ohio reservoirs. Transactions of the American Fisheries Society, 135(1):12–27, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dillman:2006:GDA

[DK06] Casey B. Dillman and Jeffrey B. Koppelman. Genetic diversity among hatchery stocks and established populations of rainbow trout in Missouri. Transactions of the American Fisheries Society, 135(2):341–347, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeKoning:2006:GAW

[DKP^{+}06] Jenefer DeKoning, Kimberly Keatley, Ruth Phillips, James Rhydderch, John Janssen, and Marc Noakes. Genetic analysis of wild lake trout embryos recovered from Lake Michigan. Transactions of the American Fisheries Society, 135(2):399–407, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeHaan:2006:GPS

[DLES06] Patrick W. DeHaan, Scot V. Libants, Robert F. Elliott, and Kim T. Scribner. Genetic population structure of remnant lake sturgeon populations in the Upper Great Lakes Basin. Transactions of the American Fisheries Society, 135(6):1478–1492, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Franco:2004:LEH

[dlHFB04] Ernesto de la Hoz Franco and Phaedra Budy. Linking environmental heterogeneity to the distribution and prevalence of Myxobolus cerebralis: a comparison across sites in a Northern Utah watershed. Transactions of the American Fisheries Society, 133(5): 1176–1189, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolan:2003:ITE

[DM03] C. R. Dolan and L. E. Miranda. Immobilization thresholds of electrofishing relative to fish size. Transactions of the American Fisheries Society, 132(5):969–976, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauwalter:2009:TVT

[DRG09] Daniel C. Dauwalter, Frank J. Rahel, and Kenneth G. Gerow. Temporal variation in trout populations: Implications for monitoring and trend detection. Transactions of the American Fisheries Society, 138(1):38–51, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daniels:2002:IBI

[DRMH^{+}02] Robert A. Daniels, Karen Riva-Murray, David B. Halliwell, David L. Vana-Miller, and Michael D. Bilger. An index of biological integrity for Northern Mid-Atlantic slope drainages. Transactions of the American Fisheries Society, 131(6):1044– 1060, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denton:2009:DMG

[DRQ09] Keith P. Denton, Harry B. Rich, Jr., and Thomas P. Quinn. Diet, movement, and growth of Dolly Varden in response to sockeye salmon subsidies. Transactions of the American Fisheries Society, 138(6):1207–1219, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denit:2004:GVS

[DS04] Kelly Denit and Su Sponaugle. Growth variation, settlement, and spawning of gray snapper across a latitudinal gradient. Transactions of the American Fisheries Society, 133(6):1339–1355, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:2005:RPH

[DS05] M. W. Davis and C. B. Schreck. Responses by Pacific halibut to air exposure: Lack of correspondence among plasma constituents and mortality. Transactions of the American Fisheries Society, 134(4):991–998, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunlop:2006:NIP

[DS06] Erin S. Dunlop and Brian J. Shuter. Native and introduced populations of smallmouth bass differ in concordance between climate and somatic growth. Transactions of the American Fisheries Society, 135(5):1175–1190, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunlop:2007:DEC

[DSD07] Erin S. Dunlop, Brian J. Shuter, and Ulf Dieckmann. Demographic and evolutionary consequences of selective mortality: Predictions from an eco-genetic model for smallmouth bass. Transactions of the American Fisheries Society, 136(3):749–765, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauer:2009:ERH

[DSH^{+}09] M. B. Dauer, T. R. Seamons, L. Hauser, T. P. Quinn, and K. A. Naish. Estimating the ratio of hatchery-produced to wild adult steelhead on the spawning grounds using scale pattern analyses. Transactions of the American Fisheries Society, 138(1):15– 22, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diaz:2001:AGP

[DSSA01] Guillermo A. Diaz, Steven G. Smith, Joseph E. Serafy, and Jerald S. Ault. Allometry of the growth of pink shrimp Farfantepenaeus duorarum in a subtropical bay. Transactions of the American Fisheries Society, 130(2):328–335, March 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Danylchuk:2003:NDF

[DT03] Andy J. Danylchuk and William M. Tonn. Natural disturbances and fish: Local and regional influences on winterkill of fathead minnows in boreal lakes. Transactions of the American Fisheries Society, 132(2):289–298, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Durham:2006:ISD

[DW06] Bart W. Durham and Gene R. Wilde. Influence of stream discharge on reproductive success of a prairie stream fish assemblage. Transactions of the American Fisheries Society, 135(6): 1644–1653, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DAmelio:2008:GPS

[DW08a] Silvia D’Amelio and Chris Wilson. Genetic population structure among source populations for coaster brook trout in Nipigon Bay, Lake Superior. Transactions of the American Fisheries Society, 137(4):1213–1228, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeRobertis:2008:WLR

[DW08b] Alex De Robertis and Kresimir Williams. Weight–length relationships in fisheries studies: The standard allometric model should be applied with caution. Transactions of the American Fisheries Society, 137(3):707–719, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Durham:2009:PDS

[DW09] Bart W. Durham and Gene R. Wilde. Population dynamics of the smalleye shiner, an imperiled cyprinid fish endemic to the Brazos River, Texas. Transactions of the American Fisheries Society, 138(3):666–674, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diaz:2000:EHE

[DWBE00] Marilyn Diaz, David Wethey, James Bulak, and Bert Ely. Effect of harvest and effective population size on genetic diversity in a striped bass population. Transactions of the American Fisheries Society, 129(6):1367–1372, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickerson:2005:HLH

[DWBQ05] Bobette R. Dickerson, Mary F. Willson, Paul Bentzen, and Thomas P. Quinn. Heritability of life history and morphological traits in a wild pink salmon population assessed by DNA parentage analysis. Transactions of the American Fisheries Society, 134(5):1323–1328, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deeds:2008:CST

[DWEL08] Jonathan R. Deeds, Kevin D. White, Stacey M. Etheridge, and Jan H. Landsberg. Concentrations of saxitoxin and tetrodotoxin in three species of puffers from the Indian River lagoon, Florida, the location for multiple cases of saxitoxin puffer poisoning from 2002 to 2004. Transactions of the American Fisheries Society, 137 (5):1317–1326, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:2004:DBH

[DYWA^{+}04] Jana L. D. Davis, Alicia C. Young-Williams, Robert Aguilar, Benjamin L. Carswell, Michael R. Goodison, Anson H. Hines, Margaret A. Kramer, Yonathan Zohar, and Oded Zmora. Differences between hatchery-raised and wild blue crabs: Implications for stock enhancement potential. Transactions of the American Fisheries Society, 133(1):1–14, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:2001:CVM

[EBC01] Randy L. Eshenroder and Mary K. Burnham-Curtis. Comment: Vertical migration of bloaters. Transactions of the American Fisheries Society, 130(1):166–167, January 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [TF01].

Evans:2004:IES

[EBFC04] Allen F. Evans, Roy E. Beaty, Martin S. Fitzpatrick, and Ken Collis. Identification and enumeration of steelhead kelts at a Snake River hydroelectric dam. Transactions of the American Fisheries Society, 133(5):1089–1099, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eckmann:2002:EFC

[EBS02] Reiner Eckmann, Michael Becker, and Myriam Schmid. Estimating food consumption by a heavily fished stock of zooplanktivorous Coregonus lavaretus. Transactions of the American Fisheries Society, 131(5):946–955, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eggleston:2009:DBC

[EBS09] David B. Eggleston, Geoffrey W. Bell, and Steven P. Searcy. Do blue crab spawning sanctuaries in North Carolina protect the spawning stock. Transactions of the American Fisheries Society, 138(3):581–592, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ebersole:2009:HML

[ECW^{+}09a] Joseph L. Ebersole, Michael E. Colvin, Parker J. Wigington, Jr., Scott G. Leibowitz, Joan P. Baker, M. Robbins Church, Jana E. Compton, and Michael A. Cairns. Hierarchical modeling of latesummer weight and summer abundance of juvenile Coho salmon across a stream network. Transactions of the American Fisheries Society, 138(5):1138–1156, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ebersole:2009:MSN

[ECW^{+}09b] Joseph L. Ebersole, Mike E. Colvin, Parker J. Wigington, Jr., Scott G. Leibowitz, Joan P. Baker, M. Robbins Church, Jana E. Compton, Bruce A. Miller, Michael A. Cairns, Bruce P. Hansen, and Henry R. La Vigne. Modeling stream network-scale variation in Coho salmon overwinter survival and smolt size. Transactions of the American Fisheries Society, 138(3):564–580, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Essington:2003:EWL

[EH03] Timothy E. Essington and Jeffrey N. Houser. The effect of wholelake nutrient enrichment on mercury concentration in age-1 yellow perch. Transactions of the American Fisheries Society, 132(1): 57–68, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Emmett:2008:NFP

[EK08] Robert L. Emmett and Gregory K. Krutzikowsky. Nocturnal feeding of Pacific hake and Jack mackerel off the mouth of the Columbia River, 1998–2004: Implications for juvenile salmon predation. Transactions of the American Fisheries Society, 137(3): 657–676, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

English:2005:MTR

[EKS^{+}05] Karl K. English, William R. Koski, Cezary Sliwinski, Anita Blakley, Alan Cass, and James C. Woodey. Migration timing and river survival of late-run Fraser River sockeye salmon estimated using radiotelemetry techniques. Transactions of the American Fisheries Society, 134(5):1342–1365, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:2008:RBF

[EKSX08] Joyce J. Evans, Phillip H. Klesius, Richard A. Shelby, and De-Hai Xu. Relationship between frequency of skin ulceration and total plasma immunoglobulin levels in Atlantic menhaden from Delaware and Maryland inland bays. Transactions of the American Fisheries Society, 137(6):1667–1673, November 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edmundson:2001:LGJ

[EM01] J. A. Edmundson and A. Mazumder. Linking growth of juvenile sockeye salmon to habitat temperature in Alaskan lakes. Transactions of the American Fisheries Society, 130(4):644–662, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellrott:2004:LTR

[EM04a] Brian J. Ellrott and J. Ellen Marsden. Lake trout reproduction in Lake Champlain. Transactions of the American Fisheries Society, 133(2):252–264, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Esteban:2004:WME

[EM04b] Elaine M. Esteban and Michael P. Marchetti. What’s on the menu? Evaluating a food availability model with young-of-the-year Chinook salmon in the Feather River, California. Transactions of the American Fisheries Society, 133(3):777–788, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

English:2006:CAS

[ERS^{+}06] K. K. English, D. Robichaud, C. Sliwinski, R. F. Alexander, W. R. Koski, T. C. Nelson, B. L. Nass, S. A. Bickford, S. Hammond, and T. R. Mosey. Comparison of adult steelhead migrations in the Mid-Columbia hydrosystem and in large naturally flowing British Columbia Rivers. Transactions of the American Fisheries Society, 135(3):739–754, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:2009:CMD

[Esh09] Randy L. Eshenroder. Comment: Mitochondrial DNA analysis indicates sea lampreys are indigenous to Lake Ontario. Transactions of the American Fisheries Society, 138(5):1178–1189, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [WDHW09].

Emery:2003:DMI

[ESM^{+}03] Erich B. Emery, Thomas P. Simon, Frank H. McCormick, Paul L. Angermeier, Jeffrey E. Deshon, Chris O. Yoder, Randall E.

Sanders, William D. Pearson, Gary D. Hickman, Robin J. Reash, and Jeffrey A. Thomas. Development of a multimetric index for assessing the biological condition of the Ohio River. Transactions of the American Fisheries Society, 132(4):791–808, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Essington:2003:DSA

[Ess03] Timothy E. Essington. Development and sensitivity analysis of bioenergetics models for skipjack tuna and albacore: a comparison of alternative life histories. Transactions of the American Fisheries Society, 132(4):759–770, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ebersole:2006:JCS

[EWB^{+}06] Joseph L. Ebersole, Parker J. Wigington, Jr., Joan P. Baker, Michael A. Cairns, M. Robbins Church, Bruce P. Hansen, Bruce A. Miller, Henry R. LaVigne, Jana E. Compton, and Scott G. Leibowitz. Juvenile Coho salmon growth and survival across stream network seasonal habitats. Transactions of the American Fisheries Society, 135(6):1681–1697, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ferguson:2006:EDM

[FACS06] John W. Ferguson, Randall F. Absolon, Thomas J. Carlson, and Benjamin P. Sandford. Evidence of delayed mortality on juvenile Pacific salmon passing through turbines at Columbia River dams. Transactions of the American Fisheries Society, 135(1):139–150, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farrell:2001:RSS

[Far01] John M. Farrell. Reproductive success of sympatric northern pike and muskellunge in an Upper St. Lawrence River Bay. Transactions of the American Fisheries Society, 130(5):796–808, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:2008:EEC

[FB08] Dylan J. Fraser and Louis Bernatchez. Ecology, evolution, and conservation of lake-migratory brook trout: a perspective from pristine populations. Transactions of the American Fisheries Society, 137(4):1192–1202, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Floyd:2007:EWV

[FCC07] Emily Y. Floyd, Roger Churchwell, and Joseph J. Cech, Jr. Effects of water velocity and trash rack architecture on juvenile fish passage and interactions: a simulation. Transactions of the American Fisheries Society, 136(5):1177–1186, September 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Furimsky:2003:RCR

[FCS^{+}03] Marosh Furimsky, Steven J. Cooke, Cory D. Suski, Yuxiang Wang, and Bruce L. Tufts. Respiratory and circulatory responses to hypoxia in largemouth bass and smallmouth bass: Implications for “Live-Release” angling tournaments. Transactions of the American Fisheries Society, 132(6):1065–1075, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fullerton:2000:OGS

[FGWS00] Aimee H. Fullerton, James E. Garvey, Russell A. Wright, and Roy A. Stein. Overwinter growth and survival of largemouth bass: Interactions among size, food, origin, and winter severity. Transactions of the American Fisheries Society, 129(1):1–12, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Feyrer:2007:OMP

[FHB^{+}07] Frederick Feyrer, James Hobbs, Melinda Baerwald, Ted Sommer, Qing zhu Yin, Kevin Clark, Bernie May, and William Bennett. Otolith microchemistry provides information complementary to microsatellite DNA for a migratory fish. Transactions of the American Fisheries Society, 136(2):469–476, March 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Feldhaus:2008:HHS

[FHML08] Joseph W. Feldhaus, Scott A. Heppell, Matthew G. Mesa, and Hiram Li. Hepatic heat shock protein 70 and plasma cortisol levels in rainbow trout after tagging with a passive integrated transponder. Transactions of the American Fisheries Society, 137(3):690–695, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fox:2000:GSS

[FHP00] Dewayne A. Fox, Joseph E. Hightower, and Frank M. Parauka. Gulf sturgeon spawning migration and habitat in the Choctawhatchee River System, Alabama–Florida. Transactions of the American Fisheries Society, 129(3):811–826, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraboulet:2009:ELO

[FLLA09] E. Fraboulet, Y. Lambert, M. Litvak, and C. Audet. The effects of latitudinal origin and paternal contribution on larval growth of winter flounder in a northern environment. Transactions of the American Fisheries Society, 138(2):407–415, March 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fullerton:2000:PRC

[FLLG00] Aimee H. Fullerton, Gary A. Lamberti, David M. Lodge, and Frederick W. Goetz. Potential for resource competition between Eurasian ruffe and yellow perch: Growth and RNA responses in laboratory experiments. Transactions of the American Fisheries Society, 129(6):1331–1339, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fast:2008:SCS

[FNL^{+}08] David E. Fast, Douglas Neeley, David T. Lind, Mark V. Johnston, Charles R. Strom, William J. Bosch, Curtis M. Knudsen, Steve L. Schroder, and Bruce D. Watson. Survival comparison of spring Chinook salmon reared in a production hatchery under optimum conventional and seminatural conditions. Transactions of the American Fisheries Society, 137(5):1507–1518, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fritts:2004:SBP

[FP04] Anthony L. Fritts and Todd N. Pearsons. Smallmouth bass predation on hatchery and wild salmonids in the Yakima River, Washington. Transactions of the American Fisheries Society, 133(4): 880–895, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fritts:2006:EPN

[FP06] Anthony L. Fritts and Todd N. Pearsons. Effects of predation by nonnative smallmouth bass on native salmonid prey: The role of predator and prey size. Transactions of the American Fisheries Society, 135(4):853–860, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flowers:2009:SSS

[FPD^{+}09] H. Jared Flowers, W. E. Pine III, A. C. Dutterer, K. G. Johnson, J. W. Ziewitz, M. S. Allen, and Frank M. Parauka. Spawning site selection and potential implications of modified flow regimes on viability of Gulf sturgeon populations. Transactions of the American Fisheries Society, 138(6):1266–1284, November 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flebbe:2006:SMP

[FRB06] Patricia A. Flebbe, Laura D. Roghair, and Jennifer L. Bruggink. Spatial modeling to project Southern Appalachian trout distribution in a warmer climate. Transactions of the American Fisheries Society, 135(5):1371–1382, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fontenot:2001:EEH

[FRK01] Q. C. Fontenot, D. A. Rutherford, and W. E. Kelso. Effects of environmental hypoxia associated with the annual flood pulse on the distribution of larval sunfish and shad in the Atchafalaya River Basin, Louisiana. Transactions of the American Fisheries Society, 130(1):107–116, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Firehammer:2006:SMM

[FS06] Jon A. Firehammer and Dennis L. Scarnecchia. Spring migratory movements by paddlefish in natural and regulated river segments of the Missouri and Yellowstone Rivers, North Dakota and Montana. Transactions of the American Fisheries Society, 135 (1):200–217, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Feyrer:2007:LDE

[FSH07] Frederick Feyrer, Ted Sommer, and James Hobbs. Living in a dynamic environment: Variability in life history traits of age-0 splittail in tributaries of San Francisco Bay. Transactions of the American Fisheries Society, 136(5):1393–1405, September 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fonseca:2008:LCS

[FSMH08] Duane B. Fonseca, Bernard Sainte-Marie, and François Hazel. Longevity and change in shell condition of adult male snow crab Chionoecetes opilio inferred from dactyl wear and mark-recapture data. Transactions of the American Fisheries Society, 137(4): 1029–1043, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farag:2003:CAH

[FSN^{+}03] Aïda M. Farag, Don Skaar, David A. Nimick, Elizabeth Mac-Connell, and Christer Hogstrand. Characterizing aquatic health using salmonid mortality, physiology, and biomass estimates in streams with elevated concentrations of arsenic, cadmium, copper, lead, and zinc in the Boulder River watershed, Montana. Transactions of the American Fisheries Society, 132(3):450–467, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ferguson:2007:BSM

[FSR^{+}07] John W. Ferguson, Benjamin P. Sandford, Rachel E. Reagan, Lyle G. Gilbreath, Edward B. Meyer, Richard D. Ledgerwood, and Noah S. Adams. Bypass system modification at Bonneville Dam on the Columbia River improved the survival of juvenile salmon. Transactions of the American Fisheries Society, 136(6): 1487–1510, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fine:2008:UPA

[FT08] Michael L. Fine and Robert F. Thorson. Use of passive acoustics for assessing behavioral interactions in individual toadfish. Transactions of the American Fisheries Society, 137(2):627–637, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flannery:2007:EMD

[FWG07a] Blair G. Flannery, John K. Wenburg, and Anthony J. Gharrett. Evolution of mitochondrial DNA variation within and among Yukon River chum salmon populations. Transactions of the American Fisheries Society, 136(4):902–910, July 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flannery:2007:VAF

[FWG07b] Blair G. Flannery, John K. Wenburg, and Anthony J. Gharrett. Variation of amplified fragment length polymorphisms in Yukon River chum salmon: Population structure and application to mixed-stock analysis. Transactions of the American Fisheries Society, 136(4):911–925, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fontaine:2007:ETF

[FWL^{+}07] Lance P. Fontaine, Kasey W. Whiteman, Peng Li, Gary S. Burr, Kenneth A. Webb, Jonathan Goff, Delbert M. Gatlin III, William H. Neill, Kenneth B. Davis, and Robert R. Vega. Effects of temperature and feed energy on the performance of juvenile red drum. Transactions of the American Fisheries Society, 136 (5):1193–1205, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faulkner:2008:ESB

[FWTS08] Sean G. Faulkner, Marek Welz, William M. Tonn, and Douglas R. Schmitt. Effects of simulated blasting on mortality of rainbow trout eggs. Transactions of the American Fisheries Society, 137(1):1–12, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodwin:2003:DCA

[GA03] Kevin R. Goodwin and Paul L. Angermeier. Demographic characteristics of American eel in the Potomac River drainage, Virginia. Transactions of the American Fisheries Society, 132(3):524–535, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grothues:2007:SAT

[GA07] Thomas M. Grothues and Kenneth W. Able. Scaling acoustic telemetry of bluefish in an estuarine observatory: Detection and habitat use patterns. Transactions of the American Fisheries Society, 136(6):1511–1519, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geist:2000:UET

[GABC00] David R. Geist, C. Scott Abernethy, Susan L. Blanton, and Valerie I. Cullinan. The use of electromyogram telemetry to estimate energy expenditure of adult fall Chinook salmon. Transactions of the American Fisheries Society, 129(1):126–135, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geist:2006:SDG

[GAH^{+}06] David R. Geist, C. Scott Abernethy, Kristine D. Hand, Valerie I. Cullinan, James A. Chandler, and Phillip A. Groves. Survival, development, and growth of fall Chinook salmon embryos, alevins, and fry exposed to variable thermal and dissolved oxygen regimes. Transactions of the American Fisheries Society, 135(6): 1462–1477, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gannon:2008:PAT

[Gan08] Damon P. Gannon. Passive acoustic techniques in fisheries science: a review and prospectus. Transactions of the American Fisheries Society, 137(2):638–656, February 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gadomski:2001:ELH

[GBBP01] Dena M. Gadomski, Craig A. Barfoot, Jennifer M. Bayer, and Thomas P. Poe. Early life history of the northern pikeminnow in the Lower Columbia River Basin. Transactions of the American Fisheries Society, 130(2):250–262, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geist:2005:MSS

[GBC^{+}05] David R. Geist, Richard S. Brown, Valerie Cullinan, Steve R. Brink, Ken Lepla, Phil Bates, and James A. Chandler. Movement, swimming speed, and oxygen consumption of juvenile white sturgeon in response to changing flow, water temperature, and light level in the Snake River, Idaho. Transactions of the American Fisheries Society, 134(4):803–816, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffiths:2001:FAL

[GBMB01] Ronald W. Griffiths, F. W. H. Beamish, B. J. Morrison, and L. A. Barker. Factors affecting larval sea lamprey growth and length at metamorphosis in lampricide-treated streams. Transactions of the American Fisheries Society, 130(2):289–306, March 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graziano:2005:NMD

[GBN05] Sara L. Graziano, Kim H. Brown, and Jennifer L. Nielsen. Nomenclature of mitochondrial DNA haplotypes for Oncorhynchus mykiss. Transactions of the American Fisheries Society, 134(5): 1271–1273, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gruber:2001:SJL

[GdMH01] Samuel H. Gruber, Jean R. C. de Marignac, and John M. Hoenig. Survival of juvenile lemon sharks at Bimini, Bahamas, estimated by mark–depletion experiments. Transactions of the American Fisheries Society, 130(3):376–384, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gutreuter:2003:EMR

[GDW03] Steve Gutreuter, John M. Dettmers, and David H. Wahl. Estimating mortality rates of adult fish from entrainment through the propellers of river towboats. Transactions of the American Fisheries Society, 132(4):646–661, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graeb:2004:FSP

[GDWC04] Brian D. S. Graeb, John M. Dettmers, David H. Wahl, and Carla E. Cáceres. Fish size and prey availability affect growth, survival, prey selection, and foraging behavior of larval yellow perch. Transactions of the American Fisheries Society, 133(3): 504–514, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffin:2007:RBF

[GF07] Joanna D. Tober Griffin and Andrew H. Fayram. Relationships between a fish index of biotic integrity and mean length and density of brook trout and brown trout in Wisconsin streams. Transactions of the American Fisheries Society, 136(6):1728– 1735, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gatt:2000:CCR

[GFL00] Michael H. Gatt, Moira M. Ferguson, and Arunas P. Liskauskas. Comparison of control region sequencing and fragment RFLP analysis for resolving mitochondrial DNA variation and phylogenetic relationships among Great Lakes walleyes. Transactions of the American Fisheries Society, 129(6):1288–1299, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gatt:2002:MDV

[GFLF02] Michael H. Gatt, Dylan J. Fraser, Arunas P. Liskauskas, and Moira M. Ferguson. Mitochondrial DNA variation and stock structure of walleyes from Eastern Lake Huron: an analysis of contemporary and historical samples. Transactions of the American Fisheries Society, 131(1):99–108, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gallaway:2007:EPI

[GGCF07] Benny J. Gallaway, William J. Gazey, John G. Cole, and Robert G. Fechhelm. Estimation of potential impacts from offshore liquefied natural gas terminals on red snapper and red drum fisheries in the Gulf of Mexico: an alternative approach. Transactions of the American Fisheries Society, 136(3):655–677, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gerrity:2006:JPS

[GGG06] Paul C. Gerrity, Christopher S. Guy, and William M. Gardner. Juvenile pallid sturgeon are piscivorous: a call for conserving native cyprinids. Transactions of the American Fisheries Society, 135(3):604–609, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gillig:2001:BAG

[GGO01] Dhazn Gillig, Wade L. Griffin, and Teofilo Ozuna, Jr. A bioeconomic assessment of Gulf of Mexico red snapper management policies. Transactions of the American Fisheries Society, 130(1):117– 129, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gedamke:2006:EMM

[GH06] Todd Gedamke and John M. Hoenig. Estimating mortality from mean length data in nonequilibrium situations, with application to the assessment of goosefish. Transactions of the American Fisheries Society, 135(2):476–487, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gresswell:2007:PSM

[GH07] Robert E. Gresswell and Steven R. Hendricks. Population-scale movement of coastal cutthroat trout in a naturally isolated stream network. Transactions of the American Fisheries Society, 136 (1):238–253, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunckel:2002:EBT

[GHL02] Stephanie L. Gunckel, Alan R. Hemmingsen, and Judith L. Li. Effect of bull trout and brook trout interactions on foraging habitat, feeding behavior, and growth. Transactions of the American Fisheries Society, 131(6):1119–1130, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grabowski:2006:SDM

[GI06] Timothy B. Grabowski and J. Jeffery Isely. Seasonal and diel movements and habitat use of robust redhorses in the Lower Savannah River, Georgia and South Carolina. Transactions of the American Fisheries Society, 135(5):1145–1155, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gido:2002:ICP

[Gid02] Keith B. Gido. Interspecific comparisons and the potential importance of nutrient excretion by benthic fishes in a large reservoir. Transactions of the American Fisheries Society, 131(2):260–270, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greene:2005:EEC

[GJP^{+}05] Correigh M. Greene, David W. Jensen, George R. Pess, E. Ashley Steel, and Eric Beamer. Effects of environmental conditions during stream, estuary, and ocean residency on Chinook salmon return rates in the Skagit River, Washington. Transactions of the American Fisheries Society, 134(6):1562–1581, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Ano06b].

Gregalis:2009:ROR

[GJP09] Kevan C. Gregalis, Matthew W. Johnson, and Sean P. Powers. Restored oyster reef location and design affect responses of resident and transient fish, crab, and shellfish species in Mobile Bay, Alabama. Transactions of the American Fisheries Society, 138(2): 314–327, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Godinho:2006:MSR

[GK06] Alexandre L. Godinho and Boyd Kynard. Migration and spawning of radio-tagged zulega Prochilodus argenteus in a dammed Brazilian River. Transactions of the American Fisheries Society, 135(3):811–824, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goniea:2006:BTS

[GKB^{+}06] Thomas M. Goniea, Matthew L. Keefer, Theodore C. Bjornn, Christopher A. Peery, David H. Bennett, and Lowell C. Stuehrenberg. Behavioral thermoregulation and slowed migration by adult fall Chinook salmon in response to high Columbia River water temperatures. Transactions of the American Fisheries Society, 135(2):408–419, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gray:2006:IFD

[GKW^{+}06] Andrew K. Gray, Arthur W. Kendall, Jr., Bruce L. Wing, Mark G. Carls, Jonathan Heifetz, Zhuozhuo Li, and Anthony J. Gharrett. Identification and first documentation of larval rockfishes in Southeast Alaskan waters was possible using mitochondrial markers but not pigmentation patterns. Transactions of the American Fisheries Society, 135(1):1–11, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gartland:2006:DCY

[GLHA06] James Gartland, Robert J. Latour, Aimee D. Halvorson, and Herbert M. Austin. Diet composition of young-of-the-year bluefish in the Lower Chesapeake Bay and the coastal ocean of Virginia. Transactions of the American Fisheries Society, 135(2):371–378, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gibson:2002:LRM

[GM02] A. Jamie F. Gibson and Ransom A. Myers. A logistic regression model for estimating turbine mortality at hydroelectric generating stations. Transactions of the American Fisheries Society, 131(4): 623–633, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goldstein:2004:CFS

[GM04] Robert M. Goldstein and Michael R. Meador. Comparisons of fish species traits from small streams to large rivers. Transactions of the American Fisheries Society, 133(4):971–983, July 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gorman:2008:NHF

[GMCQ08] Owen T. Gorman, Seth A. Moore, Andrew J. Carlson, and Henry R. Quinlan. Nearshore habitat and fish community associations of coaster brook trout in Isle Royale, Lake Superior. Transactions of the American Fisheries Society, 137(4):1252–1267, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gephard:2000:ESN

[GMGV00] Stephen Gephard, Paloma Moran, and Eva Garcia-Vazquez. Evidence of successful natural reproduction between brown trout and mature male Atlantic salmon parr. Transactions of the American Fisheries Society, 129(1):301–306, January 2000. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graeb:2006:OCP

[GMJ^{+}06] Brian D. S. Graeb, Matthew T. Mangan, Jeffrey C. Jolley, David H. Wahl, and John M. Dettmers. Ontogenetic changes in prey preference and foraging ability of yellow perch: Insights based on relative energetic return of prey. Transactions of the American Fisheries Society, 135(6):1493–1498, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gharrett:2005:TGD

[GMP^{+}05] Anthony J. Gharrett, Andrew P. Matala, Eric L. Peterson, Andrew K. Gray, Zhouzhou Li, and Jonathan Heifetz. Two genetically distinct forms of rougheye rockfish are different species. Transactions of the American Fisheries Society, 134(1):242–260, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gharrett:2006:DGD

[GMS^{+}06] A. J. Gharrett, C. W. Mecklenburg, L. W. Seeb, Z. Li, A. P. Matala, A. K. Gray, and J. Heifetz. Do genetically distinct rougheye rockfish sibling species differ phenotypically? Transactions of the American Fisheries Society, 135(3):792–800, May 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gold:2008:GES

[GMS^{+}08] John R. Gold, Liang Ma, Eric Saillant, Paul S. Silva, and R. R. Vega. Genetic effective size in populations of hatchery-raised red drum released for stock enhancement. Transactions of the American Fisheries Society, 137(5):1327–1334, September 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gulseth:2000:BPS

[GN00] Odd A. Gulseth and Kjell J. Nilssen. The brief period of spring migration, short marine residence, and high return rate of a Northern Svalbard population of Arctic char. Transactions of the American Fisheries Society, 129(3):782–796, May 2000. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gadomski:2005:ETL

[GP05] Dena M. Gadomski and Michael J. Parsley. Effects of turbidity, light level, and cover on predation of white sturgeon larvae by prickly sculpins. Transactions of the American Fisheries Society, 134(2):369–374, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glowacki:2008:NSR

[GP08] Łukasz Głowacki and Tadeusz Penczak. New species richness predictors first tested on fish in a small tropical stream. Transactions of the American Fisheries Society, 137(2):431–437, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garrett:2007:HSI

[GPUH07] Daniel L. Garrett, Theodore W. Pietsch, Fred M. Utter, and Lorenz Hauser. The hybrid sole Inopsetta ischyra (Teleostei: Pleuronectiformes: Pleuronectidae): Hybrid or biological species? Transactions of the American Fisheries Society, 136(2):460–468, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grover:2006:ENR

[Gro06] Mark C. Grover. Evaluation of a negative relationship between abundance during spawning and size at maturity in Kokanee.

Transactions of the American Fisheries Society, 135(4):970–978, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:2005:CPT

[GRR05] David J. Graham, Stephen P. Rice, and Ian Reid. Comment: Photographic techniques for characterizing streambed particle sizes. Transactions of the American Fisheries Society, 134(6): 1599–1603, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [OM05].

Guenther:2006:CFA

[GS06] Cameron B. Guenther and Anne Spacie. Changes in fish assemblage structure upstream of impoundments within the Upper Wabash River Basin, Indiana. Transactions of the American Fisheries Society, 135(3):570–583, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greenberg:2001:EDS

[GSP01] Larry A. Greenberg, Thomas Steinwall, and Henrik Persson. Effect of depth and substrate on use of stream pools by brown trout. Transactions of the American Fisheries Society, 130(4):699–705, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glover:2003:SSG

[GSS03] K. A. Glover, O. T. Skilbrei, and Ø Skaala. Stock-specific growth and length frequency bimodality in brown trout. Transactions of the American Fisheries Society, 132(2):307–315, March 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gurtin:2003:UER

[GSSB03] Scott D. Gurtin, Joe E. Slaughter IV, Schuyler J. Sampson, and Richard H. Bradford. Use of existing and reconnected backwater habitats by hatchery-reared adult razorback suckers: a predictive model for the Imperial Division, Lower Colorado River. Transactions of the American Fisheries Society, 132(6):1125– 1137, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gobalet:2004:APN

[GSWS04] Kenneth W. Gobalet, Peter D. Schulz, Thomas A. Wake, and Nelson Siefkin. Archaeological perspectives on native American fisheries of California, with emphasis on steelhead and salmon. Transactions of the American Fisheries Society, 133(4):801–833, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garcia-Vazquez:2009:MTC

[GVHC^{+}09] Eva Garcia-Vazquez, Jose L. Horreo, Daniel Campo, Gonzalo Machado-Schiaffino, Iliana Bista, Alexandros Triantafyllidis, and Francis Juanes. Mislabeling of two commercial North American hake species suggests underreported exploitation of offshore hake. Transactions of the American Fisheries Society, 138(4):790–796, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grant:2002:SIB

[GVS02] Gerold C. Grant, Bruce Vondracek, and Peter W. Sorensen. Spawning interactions between sympatric brown and brook trout may contribute to species replacement. Transactions of the American Fisheries Society, 131(4):569–576, May 2002. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galarowicz:2003:DGC

[GW03] Tracy L. Galarowicz and David H. Wahl. Differences in growth, consumption, and metabolism among walleyes from different latitudes. Transactions of the American Fisheries Society, 132(3): 425–437, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geiger:2007:WCV

[GWM^{+}07] Harold J. Geiger, Ivan Wang, Pat Malecha, Kyle Hebert, William W. Smoker, and Anthony J. Gharrett. What causes variability in pink salmon family size? Transactions of the American Fisheries Society, 136(6):1688–1698, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garvey:2000:EHL

[GWSF00] James E. Garvey, Russell A. Wright, Roy A. Stein, and Kristen H. Ferry. Evaluating how local- and regional-scale processes interact to regulate growth of age-0 largemouth bass. Transactions of the American Fisheries Society, 129(4):1044–1059, July 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2000:VDR

[Har00] K. J. Hartman. Variability in daily ration estimates of age-0 striped bass in the Chesapeake Bay. Transactions of the American Fisheries Society, 129(5):1181–1186, September 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heidarsson:2006:RBB

[HAS06] Thorkell Heidarsson, Thorolfur Antonsson, and Sigurdur S. Snorrason. The relationship between body and scale growth proportions and validation of two back-calculation methods using individually tagged and recaptured wild Atlantic salmon. Transactions of the American Fisheries Society, 135(5):1156–1164, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoxmeier:2009:EIV

[HAW09] R. John H. Hoxmeier, D. Derek Aday, and David H. Wahl. Examining interpopulation variation in bluegill growth rates and size structure: Effects of harvest, maturation, and environmental variables. Transactions of the American Fisheries Society, 138(2): 423–432, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinch:2000:ESS

[HB00] Scott G. Hinch and Jessica Bratty. Effects of swim speed and activity pattern on success of adult sockeye salmon migration through an area of difficult passage. Transactions of the American Fisheries Society, 129(2):598–606, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hata:2004:FAH

[HB04] David Hata and Jim Berkson. Factors affecting horseshoe crab Limulus polyphemus trawl survey design. Transactions of the American Fisheries Society, 133(2):292–299, March 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houston:2006:GVS

[HB06] Derek D. Houston and Mark C. Belk. Geographic variation in somatic growth of redside shiner. Transactions of the American Fisheries Society, 135(3):801–810, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

He:2007:MAG

[HB07] J. X. He and J. R. Bence. Modeling annual growth variation using a hierarchical Bayesian approach and the von Bertalanffy growth function, with application to lake trout in Southern Lake Huron. Transactions of the American Fisheries Society, 136(2): 318–330, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Homel:2008:TSV

[HB08a] Kristen Homel and Phaedra Budy. Temporal and spatial variability in the migration patterns of juvenile and subadult bull trout in Northeastern Oregon. Transactions of the American Fisheries Society, 137(3):869–880, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huckins:2008:MBC

[HB08b] Casey J. Huckins and Edward A. Baker. Migrations and biological characteristics of adfluvial coaster brook trout in a south shore Lake Superior tributary. Transactions of the American Fisheries Society, 137(4):1229–1243, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heery:2009:SEL

[HB09] Eliza C. Heery and Jim Berkson. Systematic errors in length frequency data and their effect on age-structured stock assessment models and management. Transactions of the American Fisheries Society, 138(1):218–232, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hobbs:2007:CLA

[HBBG07] James A. Hobbs, William A. Bennett, Jessica Burton, and Michelle Gras. Classification of larval and adult delta smelt to nursery areas by use of trace elemental fingerprinting. Transactions of the American Fisheries Society, 136(2):518–527, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hymel:2002:SPJ

[HBC^{+}02] M. K. Hymel, D. M. Baltz, E. J. Chesney, M. A. Tarr, and A. S. Kolok. Swimming performance of juvenile Florida pompano exposed to ethylene glycol. Transactions of the American Fisheries Society, 131(6):1152–1163, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:2008:SGS

[HBH^{+}08] Sean A. Hayes, Morgan H. Bond, Chad V. Hanson, Ellen V. Freund, Jerry J. Smith, Eric C. Anderson, Arnold J. Ammann, and R. Bruce MacFarlane. Steelhead growth in a small Central California watershed: Upstream and estuarine rearing patterns. Transactions of the American Fisheries Society, 137(1):114–128, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

He:2008:MVM

[HBJ^{+}08a] Ji X. He, James R. Bence, James E. Johnson, David F. Clapp, and Mark P. Ebener. Modeling variation in mass-length relations and condition indices of lake trout and Chinook salmon in Lake Huron: a hierarchical Bayesian approach. Transactions of the American Fisheries Society, 137(3):801–817, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heath:2008:GDA

[HBJ^{+}08b] Daniel D. Heath, Corwyn M. Bettles, Sara Jamieson, Iga Stasiak, and Margaret F. Docker. Genetic differentiation among sympatric migratory and resident life history forms of rainbow trout in British Columbia. Transactions of the American Fisheries Society, 137(4):1268–1278, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffman:2007:TNH

[HBO07] J. C. Hoffman, D. A. Bronk, and J. E. Olney. Tracking nursery habitat use in the York River Estuary, Virginia, by young American shad using stable isotopes. Transactions of the American Fisheries Society, 136(5):1285–1297, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heise:2009:PVE

[HBP^{+}09] Ryan J. Heise, Robert B. Bringolf, Rhonda Patterson, W. Gregory Cope, and Stephen T. Ross. Plasma vitellogenin and estradiol concentrations in adult Gulf sturgeon from the Pascagoula River Drainage, Mississippi. Transactions of the American Fisheries Society, 138(5):1028–1035, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havens:2009:LLL

[HBSA09] Kirk J. Havens, Donna Marie Bilkovic, Dave Stanhope, and Kory Angstadt. Location, location, location: The importance of cull ring placement in blue crab traps. Transactions of the American Fisheries Society, 138(4):720–724, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heggenes:2006:GDS

[HBTT06] Jan Heggenes, Mark Beere, Patrick Tamkee, and Eric B. Taylor. Genetic diversity in steelhead before and after conservation hatchery operation in a coastal, boreal river. Transactions of the American Fisheries Society, 135(1):251–267, January 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2005:ESS

[HC05] S. Dale Hanson and R. Allen Curry. Effects of size structure on trophic interactions between age-0 smallmouth bass and juvenile anadromous alewives. Transactions of the American Fisheries Society, 134(2):356–368, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2008:RTB

[HC08] Kyle J. Hartman and M. Keith Cox. Refinement and testing of a brook trout bioenergetics model. Transactions of the American Fisheries Society, 137(1):357–363, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffnagle:2006:EMI

[HCK06] Timothy L. Hoffnagle, Richard W. Carmichael, and Patrick J. Keniry. The effect of moderately increased and variable raceway flow rates on juvenile physiology, survival, and adult return of hatchery-reared Chinook salmon. Transactions of the American Fisheries Society, 135(6):1567–1577, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:2009:EGS

[HCK^{+}09a] J. E. Hall, J. Chamberlin, A. N. Kagley, C. Greene, and K. L. Fresh. Effects of gastric and surgical insertions of dummy ultrasonic transmitters on juvenile Chinook salmon in seawater. Transactions of the American Fisheries Society, 138(1):52–57, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holecek:2009:JCS

[HCK09b] Dean E. Holecek, Kara J. Cromwell, and Brian P. Kennedy. Juvenile Chinook salmon summer microhabitat availability, use, and selection in a Central Idaho wilderness stream. Transactions of the American Fisheries Society, 138(3):633–644, May 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hrabik:2001:IBY

[HCW01] Thomas R. Hrabik, Michael P. Carey, and Michael S. Webster. Interactions between young-of-the-year exotic rainbow smelt and native yellow perch in a northern temperate lake. Transactions of the American Fisheries Society, 130(4):568–582, July 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hendry:2000:PCR

[HDH00] Andrew P. Hendry, Andrew H. Dittman, and Ronald W. Hardy. Proximate composition, reproductive development, and a test for trade-offs in captive sockeye salmon. Transactions of the American Fisheries Society, 129(5):1082–1095, September 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horton:2007:MEP

[HDL07] Gregg E. Horton, Todd L. Dubreuil, and Benjamin H. Letcher. A model for estimating Passive Integrated Transponder (PIT) tag antenna efficiencies for interval-specific emigration rates. Transactions of the American Fisheries Society, 136(5):1165–1176, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heimbuch:2007:APE

[HDRB07] Douglas G. Heimbuch, Dennis J. Dunning, Quentin E. Ross, and Alan F. Blumberg. Assessing potential effects of entrainment and impingement on fish stocks of the New York–New Jersey Harbor Estuary and Long Island Sound. Transactions of the American Fisheries Society, 136(2):492–508, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilborn:2000:RHP

[HE00] Ray Hilborn and Doug Eggers. A review of the hatchery programs for pink salmon in Prince William Sound and Kodiak Island, Alaska. Transactions of the American Fisheries Society, 129(2):333–350, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilborn:2001:RHP

[HE01] Ray Hilborn and Doug Eggers. A review of the hatchery programs for pink salmon in Prince William Sound and Kodiak Island, Alaska: Response to comment. Transactions of the American Fisheries Society, 130(4):720–724, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [WSJH01].

Healey:2001:BR

[Hea01] Michael Healey. Book reviews. Transactions of the American Fisheries Society, 130(3):531–532, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heggenes:2002:FSH

[Heg02] Jan Heggenes. Flexible summer habitat selection by wild, allopatric brown trout in lotic environments. Transactions of the American Fisheries Society, 131(2):287–298, March 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heimbuch:2008:PES

[Hei08] Douglas G. Heimbuch. Potential effects of striped bass predation on juvenile fish in the Hudson River. Transactions of the American Fisheries Society, 137(6):1591–1605, November 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hershberger:2005:ALO

[HEW^{+}05] P. K. Hershberger, N. E. Elder, J. Wittouck, K. Stick, and R. M. Kocan. Abnormalities in larvae from the once-largest Pacific herring population in Washington state result primarily from factors independent of spawning location. Transactions of the American Fisheries Society, 134(2):326–337, March 2005. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hofmann:2002:TPC

[HF02] N. Hofmann and P. Fischer. Temperature preferences and critical thermal limits of burbot: Implications for habitat selection and ontogenetic habitat shift. Transactions of the American Fisheries Society, 131(6):1164–1172, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harper:2004:WHU

[HF04] David D. Harper and Aïda M. Farag. Winter habitat use by cutthroat trout in the Snake River near Jackson, Wyoming. Transactions of the American Fisheries Society, 133(1):15–25, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinrichsen:2009:ILM

[HF09] Richard A. Hinrichsen and Timothy R. Fisher. Inferences on the latent mortality of Snake River spring–summer-run Chinook salmon using spawner–recruit models. Transactions of the American Fisheries Society, 138(6):1232–1239, November 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamon:2000:SMS

[HFHR00] Troy R. Hamon, Chris J. Foote, Ray Hilborn, and Donald E. Rogers. Selection on morphology of spawning wild sockeye salmon by a gill-net fishery. Transactions of the American Fisheries Society, 129(6):1300–1315, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heist:2000:DML

[HG00] Edward J. Heist and John R. Gold. DNA microsatellite loci and genetic structure of red snapper in the Gulf of Mexico. Transactions of the American Fisheries Society, 129(2):469–475, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoenig:2007:SME

[HG07] John M. Hoenig and Todd Gedamke. A simple method for estimating survival rate from catch rates from multiple years. Transactions of the American Fisheries Society, 136(5):1245– 1251, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henry:2004:STF

[HGJO04] Theodore B. Henry, John M. Grizzle, Carol E. Johnston, and Jason A. Osborne. Susceptibility of ten fish species to electroshockinduced mortality. Transactions of the American Fisheries Society, 133(3):649–654, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henry:2003:EIM

[HGM03] Theodore B. Henry, John M. Grizzle, and Michael J. MacEina. Electroshocking-induced mortality of four fish species during posthatching development. Transactions of the American Fisheries Society, 132(2):299–306, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2008:LVP

[HGR^{+}08] Kyle C. Hanson, Marie-Ange Gravel, Tara Redpath, Steven J. Cooke, and Michael J. Siepker. Latitudinal variation in physiological and behavioral responses of nest-guarding smallmouth bass to common recreational angling practices. Transactions of the American Fisheries Society, 137(5):1558–1566, September 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huff:2004:ECS

[HGW04] David D. Huff, Gabriella Grad, and Craig E. Williamson. Environmental constraints on spawning depth of yellow perch: The roles of low temperature and high solar ultraviolet radiation. Transactions of the American Fisheries Society, 133(3):718–726, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hicks:2003:RTC

[HH03] Brendan J. Hicks and James D. Hall. Rock type and channel gradient structure salmonid populations in the Oregon Coast Range. Transactions of the American Fisheries Society, 132(3):468–482, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoenig:2005:WCW

[HH05] John M. Hoenig and David A. Hewitt. What can we learn about mortality from sex ratio data? A look at lumpfish in Newfoundland. Transactions of the American Fisheries Society, 134(3): 754–761, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [HH06].

Hoenig:2006:EWC

[HH06] J. M. Hoenig and D. A. Hewitt. Erratum: What Can We Learn about Mortality from Sex Ratio Data? A Look at Lumpfish in Newfoundland. Transactions of the American Fisheries Society, 135(3):e1, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [HH05].

Hughes:2007:EDN

[HH07] Robert M. Hughes and Alan T. Herlihy. Electrofishing distance needed to estimate consistent index of biotic integrity (IBI) scores in raftable Oregon Rivers. Transactions of the American Fisheries Society, 136(1):135–141, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Honeyfield:2002:ITP

[HHB02] Dale C. Honeyfield, Joy P. Hinterkopf, and Scott B. Brown. Isolation of thiaminase-positive bacteria from alewife. Transactions of the American Fisheries Society, 131(1):171–175, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hale:2003:PPA

[HHDC03] R. Scott Hale, John K. Horne, Donald J. Degan, and M. Elizabeth Conners. Paddlefish as potential acoustic targets for abundance estimates. Transactions of the American Fisheries Society, 132(4):746–758, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrington:2008:HBL

[HHHK08] Steven J. Herrington, Kurt N. Hettiger, Edward J. Heist, and Devon B. Keeney. Hybridization between longnose and alligator gars in captivity, with comments on possible gar hybridization in nature. Transactions of the American Fisheries Society, 137 (1):158–164, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:2004:BII

[HHK04] Robert M. Hughes, Shay Howlin, and Philip R. Kaufmann. A biointegrity index (IBI) for coldwater streams of Western Oregon and Washington. Transactions of the American Fisheries Society, 133(6):1497–1515, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2004:DDR

[HHS04] K. J. Hartman, J. Howell, and J. A. Sweka. Diet and daily ration of bay anchovy in the Hudson River, New York. Transactions of the American Fisheries Society, 133(3):762–771, May 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayer:2008:IGM

[HI08] Cari-Ann Hayer and Elise R. Irwin. Influence of gravel mining and other factors on detection probabilities of coastal plain fishes in the Mobile River Basin, Alabama. Transactions of the American Fisheries Society, 137(6):1606–1620, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilborn:2000:BR

[Hil00] Ray Hilborn. Book review. Transactions of the American Fisheries Society, 129(3):887, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hitt:2004:BRS

[Hit04] Nathaniel P. Hitt. Biological response signatures. Transactions of the American Fisheries Society, 133(1):232–233, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hardiman:2004:DPI

[HJM04] Jill M. Hardiman, Brett M. Johnson, and Patrick J. Martinez. Do predators influence the distribution of age-0 Kokanee in a Colorado reservoir? Transactions of the American Fisheries Society, 133(6):1366–1378, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hightower:2001:UTM

[HJP01] Joseph E. Hightower, James R. Jackson, and Kenneth H. Pollock. Use of telemetry methods to estimate natural and fishing mortality of striped bass in Lake Gaston, North Carolina. Transactions of the American Fisheries Society, 130(4):557–567, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilderbrand:2000:MPS

[HK00] Robert H. Hilderbrand and Jeffrey L. Kershner. Movement patterns of stream-resident cutthroat trout in Beaver Creek, Idaho– Utah. Transactions of the American Fisheries Society, 129(5): 1160–1170, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hickey:2004:CBC

[HK04a] Christopher W. Hickey and Christopher C. Kohler. Comparison of bluegill consumption rates by largemouth bass and sunshine bass in structured and nonstructured artificial environments. Transactions of the American Fisheries Society, 133(6): 1524–1528, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilderbrand:2004:TDG

[HK04b] Robert H. Hilderbrand and Jeffrey L. Kershner. Are there differences in growth and condition between mobile and resident cutthroat trout? Transactions of the American Fisheries Society, 133(4):1042–1046, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2008:BMP

[HK08] Kyle J. Hartman and James F. Kitchell. Bioenergetics modeling: Progress since the 1992 Symposium. Transactions of the American Fisheries Society, 137(1):216–223, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holliman:2007:EUM

[HKCL07] F. Michael Holliman, Thomas J. Kwak, W. Gregory Cope, and Jay F. Levine. Exposure of unionid mussels to electric current: Assessing risks associated with electrofishing. Transactions of the American Fisheries Society, 136(6):1593–1606, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hewitt:2009:PDI

[HKCP09] Amanda H. Hewitt, Thomas J. Kwak, W. Gregory Cope, and Kenneth H. Pollock. Population density and instream habitat suitability of the endangered Cape Fear shiner. Transactions of the American Fisheries Society, 138(6):1439–1457, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horan:2000:EHA

[HKHC00] Dona L. Horan, Jeffrey L. Kershner, Charles P. Hawkins, and Todd A. Crowl. Effects of habitat area and complexity on Colorado River cutthroat trout density in Uinta Mountain streams.

Transactions of the American Fisheries Society, 129(6):1250– 1263, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:2000:EAC

[HL00a] John Mark Hanson and Marc Lanteigne. Evaluation of Atlantic cod predation on American lobster in the Southern Gulf of St. Lawrence, with comments on other potential fish predators. Transactions of the American Fisheries Society, 129(1):13– 29, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Healy:2000:MUB

[HL00b] Brian D. Healy and David G. Lonzarich. Microhabitat use and behavior of overwintering juvenile Coho salmon in a Lake Superior tributary. Transactions of the American Fisheries Society, 129(3): 866–872, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hewitt:2007:DIE

[HLH^{+}07] David A. Hewitt, Debra M. Lambert, John M. Hoenig, Romuald N. Lipcius, David B. Bunnell, and Thomas J. Miller. Direct and indirect estimates of natural mortality for Chesapeake Bay blue crab. Transactions of the American Fisheries Society, 136 (4):1030–1040, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hannah:2007:BNS

[HM07] Robert W. Hannah and Keith M. Matteson. Behavior of nine species of Pacific rockfish after hook-and-line capture, recompression, and release. Transactions of the American Fisheries Society, 136(1):24–33, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heist:2008:RGS

[HM08] Edward J. Heist and Akeem Mustapha. Rangewide genetic structure in paddlefish inferred from DNA microsatellite loci. Transactions of the American Fisheries Society, 137(3):909–915, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2009:ASF

[HM09] Julianne E. Harris and Richard S. McBride. American shad feeding on spawning grounds in the St. Johns River, Florida. Transactions of the American Fisheries Society, 138(4):888–898, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:2002:MDA

[HMD02] Matthew R. Hall, Raymond P. Morgan II, and Roy G. Danzmann. Mitochondrial DNA analysis of Mid-Atlantic populations of brook trout: The zone of contact for major historical lineages. Transactions of the American Fisheries Society, 131(6): 1140–1151, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hueter:2006:AMR

[HMT^{+}06] Robert E. Hueter, Charles A. Manire, John P. Tyminski, John M. Hoenig, and Daniel A. Hepworth. Assessing mortality of released or discarded fish using a logistic model of relative survival derived from tagging data. Transactions of the American Fisheries Society, 135(2):500–508, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2007:LHH

[HMW07] Julianne E. Harris, Richard S. McBride, and Roy O. Williams. Life history of hickory shad in the St. Johns River, Florida. Transactions of the American Fisheries Society, 136(6):1463– 1471, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henson:2000:ETG

[HN00] Fred G. Henson and Raymond M. Newman. Effect of temperature on growth at ration and gastric evacuation rate of ruffe. Transactions of the American Fisheries Society, 129(2):552–560, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:2005:TSL

[HN05] K. J. Hartman and B. W. Nagy. A target strength and length relationship for striped bass and white perch. Transactions of the American Fisheries Society, 134(2):375–380, March 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heintz:2004:MSF

[HNH^{+}04] Ron A. Heintz, Bonita D. Nelson, John Hudson, Marie Larsen, Larry Holland, and Mark Wipfli. Marine subsidies in freshwater: Effects of salmon carcasses on lipid class and fatty acid composition of juvenile Coho salmon. Transactions of the American Fisheries Society, 133(3):559–567, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:2006:RSL

[HNW06] Bret C. Harvey, Rodney J. Nakamoto, and Jason L. White. Reduced streamflow lowers dry-season growth of rainbow trout in a small stream. Transactions of the American Fisheries Society, 135(4):998–1005, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffman:2005:CSG

[HO05] Joel C. Hoffman and John E. Olney. Cohort-specific growth and mortality of juvenile American shad in the Pamunkey River, Virginia. Transactions of the American Fisheries Society, 134(1): 1–18, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Humston:2004:BAM

[HOA04] Robert Humston, Donald B. Olson, and Jerald S. Ault. Behavioral assumptions in models of fish movement and their influence on population dynamics. Transactions of the American Fisheries Society, 133(6):1304–1328, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holt:2008:DRD

[Hol08] Scott A. Holt. Distribution of red drum spawning sites identified by a towed hydrophone array. Transactions of the American Fisheries Society, 137(2):551–561, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hondorp:2005:IDD

[HPB05] Darryl W. Hondorp, Steven A. Pothoven, and Stephen B. Brandt. Influence of Diporeia density on diet composition, relative abundance, and energy density of planktivorous fishes in Southeast Lake Michigan. Transactions of the American Fisheries Society, 134(3):588–601, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

High:2006:TSC

[HPB06] Brett High, Christopher A. Peery, and David H. Bennett. Temporary staging of Columbia River summer steelhead in coolwater areas and its effect on migration rates. Transactions of the American Fisheries Society, 135(2):519–528, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2005:DGM

[HPM05] Julianne E. Harris, Daryl C. Parkyn, and Debra J. Murie. Distribution of Gulf of Mexico sturgeon in relation to benthic invertebrate prey resources and environmental parameters in the Suwannee River Estuary, Florida. Transactions of the American Fisheries Society, 134(4):975–990, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamel:2008:RDM

[HPW08] Owen S. Hamel, Kevin R. Piner, and John R. Wallace. A robust deterministic model describing the bomb radiocarbon signal for use in fish age validation. Transactions of the American Fisheries Society, 137(3):852–859, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:2009:EPS

[HR09] Bret C. Harvey and Steven F. Railsback. Exploring the persistence of stream-dwelling trout populations under alternative realworld turbidity regimes with an individual-based model. Transactions of the American Fisheries Society, 138(2):348–360, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hook:2007:HDG

[HRMC07] Tomas O. Höök, Edward S. Rutherford, Doran M. Mason, and Glenn S. Carter. Hatch dates, growth, survival, and overwinter mortality of age-0 alewives in Lake Michigan: Implications for habitat-specific recruitment success. Transactions of the American Fisheries Society, 136(5):1298–1312, September 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Higgins:2008:MSF

[HS08] Christopher L. Higgins and Richard E. Strauss. Modeling stream fish assemblages with niche apportionment models: Patterns, processes, and scale dependence. Transactions of the American Fisheries Society, 137(3):696–706, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hiscock:2002:WMR

[HSBC02] Martha J. Hiscock, David A. Scruton, Joseph A. Brown, and Keith D. Clarke. Winter movement of radio-tagged juvenile Atlantic salmon in Northeast Brook, Newfoundland. Transactions of the American Fisheries Society, 131(4):577–581, May 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hurst:2000:SED

[HSC00] Thomas P. Hurst, Eric T. Schultz, and David O. Conover. Seasonal energy dynamics of young-of-the-year Hudson River striped bass. Transactions of the American Fisheries Society, 129(1):145– 157, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hurley:2004:HUM

[HSH^{+}04] Keith L. Hurley, Robert J. Sheehan, Roy C. Heidinger, Paul S. Wills, and Bob Clevenstine. Habitat use by Middle Mississippi River pallid sturgeon. Transactions of the American Fisheries Society, 133(4):1033–1041, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [SMZB05].

Hurley:2005:ELT

[HSH^{+}05] Keith L. Hurley, Robert J. Sheehan, Roy C. Heidinger, Paul S. Wills, and Bob Clevenstine. Erratum: Long-Term Residence of Red Snapper on Artificial Reefs in the Northeastern Gulf of Mexico. Transactions of the American Fisheries Society, 134(2):1, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [SS05].

Harvey:2003:TRA

[HSK03] Chris J. Harvey, Stephen T. Schram, and James F. Kitchell. Trophic relationships among lean and Siscowet lake trout in Lake Superior. Transactions of the American Fisheries Society, 132(2): 219–228, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heise:2004:SAM

[HSRD04] Ryan J. Heise, William T. Slack, Stephen T. Ross, and Mark A. Dugo. Spawning and associated movement patterns of Gulf sturgeon in the Pascagoula River Drainage, Mississippi. Transactions of the American Fisheries Society, 133(1):221–230, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:2000:DTW

[HSS00] John W. Hayes, John D. Stark, and Karen A. Shearer. Development and test of a whole-lifetime foraging and bioenergetics growth model for drift-feeding brown trout. Transactions of the American Fisheries Society, 129(2):315–332, March 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Habicht:2007:GED

[HSS07] Christopher Habicht, Lisa W. Seeb, and James E. Seeb. Genetic and ecological divergence defines population structure of sockeye salmon populations returning to Bristol Bay, Alaska, and provides a tool for admixture analysis. Transactions of the American Fisheries Society, 136(1):82–94, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howland:2001:IFA

[HTBT01] Kimberly L. Howland, William M. Tonn, John A. Babaluk, and Ross F. Tallman. Identification of freshwater and anadromous inconnu in the Mackenzie River system by analysis of otolith strontium. Transactions of the American Fisheries Society, 130(5): 725–741, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howland:2000:MPF

[HTT00] Kimberly L. Howland, Ross F. Tallman, and William M. Tonn. Migration patterns of freshwater and anadromous inconnu in the Mackenzie River system. Transactions of the American Fisheries Society, 129(1):41–59, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:2004:HFA

[HW04] Julie L. Hall and Robert C. Wissmar. Habitat factors affecting sockeye salmon redd site selection in off-channel ponds of a river floodplain. Transactions of the American Fisheries Society, 133 (6):1480–1496, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haag:2008:ESD

[HW08a] Wendell R. Haag and Melvin L. Warren, Jr. Effects of severe drought on freshwater mussel assemblages. Transactions of the American Fisheries Society, 137(4):1165–1178, July 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:2008:UBP

[HW08b] Bret C. Harvey and Jason L. White. Use of benthic prey by salmonids under turbid conditions in a laboratory stream. Transactions of the American Fisheries Society, 137(6):1756–1763, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Halbisen:2009:VIS

[HW09] Michael A. Halbisen and Chris C. Wilson. Variable introgression from supplemental stocking in Southern Ontario populations of lake trout. Transactions of the American Fisheries Society, 138(4):699–719, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoxmeier:2004:GSL

[HWHP04] R. John H. Hoxmeier, David H. Wahl, Mike L. Hooe, and Clay L. Pierce. Growth and survival of larval walleyes in response to prey availability. Transactions of the American Fisheries Society, 133(1):45–54, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:2002:ECE

[HWLC02] J. A. Hansen, P. G. Welsh, J. Lipton, and D. Cacela. Effects of copper exposure on growth and survival of juvenile bull trout. Transactions of the American Fisheries Society, 131(4):690–697, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2004:AGR

[HWM04] Patrick J. Harris, David M. Wyanski, and Paulette T. Powers Mikell. Age, growth, and reproductive biology of blueline tilefish along the Southeastern Coast of the United States, 1982– 1999. Transactions of the American Fisheries Society, 133(5): 1190–1204, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hendry:2002:GPV

[HWMH02] Michael A. Hendry, John K. Wenburg, Katherine W. Myers, and Andrew P. Hendry. Genetic and phenotypic variation through the migratory season provides evidence for multiple populations of wild steelhead in the Dean River, British Columbia. Transactions of the American Fisheries Society, 131(4):418–434, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:2002:HRL

[HWN02] Bret C. Harvey, Jason L. White, and Rodney J. Nakamoto. Habitat relationships and larval drift of native and nonindigenous fishes in neighboring tributaries of a coastal California River. Transactions of the American Fisheries Society, 131(1):159–170, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Halvorsen:2009:BES

[HWS^{+}09] Michele B. Halvorsen, Lidia E. Wysocki, Carla M. Stehr, David H. Baldwin, David R. Chicoine, Nathaniel L. Scholz, and Arthur N. Popper. Barging effects on sensory systems of Chinook salmon smolts. Transactions of the American Fisheries Society, 138(4): 777–789, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henne:2006:SHS

[HWW^{+}06] James P. Henne, Kent M. Ware, William R. Wayman, Robert S. Bakal, and Ákos Horváth. Synchronous hermaphroditism and self-fertilization in a captive shortnose sturgeon. Transactions of the American Fisheries Society, 135(1):55–60, January 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:2007:AGR

[HWW^{+}07] Patrick J. Harris, David M. Wyanski, D. Byron White, Paulette P. Mikell, and Pamela B. Eyo. Age, growth, and reproduction of greater amberjack off the Southeastern U.S. Atlantic Coast. Transactions of the American Fisheries Society, 136(6):1534– 1545, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ihde:2002:CCS

[IC02] Thomas F. Ihde and Mark E. Chittenden, Jr. Comparison of calcified structures for aging spotted seatrout. Transactions of the American Fisheries Society, 131(4):634–642, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irwin:2003:EPP

[IDW03] Brian J. Irwin, Dennis R. DeVries, and Russell A. Wright. Evaluating the potential for predatory control of gizzard shad by largemouth bass in small impoundments: a bioenergetics approach. Transactions of the American Fisheries Society, 132(5):913–924, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaak:2004:NRT

[IH04] Daniel J. Isaak and Wayne A. Hubert. Nonlinear response of trout abundance to summer stream temperatures across a thermally diverse montane landscape. Transactions of the American Fisheries Society, 133(5):1254–1259, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ihde:2008:IRA

[IHF08] Thomas F. Ihde, John M. Hoenig, and Stewart D. Frusher. An index-removal abundance estimator that allows for seasonal change in catchability, with application to southern rock lobster Jasus edwardsii. Transactions of the American Fisheries Society, 137(3):720–735, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isely:2000:ICG

[IK00] J. Jeffery Isely and Chris Kempton. Influence of costocking on growth of young-of-year brook trout and rainbow trout. Transactions of the American Fisheries Society, 129(2):613–617, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iguchi:2004:PIN

[IMM^{+}04] Kei’ichiro Iguchi, Keiichi Matsuura, Kristina M. McNyset, A. Townsend Peterson, Ricardo Scachetti-Pereira, Katherine A. Powers, Dave A. Vieglais, E. O. Wiley, and Taiga Yodo. Predicting invasions of North American basses in Japan using native range data and a genetic algorithm. Transactions of the American Fisheries Society, 133(4):845–854, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irwin:2009:DMD

[IRJ^{+}09] Brian J. Irwin, Lars G. Rudstam, James R. Jackson, Anthony J. VanDeValk, John L. Forney, and Dean G. Fitzgerald. Depensatory mortality, density-dependent growth, and delayed compensation: Disentangling the interplay of mortality, growth, and density during early life stages of yellow perch. Transactions of the American Fisheries Society, 138(1):99–110, January 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irvine:2007:PSW

[ISH07] Robyn L. Irvine, Dana C. Schmidt, and Larry R. Hildebrand. Population status of white sturgeon in the Lower Columbia River within Canada. Transactions of the American Fisheries Society, 136(6):1472–1479, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tsuboi:2008:RBC

[iTE08] Jun ichi Tsuboi and Shinsuke Endou. Relationships between catch per unit effort, catchability, and abundance based on actual measurements of salmonids in a mountain stream. Transactions of the American Fisheries Society, 137(2):496–502, February 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2000:EPS

[JAJ^{+}00] Gary E. Johnson, Noah S. Adams, Robert L. Johnson, Dennis W. Rondorf, Dennis D. Dauble, and Theresa Y. Barila. Evaluation of the prototype surface bypass for salmonid smolts in spring 1996 and 1997 at Lower Granite Dam on the Snake River, Washington. Transactions of the American Fisheries Society, 129(2):381–397, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2005:MDA

[JCB^{+}05] Eric L. Johnson, Tami S. Clabough, David H. Bennett, Theodore C. Bjornn, Christopher A. Peery, Christopher C. Caudill, and Lowell C. Stuehrenberg. Migration depths of adult spring and summer Chinook salmon in the Lower Columbia and Snake Rivers in relation to dissolved gas supersaturation. Transactions of the American Fisheries Society, 134(5):1213–1227, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jowett:2007:CCH

[JD07] Ian G. Jowett and Andrew J. H. Davey. A comparison of composite habitat suitability indices and generalized additive models of invertebrate abundance and fish presence–habitat availability. Transactions of the American Fisheries Society, 136(2):428–444, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jepsen:2001:PRC

[JDSS01] N. Jepsen, L. E. Davis, C. B. Schreck, and B. Siddens. The physiological response of Chinook salmon smolts to two methods of radio-tagging. Transactions of the American Fisheries Society, 130(3):495–500, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jessop:2000:EPS

[Jes00] B. M. Jessop. Estimates of population size and instream mortality rate of American eel elvers in a Nova Scotia river. Transactions of the American Fisheries Society, 129(2):514–526, March 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jardine:2005:SIV

[JGMC05] Timothy D. Jardine, Michelle A. Gray, Sherisse M. McWilliam, and Richard A. Cunjak. Stable isotope variability in tissues of temperate stream fishes. Transactions of the American Fisheries Society, 134(5):1103–1110, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jackson:2002:EBT

[JHS02] John K. Jackson, Richard J. Horwitz, and Bernard W. Sweeney. Effects of Bacillus thuringiensis israelensis on black flies and nontarget macroinvertebrates and fish in a large river. Transactions of the American Fisheries Society, 131(5):910–930, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jacobson:2008:FEL

[JJRP08] Peter C. Jacobson, Thomas S. Jones, Pat Rivers, and Donald L. Pereira. Field estimation of a lethal oxythermal niche boundary for adult ciscoes in Minnesota lakes. Transactions of the American Fisheries Society, 137(5):1464–1474, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jordan:2006:PMH

[JKWS06] George R. Jordan, Robert A. Klumb, Greg A. Wanner, and Wayne J. Stancill. Poststocking movements and habitat use of hatchery-reared juvenile pallid sturgeon in the Missouri River below Fort Randall Dam, South Dakota and Nebraska. Transactions of the American Fisheries Society, 135(6):1499–1511, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:2005:GAA

[JM05] Olaf P. Jensen and Thomas J. Miller. Geostatistical analysis of the abundance and winter distribution patterns of the blue crab Callinectes sapidus in Chesapeake Bay. Transactions of the American Fisheries Society, 134(6):1582–1598, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jastrebski:2009:ESV

[JM09] C. J. Jastrebski and Y. E. Morbey. Egg size variation in lake trout: Phenotype–habitat correlations show an effect of rearing environment. Transactions of the American Fisheries Society, 138 (6):1342–1351, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2002:TTI

[JMS02] Brett M. Johnson, Patrick J. Martinez, and Jason D. Stockwell. Tracking trophic interactions in coldwater reservoirs using naturally occurring stable isotopes. Transactions of the American Fisheries Society, 131(1):1–13, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jackson:2000:FYC

[JN00a] James R. Jackson and Richard L. Noble. First-year cohort dynamics and overwinter mortality of juvenile largemouth bass. Transactions of the American Fisheries Society, 129(3):716–726, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jackson:2000:RBA

[JN00b] James R. Jackson and Richard L. Noble. Relationships between annual variations in reservoir conditions and age-0 largemouth bass year-class strength. Transactions of the American Fisheries Society, 129(3):699–715, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2009:RSL

[JPLSH09] Donald R. Johnson, Harriet M. Perry, Joanne Lyczkowski-Shultz, and David Hanisko. Red snapper larval transport in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 138(3):458–470, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnstone:2003:ATT

[JR03] Helene C. Johnstone and Frank J. Rahel. Assessing temperature tolerance of Bonneville cutthroat trout based on constant and cycling thermal regimes. Transactions of the American Fisheries Society, 132(1):92–99, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jacobs:2008:TSS

[JRM^{+}08] John M. Jacobs, Matt R. Rhodes, Brandon Martin, Dennis McIntosh, William F. Van Heukelem, and Reginal M. Harrell. Tissuespecific sampling for the estimation of striped bass whole-body proximate composition. Transactions of the American Fisheries Society, 137(2):386–392, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Janney:2008:DAL

[JSH^{+}08] Eric C. Janney, Rip S. Shively, Brian S. Hayes, Patrick M. Barry, and David Perkins. Demographic analysis of Lost River sucker and shortnose sucker populations in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 137(6): 1812–1825, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jessop:2009:LHA

[JSI09] B. M. Jessop, J. C. Shiao, and Y. Iizuka. Life history of American eels from Western Newfoundland. Transactions of the American Fisheries Society, 138(4):861–871, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:2004:EPC

[JT04] Nicholas E. Jones and William M. Tonn. Enhancing productive capacity in the Canadian Arctic: Assessing the effectiveness of instream habitat structures in habitat compensation. Transactions of the American Fisheries Society, 133(6):1356–1365, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2008:DUS

[JTM08] Scott W. Johnson, John F. Thedinga, and Kristen M. Munk. Distribution and use of shallow-water habitats by Pacific sand lances in Southeastern Alaska. Transactions of the American Fisheries Society, 137(5):1455–1463, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2009:EDE

[JTPW09] Christopher L. Johnson, Gabriel M. Temple, Todd N. Pearsons, and Timothy D. Webster. An evaluation of data entry error and proofing methods for fisheries data. Transactions of the American Fisheries Society, 138(3):593–601, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kanda:2001:GPS

[KA01] Naohisa Kanda and Fred W. Allendorf. Genetic population structure of bull trout from the Flathead River Basin as shown by microsatellites and mitochondrial DNA markers. Transactions of the American Fisheries Society, 130(1):92–106, January 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koehle:2007:ETD

[KA07] Jessica J. Koehle and Ira R. Adelman. The effects of temperature, dissolved oxygen, and Asian tapeworm infection on growth and survival of the Topeka shiner. Transactions of the American Fisheries Society, 136(6):1607–1613, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kimmerer:2005:VLW

[KAB^{+}05] Wim Kimmerer, Sean R. Avent, Steven M. Bollens, Fred Feyrer, Lenny F. Grimaldo, Peter B. Moyle, Matt Nobriga, and Tammie Visintainer. Variability in length–weight relationships used to estimate biomass of estuarine fish from survey data. Transactions of the American Fisheries Society, 134(2):481–495, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocovsky:2009:ESS

[KAB09] Patrick M. Kocovsky, Jean V. Adams, and Charles R. Bronte. The effect of sample size on the stability of principal components analysis of truss-based fish morphometrics. Transactions of the American Fisheries Society, 138(3):487–496, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kandl:2001:EIS

[Kan01] Karen L. Kandl. Effects of inbreeding and salinity stress on population dynamics of Eastern mosquitofish. Transactions of the American Fisheries Society, 130(6):1224–1232, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Korstrom:2006:ESS

[KB06] Jillian S. Korstrom and Ian K. Birtwell. Effects of suspended sediment on the escape behavior and cover-seeking response of juvenile Chinook salmon in freshwater. Transactions of the American Fisheries Society, 135(4):1006–1016, July 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:2008:IBF

[KBJ08] Matthew S. Kendall, Laurie J. Bauer, and Christopher F. G. Jeffrey. Influence of benthic features and fishing pressure on size and distribution of three exploited reef fishes from the Southeastern United States. Transactions of the American Fisheries Society, 137(4):1134–1146, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keeler:2007:ASD

[KBPC07] Rachel A. Keeler, AndréR. Breton, Douglas P. Peterson, and Richard A. Cunjak. Apparent survival and detection estimates for PIT-Tagged slimy sculpin in five small New Brunswick streams. Transactions of the American Fisheries Society, 136(1):281–292, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ketola:2000:ETR

[KBW^{+}00] H. George Ketola, Paul R. Bowser, Gregory A. Wooster, Leslie R. Wedge, and Steven S. Hurst. Effects of thiamine on reproduction of Atlantic salmon and a new hypothesis for their extirpation in Lake Ontario. Transactions of the American Fisheries Society, 129(2):607–612, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocovsky:2005:SPA

[KC05] Patrick M. Kocovsky and Robert F. Carline. Stream pH as an abiotic gradient influencing distributions of trout in Pennsylvania streams. Transactions of the American Fisheries Society, 134(5): 1299–1312, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocovsky:2006:ILS

[KC06] Patrick M. Kocovsky and Robert F. Carline. Influence of landscape-scale factors in limiting brook trout populations in Pennsylvania streams. Transactions of the American Fisheries Society, 135(1):76–88, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keeler:2007:RES

[KC07] Rachel A. Keeler and Richard A. Cunjak. Reproductive ecology of slimy sculpin in small New Brunswick streams. Transactions of the American Fisheries Society, 136(6):1762–1768, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Korman:2009:EHN

[KC09] Josh Korman and Steven E. Campana. Effects of hydropeaking on nearshore habitat use and growth of age-0 rainbow trout in a large regulated river. Transactions of the American Fisheries Society, 138(1):76–87, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaplan:2007:CHP

[KCK07] Isaac C. Kaplan, Sean P. Cox, and James F. Kitchell. Circle hooks for Pacific longliners: Not a panacea for marlin and shark bycatch, but part of the solution. Transactions of the American Fisheries Society, 136(2):392–401, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kozfkay:2007:IHB

[KCY^{+}07] Christine C. Kozfkay, Matthew R. Campbell, Steven P. Yundt, Michael P. Peterson, and Madison S. Powell. Incidence of hybridization between naturally sympatric westslope cutthroat trout and rainbow trout in the Middle Fork Salmon River drainage, Idaho. Transactions of the American Fisheries Society, 136(3): 624–638, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kim:2001:AFP

[KD01] Gene W. Kim and Dennis R. DeVries. Adult fish predation on freshwater limnetic fish larvae: a mesocosm experiment. Transactions of the American Fisheries Society, 130(2):189–203, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:2005:RDS

[KESB05] Tim L. King, Michael S. Eackles, Adrian P. Spidle, and H. Jane Brockmann. Regional differentiation and sex-biased dispersal among populations of the horseshoe crab Limulus polyphemus. Transactions of the American Fisheries Society, 134(2):441–465, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kuligowski:2005:BSS

[KFB05] David R. Kuligowski, Michael J. Ford, and Barry A. Berejikian. Breeding structure of steelhead inferred from patterns of genetic relatedness among nests. Transactions of the American Fisheries Society, 134(5):1202–1212, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koehler:2006:DBL

[KFB^{+}06] Michele E. Koehler, Kurt L. Fresh, David A. Beauchamp, Jeffery R. Cordell, Charles A. Simenstad, and David E. Seiler. Diet and bioenergetics of lake-rearing juvenile Chinook salmon in Lake Washington. Transactions of the American Fisheries Society, 135 (6):1580–1591, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kappenman:2009:ETG

[KFT^{+}09] Kevin M. Kappenman, William C. Fraser, Matt Toner, Jan Dean, and Molly A. H. Webb. Effect of temperature on growth, condition, and survival of juvenile shovelnose sturgeon. Transactions of the American Fisheries Society, 138(4):927–937, July 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koppelman:2000:AMV

[KGS00] Jeffrey B. Koppelman, Craig M. Gale, and John S. Stanovick. Allozyme and morphological variation among three nominal species of Ambloplites (Centrarchidae) inhabiting theOzarksRegion. Transactions of the American Fisheries Society, 129(5): 1134–1149, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kinziger:2007:MDV

[KGS07] Andrew P. Kinziger, Damon H. Goodman, and Rebecca S. Studebaker. Mitochondrial DNA variation in the Ozark highland members of the banded sculpin Cottus carolinae complex. Transactions of the American Fisheries Society, 136(6):1742–1749, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kemp:2005:FSB

[KGW05] Paul S. Kemp, Mike H. Gessel, and John G. Williams. Fine-scale behavioral responses of Pacific salmonid smolts as they encounter divergence and acceleration of flow. Transactions of the American Fisheries Society, 134(2):390–398, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kynard:2000:HUS

[KHKS00] Boyd Kynard, Martin Horgan, Micah Kieffer, and David Seibel. Habitats used by shortnose sturgeon in two Massachusetts Rivers, with notes on estuarine Atlantic sturgeon: a hierarchical approach. Transactions of the American Fisheries Society, 129(2): 487–503, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kassler:2008:SRH

[KHT08] Todd W. Kassler, Denise K. Hawkins, and Jack M. Tipping. Summer-run hatchery steelhead have naturalized in the South Fork Skykomish River, Washington. Transactions of the American Fisheries Society, 137(3):763–771, May 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaufman:2007:RBB

[KJL^{+}07] S. D. Kaufman, T. A. Johnston, W. C. Leggett, M. D. Moles, J. M. Casselman, and A. I. Schulte-Hostedde. Relationships between body condition indices and proximate composition in adult walleyes. Transactions of the American Fisheries Society, 136(6): 1566–1576, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keskinen:2008:BMZ

[KJM^{+}08] Tapio Keskinen, Juha Jääskeläinen, Timo J. Marjomäki, Tero Matilainen, and Juha Karjalainen. A bioenergetics model for zander: Construction, validation, and evaluation of uncertainty caused by multiple input parameters. Transactions of the American Fisheries Society, 137(6):1741–1755, November 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kanda:2002:EIH

[KLA02] Naohisa Kanda, Robb F. Leary, and Fred W. Allendorf. Evidence of introgressive hybridization between bull trout and brook trout. Transactions of the American Fisheries Society, 131(4):772–782, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kanda:2002:MGM

[KLSA02] Naohisa Kanda, Robb F. Leary, Paul Spruell, and Fred W. Allendorf. Molecular genetic markers identifying hybridization between the Colorado River-greenback cutthroat trout complex and Yellowstone cutthroat trout or rainbow trout. Transactions of the American Fisheries Society, 131(2):312–319, March 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kosa:2001:PCV

[KM01] Jarrad T. Kosa and Martha E. Mather. Processes contributing to variability in regional patterns of juvenile river herring abundance across small coastal systems. Transactions of the American Fisheries Society, 130(4):600–619, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kondratieff:2005:TAW

[KM05] Matthew C. Kondratieff and Christopher A. Myrick. Two adjustable waterfalls for evaluating fish jumping performance. Transactions of the American Fisheries Society, 134(2):503–508, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kondratieff:2006:EHH

[KM06a] M. C. Kondratieff and C. A. Myrick. Erratum: How High Can Brook Trout Jump? A Laboratory Evaluation of Brook Trout Jumping Performance. Transactions of the American Fisheries Society, 134(3):e2, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [KM06b].

Kondratieff:2006:HHC

[KM06b] Matthew C. Kondratieff and Christopher A. Myrick. How high can brook trout jump? A laboratory evaluation of brook trout jumping performance. Transactions of the American Fisheries Society, 135(2):361–370, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [KM06a].

Kostow:2003:NSH

[KMP03] Kathryn E. Kostow, Anne R. Marshall, and Stevan R. Phelps. Naturally spawning hatchery steelhead contribute to smolt production but experience low reproductive success. Transactions of the American Fisheries Society, 132(4):780–790, July 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knudsen:2002:GSC

[KMSL02] Kathy L. Knudsen, Clint C. Muhlfeld, George K. Sage, and Robb F. Leary. Genetic structure of Columbia River redband trout populations in the Kootenai River drainage, Montana, revealed by microsatellite and allozyme loci. Transactions of the American Fisheries Society, 131(6):1093–1105, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krkosek:2005:NAJ

[KMV05] Martin Krkošek, Alexandra Morton, and John P. Volpe. Nonlethal assessment of juvenile pink and chum salmon for parasitic sea lice infections and fish health. Transactions of the American Fisheries Society, 134(3):711–716, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koel:2004:SVF

[Koe04] Todd M. Koel. Spatial variation in fish species richness of the Upper Mississippi River system. Transactions of the American Fisheries Society, 133(4):984–1003, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kolz:2006:ECA

[Kol06] A. Lawrence Kolz. Electrical conductivity as applied to electrofishing. Transactions of the American Fisheries Society, 135 (2):509–518, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kondolf:2000:ASS

[Kon00] G. Mathias Kondolf. Assessing salmonid spawning gravel quality. Transactions of the American Fisheries Society, 129(1):262–281, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keefer:2004:HDR

[KPB^{+}04] Matthew L. Keefer, Christopher A. Peery, Theodore C. Bjornn, Michael A. Jepson, and Lowell C. Stuehrenberg. Hydrosystem, dam, and reservoir passage rates of adult Chinook salmon and steelhead in the Columbia and Snake Rivers. Transactions of the American Fisheries Society, 133(6):1413–1439, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keefer:2008:MTC

[KPC08] Matthew L. Keefer, Christopher A. Peery, and Christopher C. Caudill. Migration timing of Columbia River spring Chinook salmon: Effects of temperature, river discharge, and ocean environment. Transactions of the American Fisheries Society, 137 (4):1120–1133, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kiffney:2007:RBP

[KR07] P. M. Kiffney and P. Roni. Relationships between productivity, physical habitat, and aquatic invertebrate and vertebrate populations of forest streams: an information-theoretic approach. Transactions of the American Fisheries Society, 136(4):1088–1103, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klumb:2003:CAY

[KRM03] Robert A. Klumb, Lars G. Rudstam, and Edward L. Mills. Comparison of alewife young-of-the-year and adult respiration and swimming speed bioenergetics model parameters: Implications of extrapolation. Transactions of the American Fisheries Society, 132(6):1089–1103, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaeser:2001:IAR

[KS01] Adam J. Kaeser and William E. Sharpe. The influence of acidic runoff episodes on slimy sculpin reproduction in stone run. Transactions of the American Fisheries Society, 130(6):1106–1115, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knudsen:2006:CLH

[KSB^{+}06] Curtis M. Knudsen, Steve L. Schroder, Craig A. Busack, Mark V. Johnston, Todd N. Pearsons, William J. Bosch, and David E. Fast. Comparison of life history traits between first-generation hatchery and wild Upper Yakima River spring Chinook salmon. Transactions of the American Fisheries Society, 135(4):1130– 1144, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knudsen:2008:CFR

[KSB^{+}08] Curtis M. Knudsen, Steve L. Schroder, Craig Busack, Mark V. Johnston, Todd N. Pearsons, and Charles R. Strom. Comparison of female reproductive traits and progeny of first-generation hatchery and wild Upper Yakima River spring Chinook salmon. Transactions of the American Fisheries Society, 137(5):1433– 1445, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Killen:2006:FCP

[KSC^{+}06] Shaun S. Killen, Cory D. Suski, Steven J. Cooke, David P. Philipp, and Bruce L. Tufts. Factors contributing to the physiological disturbance in walleyes during simulated live-release angling tournaments. Transactions of the American Fisheries Society, 135(3): 557–569, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelsey:2002:EJS

[KSCD02] D. A. Kelsey, C. B. Schreck, J. L. Congleton, and L. E. Davis. Effects of juvenile steelhead on juvenile Chinook salmon behavior and physiology. Transactions of the American Fisheries Society, 131(4):676–689, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kruzic:2001:CMS

[KSR01] Lance M. Kruzic, Dennis L. Scarnecchia, and Brett B. Roper. Comparison of midsummer survival and growth of age-0 hatchery Coho salmon held in pools and riffles. Transactions of the American Fisheries Society, 130(1):147–154, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kondou:2001:ESF

[KTNK01] Takuya Kondou, Naohiko Takeshita, Akinobu Nakazono, and Seirô Kimura. Egg survival in a fluvial population of masu salmon in relation to intragravel conditions in spawning redds. Transactions of the American Fisheries Society, 130(5):969–974, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kanno:2009:IRS

[KVD^{+}09] Yoichiro Kanno, Jason C. Vokoun, Daniel C. Dauwalter, Robert M. Hughes, Alan T. Herlihy, Terry R. Maret, and Tim M. Patton. Influence of rare species on electrofishing distance when estimating species richness of stream and river reaches. Transactions of the American Fisheries Society, 138(6):1240–1251, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knights:2002:HML

[KVZD02] Brent C. Knights, Jonathan M. Vallazza, Steven J. Zigler, and Michael R. Dewey. Habitat and movement of lake sturgeon in the Upper Mississippi River system, USA. Transactions of the American Fisheries Society, 131(4):507–522, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kolar:2003:PJW

[KWH03] Cynthia S. Kolar, David H. Wahl, and Michael L. Hooe. Piscivory in juvenile walleyes: Relative importance of prey species, timing of spawning of prey fish, and density on growth and survival. Transactions of the American Fisheries Society, 132(4):679–690, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Korman:2009:EFS

[KYWC09] Josh Korman, Mike Yard, Carl Walters, and Lewis G. Coggins. Effects of fish size, habitat, flow, and density on capture probabilities of age-0 rainbow trout estimated from electrofishing at discrete sites in a large river. Transactions of the American Fisheries Society, 138(1):58–75, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kostow:2006:EIS

[KZ06] Kathryn E. Kostow and Shijie Zhou. The effect of an introduced summer steelhead hatchery stock on the productivity of a wild winter steelhead population. Transactions of the American Fisheries Society, 135(3):825–841, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lauer:2004:CMS

[LAM04] Thomas E. Lauer, Paul J. Allen, and Thomas S. McComish. Changes in mottled sculpin and Johnny darter trawl catches after the appearance of round gobies in the Indiana waters of Lake Michigan. Transactions of the American Fisheries Society, 133 (1):185–189, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LeBreton:2000:SLS

[LB00] Gregory T. O. LeBreton and F. William H. Beamish. Suitability of lake sturgeon growth rings in chronology development. Transactions of the American Fisheries Society, 129(4):1018–1030, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lowerre-Barbieri:2008:UPA

[LBBF^{+}08] Susan K. Lowerre-Barbieri, Luiz R. Barbieri, J. R. Flanders, A. G. Woodward, C. F. Cotton, and M. Katheryn Knowlton. Use of passive acoustics to determine red drum spawning in Georgia waters. Transactions of the American Fisheries Society, 137(2):562–575, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larsen:2004:AHR

[LBC^{+}04] Donald A. Larsen, Brian R. Beckman, Kathleen A. Cooper, Dan Barrett, Mark Johnston, Penny Swanson, and Walton W. Dickhoff. Assessment of high rates of precocious male maturation in a spring Chinook salmon supplementation hatchery program. Transactions of the American Fisheries Society, 133(1):98–120, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leips:2001:SCM

[LBR^{+}01] J. Leips, C. T. Baril, F. H. Rodd, D. N. Reznick, F. Bashey, G. J. Visser, and J. Travis. The suitability of calcein to mark poeciliid fish and a new method of detection. Transactions of the American Fisheries Society, 130(3):501–507, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larsen:2006:GMA

[LBS^{+}06] Donald A. Larsen, Brian R. Beckman, Charles R. Strom, Paul J. Parkins, Kathleen A. Cooper, David E. Fast, and Walton W. Dickhoff. Growth modulation alters the incidence of early male maturation and physiological development of hatchery-reared spring Chinook salmon: a comparison with wild fish. Transactions of the American Fisheries Society, 135(4):1017–1032, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lowerre-Barbieri:2003:CRF

[LBVW03] Susan K. Lowerre-Barbieri, Frederic E. Vose, and James A. Whittington. Catch-and-release fishing on a spawning aggregation of common snook: Does it affect reproductive output? Transactions of the American Fisheries Society, 132(5):940–952, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lobon-Cervia:2003:SDB

[LC03] Javier Lobón-Cerviá. Spatiotemporal dynamics of brown trout production in a Cantabrian stream: Effects of density and habitat quality. Transactions of the American Fisheries Society, 132(4): 621–637, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lapointe:2007:SOS

[LCM07] Nicolas W. R. Lapointe, Lynda D. Corkum, and Nicholas E. Mandrak. Seasonal and ontogenic shifts in microhabitat selection by fishes in the shallow waters of the Detroit River, a large connecting channel. Transactions of the American Fisheries Society, 136 (1):155–166, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lutz-Carrillo:2006:AAF

[LCNB^{+}06] Dijar J. Lutz-Carrillo, Chris C. Nice, Timothy H. Bonner, Michael R. J. Forstner, and Loraine T. Fries. Admixture analysis of Florida largemouth bass and northern largemouth bass using microsatellite loci. Transactions of the American Fisheries Society, 135(3):779–791, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lachance:2008:TSQ

[LDDB08] Stephanie Lachance, Maryse Dubé, Renaud Dostie, and Pierre Bérubé. Temporal and spatial quantification of fine-sediment accumulation downstream of culverts in brook trout habitat. Transactions of the American Fisheries Society, 137(6):1826– 1838, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lester:2004:LTK

[LDK^{+}04] Nigel P. Lester, Alan J. Dextrase, Robert S. Kushneriuk, Michael R. Rawson, and Phil A. Ryan. Light and temperature: Key factors affecting walleye abundance and production. Transactions of the American Fisheries Society, 133(3):588–605, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lupes:2006:CRS

[LDOS06] Sarah C. Lupes, Michael W. Davis, Bori L. Olla, and Carl B. Schreck. Capture-related stressors impair immune system function in sablefish. Transactions of the American Fisheries Society, 135(1):129–138, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Legault:2005:PVA

[Leg05] Christopher M. Legault. Population viability analysis of Atlantic salmon in Maine, USA. Transactions of the American Fisheries Society, 134(3):549–562, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lonzarich:2009:SMJ

[LFA09] David George Lonzarich, Ryan Patrick Franckowiak, and Matthew David Allen. Summer movements of juvenile Coho salmon under variable stream flow conditions. Transactions of the American Fisheries Society, 138(2):397–406, March 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ludsin:2006:CSB

[LFG06] Stuart A. Ludsin, Brian J. Fryer, and Joel E. Gagnon. Comparison of solution-based versus laser ablation inductively coupled plasma mass spectrometry for analysis of larval fish otolith microelemental composition. Transactions of the American Fisheries Society, 135(1):218–231, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Letcher:2002:SSD

[LGJ02] Benjamin H. Letcher, Gabe Gries, and Francis Juanes. Survival of stream-dwelling Atlantic salmon: Effects of life history variation, season, and age. Transactions of the American Fisheries Society, 131(5):838–854, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Levin:2003:SEH

[LH03] Phillip S. Levin and Mark E. Hay. Selection of estuarine habitats by juvenile gags in experimental mesocosms. Transactions of the American Fisheries Society, 132(1):76–83, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lambert:2006:TRE

[LHL06] Debra M. Lambert, John M. Hoenig, and Romuald N. Lipcius. Tag return estimation of annual and semiannual survival rates of adult female blue crabs. Transactions of the American Fisheries Society, 135(6):1592–1603, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:2002:ODF

[LHM02] Gary L. Larson, Robert L. Hoffman, and Stephen E. Moore. Observations of the distributions of five fish species in a small Appalachian stream. Transactions of the American Fisheries Society, 131(4):791–796, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Latour:2001:STN

[LHOP01] Robert J. Latour, John M. Hoenig, John E. Olney, and Kenneth H. Pollock. A simple test for nonmixing in multiyear tagging studies: Application to striped bass tagged in the Rappahannock River, Virginia. Transactions of the American Fisheries Society, 130(5):848–856, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linley:2001:CFF

[Lin01] Timothy J. Linley. A comparison of first feeding characteristics in two populations of Chinook salmon. Transactions of the American Fisheries Society, 130(3):519–525, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Link:2004:GMS

[Lin04] Jason S. Link. A general model of selectivity for fish feeding: a rank proportion algorithm. Transactions of the American Fisheries Society, 133(3):655–673, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:2007:SFC

[LJC^{+}07] Michael Lewis, Stephen Jordan, Cynthia Chancy, Linda Harwell, Larry Goodman, and Robert Quarles. Summer fish community of the coastal Northern Gulf of Mexico: Characterization of a large-scale trawl survey. Transactions of the American Fisheries Society, 136(3):829–845, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lubinski:2008:RBF

[LJE08] Benjamin J. Lubinski, John R. Jackson, and Michael A. Eggleton. Relationships between floodplain lake fish communities and environmental variables in a large river-floodplain ecosystem. Transactions of the American Fisheries Society, 137(3):895–908, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lacroix:2004:EID

[LKM04] Gilles L. Lacroix, Derek Knox, and Paul McCurdy. Effects of implanted dummy acoustic transmitters on juvenile Atlantic salmon. Transactions of the American Fisheries Society, 133(1):211–220, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LAbee-Lund:2004:STV

[LLVB04] J. H. L’Abée-Lund, L. A. Vøllestad, and S. Beldring. Spatial and temporal variation in the grilse proportion of Atlantic salmon in Norwegian Rivers. Transactions of the American Fisheries Society, 133(3):743–761, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Locascio:2008:DPF

[LM08a] James V. Locascio and David A. Mann. Diel periodicity of fish sound production in Charlotte Harbor, Florida. Transactions of the American Fisheries Society, 137(2):606–615, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luecke:2008:LEG

[LM08b] Chris Luecke and Peter MacKinnon. Landscape effects on growth of age-0 Arctic grayling in tundra streams. Transactions of the American Fisheries Society, 137(1):236–243, January 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Logsdon:2009:ELT

[LMA09] Dale E. Logsdon, Loren M. Miller, and Charles S. Anderson. Evaluation of long-term retention and detection of oxytetracycline marks in walleye otoliths using genetic methodology. Transactions of the American Fisheries Society, 138(4):872–887, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LaHood:2008:REF

[LMAF08] Eric S. LaHood, Jason J. Miller, Chase Apland, and Michael J. Ford. A rapid, ethanol-free fish tissue collection method for molecular genetic analyses. Transactions of the American Fisheries Society, 137(4):1104–1107, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindley:2008:MMN

[LME^{+}08] Steven T. Lindley, Mary L. Moser, Daniel L. Erickson, Michael Belchik, David W. Welch, Erin L. Rechisky, John T. Kelly, Joseph Heublein, and A. Peter Klimley. Marine migration of North American green sturgeon. Transactions of the American Fisheries Society, 137(1):182–194, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lacroix:2004:MAS

[LMK04] Gilles L. Lacroix, Paul McCurdy, and Derek Knox. Migration of Atlantic salmon postsmolts in relation to habitat use in a coastal system. Transactions of the American Fisheries Society, 133(6): 1455–1471, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luczkovich:2008:PAT

[LMR08] Joseph J. Luczkovich, David A. Mann, and Rodney A. Rountree. Passive acoustics as a tool in fisheries science. Transactions of the American Fisheries Society, 137(2):533–541, February 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Logerwell:2001:MRC

[Log01] Elizabeth A. Logerwell. Metabolic rate of California Pacific sardine estimated from energy losses during starvation. Transactions of the American Fisheries Society, 130(3):526–530, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lonzarich:2004:SRR

[Lon04] David Lonzarich. Strategies for restoring river ecosystems: Sources of variability and uncertainty in natural and managed systems. Transactions of the American Fisheries Society, 133(5): 1292–1293, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luczkovich:2008:ISC

[LPJS08] Joseph J. Luczkovich, R. Christopher Pullinger, Stephen E. Johnson, and Mark W. Sprague. Identifying sciaenid critical spawning habitats by the use of passive acoustics. Transactions of the American Fisheries Society, 137(2):576–605, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liao:2001:EAI

[LPL01] Hongsheng Liao, Clay L. Pierce, and Joe G. Larscheid. Empirical assessment of indices of prey importance in the diets of predacious fish. Transactions of the American Fisheries Society, 130(4):583– 591, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lyons:2001:DVA

[LPN01] John Lyons, Randal R. Piette, and Kent W. Niermeyer. Development, validation, and application of a fish-based index of biotic integrity for Wisconsin’s large warmwater rivers. Transactions of the American Fisheries Society, 130(6):1077–1094, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lantry:2008:CBC

[LRF^{+}08] Brian F. Lantry, Lars G. Rudstam, John L. Forney, Anthony J. VanDeValk, Edward L. Mills, Donald J. Stewart, and Jean V. Adams. Comparisons between consumption estimates from bioenergetics simulations and field measurements for walleyes from Oneida Lake, New York. Transactions of the American Fisheries Society, 137(5):1406–1421, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lacroix:2004:ARF

[LS04] Gilles L. Lacroix and Michael J. W. Stokesbury. Adult return of farmed Atlantic salmon escaped as juveniles into freshwater. Transactions of the American Fisheries Society, 133(2):484–490, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LaPlante:2007:AFT

[LS07] Lori Hosaka LaPlante and Eric T. Schultz. Annual fecundity of tautog in Long Island Sound: Size effects and long-term changes in a harvested population. Transactions of the American Fisheries Society, 136(6):1520–1533, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lankford:2001:LTT

[LT01a] Thomas E. Lankford, Jr. and Timothy E. Targett. Lowtemperature tolerance of age-0 Atlantic croakers: Recruitment implications for U.S. Mid-Atlantic Estuaries. Transactions of the American Fisheries Society, 130(2):236–249, March 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lankford:2001:PPY

[LT01b] Thomas E. Lankford, Jr. and Timothy E. Targett. Physiological performance of young-of-the-year Atlantic croakers from different Atlantic Coast Estuaries: Implications for stock structure. Transactions of the American Fisheries Society, 130(3):367–375, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LeClair:2006:AMD

[LYS06] Larry L. LeClair, Sewall F. Young, and James B. Shaklee. Allozyme and microsatellite DNA analyses of lingcod from Puget Sound, Washington, and adjoining waters. Transactions of the American Fisheries Society, 135(6):1631–1643, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGrath:2009:SFH

[MA09] Patrick McGrath and Herbert A. Austin. Site fidelity, home range, and tidal movements of white perch during the summer in two small tributaries of the York River, Virginia. Transactions of the American Fisheries Society, 138(5):966–974, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matern:2001:UTS

[Mat01] Scott A. Matern. Using temperature and salinity tolerances to predict the success of the shimofuri goby, a recent invader into California. Transactions of the American Fisheries Society, 130 (4):592–599, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Malick:2009:VEB

[MAW09] Michael J. Malick, Milo D. Adkison, and Alex C. Wertheimer. Variable effects of biological and environmental processes on Coho salmon marine survival in Southeast Alaska. Transactions of the American Fisheries Society, 138(4):846–860, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:2001:PIW

[MB01a] Douglas J. Martin and Lee E. Benda. Patterns of instream wood recruitment and transport at the watershed scale. Transactions of the American Fisheries Society, 130(5):940–958, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:2001:DSM

[MB01b] Seth A. Moore and Charles R. Bronte. Delineation of sympatric morphotypes of lake trout in Lake Superior. Transactions of the American Fisheries Society, 130(6):1233–1240, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moran:2002:PCR

[MB02] Paul Moran and Jon Baker. Polymerase chain reaction inhibition in historical fish scale mounting cards. Transactions of the American Fisheries Society, 131(1):109–119, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McHugh:2004:PSH

[MB04] Peter McHugh and Phaedra Budy. Patterns of spawning habitat selection and suitability for two populations of spring Chinook salmon, with an evaluation of generic versus site-specific suitability criteria. Transactions of the American Fisheries Society, 133(1):89–97, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:2005:AGS

[MB05a] Paul H. Michaletz and Joseph L. Bonneau. Age-0 gizzard shad abundance is reduced in the presence of macrophytes: Implications for interactions with bluegills. Transactions of the American Fisheries Society, 134(1):149–159, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mitchell:2005:TTR

[MB05b] Andrew J. Mitchell and Thomas M. Brandt. Temperature tolerance of red-rim melania Melanoides tuberculatus, an exotic aquatic snail established in the United States. Transactions of the American Fisheries Society, 134(1):126–131, January 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McHugh:2006:EEN

[MB06] Peter McHugh and Phaedra Budy. Experimental effects of nonnative brown trout on the individual- and population-level performance of native Bonneville cutthroat trout. Transactions of the American Fisheries Society, 135(6):1441–1455, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2008:DDD

[MB08a] Charles P. Madenjian and David B. Bunnell. Depth distribution dynamics of the sculpin community in Lake Michigan. Transactions of the American Fisheries Society, 137(5):1346–1357, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matzen:2008:ASS

[MB08b] Deanna Akre Matzen and Hans B. Berge. Assessing small-stream biotic integrity using fish assemblages across an urban landscape in the Puget Sound lowlands of Western Washington. Transactions of the American Fisheries Society, 137(3):677–689, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marshall:2000:GCN

[MBC00] Anne R. Marshall, H. Lee Blankenship, and William P. Connor. Genetic characterization of naturally spawned Snake River fallrun Chinook salmon. Transactions of the American Fisheries Society, 129(3):680–698, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moss:2005:ESS

[MBC^{+}05] Jamal H. Moss, David A. Beauchamp, Alison D. Cross, Katherine W. Myers, Edward V. Farley, Jr., James M. Murphy, and John H. Helle. Evidence for size-selective mortality after the first summer of ocean growth by pink salmon. Transactions of the American Fisheries Society, 134(5):1313–1322, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Modde:2005:STU

[MBK05] Timothy Modde, Zachary H. Bowen, and David C. Kitcheyan. Spatial and temporal use of a spawning site in the Middle Green River by wild and hatchery-reared razorback suckers. Transactions of the American Fisheries Society, 134(4):937–944, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McIntyre:2006:OTI

[MBMO06] J. K. McIntyre, D. A. Beauchamp, M. M. Mazur, and N. C. Overman. Ontogenetic trophic interactions and benthopelagic coupling in Lake Washington: Evidence from stable isotopes and diet analysis. Transactions of the American Fisheries Society, 135(5): 1312–1328, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCollum:2003:CNW

[MBS03a] Arthur B. McCollum, David B. Bunnell, and Roy A. Stein. Cold, northern winters: The importance of temperature to overwinter mortality of age-0 white crappies. Transactions of the American Fisheries Society, 132(5):977–987, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:2003:SPP

[MBS03b] Matthew G. Mesa, Jennifer M. Bayer, and James G. Seelye. Swimming performance and physiological responses to exhaustive exercise in radio-tagged and untagged Pacific lampreys. Transactions of the American Fisheries Society, 132(3):483–492, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McHugh:2004:MBA

[MBS04] Peter McHugh, Phaedra Budy, and Howard Schaller. A modelbased assessment of the potential response of Snake River spring– summer Chinook salmon to habitat improvements. Transactions of the American Fisheries Society, 133(3):622–638, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meeuwig:2005:ETS

[MBS05] Michael H. Meeuwig, Jennifer M. Bayer, and James G. Seelye. Effects of temperature on survival and development of early life stage Pacific and western brook lampreys. Transactions of the American Fisheries Society, 134(1):19–27, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacInnis:2000:AGR

[MC00a] Andrew J. MacInnis and Lynda D. Corkum. Age and growth of round goby Neogobius melanostomus in the Upper Detroit River. Transactions of the American Fisheries Society, 129(3):852–858, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacInnis:2000:FRS

[MC00b] Andrew J. MacInnis and Lynda D. Corkum. Fecundity and reproductive season of the round goby Neogobius melanostomus in the Upper Detroit River. Transactions of the American Fisheries Society, 129(1):136–144, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayfield:2004:TEG

[MC04] Ryan B. Mayfield and Joseph J. Cech, Jr. Temperature effects on green sturgeon bioenergetics. Transactions of the American Fisheries Society, 133(4):961–970, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:2009:PMF

[MC09] Michael R. Meador and Daren M. Carlisle. Predictive models for fish assemblages in Eastern U.S. streams: Implications for assessing biodiversity. Transactions of the American Fisheries Society, 138(4):725–740, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDowall:2003:IIS

[McD03] R. M. McDowall. Impacts of introduced salmonids on native galaxiids in New Zealand upland streams: a new look at an old problem. Transactions of the American Fisheries Society, 132(2): 229–238, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKenna:2005:ANN

[McK05] James E. McKenna, Jr. Application of neural networks to prediction of fish diversity and salmonid production in the Lake Ontario Basin. Transactions of the American Fisheries Society, 134(1): 28–43, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:2005:RBJ

[MCR05] Lesa Meng, Giancarlo Cicchetti, and Steve Raciti. Relationships between juvenile winter flounder and multiple-scale habitat variation in Narragansett Bay, Rhode Island. Transactions of the American Fisheries Society, 134(6):1509–1519, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:2006:EWS

[MCS06] Margaret H. Murphy, Michael J. Connerton, and Donald J. Stewart. Evaluation of winter severity on growth of young-of-the-year Atlantic salmon. Transactions of the American Fisheries Society, 135(2):420–430, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacLean:2008:ERG

[MCS08] Sharon A. MacLean, Elaine M. Caldarone, and Jeanne M. St. Onge-Burns. Estimating recent growth rates of Atlantic salmon smolts using RNA–DNA ratios from nonlethally sampled tissues. Transactions of the American Fisheries Society, 137(5): 1279–1284, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2003:TPT

[MD03] L. E. Miranda and C. R. Dolan. Test of a power transfer model for standardized electrofishing. Transactions of the American Fisheries Society, 132(6):1179–1185, November 2003. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McRae:2005:FID

[MD05] Brian J. McRae and James S. Diana. Factors influencing density of age-0 brown trout and brook trout in the Au Sable River, Michigan. Transactions of the American Fisheries Society, 134 (1):132–140, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markle:2007:EHV

[MD07] Douglas F. Markle and Larry K. Dunsmoor. Effects of habitat volume and fathead minnow introduction on larval survival of two endangered sucker species in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 136(3):567–579, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarthy:2009:LHQ

[MDE^{+}09] Sarah G. McCarthy, Jeffrey J. Duda, John M. Emlen, Garth R. Hodgson, and David A. Beauchamp. Linking habitat quality with trophic performance of steelhead along forest gradients in the South Fork Trinity River watershed, California. Transactions of the American Fisheries Society, 138(3):506–521, May 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Milston:2006:CPS

[MDP^{+}06] Ruth H. Milston, Michael W. Davis, Steven J. Parker, Bori L. Olla, Shaun Clements, and Carl B. Schreck. Characterization of the physiological stress response in lingcod. Transactions of the American Fisheries Society, 135(5):1165–1174, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2007:UCS

[ME07] Rafael Miranda and M. Carmen Escala. On the use of cyprinid scales in the diet analysis of piscivorous species: How much information is hidden in a fish scale? Transactions of the American Fisheries Society, 136(4):1008–1017, July 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:2005:SPV

[Mea05] Michael R. Meador. Single-pass versus two-pass boat electrofishing for characterizing river fish assemblages: Species richness estimates and sampling distance. Transactions of the American Fisheries Society, 134(1):59–67, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moorman:2009:IBC

[MEA^{+}09] Michelle C. Moorman, David B. Eggleston, Christopher B. Anderson, Andres Mansilla, and Paul Szejner. Implications of beaver Castor canadensis and trout introductions on native fish in the Cape Horn Biosphere Reserve, Chile. Transactions of the American Fisheries Society, 138(2):306–313, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Myrick:2004:ACA

[MFC04] Christopher A. Myrick, D. Kurtis Folgner, and Joseph J. Cech, Jr. An annular chamber for aquatic animal preference studies. Transactions of the American Fisheries Society, 133(2):427–433, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moss:2009:SDE

[MFFI09] Jamal H. Moss, Edward V. Farley, Jr., Angela M. Feldmann, and James N. Ianelli. Spatial distribution, energetic status, and food habits of Eastern Bering Sea age-0 walleye pollock. Transactions of the American Fisheries Society, 138(3):497–505, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyes:2006:PPS

[MFMB06] Christopher D. Moyes, Nuno Fragoso, Michael K. Musyl, and Richard W. Brill. Predicting postrelease survival in large pelagic fish. Transactions of the American Fisheries Society, 135(5):1389– 1397, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matala:2009:EDA

[MFOA09] A. P. Matala, R. French, E. Olsen, and W. R. Ardren. Ecotype distinctions among steelhead in Hood River, Oregon, allow real-time genetic assignment of conservation broodstocks. Transactions of the American Fisheries Society, 138(6):1490– 1509, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minakawa:2003:ECS

[MG03] Noboru Minakawa and Robert I. Gara. Effects of chum salmon redd excavation on benthic communities in a stream in the Pacific Northwest. Transactions of the American Fisheries Society, 132(3):598–604, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maceina:2006:RLB

[MG06] Michael J. Maceina and John M. Grizzle. The relation of largemouth bass virus to largemouth bass population metrics in five Alabama reservoirs. Transactions of the American Fisheries Society, 135(3):545–555, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:2006:MER

[MGG^{+}06] William J. Matthews, Keith B. Gido, Gary P. Garrett, Frances P. Gelwick, Jeffrey G. Stewart, and Jacob Schaefer. Modular experimental riffle–pool stream system. Transactions of the American Fisheries Society, 135(6):1559–1566, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martinez:2001:GLH

[MGJ^{+}01] J. L. Martinez, S. Gephard, F. Juanes, J. Perez, and E. Garcia-Vazquez. Genetic and life history differentiation between donor and derivative populations of Atlantic salmon. Transactions of the American Fisheries Society, 130(3):508–515, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McQuown:2002:CIS

[MGM02] Eve McQuown, Graham A. E. Gall, and Bernie May. Characterization and inheritance of six microsatellite loci in lake sturgeon. Transactions of the American Fisheries Society, 131(2):299–307, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:2001:DDO

[MGMM01] William J. Matthews, Keith B. Gido, and Edie Marsh-Matthews. Density-dependent overwinter survival and growth of red shiners from a southwestern river. Transactions of the American Fisheries Society, 130(3):478–488, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mackenzie-Grieve:2006:THU

[MGP06] Jody L. Mackenzie-Grieve and John R. Post. Thermal habitat use by lake trout in two contrasting Yukon Territory lakes. Transactions of the American Fisheries Society, 135(3):727–738, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2002:STL

[MH02] Joseph S. Meyer and James A. Hansen. Subchronic toxicity of low dissolved oxygen concentrations, elevated pH, and elevated ammonia concentrations to Lost River suckers. Transactions of the American Fisheries Society, 131(4):656–666, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2005:TES

[MH05] Carl G. Meyer and Randy R. Honebrink. Transintestinal expulsion of surgically implanted dummy transmitters by bluefin trevally — implications for long-term movement studies. Transactions of the American Fisheries Society, 134(3):602–606, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2003:GCA

[MHD03] Charles P. Madenjian, Jeffrey D. Holuszko, and Timothy J. Desorcie. Growth and condition of alewives in Lake Michigan, 1984– 2001. Transactions of the American Fisheries Society, 132(6): 1104–1116, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manire:2001:SCA

[MHHS01] C. Manire, R. Hueter, E. Hull, and R. Spieler. Serological changes associated with gill-net capture and restraint in three species of sharks. Transactions of the American Fisheries Society, 130(6): 1038–1048, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:2001:DIB

[MHK^{+}01] Frank H. McCormick, Robert M. Hughes, Philip R. Kaufmann, David V. Peck, John L. Stoddard, and Alan T. Herlihy. Development of an index of biotic integrity for the Mid-Atlantic highlands region. Transactions of the American Fisheries Society, 130(5): 857–877, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2008:LGA

[MHM08] L. E. Miranda, M. D. Habrat, and S. Miyazono. Longitudinal gradients along a reservoir cascade. Transactions of the American Fisheries Society, 137(6):1851–1865, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mellina:2005:SMP

[MHMP05] Eric Mellina, Scott G. Hinch, Kirsten D. MacKenzie, and Greg Pearson. Seasonal movement patterns of stream-dwelling rainbow trout in North-Central British Columbia, Canada. Transactions of the American Fisheries Society, 134(4):1021–1037, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2005:RVA

[MHR^{+}05] Charles P. Madenjian, Tomas O. Höök, Edward S. Rutherford, Doran M. Mason, Thomas E. Croley II, Emily B. Szalai, and James R. Bence. Recruitment variability of alewives in Lake Michigan. Transactions of the American Fisheries Society, 134 (1):218–230, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:2006:DCC

[Mic06] Paul H. Michaletz. Do channel catfish stockings affect growth and size structure of bluegills in small impoundments? Transactions of the American Fisheries Society, 135(5):1254–1265, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:2003:MDL

[Mil03] Loren M. Miller. Microsatellite DNA loci reveal genetic structure of yellow perch in Lake Michigan. Transactions of the American Fisheries Society, 132(3):503–513, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mills:2004:RSL

[Mil04] Edward Mills. Regime shifts in lake ecosystems: Pattern and variation. Transactions of the American Fisheries Society, 133 (6):1540, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2005:CRR

[Mir05a] L. E. Miranda. Catch rates relative to angler party size with implications for monitoring angler success. Transactions of the American Fisheries Society, 134(4):1005–1010, July 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2005:FAO

[Mir05b] L. E. Miranda. Fish assemblages in Oxbow Lakes relative to connectivity with the Mississippi River. Transactions of the American Fisheries Society, 134(6):1480–1489, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2007:ASS

[Mir07] L. E. Miranda. Approximate sample sizes required to estimate length distributions. Transactions of the American Fisheries Society, 136(2):409–415, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDermid:2007:GEI

[MISS07] Jenni L. McDermid, Peter E. Ihssen, William N. Sloan, and Brian J. Shuter. Genetic and environmental influences on life history traits in lake trout. Transactions of the American Fisheries Society, 136(4):1018–1029, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Monzyk:2009:MCH

[MJH^{+}09] Fred R. Monzyk, Brian C. Jonasson, Timothy L. Hoffnagle, Patrick J. Keniry, Richard W. Carmichael, and Peter J. Cleary.

Migration characteristics of hatchery and natural spring Chinook salmon smolts from the Grande Ronde River Basin, Oregon, to Lower Granite Dam on the Snake River. Transactions of the American Fisheries Society, 138(5):1093–1108, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meise:2003:GRJ

[MJS^{+}03] C. J. Meise, D. L. Johnson, L. L. Stehlik, J. Manderson, and P. Shaheen. Growth rates of juvenile winter flounder under varying environmental conditions. Transactions of the American Fisheries Society, 132(2):335–345, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Myers:2009:RPE

[MJSY09] Jared T. Myers, Michael L. Jones, Jason D. Stockwell, and Daniel L. Yule. Reassessment of the predatory effects of rainbow smelt on ciscoes in Lake Superior. Transactions of the American Fisheries Society, 138(6):1352–1368, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mogen:2005:ICM

[MK05] Jim T. Mogen and Lynn R. Kaeding. Identification and characterization of migratory and nonmigratory bull trout populations in the St. Mary River Drainage, Montana. Transactions of the American Fisheries Society, 134(4):841–852, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2008:BER

[MK08] L. E. Miranda and M. Kratochvíl. Boat electrofishing relative to anode arrangement. Transactions of the American Fisheries Society, 137(5):1358–1362, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2000:RRW

[MKBF00] Charles P. Madenjian, Roger L. Knight, Michael T. Bur, and John L. Forney. Reduction in recruitment of white bass in Lake Erie after invasion of white perch. Transactions of the American Fisheries Society, 129(6):1340–1353, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meka:2003:VMP

[MKDB03] Julie M. Meka, E. Eric Knudsen, David C. Douglas, and Robert B. Benter. Variable migratory patterns of different adult rainbow trout life history types in a Southwest Alaska watershed. Transactions of the American Fisheries Society, 132(4):717–732, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMillan:2007:OES

[MKP07] John R. McMillan, Stephen L. Katz, and George R. Pess. Observational evidence of spatial and temporal structure in a sympatric anadromous (winter steelhead) and resident rainbow trout mating system on the Olympic Peninsula, Washington. Transactions of the American Fisheries Society, 136(3):736–748, May 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:2001:SSN

[MKS01] Loren M. Miller, Larry Kallemeyn, and Wansuk Senanan. Spawning-site and natal-site fidelity by northern pike in a large lake: Mark–recapture and genetic evidence. Transactions of the American Fisheries Society, 130(2):307–316, March 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:2004:DFA

[ML04] L. E. Miranda and G. M. Lucas. Determinism in fish assemblages of floodplain lakes of the vastly disturbed Mississippi alluvial valley. Transactions of the American Fisheries Society, 133(2):358– 370, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGrath:2007:CPM

[ML07] C. C. McGrath and W. M. Lewis, Jr. Competition and predation as mechanisms for displacement of greenback cutthroat trout by brook trout. Transactions of the American Fisheries Society, 136 (5):1381–1392, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNabb:2003:PJC

[MLB03] C. D. McNabb, C. R. Liston, and S. M. Borthwick. Passage of juvenile Chinook salmon and other fish species through Archimedes lifts and a hidrostal pump at Red Bluff, California. Transactions of the American Fisheries Society, 132(2):326–334, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Machut:2007:AIA

[MLSD07] Leonard S. Machut, Karin E. Limburg, Robert E. Schmidt, and Dawn Dittman. Anthropogenic impacts on American eel demographics in Hudson River tributaries, New York. Transactions of the American Fisheries Society, 136(6):1699–1713, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:2001:NIL

[MM01] Lesa Meng and Scott A. Matern. Native and introduced larval fishes of Suisun Marsh, California: The effects of freshwater flow. Transactions of the American Fisheries Society, 130(5):750–765, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maret:2002:FAE

[MM02] T. R. Maret and D. E. MacCoy. Fish assemblages and environmental variables associated with hard-rock mining in the Coeur d’Alene River Basin, Idaho. Transactions of the American Fisheries Society, 131(5):865–884, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:2003:EEG

[MM03] Michael R. Meador and Julie P. McIntyre. Effects of electrofishing gear type on spatial and temporal variability in fish community sampling. Transactions of the American Fisheries Society, 132(4): 709–716, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mock:2005:PMD

[MM05] K. E. Mock and M. P. Miller. Patterns of molecular diversity in naturally occurring and refugial populations of the least chub. Transactions of the American Fisheries Society, 134(1):267–278, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mucha:2008:HUM

[MM08] Jamie M. Mucha and Robert W. Mackereth. Habitat use and movement patterns of brook trout in Nipigon Bay, Lake Superior. Transactions of the American Fisheries Society, 137(4):

1203–1212, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muhlfeld:2009:LHW

[MMBG09] Clint C. Muhlfeld, Thomas E. McMahon, Matthew C. Boyer, and Robert E. Gresswell. Local habitat, watershed, and biotic factors influencing the spread of hybridization between native westslope cutthroat trout and introduced rainbow trout. Transactions of the American Fisheries Society, 138(5):1036–1051, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minello:2008:PPE

[MMCR08] Thomas J. Minello, Geoffrey A. Matthews, Philip A. Caldwell, and Lawrence P. Rozas. Population and production estimates for decapod crustaceans in wetlands of Galveston Bay, Texas. Transactions of the American Fisheries Society, 137(1):129–146, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Modde:2001:FWS

[MMH01] Timothy Modde, Robert T. Muth, and G. Bruce Haines. Floodplain wetland suitability, access, and potential use by juvenile razorback suckers in the Middle Green River, Utah. Transactions of the American Fisheries Society, 130(6):1095–1105, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mebane:2003:IBI

[MMH03] Christopher A. Mebane, Terry R. Maret, and Robert M. Hughes. An Index of Biological Integrity (IBI) for Pacific Northwest rivers. Transactions of the American Fisheries Society, 132(2):239–261, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:2007:ERS

[MMM07] William J. Matthews and Edie Marsh-Matthews. Extirpation of red shiner in direct tributaries of Lake Texoma (Oklahoma– Texas): a cautionary case history from a fragmented river– reservoir system. Transactions of the American Fisheries Society, 136(4):1041–1062, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matern:2002:NAF

[MMP02] Scott A. Matern, Peter B. Moyle, and Leslie C. Pierce. Native and alien fishes in a California estuarine Marsh: Twenty-one years of changing assemblages. Transactions of the American Fisheries Society, 131(5):797–816, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:2003:AES

[MMP03] Michael R. Meador, Julie P. McIntyre, and Kenneth H. Pollock. Assessing the efficacy of single-pass backpack electrofishing to characterize fish community structure. Transactions of the American Fisheries Society, 132(1):39–46, January 2003. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:2000:IIR

[MMS00] Matthew G. Mesa, Alec G. Maule, and Carl B. Schreck. Interaction of infection with Renibacterium salmoninarum and physical stress in juvenile Chinook salmon: Physiological responses, disease progression, and mortality. Transactions of the American Fisheries Society, 129(1):158–173, January 2000. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muir:2006:PHS

[MMS^{+}06] William D. Muir, Douglas M. Marsh, Benjamin P. Sandford, Steven G. Smith, and John G. Williams. Post-hydropower system delayed mortality of transported Snake River stream-type Chinook salmon: Unraveling the mystery. Transactions of the American Fisheries Society, 135(6):1523–1534, November 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muhlfeld:2005:GSS

[MMTF05] Clint C. Muhlfeld, Brian Marotz, Simon R. Thorrold, and Jennifer L. FitzGerald. Geochemical signatures in scales record stream of origin in westslope cutthroat trout. Transactions of the American Fisheries Society, 134(4):945–959, July 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:2009:GLS

[MMY09] Loren M. Miller, Steven W. Mero, and Jerry A. Younk. The genetic legacy of stocking muskellunge in a Northern Minnesota Lake. Transactions of the American Fisheries Society, 138(3):

602–615, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McClelland:2007:CGR

[MN07] Erin K. McClelland and Kerry A. Naish. Comparisons of F_{ST} and Q_{ST} of growth-related traits in two populations of Coho salmon. Transactions of the American Fisheries Society, 136(5): 1276–1284, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:2002:UFI

[MOP02] Lesa Meng, Christopher D. Orphanides, and J. Christopher Powell. Use of a fish index to assess habitat quality in Narragansett Bay, Rhode Island. Transactions of the American Fisheries Society, 131(4):731–742, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:2006:ELW

[MOP^{+}06] Charles P. Madenjian, Daniel V. O’Connor, Steven A. Pothoven, Philip J. Schneeberger, Richard R. Rediske, James P. O’Keefe, Roger A. Bergstedt, Ray L. Argyle, and Stephen B. Brandt. Evaluation of a lake whitefish bioenergetics model. Transactions of the American Fisheries Society, 135(1):61–75, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moser:2002:PEA

[MOSB02] Mary L. Moser, Paul A. Ocker, Lowell C. Stuehrenberg, and Theodore C. Bjornn. Passage efficiency of adult Pacific lampreys at hydropower dams on the Lower Columbia River, USA. Transactions of the American Fisheries Society, 131(5):956–965, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McPherson:2004:GDT

[MOT04] Arran A. McPherson, Patrick T. O’Reilly, and Christopher T. Taggart. Genetic differentiation, temporal stability, and the absence of isolation by distance among Atlantic herring populations. Transactions of the American Fisheries Society, 133(2):434–446, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:2003:DRS

[MPK03] Paul C. Marsh, Carol A. Pacey, and Brian R. Kesner. Decline of the razorback sucker in Lake Mohave, Colorado River, Arizona and Nevada. Transactions of the American Fisheries Society, 132 (6):1251–1256, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mueter:2005:RBC

[MPP05] Franz J. Mueter, Brian J. Pyper, and Randall M. Peterman. Relationships between coastal ocean conditions and survival rates of Northeast Pacific salmon at multiple lags. Transactions of the American Fisheries Society, 134(1):105–119, January 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maki-Petays:2000:MUA

[MPVM00] Aki Mäki-Petäys, Teppo Vehanen, and Timo Muotka. Microhabitat use by age-0 brown trout and grayling: Seasonal responses to streambed restoration under different flows. Transactions of the American Fisheries Society, 129(3):771–781, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Myers:2000:UBH

[MR00] Marilyn J. Myers and Vincent H. Resh. Undercut banks: a habitat for more than just trout. Transactions of the American Fisheries Society, 129(2):594–597, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Markle:2009:LFT

[MRC^{+}09] Douglas F. Markle, Susan A. Reithel, John Crandall, Tammy Wood, Torrey J. Tyler, Mark Terwilliger, and David C. Simon. Larval fish transport and retention and the importance of location for juvenile fish recruitment in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 138(2):328–347, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:2003:RTS

[MS03] Bruce A. Miller and Steve Sadro. Residence time and seasonal movements of juvenile Coho salmon in the ecotone and lower estuary of Winchester Creek, South Slough, Oregon. Transactions of the American Fisheries Society, 132(3):546–559, May 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morrison:2004:AYP

[MS04] W. E. Morrison and D. H. Secor. Abundance of yellow-phase American eels in the Hudson River Estuary. Transactions of the American Fisheries Society, 133(4):896–910, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macdonald:2008:MIE

[MSC08] Jed I. Macdonald, J. Michael G. Shelley, and David A. Crook. A method for improving the estimation of natal chemical signatures in otoliths. Transactions of the American Fisheries Society, 137 (6):1674–1682, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCaffery:2007:ERD

[MSE07] Magnus McCaffery, T. Adam Switalski, and Lisa Eby. Effects of road decommissioning on stream habitat characteristics in the South Fork Flathead River, Montana. Transactions of the American Fisheries Society, 136(3):553–561, May 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2003:RDF

[MSEL03] Kevin A. Meyer, Daniel J. Schill, F. Steven Elle, and James A. Lamansky, Jr. Reproductive demographics and factors that influence length at sexual maturity of Yellowstone cutthroat trout in Idaho. Transactions of the American Fisheries Society, 132(2): 183–195, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morrissey:2005:IPC

[MSET05] Michael B. Morrissey, Cory D. Suski, Kevin R. Esseltine, and Bruce L. Tufts. Incidence and physiological consequences of decompression in smallmouth bass after live-release angling tournaments. Transactions of the American Fisheries Society, 134(4): 1038–1047, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2006:SYC

[MSL^{+}06] Kevin A. Meyer, Daniel J. Schill, James A. Lamansky, Jr., Matthew R. Campbell, and Christine C. Kozfkay. Status of Yellowstone cutthroat trout in Idaho. Transactions of the American Fisheries Society, 135(5):1329–1347, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mueller:2008:JCS

[MSM^{+}08] Robert P. Mueller, Susan S. Southard, Christopher W. May, Walter H. Pearson, and Valerie I. Cullinan. Juvenile Coho salmon leaping ability and behavior in an experimental culvert test bed. Transactions of the American Fisheries Society, 137(4):941–950, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McQuown:2000:MAG

[MSS^{+}00] Eve C. McQuown, Brian L. Sloss, Robert J. Sheehan, Jeff Rodzen, Gregory J. Tranah, and Bernie May. Microsatellite analysis of genetic variation in sturgeon: New primer sequences for Scaphirhynchus and Acipenser. Transactions of the American Fisheries Society, 129(6):1380–1388, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muhlfeld:2006:OES

[MTSS06] Clint C. Muhlfeld, Mark L. Taper, David F. Staples, and Bradley B. Shepard. Observer error structure in bull trout redd counts in Montana streams: Implications for inference on true redd numbers. Transactions of the American Fisheries Society, 135(3):643–654, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McPhee:2009:GDP

[MTWS09] Megan V. McPhee, Tyler H. Tappenbeck, Diane C. Whited, and Jack A. Stanford. Genetic diversity and population structure in the Kuskokwim River drainage support the recurrent evolution hypothesis for sockeye salmon life histories. Transactions of the American Fisheries Society, 138(6):1481–1489, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meretsky:2000:SVL

[MVD^{+}00] Vicky J. Meretsky, Richard A. Valdez, Michael E. Douglas, Mark J. Brouder, Owen T. Gorman, and Paul C. Marsh. Spatiotemporal variation in length–weight relationships of endangered humpback chub: Implications for conservation and management. Transactions of the American Fisheries Society, 129(2): 419–428, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morbey:2007:DPL

[MVS07] Yolanda E. Morbey, Kris Vascotto, and Brian J. Shuter. Dynamics of piscivory by lake trout following a smallmouth bass invasion: a historical reconstruction. Transactions of the American Fisheries Society, 136(2):477–483, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKelvey:2006:DCF

[MW06] Denise R. McKelvey and Christopher D. Wilson. Discriminant classification of fish and zooplankton backscattering at 38 and 120 kHz. Transactions of the American Fisheries Society, 135(2): 488–499, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:2008:EIB

[MWG^{+}08] Michael R. Meador, Thomas R. Whittier, Robert M. Goldstein, Robert M. Hughes, and David V. Peck. Evaluation of an index of biotic integrity approach used to assess biological condition in Western U.S. streams and rivers at varying spatial scales. Transactions of the American Fisheries Society, 137(1):13–22, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:2000:PSG

[MWM00] Matthew G. Mesa, Lisa K. Weiland, and Alec G. Maule. Progression and severity of gas bubble trauma in juvenile salmonids. Transactions of the American Fisheries Society, 129(1):174–185, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mortensen:2002:SSA

[MWMT02] Donald G. Mortensen, Alex C. Wertheimer, Jacek M. Maselko, and S. G. Taylor. Survival and straying of Auke Creek, Alaska, pink salmon marked with coded wire tags and thermally induced otolith marks. Transactions of the American Fisheries Society, 131(1):14–26, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McAdam:2005:LBW

[MWN05] Steven O. McAdam, Carl J. Walters, and Craig Nistor. Linkages between white sturgeon recruitment and altered bed substrates in the Nechako River, Canada. Transactions of the American Fisheries Society, 134(6):1448–1456, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGree:2008:LMI

[MWS08] Michelle McGree, Timothy A. Whitesel, and Jen Stone. Larval metamorphosis of individual Pacific lampreys reared in captivity. Transactions of the American Fisheries Society, 137(6):1866– 1878, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mann:2009:HSW

[MWSP09] David A. Mann, Christopher D. Wilson, Jiakun Song, and Arthur N. Popper. Hearing sensitivity of the walleye pollock. Transactions of the American Fisheries Society, 138(5):1000– 1008, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mitro:2002:SSM

[MZ02] Matthew G. Mitro and Alexander V. Zale. Seasonal survival, movement, and habitat use of age-0 rainbow trout in the Henry’s Fork of the Snake River, Idaho. Transactions of the American Fisheries Society, 131(2):271–286, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMahon:2007:TCB

[MZB^{+}07] Thomas E. McMahon, Alexander V. Zale, Frederic T. Barrows, Jason H. Selong, and Robert J. Danehy. Temperature and competition between bull trout and brook trout: a test of the elevation refuge hypothesis. Transactions of the American Fisheries Society, 136(5):1313–1326, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nicola:2004:GPS

[NA04] Graciela G. Nicola and Ana Almodóvar. Growth pattern of stream-dwelling brown trout under contrasting thermal conditions. Transactions of the American Fisheries Society, 133(1): 66–78, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ng:2007:HUS

[NAG07] Clare L. Ng, Kenneth W. Able, and Thomas M. Grothues. Habitat use, site fidelity, and movement of adult striped bass in a Southern New Jersey Estuary based on mobile acoustic telemetry. Transactions of the American Fisheries Society, 136(5):1344– 1355, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Natsumeda:2007:MPJ

[Nat07] Takaharu Natsumeda. Movement patterns of Japanese fluvial sculpin Cottus pollux in a headwater stream. Transactions of the American Fisheries Society, 136(6):1769–1777, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nitschke:2001:AGV

[NBK01] Paul Nitschke, Jay Burnett, and Brian C. Kelly. Age and growth verification for cunner in Western Cape Cod Bay, Massachusetts, using tag–recapture data. Transactions of the American Fisheries Society, 130(6):1150–1163, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narum:2006:SSG

[NBMP06] Shawn R. Narum, Steve Boe, Paul Moran, and Matt Powell. Small-scale genetic structure and variation in steelhead of the Grande Ronde River, Oregon, USA. Transactions of the American Fisheries Society, 135(4):979–986, July 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Newby:2007:PIT

[NBS07] Nathalie C. Newby, Thomas R. Binder, and E. Don Stevens. Passive integrated transponder (PIT) tagging did not negatively affect the short-term feeding behavior or swimming performance of juvenile rainbow trout. Transactions of the American Fisheries Society, 136(2):341–345, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:2005:SSP

[NC05] Jay A. Nelson and Guy Claireaux. Sprint swimming performance of juvenile European Sea bass. Transactions of the American Fisheries Society, 134(5):1274–1284, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neitzel:2004:SEJ

[NDČ^{+}04] D. A. Neitzel, D. D. Dauble, G. F. Čada, M. C. Richmond, G. R. Guensch, R. P. Mueller, C. S. Abernethy, and Brett Amidan.

Survival estimates for juvenile fish subjected to a laboratorygenerated shear environment. Transactions of the American Fisheries Society, 133(2):447–454, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neville:2009:IWH

[NDR^{+}09] Helen Neville, Jason Dunham, Amanda Rosenberger, John Umek, and Brooke Nelson. Influences of wildfire, habitat size, and connectivity on trout in headwater streams revealed by patterns of genetic diversity. Transactions of the American Fisheries Society, 138(6):1314–1327, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neves:2004:BMB

[Nev04] Richard J. Neves. Bivalve molluscs: Biology, ecology, and culture. Transactions of the American Fisheries Society, 133(5): 1290, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Newcomb:2001:BR

[New01] Tammy J. Newcomb. Book review. Transactions of the American Fisheries Society, 130(1):178–179, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neff:2000:MMF

[NFG00] Bryan D. Neff, Peng Fu, and Mart R. Gross. Microsatellite multiplexing in fish. Transactions of the American Fisheries Society, 129(2):584–593, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nislow:2000:SEB

[NFP00] K. H. Nislow, C. L. Folt, and D. L. Parrish. Spatially explicit bioenergetic analysis of habitat quality for age-0 Atlantic salmon. Transactions of the American Fisheries Society, 129(5): 1067–1081, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nestler:2002:SMP

[NGC^{+}02] John M. Nestler, R. Andrew Goodwin, Thomas M. Cole, Donald Degan, and Donald Dennerline. Simulating movement patterns of blueback herring in a stratified southern impoundment. Transactions of the American Fisheries Society, 131(1):55–69, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nico:2001:BR

[NH01] Leo G. Nico and Jeffery J. Herod. Book review. Transactions of the American Fisheries Society, 130(2):347–348, March 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norman:2009:AMM

[NHFF09] James R. Norman, Megan M. Hagler, Mary C. Freeman, and Byron J. Freeman. Application of a multistate model to estimate culvert effects on movement of small fishes. Transactions of the American Fisheries Society, 138(4):826–838, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nemeth:2003:RRT

[NKJ03] Matthew J. Nemeth, Charles C. Krueger, and Daniel C. Josephson. Rheotactic response of two strains of juvenile landlocked Atlantic salmon: Implications for population restoration. Transactions of the American Fisheries Society, 132(5):904–912, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nislow:2003:ILH

[NL03] Keith H. Nislow and Winsor H. Lowe. Influences of logging history and stream pH on brook trout abundance in first-order streams in New Hampshire. Transactions of the American Fisheries Society, 132(1):166–171, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neal:2006:BBA

[NN06] J. Wesley Neal and Richard L. Noble. A bioenergetics-based approach to explain largemouth bass size in tropical reservoirs. Transactions of the American Fisheries Society, 135(6):1535– 1545, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neal:2006:EEC

[NNL06] J. Wesley Neal, Richard L. Noble, and Craig G. Lilyestrom. Evaluation of the ecological compatibility of butterfly peacock cichlids and largemouth bass in Puerto Rico reservoirs. Transactions of the American Fisheries Society, 135(2):288–296, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narum:2004:DML

[NPT04] Shawn R. Narum, Madison S. Powell, and AndréJ. Talbot. A distinctive microsatellite locus that differentiates ocean-type from stream-type Chinook salmon in the interior Columbia River Basin. Transactions of the American Fisheries Society, 133(4): 1051–1055, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nowak:2002:DSP

[NQ02] Gretchen M. Nowak and Thomas P. Quinn. Diel and seasonal patterns of horizontal and vertical movements of telemetered cutthroat trout in Lake Washington, Washington. Transactions of the American Fisheries Society, 131(4):452–462, May 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nielsen:2002:PGS

[NS02] Jennifer L. Nielsen and George K. Sage. Population genetic structure in Lahontan cutthroat trout. Transactions of the American Fisheries Society, 131(4):376–388, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neufeld:2004:ESD

[NS04] Matthew D. Neufeld and Colin R. Spence. Evaluation of a simple decompression procedure to reduce decompression trauma in trap-caught burbot. Transactions of the American Fisheries Society, 133(5):1260–1263, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narum:2007:GVS

[NSC07] Shawn R. Narum, Jeffrey J. Stephenson, and Matthew R. Campbell. Genetic variation and structure of Chinook salmon life history types in the Snake River. Transactions of the American Fisheries Society, 136(5):1252–1262, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nislow:2004:MLH

[NSF04] K. H. Nislow, A. J. Sepulveda, and C. L. Folt. Mechanistic linkage of hydrologic regime to summer growth of age-0 Atlantic salmon. Transactions of the American Fisheries Society, 133(1): 79–88, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neilson:2003:AGC

[NSP03] John D. Neilson, Wayne T. Stobo, and Peter Perley. Age and growth of Canadian East Coast pollock: Comparison of results from otolith examination and mark–recapture studies. Transactions of the American Fisheries Society, 132(3):536–545, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narum:2008:LGS

[NSVT08] Shawn R. Narum, Terra L. Schultz, Donald M. Van Doornik, and David Teel. Localized genetic structure persists in wild populations of Chinook salmon in the John Day River despite gene flow from outside sources. Transactions of the American Fisheries Society, 137(6):1650–1656, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nilo:2006:FEJ

[NTB^{+}06] Pedro Nilo, Serge Tremblay, Aline Bolon, Julian Dodson, Pierre Dumont, and Réjean Fortin. Feeding ecology of juvenile lake sturgeon in the St. Lawrence River system. Transactions of the American Fisheries Society, 135(4):1044–1055, July 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nieland:2002:AGR

[NTW02] David L. Nieland, R. Glenn Thomas, and Charles A. Wilson. Age, growth, and reproduction of spotted seatrout in Barataria Bay, Louisiana. Transactions of the American Fisheries Society, 131 (2):245–259, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oyadomari:2007:IRT

[OA07] Jason K. Oyadomari and Nancy N. Auer. Influence of rearing temperature and feeding regime on otolith increment deposition in larval ciscoes. Transactions of the American Fisheries Society, 136(3):766–777, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:2008:PUR

[OADC^{+}08] Mark H. Olson, Tristan Adams-Deutsch, Kelly J. Cassels, Amy E. Oliver, and David L. Mitchell. Patterns of ultraviolet radiation exposure in bluegill nests over the course of the spawning season. Transactions of the American Fisheries Society, 137(5):1446– 1454, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:2000:MRP

[OBB^{+}00] Jeffrey B. Olsen, Paul Bentzen, Michael A. Banks, James B. Shaklee, and Sewall Young. Microsatellites reveal population identity of individual pink salmon to allow supportive breeding of a population at risk of extinction. Transactions of the American Fisheries Society, 129(1):232–242, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Overman:2009:DFF

[OBB^{+}09] Nathanael C. Overman, David A. Beauchamp, Hans B. Berge, Michael M. Mazur, and Jenifer K. McIntyre. Differing forage fish assemblages influence trophic structure in neighboring urban lakes. Transactions of the American Fisheries Society, 138(4): 741–755, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ostrand:2004:CVD

[OBW04] Kenneth G. Ostrand, Ben J. Braeutigam, and David H. Wahl. Consequences of vegetation density and prey species on spotted gar foraging. Transactions of the American Fisheries Society, 133(3):794–800, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OGorman:2000:SDD

[OEO^{+}00] Robert O’Gorman, Joseph H. Elrod, Randall W. Owens, Clifford P. Schneider, Thomas H. Eckert, and Brian F. Lantry. Shifts in depth distributions of alewives, rainbow smelt, and age-2 lake trout in Southern Lake Ontario following establishment of dreissenids. Transactions of the American Fisheries Society, 129(5): 1096–1106, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olden:2001:FHR

[OJ01] Julian D. Olden and Donald A. Jackson. Fish–habitat relationships in lakes: Gaining predictive and explanatory insight by using artificial neural networks. Transactions of the American Fisheries Society, 130(5):878–897, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olden:2002:PMF

[OJPN02] Julian D. Olden, Donald A. Jackson, and Pedro R. Peres-Neto. Predictive models of fish species distributions: a note on proper validation and chance predictions. Transactions of the American Fisheries Society, 131(2):329–336, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OGorman:2004:ESS

[OLS04] Robert O’Gorman, Brian F. Lantry, and Clifford P. Schneider. Effect of stock size, climate, predation, and trophic status on recruitment of alewives in Lake Ontario, 1978–2000. Transactions of the American Fisheries Society, 133(4):855–867, July 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olney:2006:MBA

[OLWC06] John E. Olney, Robert J. Latour, Brian E. Watkins, and Douglas G. Clarke. Migratory behavior of American shad in the York River, Virginia, with implications for estimating in-river exploitation from tag recovery data. Transactions of the American Fisheries Society, 135(4):889–896, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oliveira:2002:SDG

[OM02] Kenneth Oliveira and James D. McCleave. Sexually different growth histories of the American eel in four rivers in Maine. Transactions of the American Fisheries Society, 131(2):203–211, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ourso:2005:PTC

[OM05] R. T. Ourso and E. H. Moran. Photographic techniques for characterizing streambed particle sizes: Response to comment. Transactions of the American Fisheries Society, 134(6):1604, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [GRR05].

Overton:2009:STP

[OMM09] Anthony S. Overton, F. Joseph Margraf, and Eric B. May. Spatial and temporal patterns in the diet of striped bass in Chesapeake Bay. Transactions of the American Fisheries Society, 138(4):915– 926, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Odenkirk:2005:NST

[OO05] John Odenkirk and Steve Owens. Northern snakeheads in the tidal Potomac River system. Transactions of the American Fisheries Society, 134(6):1605–1609, November 2005. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Odenkirk:2007:ENS

[OO07] John Odenkirk and Steve Owens. Expansion of a northern snakehead population in the Potomac River system. Transactions of the American Fisheries Society, 136(6):1633–1639, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ostberg:2009:ERA

[OPH09] C. O. Ostberg, S. D. Pavlov, and L. Hauser. Evolutionary relationships among sympatric life history forms of Dolly Varden inhabiting the landlocked Kronotsky Lake, Kamchatka, and a neighboring anadromous population. Transactions of the American Fisheries Society, 138(1):1–14, January 2009. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ostberg:2006:HCA

[OR06] Carl O. Ostberg and Rusty J. Rodriguez. Hybridization and cytonuclear associations among native westslope cutthroat trout, introduced rainbow trout, and their hybrids within the Stehekin River Drainage, North Cascades National Park. Transactions of the American Fisheries Society, 135(4):924–942, July 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ouzts:2003:DFP

[OS03] Allyson Clair Ouzts and Stephen T. Szedlmayer. Diel feeding patterns of red snapper on artificial reefs in the North-Central Gulf of Mexico. Transactions of the American Fisheries Society, 132(6):1186–1193, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:2004:VPS

[OSS^{+}04] Jeffrey B. Olsen, William J. Spearman, George K. Sage, Steve J. Miller, Blair G. Flannery, and John K. Wenburg. Variation in the population structure of Yukon River chum and Coho salmon: Evaluating the potential impact of localized habitat degradation. Transactions of the American Fisheries Society, 133(2):476–483, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ostrand:2001:TDO

[OW01] Kenneth G. Ostrand and Gene R. Wilde. Temperature, dissolved oxygen, and salinity tolerances of five prairie stream fishes and their role in explaining fish assemblage patterns. Transactions of the American Fisheries Society, 130(5):742–749, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ostrand:2004:CPS

[OW04] Kenneth G. Ostrand and Gene R. Wilde. Changes in prairie stream fish assemblages restricted to isolated streambed pools. Transactions of the American Fisheries Society, 133(6):1329– 1338, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

ONeill:2009:MDL

[OW09] Sandra M. O’Neill and James E. West. Marine distribution, life history traits, and the accumulation of polychlorinated biphenyls in Chinook salmon from Puget Sound, Washington. Transactions of the American Fisheries Society, 138(3):616–632, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:2003:PDG

[OY03] Mark H. Olson and Brian P. Young. Patterns of diet and growth in co-occurring populations of largemouth bass and smallmouth bass. Transactions of the American Fisheries Society, 132(6): 1207–1213, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:2003:MUH

[OYB03] Mark H. Olson, Brian P. Young, and Kevin D. Blinkoff. Mechanisms underlying habitat use of juvenile largemouth bass and smallmouth bass. Transactions of the American Fisheries Society, 132(2):398–405, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pine:2001:DGS

[PA01] William E. Pine III and Micheal S. Allen. Differential growth and survival of weekly age-0 black crappie cohorts in a Florida Lake. Transactions of the American Fisheries Society, 130(1): 80–91, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pondella:2001:DPG

[PACH01] Daniel J. Pondella II, Larry G. Allen, Jorge A. Rosales Casian, and Tim E. Hovey. Demographic parameters of golden spotted rock bass Paralabrax auroguttatus from the Northern Gulf of California. Transactions of the American Fisheries Society, 130(4): 686–691, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pine:2001:PVG

[PAD01] William E. Pine III, Mike S. Allen, and Victoria J. Dreitz. Population viability of the Gulf of Mexico sturgeon: Inferences from capture–recapture and age-structured models. Transactions of the American Fisheries Society, 130(6):1164–1174, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Papadopoulos:2008:HBP

[Pap08] Anthony Papadopoulos. On the hydrodynamics-based power-law function and its application in fish swimming energetics. Transactions of the American Fisheries Society, 137(4):997–1006, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perry:2001:BCJ

[PAR01] Russell W. Perry, Noah S. Adams, and Dennis W. Rondorf. Buoyancy compensation of juvenile Chinook salmon implanted with two different size dummy transmitters. Transactions of the American Fisheries Society, 130(1):46–52, January 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Penaluna:2009:EIS

[PAS09] Brooke E. Penaluna, Ivan Arismendi, and Doris Soto. Evidence of interactive segregation between introduced trout and native fishes in Northern Patagonian rivers, Chile. Transactions of the American Fisheries Society, 138(4):839–845, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paragamian:2003:GET

[PB03] Vaughn L. Paragamian and Raymond C. P. Beamesderfer. Growth estimates from tagged white sturgeon suggest that ages from fin rays underestimate true age in the Kootenai River, USA and Canada. Transactions of the American Fisheries Society, 132 (5):895–903, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paragamian:2005:SPD

[PBI05] Vaughn L. Paragamian, Raymond C. P. Beamesderfer, and Susan C. Ireland. Status, population dynamics, and future prospects of the endangered Kootenai River white sturgeon population with and without hatchery intervention. Transactions of the American Fisheries Society, 134(2):518–532, March 2005. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pascual:2001:FDC

[PBR^{+}01] Miguel Pascual, Paul Bentzen, Carla Riva Rossi, Greg Mackey, Michael T. Kinnison, and Robert Walker. First documented case of anadromy in a population of introduced rainbow trout in Patagonia, Argentina. Transactions of the American Fisheries Society, 130(1):53–67, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Piller:2004:DFM

[PBT04] Kyle R. Piller, Henry L. Bart, Jr., and Jason A. Tipton. Decline of the frecklebelly madtom in the Pearl River based on contemporary and historical surveys. Transactions of the American Fisheries Society, 133(4):1004–1013, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Philipp:2009:SVA

[PCC^{+}09] David P. Philipp, Steven J. Cooke, Julie E. Claussen, Jeffrey B. Koppelman, Cory D. Suski, and Dale P. Burkett. Selection for vulnerability to angling in largemouth bass. Transactions of the American Fisheries Society, 138(1):189–199, January 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paukert:2006:DMH

[PCF06] Craig P. Paukert, Lewis G. Coggins, Jr., and Christopher E. Flaccus. Distribution and movement of humpback chub in the Colorado River, Grand Canyon, based on recaptures. Transactions of the American Fisheries Society, 135(3):539–544, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paukert:2001:EIT

[PCHB01] Craig P. Paukert, Phillip J. Chvala, Brian L. Heikes, and Michael L. Brown. Effects of implanted transmitter size and surgery on survival, growth, and wound healing of bluegill. Transactions of the American Fisheries Society, 130(5):975–980, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purchase:2005:PLH

[PCMS05] C. F. Purchase, N. C. Collins, G. E. Morgan, and B. J. Shuter. Predicting life history traits of yellow perch from environmental characteristics of lakes. Transactions of the American Fisheries Society, 134(5):1369–1381, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patterson:2008:TSV

[PCWC08] William F. Patterson III, James H. Cowan, Jr., Charles A. Wilson, and Zhongxing Chen. Temporal and spatial variability in juvenile red snapper otolith elemental signatures in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 137 (2):521–532, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paragamian:2005:VVL

[PD05] Vaughn L. Paragamian and Jeremy P. Duehr. Variations in vertical location of Kootenai River white sturgeon during the prespawn and spawning periods. Transactions of the American Fisheries Society, 134(1):261–266, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petersen:2008:OMD

[PDP08] James H. Petersen, Donald L. DeAngelis, and Craig P. Paukert. An overview of methods for developing bioenergetic and life history models for rare and endangered species. Transactions of the American Fisheries Society, 137(1):244–253, January 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Palmisano:2001:SSR

[PE01] Aldo N. Palmisano and Nancy E. Elder. Standardized seawater rearing of Chinook salmon smolts to evaluate hatchery practices showed low statistical power. Transactions of the American Fisheries Society, 130(3):409–416, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peake:2004:EUC

[Pea04] Stephan Peake. An evaluation of the use of critical swimming speed for determination of culvert water velocity criteria for smallmouth bass. Transactions of the American Fisheries Society, 133 (6):1472–1479, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2000:DSS

[PF00] Douglas P. Peterson and Chris J. Foote. Disturbance of smallstream habitat by spawning sockeye salmon in Alaska. Transactions of the American Fisheries Society, 129(4):924–934, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paukert:2001:CPS

[PF01a] Craig P. Paukert and William L. Fisher. Characteristics of paddlefish in a Southwestern U.S. reservoir, with comparisons of lentic and lotic populations. Transactions of the American Fisheries Society, 130(4):634–643, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulsen:2001:SRB

[PF01b] Charles M. Paulsen and Timothy R. Fisher. Statistical relationship between parr-to-smolt survival of Snake River spring– summer Chinook salmon and indices of land use. Transactions of the American Fisheries Society, 130(3):347–358, May 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pratt:2002:IPR

[PF02] Thomas C. Pratt and Michael G. Fox. Influence of predation risk on the overwinter mortality and energetic relationships of youngof-year walleyes. Transactions of the American Fisheries Society, 131(5):885–898, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulsen:2005:DHA

[PF05] Charles M. Paulsen and Timothy R. Fisher. Do habitat actions affect juvenile survival? An information-theoretic approach applied to endangered Snake River Chinook salmon. Transactions of the American Fisheries Society, 134(1):68–85, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prager:2001:EMM

[PG01] Michael H. Prager and C. Phillip Goodyear. Effects of mixedmetric data on production model estimation: Simulation study of a blue-marlin-like stock. Transactions of the American Fisheries Society, 130(5):927–939, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Propst:2004:RNN

[PG04] David L. Propst and Keith B. Gido. Responses of native and nonnative fishes to natural flow regime mimicry in the San Juan River. Transactions of the American Fisheries Society, 133(4): 922–931, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petty:2007:SDT

[PG07] J. Todd Petty and Gary D. Grossman. Size-dependent territoriality of mottled sculpin in a Southern Appalachian stream. Transactions of the American Fisheries Society, 136(6):1750– 1761, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phelps:2008:IME

[PGW08] Quinton E. Phelps, Brian D. S. Graeb, and David W. Willis. Influence of the Moran Effect on spatiotemporal synchrony in common carp recruitment. Transactions of the American Fisheries Society, 137(6):1701–1708, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulsen:2007:MTE

[PHF07] Charles M. Paulsen, Richard A. Hinrichsen, and Timothy R. Fisher. Measure twice, estimate once: Pacific salmon population viability analysis for highly variable populations. Transactions of the American Fisheries Society, 136(2):346–364, March 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pont:2009:PIB

[PHWS09] Didier Pont, Robert M. Hughes, Thomas R. Whittier, and Stefan Schmutz. A predictive index of biotic integrity model for aquaticvertebrate assemblages of Western U.S. streams. Transactions of the American Fisheries Society, 138(2):292–305, March 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pirhalla:2004:EFH

[Pir04] Douglas E. Pirhalla. Evaluating fish–habitat relationships for refining regional indexes of biotic integrity: Development of a tolerance index of habitat degradation for Maryland stream fishes. Transactions of the American Fisheries Society, 133(1):144–159, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pritchard:2007:GDW

[PJC07] V. L. Pritchard, K. Jones, and D. E. Cowley. Genetic diversity within fragmented cutthroat trout populations. Transactions of the American Fisheries Society, 136(3):606–623, May 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pollock:2004:CTF

[PJH04] Kenneth H. Pollock, Honghua Jiang, and Joseph E. Hightower. Combining telemetry and fisheries tagging models to estimate fishing and natural mortality rates. Transactions of the American Fisheries Society, 133(3):639–648, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2009:DES

[PJSL09] James T. Peterson, C. Rhett Jackson, Colin P. Shea, and Guoyuan Li. Development and evaluation of a stream channel classification for estimating fish responses to changing streamflow. Transactions of the American Fisheries Society, 138(5):1123–1137, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pender:2002:FIB

[PK02] Danielle R. Pender and Thomas J. Kwak. Factors influencing brown trout reproductive success in Ozark tailwater rivers. Transactions of the American Fisheries Society, 131(4):698–717, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peacock:2004:ACV

[PK04] Mary M. Peacock and Veronica Kirchoff. Assessing the conservation value of hybridized cutthroat trout populations in the Quinn River drainage, Nevada. Transactions of the American Fisheries Society, 133(2):309–325, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Principe:2007:GED

[PKM07] Nicholas D. Principe, Clifford E. Kraft, and Edward L. Mills. Gastric evacuation and daily ration of naturally produced age-0 Chinook salmon in Lake Ontario. Transactions of the American Fisheries Society, 136(5):1206–1215, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pascual:2002:FDC

[PKR02] Miguel Pascual, Michael Kinnison, and Carla Riva Rossi. First documented case of anadromy in a population of introduced rainbow trout in Patagonia, Argentina: Response to comment. Transactions of the American Fisheries Society, 131(4):585–588, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Beh02].

Pine:2007:MMS

[PKR07] William E. Pine III, Thomas J. Kwak, and James A. Rice. Modeling management scenarios and the effects of an introduced apex predator on a coastal riverine fish community. Transactions of the American Fisheries Society, 136(1):105–120, January 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pine:2005:DSI

[PKWR05] William E. Pine III, Thomas J. Kwak, D. Scott Waters, and James A. Rice. Diet selectivity of introduced flathead catfish in coastal rivers. Transactions of the American Fisheries Society, 134(4):901–909, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pine:2000:FSS

[PLD00] William E. Pine III, Stuart A. Ludsin, and Dennis R. DeVries. First-summer survival of largemouth bass cohorts: Is early spawning really best? Transactions of the American Fisheries Society, 129(2):504–513, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pothoven:2008:RBI

[PLH^{+}08] Steven A. Pothoven, Stuart A. Ludsin, Tomas O. Höök, David L. Fanslow, Doran M. Mason, Paris D. Collingsworth, and Jason J.

Van Tassell. Reliability of bioelectrical impedance analysis for estimating whole-fish energy density and percent lipids. Transactions of the American Fisheries Society, 137(5):1519–1529, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petty:2005:SSD

[PLM05] J. Todd Petty, Peter J. Lamothe, and Patricia M. Mazik. Spatial and seasonal dynamics of brook trout populations inhabiting a Central Appalachian watershed. Transactions of the American Fisheries Society, 134(3):572–587, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearlstein:2007:EDA

[PLO07] Jamie H. Pearlstein, Benjamin H. Letcher, and Mariska Obedzinski. Early discrimination of Atlantic salmon smolt age: Time course of the relative effectiveness of body size and shape. Transactions of the American Fisheries Society, 136(6):1622–1632, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pyper:2002:SCS

[PMP^{+}02] Brian J. Pyper, Franz J. Mueter, Randall M. Peterman, David J. Blackbourn, and Chris C. Wood. Spatial covariation in survival rates of Northeast Pacific chum salmon. Transactions of the American Fisheries Society, 131(3):343–363, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pyper:2005:ASC

[PMP05] Brian J. Pyper, Franz J. Mueter, and Randall M. Peterman. Across-species comparisons of spatial scales of environmental effects on survival rates of Northeast Pacific salmon. Transactions of the American Fisheries Society, 134(1):86–104, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:2006:BRB

[PMRH06] Steven J. Parker, Howard I. McElderry, Polly S. Rankin, and Robert W. Hannah. Buoyancy regulation and barotrauma in two species of nearshore rockfish. Transactions of the American Fisheries Society, 135(5):1213–1223, September 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pratt:2009:BAH

[POH^{+}09] T. C. Pratt, L. M. O’Connor, A. G. Hallett, R. L. McLaughlin, C. Katopodis, D. B. Hayes, and R. A. Bergstedt. Balancing aquatic habitat fragmentation and control of invasive species: Enhancing selective fish passage at sea lamprey control barriers. Transactions of the American Fisheries Society, 138(3):652–665, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Polivka:2007:UTF

[Pol07] Karl M. Polivka. Use of techniques from foraging theory to quantify the cost of predation for benthic fishes. Transactions of the American Fisheries Society, 136(6):1778–1790, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paul:2001:SDN

[PP01] Andrew J. Paul and John R. Post. Spatial distribution of native and nonnative salmonids in streams of the Eastern slopes of the Canadian rocky mountains. Transactions of the American Fisheries Society, 130(3):417–430, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pientka:2002:HSP

[PP02] Bernard Pientka and Donna L. Parrish. Habitat selection of predator and prey: Atlantic salmon and rainbow smelt overlap, based on temperature and dissolved oxygen. Transactions of the American Fisheries Society, 131(6):1180–1193, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petersen:2005:DBM

[PP05] James H. Petersen and Craig P. Paukert. Development of a bioenergetics model for humpback chub and evaluation of water temperature changes in the Grand Canyon, Colorado River. Transactions of the American Fisheries Society, 134(4):960–974, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Penne:2008:SDA

[PP08] Christopher R. Penne and Clay L. Pierce. Seasonal distribution, aggregation, and habitat selection of common carp in Clear Lake, Iowa. Transactions of the American Fisheries Society, 137(4): 1050–1062, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2009:CFR

[PPDV09] Ron Pierce, Craig Podner, Michael Davidson, and E. Richard Vincent. Correlation of fluvial rainbow trout spawning life history with severity of infection by Myxobolus cerebralis in the Blackfoot River Basin, Montana. Transactions of the American Fisheries Society, 138(2):251–263, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peters:2006:MAR

[PPR06] Lindsay M. Peters, Mark A. Pegg, and Ulrich G. Reinhardt. Movements of adult radio-tagged bighead carp in the Illinois River. Transactions of the American Fisheries Society, 135(5): 1205–1212, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsley:2008:SDM

[PPWvdL08] Michael J. Parsley, Nicholas D. Popoff, Corey D. Wright, and Bjorn K. van der Leeuw. Seasonal and diel movements of white sturgeon in the Lower Columbia River. Transactions of the American Fisheries Society, 137(4):1007–1017, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paramore:2001:DCI

[PR01a] Lee M. Paramore and Roger A. Rulifson. Dorsal coloration as an indicator of different life history patterns for striped bass within a single watershed of Atlantic Canada. Transactions of the American Fisheries Society, 130(4):663–674, July 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2001:EPC

[PR01b] James T. Peterson and Charles F. Rabeni. Evaluating the physical characteristics of channel units in an Ozark stream. Transactions of the American Fisheries Society, 130(5):898–910, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2001:RFA

[PR01c] James T. Peterson and Charles F. Rabeni. The relation of fish assemblages to channel units in an Ozark stream. Transactions of the American Fisheries Society, 130(5):911–926, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petrie:2006:LFE

[PR06] Megan E. Petrie and Clifford H. Ryer. Laboratory and field evidence for structural habitat affinity of young-of-the-year lingcod. Transactions of the American Fisheries Society, 135(6):1622– 1630, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:2001:RBS

[PRME01] S. L. Parker, L. G. Rudstam, E. L. Mills, and D. W. Einhouse. Retention of Bythotrephes spines in the stomachs of Eastern Lake Erie rainbow smelt. Transactions of the American Fisheries Society, 130(5):988–994, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perrin:2003:WSS

[PRR03] Chris J. Perrin, Laura L. Rempel, and Marvin L. Rosenau. White sturgeon spawning habitat in an unregulated river: Fraser River, Canada. Transactions of the American Fisheries Society, 132 (1):154–165, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:2003:GAS

[PSB^{+}03] Kevin S. Page, Kim T. Scribner, Kristine R. Bennett, Laura M. Garzel, and Mary K. Burnham-Curtis. Genetic assessment of strain-specific sources of lake trout recruitment in the Great Lakes. Transactions of the American Fisheries Society, 132(5): 877–894, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:2005:GEG

[PSB^{+}05] Kevin S. Page, Kim T. Scribner, Dale Bast, Mark E. Holey, and Mary K. Burnham-Curtis. Genetic evaluation of a Great Lakes Lake trout hatchery program. Transactions of the American Fisheries Society, 134(4):872–891, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:2004:GDW

[PSBC04] Kevin S. Page, Kim T. Scribner, and Mary Burnham-Curtis. Genetic diversity of wild and hatchery lake trout populations: Relevance for management and restoration in the Great Lakes. Transactions of the American Fisheries Society, 133(3):674–691, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2008:ARS

[PSD^{+}08] Douglas L. Peterson, Paul Schueller, Rob DeVries, Joel Fleming, Cheryl Grunwald, and Isaac Wirgin. Annual run size and genetic characteristics of Atlantic sturgeon in the Altamaha River, Georgia. Transactions of the American Fisheries Society, 137(2):393– 401, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Puckett:2008:VAL

[PSJ08] Brandon J. Puckett, David H. Secor, and Se-Jong Ju. Validation and application of lipofuscin-based age determination for Chesapeake Bay blue crabs Callinectes sapidus. Transactions of the American Fisheries Society, 137(6):1637–1649, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pangle:2004:OSJ

[PSKH04] Kevin L. Pangle, Trent M. Sutton, Ronald E. Kinnunen, and Michael H. Hoff. Overwinter survival of juvenile lake herring in relation to body size, physiological condition, energy stores, and food ration. Transactions of the American Fisheries Society, 133 (5):1235–1246, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patrick:2009:LSL

[PSS09] Holly K. Patrick, Trent M. Sutton, and William D. Swink. Lethality of sea lamprey parasitism on lake sturgeon. Transactions of the American Fisheries Society, 138(5):1065–1075, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:2009:IMH

[PSSV^{+}09] Aaron D. Parker, Carol A. Stepien, Osvaldo J. Sepulveda-Villet, Clifton B. Ruehl, and Donald G. Uzarski. The interplay of morphology, habitat, resource use, and genetic relationships in young yellow perch. Transactions of the American Fisheries Society, 138(4):899–914, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2003:IAP

[PT03a] Rodney B. Pierce and Cynthia M. Tomcko. Interrelationships among production, density, growth, and mortality of northern pike in seven North-Central Minnesota lakes. Transactions of the American Fisheries Society, 132(1):143–153, January 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2003:VGN

[PT03b] Rodney B. Pierce and Cynthia M. Tomcko. Variation in gillnet and angling catchability with changing density of northern pike in a small Minnesota Lake. Transactions of the American Fisheries Society, 132(4):771–779, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2005:DBN

[PT05] Rodney B. Pierce and Cynthia M. Tomcko. Density and biomass of native northern pike populations in relation to basin-scale characteristics of North-Central Minnesota lakes. Transactions of the American Fisheries Society, 134(1):231–241, January 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:2004:EME

[PTG04] James T. Peterson, Russell F. Thurow, and John W. Guzevich. An evaluation of multipass electrofishing for estimating the abundance of stream-dwelling salmonids. Transactions of the American Fisheries Society, 133(2):462–475, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pothoven:2004:DPS

[PV04] Steven A. Pothoven and Henry A. Vanderploeg. Diet and prey selection of alewives in Lake Michigan: Seasonal, depth, and interannual patterns. Transactions of the American Fisheries Society, 133(5):1068–1077, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paukert:2002:SDH

[PW02] Craig P. Paukert and David W. Willis. Seasonal and diel habitat selection by bluegills in a shallow natural lake. Transactions of the American Fisheries Society, 131(6):1131–1139, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:2009:NER

[PW09] David M. Post and Annika W. Walters. Nutrient excretion rates of anadromous alewives during their spawning migration. Transactions of the American Fisheries Society, 138(2):264–268, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearse:2007:INP

[PWDG07] Devon E. Pearse, Lisa Wooninck, Cheryl A. Dean, and John Carlos Garza. Identification of Northeastern Pacific rockfish using multilocus nuclear DNA genotypes. Transactions of the American Fisheries Society, 136(1):272–280, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Piller:2005:CGI

[PWLL05] Kyle R. Piller, Chris C. Wilson, Carol Eunmi Lee, and John Lyons. Conservation genetics of inland lake trout in the Upper Mississippi River Basin: Stocked or native ancestry? Transactions of the American Fisheries Society, 134(4):789–802, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patterson:2001:MTR

[PWSC01] William F. Patterson III, J. Carter Watterson, Robert L. Shipp, and James H. Cowan, Jr. Movement of tagged red snapper in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 130(4):533–545, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pollock:2007:FAS

[PYFS07] Kenneth H. Pollock, Jun Yoshizaki, Mary C. Fabrizio, and Stephen T. Schram. Factors affecting survival rates of a recovering lake trout population estimated by mark–recapture in Lake Superior, 1969–1996. Transactions of the American Fisheries Society, 136(1):185–194, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2001:SSS

[QB01] Thomas P. Quinn and Gregory B. Buck. Size- and sex-selective mortality of adult sockeye salmon: Bears, gulls, and fish out of water. Transactions of the American Fisheries Society, 130(6):

995–1005, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quist:2005:REB

[QH05] Michael C. Quist and Wayne A. Hubert. Relative effects of biotic and abiotic processes: a test of the biotic–abiotic constraining hypothesis as applied to cutthroat trout. Transactions of the American Fisheries Society, 134(3):676–686, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quist:2004:RAH

[QHR04] Michael C. Quist, Wayne A. Hubert, and Frank J. Rahel. Relations among habitat characteristics, exotic species, and turbidriver cyprinids in the Missouri River drainage of Wyoming. Transactions of the American Fisheries Society, 133(3):727–742, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2003:FAC

[QK03] Jeffrey W. Quinn and Thomas J. Kwak. Fish assemblage changes in an Ozark river after impoundment: a long-term perspective. Transactions of the American Fisheries Society, 132(1):110–119, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinones:2005:HUJ

[QM05] Rebecca M. Quiñones and Timothy J. Mulligan. Habitat use by juvenile salmonids in the Smith River Estuary, California. Transactions of the American Fisheries Society, 134(5):1147– 1158, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2002:ASE

[QPG^{+}02] Thomas P. Quinn, Jeramie A. Peterson, Vincent F. Gallucci, William K. Hershberger, and Ernest L. Brannon. Artificial selection and environmental change: Countervailing factors affecting the timing of spawning by Coho and Chinook salmon. Transactions of the American Fisheries Society, 131(4):591–598, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2004:CRM

[Qui04] Stephen P. Quinn. Catch and release in marine recreational fisheries. Transactions of the American Fisheries Society, 133(2): 491–492, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2005:CSC

[Qui05] Thomas P. Quinn. Comment: Sperm competition in salmon hatcheries — the need to institutionalize genetically benign spawning protocols. Transactions of the American Fisheries Society, 134(6):1490–1494, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [Cam05].

Quinn:2004:IFM

[QVPG04] Thomas P. Quinn, Leif Asbjørn Vøllestad, Jeramie Peterson, and Vincent Gallucci. Influences of freshwater and marine growth on the egg size–egg number tradeoff in Coho and Chinook salmon. Transactions of the American Fisheries Society, 133(1):55–65, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:2007:MRS

[RA07] James H. Roberts and Paul L. Angermeier. Movement responses of stream fishes to introduced corridors of complex cover. Transactions of the American Fisheries Society, 136(4):971–978, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rogers:2009:EGR

[RA09] Mark W. Rogers and Micheal S. Allen. Exploring the generality of recruitment hypotheses for largemouth bass along a latitudinal gradient of Florida lakes. Transactions of the American Fisheries Society, 138(1):23–37, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahel:2000:BR

[Rah00] Frank J. Rahel. Book review. Transactions of the American Fisheries Society, 129(3):886–887, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rummer:2005:PES

[RB05] Jodie L. Rummer and Wayne A. Bennett. Physiological effects of swim bladder overexpansion and catastrophic decompression on red snapper. Transactions of the American Fisheries Society, 134 (6):1457–1470, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosenfield:2007:PDD

[RB07] Jonathan A. Rosenfield and Randall D. Baxter. Population dynamics and distribution patterns of longfin smelt in the San Francisco Estuary. Transactions of the American Fisheries Society, 136(6):1577–1592, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Renan:2006:FEN

[RBLF06] X. Renán, T. Brulé, and R. Lecomte-Finiger. First evidence of a nursery habitat for juvenile gag in the Southern Gulf of Mexico. Transactions of the American Fisheries Society, 135(3):595–603, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rypel:2006:GFD

[RBM06] Andrew L. Rypel, David R. Bayne, and Justin B. Mitchell. Growth of freshwater drum from lotic and lentic habitats in Alabama. Transactions of the American Fisheries Society, 135(4): 987–997, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryon:2000:SDD

[RBR^{+}00] Michael G. Ryon, John J. Beauchamp, W. Kelly Roy, Elizabeth Schilling, Brian A. Carrico, and Roxanna L. Hinzman. Stream dispersal of dead fish and survey effectiveness in a simulated fish kill. Transactions of the American Fisheries Society, 129(1):89– 100, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rogers:2004:GPS

[RC04] Sean M. Rogers and R. Allen Curry. Genetic population structure of brook trout inhabiting a large river watershed. Transactions of the American Fisheries Society, 133(5):1138–1149, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robinson:2004:OMY

[RCB^{+}04] M. Robinson, S. Courtenay, T. Benfey, L. Maceda, and I. Wirgin. Origin and movements of young-of-the-year striped bass in the Southern Gulf of St. Lawrence, New Brunswick. Transactions of the American Fisheries Society, 133(2):412–426, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rakocinski:2000:REF

[RCP00] Chet F. Rakocinski, Bruce H. Comyns, and Mark S. Peterson. Relating environmental fluctuation and the early growth of estuarine fishes: Ontogenetic standardization. Transactions of the American Fisheries Society, 129(1):210–221, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roghair:2002:RBT

[RDU02] Craig N. Roghair, C. Andrew Dolloff, and Martin K. Underwood. Response of a brook trout population and instream habitat to a catastrophic flood and debris flow. Transactions of the American Fisheries Society, 131(4):718–730, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:2008:PEC

[RE08] Stephen C. Riley and Allison N. Evans. Phylogenetic and ecological characteristics associated with thiaminase activity in Laurentian Great Lakes fishes. Transactions of the American Fisheries Society, 137(1):147–157, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinhardt:2001:SSF

[Rei01] Ulrich G. Reinhardt. Selection for surface feeding in farmed and sea-ranched masu salmon juveniles. Transactions of the American Fisheries Society, 130(1):155–158, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruetz:2006:EPI

[REK06] Carl R. Ruetz III, Brendan M. Earl, and Steven L. Kohler. Evaluating passive integrated transponder tags for marking mottled sculpins: Effects on growth and mortality. Transactions of the American Fisheries Society, 135(6):1456–1461, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ramstad:2003:GPE

[RFOR03] K. M. Ramstad, C. J. Foote, J. B. Olsen, and D. Rogers. Genetic and phenotypic evidence of reproductive isolation between seasonal runs of sockeye salmon in Bear Lake, Alaska. Transactions of the American Fisheries Society, 132(5):997–1013, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rutherford:2001:PEF

[RGK01] D. Allen Rutherford, Kevin R. Gelwicks, and William E. Kelso. Physicochemical effects of the flood pulse on fishes in the Atchafalaya River Basin, Louisiana. Transactions of the American Fisheries Society, 130(2):276–288, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ramcharitar:2006:BFF

[RGP06] John Ramcharitar, Damon P. Gannon, and Arthur N. Popper. Bioacoustics of fishes of the family Sciaenidae (croakers and drums). Transactions of the American Fisheries Society, 135(5): 1409–1431, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rowe:2001:STL

[RH01] Sherrylynn Rowe and Richard L. Haedrich. Streamer tag loss from American lobsters. Transactions of the American Fisheries Society, 130(3):516–518, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reese:2002:TDI

[RH02] Carl D. Reese and Bret C. Harvey. Temperature-dependent interactions between juvenile steelhead and Sacramento pikeminnow in laboratory streams. Transactions of the American Fisheries Society, 131(4):599–606, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rowe:2006:ESP

[RH06a] Sherrylynn Rowe and Jeffrey A. Hutchings. Erratum: Sound Production by Atlantic Cod During Spawning. Transactions of the American Fisheries Society, 135(3):e4, May 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [RH06b].

Rowe:2006:SPA

[RH06b] Sherrylynn Rowe and Jeffrey A. Hutchings. Sound production by Atlantic cod during spawning. Transactions of the American Fisheries Society, 135(2):529–538, March 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [RH06a].

Reiser:2006:DFW

[RHB^{+}06] Dudley W. Reiser, Chi-Ming Huang, Stuart Beck, Michael Gagner, and Eric Jeanes. Defining flow Windows for upstream passage of adult anadromous salmonids at cascades and falls. Transactions of the American Fisheries Society, 135(3):668–679, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rasmussen:2002:EPM

[RHG^{+}02] Paul W. Rasmussen, Dennis M. Heisey, Stephen J. Gilbert, Ruth M. King, and Steven W. Hewett. Estimating postspawning movement of walleyes among interconnected lakes of Northern Wisconsin. Transactions of the American Fisheries Society, 131 (6):1020–1032, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rand:2006:ERD

[RHM^{+}06] P. S. Rand, S. G. Hinch, J. Morrison, M. G. G. Foreman, M. J. MacNutt, J. S. Macdonald, M. C. Healey, A. P. Farrell, and D. A. Higgs. Effects of river discharge, temperature, and future climates on energetics and mortality of adult migrating Fraser River sockeye salmon. Transactions of the American Fisheries Society, 135(3):655–667, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reisenbichler:2006:GPG

[RHR^{+}06] Reginald R. Reisenbichler, Michael C. Hayes, Stephen P. Rubin, Lisa A. Wetzel, and Bruce M. Baker. PEPA-1* genotype affects return rate for hatchery steelhead. Transactions of the American Fisheries Society, 135(1):28–39, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruetz:2003:IIB

[RHV03] Carl R. Ruetz III, Amy L. Hurford, and Bruce Vondracek. Interspecific interactions between brown trout and slimy sculpin in stream enclosures. Transactions of the American Fisheries Society, 132(3):611–618, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rieman:2007:ACW

[RIA^{+}07] Bruce E. Rieman, Daniel Isaak, Susan Adams, Dona Horan, David Nagel, Charles Luce, and Deborah Myers. Anticipated climate warming effects on bull trout habitats and populations across the interior Columbia River Basin. Transactions of the American Fisheries Society, 136(6):1552–1565, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rideout:2004:NSE

[Rid04] Stephen G. Rideout. Nutrients in salmonid ecosystems: Sustaining production and biodiversity. Transactions of the American Fisheries Society, 133(2):492–493, March 2004. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ridgway:2008:RCL

[Rid08] Mark S. Ridgway. A roadmap for coasters: Landscapes, life histories, and the conservation of brook trout. Transactions of the American Fisheries Society, 137(4):1179–1191, July 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruetz:2000:SPL

[RJ00] Carl R. Ruetz III and Cecil A. Jennings. Swimming performance of larval robust redhorse Moxostoma robustum and low-velocity habitat modeling in the Oconee River, Georgia. Transactions of the American Fisheries Society, 129(2):398–407, March 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rothschild:2009:SCB

[RJ09] Brian J. Rothschild and Yue Jiao. The structure of complex biological reference points and the theory of replacement. Transactions of the American Fisheries Society, 138(5):949–965, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rundio:2008:SPT

[RL08] David E. Rundio and Steven T. Lindley. Seasonal patterns of terrestrial and aquatic prey abundance and use by Oncorhynchus mykiss in a California coastal basin with a Mediterranean climate. Transactions of the American Fisheries Society, 137(2):467–480, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radchenko:2004:DGF

[RM04] Vladimir I. Radchenko and Ole A. Mathisen. Distribution, growth, and feeding of sockeye salmon in the Western Bering Sea. Transactions of the American Fisheries Society, 133(3):606–621, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raborn:2003:MPS

[RMD03] Scott W. Raborn, Leandro E. Miranda, and M. Todd Driscoll. Modeling predation as a source of mortality for piscivorous fishes in a Southeastern U.S. reservoir. Transactions of the American Fisheries Society, 132(3):560–575, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raborn:2007:PSP

[RMD07] Scott W. Raborn, Leandro E. Miranda, and M. Todd Driscoll. Prey supply and predator demand in a reservoir of the Southeastern United States. Transactions of the American Fisheries Society, 136(1):12–23, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rulifson:2008:IHU

[RMD08] Roger A. Rulifson, Sean A. McKenna, and Michael J. Dadswell. Intertidal habitat use, population characteristics, movement, and exploitation of striped bass in the Inner Bay of Fundy, Canada. Transactions of the American Fisheries Society, 137(1):23–32, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roni:2006:CSS

[RMG^{+}06] Phil Roni, Sarah A. Morley, Patsy Garcia, Chris Detrick, Dave King, and Eric Beamer. Coho salmon smolt production from constructed and natural floodplain habitats. Transactions of the American Fisheries Society, 135(5):1398–1408, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:2008:CRE

[RMH^{+}08] Kenneth A. Rose, Bernard A. Megrey, Douglas Hay, Francisco Werner, and Jake Schweigert. Climate regime effects on Pacific herring growth using coupled nutrient-phytoplankton-zooplankton and bioenergetics models. Transactions of the American Fisheries Society, 137(1):278–297, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rich:2003:LHW

[RMRT03] Cecil F. Rich, Jr., Thomas E. McMahon, Bruce E. Rieman, and William L. Thompson. Local-habitat, watershed, and biotic features associated with bull trout occurrence in Montana streams. Transactions of the American Fisheries Society, 132(6): 1053–1064, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rasmussen:2003:IGM

[ROC^{+}03] Charlotte Rasmussen, Carl O. Ostberg, David R. Clifton, James L. Holloway, and Russell J. Rodriguez. Identification of a genetic marker that discriminates ocean-type and stream-type Chinook salmon in the Columbia River Basin. Transactions of the American Fisheries Society, 132(1):131–142, January 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roni:2002:HUF

[Ron02] Philip Roni. Habitat use by fishes and Pacific giant salamanders in small Western Oregon and Washington streams. Transactions of the American Fisheries Society, 131(4):743–761, July 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosenfeld:2003:AHR

[Ros03a] Jordan Rosenfeld. Assessing the habitat requirements of stream fishes: an overview and evaluation of different approaches. Transactions of the American Fisheries Society, 132(5):953–968, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:2003:FFD

[Ros03b] Stephen T. Ross. Freshwater fish distribution. Transactions of the American Fisheries Society, 132(2):406–407, March 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rothschild:2007:CAC

[Rot07] Brian J. Rothschild. Coherence of Atlantic cod stock dynamics in the Northwest Atlantic Ocean. Transactions of the American Fisheries Society, 136(3):858–874, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rexroad:2003:DNS

[RP03] C. E. Rexroad III and Y. Palti. Development of ninety-seven polymorphic microsatellite markers for rainbow trout. Transactions of the American Fisheries Society, 132(6):1214–1221, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruiz:2007:ERI

[RP07] John C. Ruiz and James T. Peterson. An evaluation of the relative influence of spatial, statistical, and biological factors on the accuracy of stream fish species presence models. Transactions of the American Fisheries Society, 136(6):1640–1653, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roni:2001:EWP

[RQ01] Philip Roni and Thomas P. Quinn. Effects of wood placement on movements of trout and juvenile Coho salmon in natural and artificial stream channels. Transactions of the American Fisheries Society, 130(4):675–685, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robards:2002:MTA

[RQ02] Martin D. Robards and Thomas P. Quinn. The migratory timing of adult summer-run steelhead in the Columbia River over six decades of environmental change. Transactions of the American Fisheries Society, 131(4):523–536, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:2008:ICS

[RR08] James J. Roberts and Frank J. Rahel. Irrigation canals as sink habitat for trout and other fishes in a Wyoming drainage. Transactions of the American Fisheries Society, 137(4):951–961, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:2008:DDF

[RRN^{+}08] Stephen C. Riley, Edward F. Roseman, S. Jerrine Nichols, Timothy P. O’Brien, Courtney S. Kiley, and Jeffrey S. Schaeffer. Deepwater demersal fish community collapse in Lake Huron. Transactions of the American Fisheries Society, 137(6):1879–1890, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodnick:2008:HSI

[RSB^{+}08] Kenneth J. Rodnick, Sophie St.-Hilaire, Pavan K. Battiprolu, Steven M. Seiler, Michael L. Kent, Madison S. Powell, and Joseph L. Ebersole. Habitat selection influences sex distribution, morphology, tissue biochemistry, and parasite load of juvenile Coho salmon in the West Fork Smith River, Oregon. Transactions of the American Fisheries Society, 137(6):1571–1590, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryan:2003:RVA

[RSBF03] Brad A. Ryan, Steven G. Smith, Jo Anne M. Butzerin, and John W. Ferguson. Relative vulnerability to avian predation of juvenile salmonids tagged with passive integrated transponders in the Columbia River Estuary, 1998–2000. Transactions of the American Fisheries Society, 132(2):275–288, March 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robinson:2009:OSP

[RSBS09] T. Craig Robinson, Peter W. Sorensen, Jennifer M. Bayer, and James G. Seelye. Olfactory sensitivity of Pacific lampreys to lamprey bile acids. Transactions of the American Fisheries Society, 138(1):144–152, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:2007:SES

[RSC07] Martha J. Robertson, David A. Scruton, and Keith D. Clarke. Seasonal effects of suspended sediment on the behavior of juvenile Atlantic salmon. Transactions of the American Fisheries Society, 136(3):822–828, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruetz:2009:EDI

[RSP09] Carl R. Ruetz III, Dana L. Strouse, and Steven A. Pothoven. Energy density of introduced round goby compared with four native fishes in a Lake Michigan tributary. Transactions of the American Fisheries Society, 138(4):938–947, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robison:2004:PRC

[RT04] Barrie D. Robison and Gary H. Thorgaard. The phenotypic relationship of a clonal line to its population of origin: Rapid embryonic development in an Alaskan population of rainbow trout. Transactions of the American Fisheries Society, 133(2):455–461, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rudershausen:2005:PSD

[RTBH05] Paul J. Rudershausen, Jack E. Tuomikoski, Jeffrey A. Buckel, and Joseph E. Hightower. Prey selectivity and diet of striped bass in Western Albemarle Sound, North Carolina. Transactions of the American Fisheries Society, 134(5):1059–1074, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosenfield:2000:AHI

[RTG00] Jonathan A. Rosenfield, Thomas Todd, and Roger Greil. Asymmetric hybridization and introgression between pink salmon and Chinook salmon in the Laurentian Great Lakes. Transactions of the American Fisheries Society, 129(3):670–679, May 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roseman:2001:RWE

[RTH^{+}01] Edward F. Roseman, William W. Taylor, Daniel B. Hayes, Roger L. Knight, and Robert C. Haas. Removal of walleye eggs from reefs in Western Lake Erie by a catastrophic storm. Transactions of the American Fisheries Society, 130(2):341–346, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rulifson:2003:MFF

[Rul03] Roger A. Rulifson. Migration of freshwater fishes. Transactions of the American Fisheries Society, 132(2):406–407, March 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodtka:2007:EWT

[RV07] Mike C. Rodtka and John P. Volpe. Effects of water temperature on interspecific competition between juvenile bull trout and brook trout in an artificial stream. Transactions of the American Fisheries Society, 136(6):1714–1727, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Runstrom:2001:PSP

[RVJ01] Ann L. Runstrom, Bruce Vondracek, and Cecil A. Jennings. Population statistics for paddlefish in the Wisconsin River. Transactions of the American Fisheries Society, 130(4):546–556, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rudershausen:2008:CRF

[RWB^{+}08] Paul J. Rudershausen, Erik H. Williams, Jeffrey A. Buckel, Jennifer C. Potts, and Charles S. Manooch III. Comparison of reef fish catch per unit effort and total mortality between the 1970s and 2005–2006 in Onslow Bay, North Carolina. Transactions of the American Fisheries Society, 137(5):1389–1405, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruzycki:2001:SCC

[RWL01] James R. Ruzycki, Wayne A. Wurtsbaugh, and Chris Luecke. Salmonine consumption and competition for endemic prey fishes in Bear Lake, Utah–Idaho. Transactions of the American Fisheries Society, 130(6):1175–1189, November 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smogor:2001:DRF

[SA01] Roy A. Smogor and Paul L. Angermeier. Determining a regional framework for assessing biotic integrity of Virginia streams. Transactions of the American Fisheries Society, 130(1):18–35, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shotwell:2004:EIA

[SA04a] S. Kalei Shotwell and Milo D. Adkison. Estimating indices of abundance and escapement of Pacific salmon for data-limited situations. Transactions of the American Fisheries Society, 133(3):

538–558, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spier:2004:ETI

[SA04b] Timothy W. Spier and Jonathan R. Ackerson. Effect of temperature on the identification of larval black crappies, white crappies, and F_{1} hybrid crappies. Transactions of the American Fisheries Society, 133(3):789–793, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [Ano05].

Salinger:2006:EWT

[SA06] David H. Salinger and James J. Anderson. Effects of water temperature and flow on adult salmon migration swim speed and delay. Transactions of the American Fisheries Society, 135(1):188– 199, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sackett:2009:SAM

[SARC09] Dana K. Sackett, D. Derek Aday, James A. Rice, and W. Gregory Cope. A statewide assessment of mercury dynamics in North Carolina water bodies and fish. Transactions of the American Fisheries Society, 138(6):1328–1341, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2007:IMC

[SAT^{+}07] Christian T. Smith, Anton Antonovich, William D. Templin, Carita M. Elfstrom, Shawn R. Narum, and Lisa W. Seeb. Impacts of marker class bias relative to locus-specific variability on population inferences in Chinook salmon: a comparison of single-nucleotide polymorphisms with short tandem repeats and allozymes. Transactions of the American Fisheries Society, 136 (6):1674–1687, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savitz:2004:BBE

[Sav04] Jan Savitz. Black bass: Ecology, conservation, and management. Transactions of the American Fisheries Society, 133(1):231–234, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2003:RDR

[SB03a] Todd R. Smith and Lawrence J. Buckley. RNA–DNA ratio in scales from juvenile cod provides a nonlethal measure of feeding condition. Transactions of the American Fisheries Society, 132(1):9–17, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sutton:2003:IET

[SB03b] Trent M. Sutton and Angela C. Benson. Influence of external transmitter shape and size on tag retention and growth of juvenile lake sturgeon. Transactions of the American Fisheries Society, 132(6):1257–1263, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scholten:2005:PCA

[SB05] George D. Scholten and Phillip W. Bettoli. Population characteristics and assessment of overfishing for an exploited paddlefish population in the Lower Tennessee River. Transactions of the American Fisheries Society, 134(5):1285–1298, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sergeant:2006:EPH

[SB06] Christopher J. Sergeant and David A. Beauchamp. Effects of physical habitat and ontogeny on lentic habitat preferences of juvenile Chinook salmon. Transactions of the American Fisheries Society, 135(5):1191–1204, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2008:GHO

[SB08] David L. Smith and Ernest L. Brannon. Growth and habitat occupancy of hatchery Coho salmon in an engineered stream. Transactions of the American Fisheries Society, 137(4):1108–1119, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schlechte:2005:LEP

[SBB05] J. Warren Schlechte, Robert K. Betsill, and David L. Buckmeier. A laboratory evaluation of poststocking predatory losses for cultured largemouth bass. Transactions of the American Fisheries Society, 134(1):141–148, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stead:2005:SPJ

[SBC^{+}05] M. A. Stead, D. M. Baltz, E. J. Chesney, M. A. Tarr, A. S. Kolok, and B. D. Marx. Swimming performance of juvenile Florida pompano after sublethal exposure to ethylene glycol and methanol: Synergistic effects. Transactions of the American Fisheries Society, 134(6):1438–1447, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Satterthwaite:2009:SLH

[SBC^{+}09] William H. Satterthwaite, Michael P. Beakes, Erin M. Collins, David R. Swank, Joseph E. Merz, Robert G. Titus, Susan M. Sogard, and Marc Mangel. Steelhead life history on California’s Central Coast: Insights from a state-dependent model. Transactions of the American Fisheries Society, 138(3):532–548, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sammons:2001:ELH

[SBG01a] Steve M. Sammons, Phillip W. Bettoli, and Veronica A. Greear. Early life history characteristics of age-0 white crappies in response to hydrology and zooplankton densities in Normandy Reservoir, Tennessee. Transactions of the American Fisheries Society, 130(3):442–449, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schrank:2001:SVD

[SBG01b] Sally J. Schrank, Patrick J. Braaten, and Christopher S. Guy. Spatiotemporal variation in density of larval bighead carp in the Lower Missouri River. Transactions of the American Fisheries Society, 130(5):809–814, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sutton:2000:RAF

[SBH00] Stephen G. Sutton, Tammo P. Bult, and Richard L. Haedrich. Relationships among fat weight, body weight, water weight, and condition factors in wild Atlantic salmon parr. Transactions of the American Fisheries Society, 129(2):527–538, March 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Szalanski:2000:PGS

[SBM00] Allen L. Szalanski, Richard Bischof, and Gerald Mestl. Population genetic structure of Nebraska paddlefish based on mitochondrial DNA variation. Transactions of the American Fisheries Society, 129(4):1060–1065, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2005:RJR

[SBO05] David L. Smith, Ernest L. Brannon, and Mufeed Odeh. Response of juvenile rainbow trout to turbulence produced by prismatoidal shapes. Transactions of the American Fisheries Society, 134(3): 741–753, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serrano:2008:TSB

[SBP^{+}08] Ignacio Serrano, Ishi Buffam, Daniel Palm, Eva Brännäs, and Hjalmar Laudon. Thresholds for survival of brown trout during the spring flood acid pulse in streams high in dissolved organic carbon. Transactions of the American Fisheries Society, 137(5): 1363–1377, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scharf:2006:EVP

[SBR^{+}06] Frederick S. Scharf, Jeffrey A. Buckel, Kenneth A. Rose, Francis Juanes, and James H. Cowan, Jr. Effects of variable prey and cohort dynamics on growth of young-of-the-year estuarine bluefish: Evidence for interactions between spring- and summer-spawned cohorts. Transactions of the American Fisheries Society, 135(5): 1266–1289, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2006:UAF

[SBSO06] David L. Smith, Ernest L. Brannon, Bahman Shafii, and Mufeed Odeh. Use of the average and fluctuating velocity components for estimation of volitional rainbow trout density. Transactions of the American Fisheries Society, 135(2):431–441, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schisler:2000:EMS

[SBW00] George J. Schisler, Eric P. Bergersen, and Peter G. Walker. Effects of multiple stressors on morbidity and mortality of fingerling rainbow trout infected with Myxobolus cerebralis. Transactions of the American Fisheries Society, 129(3):859–865, May 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sipe:2002:CCS

[SC02] Ann M. Sipe and Mark E. Chittenden, Jr. A comparison of calcified structures for aging bluefish in the Chesapeake Bay region. Transactions of the American Fisheries Society, 131(4):783–790, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoenebeck:2008:IEB

[SCB08] Casey W. Schoenebeck, Steven R. Chipps, and Michael L. Brown. Improvement of an esocid bioenergetics model for juvenile fish. Transactions of the American Fisheries Society, 137(6):1891– 1897, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scharf:2000:PAG

[Sch00a] Frederick S. Scharf. Patterns in abundance, growth, and mortality of juvenile red drum across estuaries on the Texas Coast with implications for recruitment and stock enhancement. Transactions of the American Fisheries Society, 129(6):1207–1222, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schleiger:2000:UIB

[Sch00b] Steve L. Schleiger. Use of an index of biotic integrity to detect effects of land uses on stream fish communities in West-Central Georgia. Transactions of the American Fisheries Society, 129(5): 1118–1133, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scheerer:2002:IFI

[Sch02] Paul D. Scheerer. Implications of floodplain isolation and connectivity on the conservation of an endangered minnow, Oregon chub, in the Willamette River, Oregon. Transactions of the American Fisheries Society, 131(6):1070–1080, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schuster:2003:BFC

[Sch03a] Guenter A. Schuster. Biology of freshwater crayfish. Transactions of the American Fisheries Society, 132(2):406–407, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2003:EPR

[SCH03b] Daniel E. Shoup, Robert E. Carlson, and Robert T. Heath. Effects of predation risk and foraging return on the diel use of vegetated habitat by two size-classes of bluegills. Transactions of the American Fisheries Society, 132(3):590–597, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2004:DAL

[SCH04a] Daniel E. Shoup, Robert E. Carlson, and Robert T. Heath. Diel activity levels of centrarchid fishes in a small Ohio Lake. Transactions of the American Fisheries Society, 133(5):1264– 1269, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sweka:2004:GER

[SCH04b] John A. Sweka, M. Keith Cox, and Kyle J. Hartman. Gastric evacuation rates of brook trout. Transactions of the American Fisheries Society, 133(1):204–210, January 2004. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scheuerell:2005:IJS

[Sch05] Mark D. Scheuerell. Influence of juvenile size on the age at maturity of individually marked wild Chinook salmon. Transactions of the American Fisheries Society, 134(4):999–1004, July 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreer:2001:CRV

[SCM01] Jason F. Schreer, Steven J. Cooke, and R. S. McKinley. Cardiac response to variable forced exercise at different temperatures: an angling simulation for smallmouth bass. Transactions of the American Fisheries Society, 130(5):783–795, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Studebaker:2009:RHS

[SCM09] Rebecca S. Studebaker, Karah N. Cox, and Timothy J. Mulligan. Recent and historical spatial distributions of juvenile rockfish species in rocky intertidal tide pools, with emphasis on black rockfish. Transactions of the American Fisheries Society, 138(3): 645–651, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sommer:2002:SRS

[SCO^{+}02] Ted R. Sommer, Louise Conrad, Gavin O’Leary, Frederick Feyrer, and William C. Harrell. Spawning and rearing of splittail in a model floodplain wetland. Transactions of the American Fisheries Society, 131(5):966–974, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simonin:2005:ASF

[SCP^{+}05] Howard A. Simonin, James R. Colquhoun, Eric A. Paul, John Symula, and Howard J. Dean. Have Adirondack stream fish populations changed in response to decreases in sulfate deposition? Transactions of the American Fisheries Society, 134(2):338–345, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siepker:2009:IRS

[SCWP09] Michael J. Siepker, Steven J. Cooke, David H. Wahl, and David P. Philipp. Individual reproductive success of largemouth bass and smallmouth bass subjected to different components of competitive angling events. Transactions of the American Fisheries Society, 138(4):818–825, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stelzer:2008:CSL

[SDSB08] Robert S. Stelzer, H. Gene Drecktrah, Michael P. Shupryt, and Ronald M. Bruch. Carbon sources for lake sturgeon in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 137(4):1018–1028, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Secor:2008:ISS

[Sec08] David H. Secor. Influence of skipped spawning and misspecified reproductive schedules on biological reference points in sustainable fisheries. Transactions of the American Fisheries Society, 137(3):782–789, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickler:2008:SGR

[SEP^{+}08] Morten Stickler, Eva C. Enders, Curtis J. Pennell, David Cote, Knut Alfredsen, and David A. Scruton. Stream gradient-related movement and growth of Atlantic salmon parr during winter. Transactions of the American Fisheries Society, 137(2):371–385, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saunders:2007:IGM

[SF07] W. Carl Saunders and Kurt D. Fausch. Improved grazing management increases terrestrial invertebrate inputs that feed trout in Wyoming rangeland streams. Transactions of the American Fisheries Society, 136(5):1216–1230, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scheurer:2003:MPR

[SFB03] Julie A. Scheurer, Kurt D. Fausch, and Kevin R. Bestgen. Multiscale processes regulate brassy minnow persistence in a Great Plains river. Transactions of the American Fisheries Society, 132 (5):840–855, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stefan:2001:SFH

[SFE01] Heinz G. Stefan, Xing Fang, and John G. Eaton. Simulated fish habitat changes in North American lakes in response to projected climate warming. Transactions of the American Fisheries Society, 130(3):459–477, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stein:2004:ASM

[SFS04] Andrew B. Stein, Kevin D. Friedland, and Michael Sutherland. Atlantic sturgeon marine distribution and habitat use along the Northeastern Coast of the United States. Transactions of the American Fisheries Society, 133(3):527–537, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shervette:2007:HSG

[SG07] Virginia R. Shervette and Frances Gelwick. Habitat-specific growth in juvenile pinfish. Transactions of the American Fisheries Society, 136(2):445–451, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schrank:2003:CIB

[SGF03] Sally J. Schrank, Christopher S. Guy, and James F. Fairchild. Competitive interactions between age-0 bighead carp and paddlefish. Transactions of the American Fisheries Society, 132(6): 1222–1228, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strakosh:2009:EAW

[SGG09] Timothy R. Strakosh, Keith B. Gido, and Christopher S. Guy. Effects of American water willow establishment on density, growth, diet, and condition of age-0 largemouth bass in Kansas reservoirs. Transactions of the American Fisheries Society, 138(2):269–279, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2005:RHT

[SGS05] Christian T. Smith, W. Stewart Grant, and Lisa W. Seeb. A rapid, high-throughput technique for detecting Tanner crabs Chionoecetes bairdi illegally taken in Alaska’s snow crab fishery. Transactions of the American Fisheries Society, 134(3):620–623, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sargent:2006:DRS

[SGS06a] Philip S. Sargent, Robert S. Gregory, and David C. Schneider. Density responses of subarctic coastal marine fish and crabs to artificial reefs. Transactions of the American Fisheries Society, 135(2):348–360, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [SGS06b].

Sargent:2006:EDR

[SGS06b] Philip S. Sargent, Robert S. Gregory, and David C. Schneider. Erratum: Density Responses of Subarctic Coastal Marine Fish and Crabs to Artificial Reefs. Transactions of the American Fisheries Society, 135(2):e3, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [SGS06a].

Salek:2001:CTS

[SGSS01] Stephen J. Salek, John Godwin, Craig V. Sullivan, and Norman E. Stacey. Courtship and tank spawning behavior of temperate basses (genus Morone). Transactions of the American Fisheries Society, 130(5):833–847, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simpkins:2000:DIS

[SH00a] Darin G. Simpkins and Wayne A. Hubert. Drifting invertebrates, stomach contents, and body conditions of juvenile rainbow trout from fall through winter in a Wyoming tailwater. Transactions of the American Fisheries Society, 129(5):1187–1195, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sutela:2000:VRZ

[SH00b] T. Sutela and A. Huusko. Varying resistance of zooplankton prey to digestion: Implications for quantifying larval fish diets. Transactions of the American Fisheries Society, 129(2):545–551, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sweka:2001:ITB

[SH01] John A. Sweka and Kyle J. Hartman. Influence of turbidity on brook trout reactive distance and foraging success. Transactions of the American Fisheries Society, 130(1):138–146, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sitar:2006:GMH

[SH06] Shawn P. Sitar and Ji X. He. Growth and maturity of hatchery and wild lean lake trout during population recovery in Michigan waters of Lake Superior. Transactions of the American Fisheries Society, 135(4):915–923, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sweka:2008:CTI

[SH08] John A. Sweka and Kyle J. Hartman. Contribution of terrestrial invertebrates to yearly brook trout prey consumption and growth. Transactions of the American Fisheries Society, 137(1):224–235, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schleuter:2007:RRE

[SHBFE07] Diana Schleuter, Susanne Haertel-Borer, Philipp Fischer, and Reiner Eckmann. Respiration rates of Eurasian perch Perca fluviatilis and ruffe: Lower energy costs in groups. Transactions of the American Fisheries Society, 136(1):43–55, January 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schulz:2001:TIB

[SHC01] Eric J. Schulz, Mark V. Hoyer, and Daniel E. Canfield, Jr. Test of an index of biotic integrity: Response to comment. Transactions of the American Fisheries Society, 130(1):172–173, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [WHP01].

Simpkins:2003:PRJ

[SHDR03] Darin G. Simpkins, Wayne A. Hubert, Carlos Martinez Del Rio, and Daniel C. Rule. Physiological responses of juvenile rainbow trout to fasting and swimming activity: Effects on body composition and condition indices. Transactions of the American Fisheries Society, 132(3):576–589, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stafford:2004:MFT

[SHS04] Craig P. Stafford, Barry Hansen, and Jack A. Stanford. Mercury in fishes and their diet items from Flathead Lake, Montana. Transactions of the American Fisheries Society, 133(2):349–357, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seeb:2000:GDS

[SHT^{+}00] Lisa W. Seeb, Chris Habicht, William D. Templin, Kenneth E. Tarbox, Randall Z. Davis, Linda K. Brannian, and James E. Seeb. Genetic diversity of sockeye salmon of Cook Inlet, Alaska, and its application to management of populations affected by the Exxon Valdez oil spill. Transactions of the American Fisheries Society, 129(6):1223–1249, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simpkins:2000:EFW

[SHW00] Darin G. Simpkins, Wayne A. Hubert, and Thomas A. Wesche. Effects of fall-to-winter changes in habitat and frazil ice on the movements and habitat use of juvenile rainbow trout in a Wyoming tailwater. Transactions of the American Fisheries Society, 129(1):101–118, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2005:DMI

[SHWS05] Chris D. Smith, Christopher L. Higgins, Gene R. Wilde, and Richard E. Strauss. Development of a morphological index of the nutritional status of juvenile largemouth bass. Transactions of the American Fisheries Society, 134(1):120–125, January 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaffler:2002:HUS

[SIH02] Jason J. Schaffler, J. Jeffery Isely, and Willard E. Hayes. Habitat use by striped bass in relation to seasonal changes in water quality in a southern reservoir. Transactions of the American Fisheries Society, 131(5):817–827, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stoneman:2000:IHF

[SJ00] C. L. Stoneman and M. L. Jones. The influence of habitat features on the biomass and distribution of three species of Southern Ontario stream salmonines. Transactions of the American Fisheries Society, 129(3):639–657, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shephard:2004:SSC

[SJ04] Samuel Shephard and Donald C. Jackson. Size selection of channel catfish in slat traps of different interslat space widths. Transactions of the American Fisheries Society, 133(1):197–203, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shephard:2006:DCC

[SJ06] Samuel Shephard and Donald C. Jackson. Difference in channel catfish growth among Mississippi stream basins. Transactions of the American Fisheries Society, 135(5):1224–1229, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2008:ASE

[SJ08] Katherine L. Smith and Michael L. Jones. Allocation of sampling effort to optimize efficiency of watershed-level ichthyofaunal inventories. Transactions of the American Fisheries Society, 137 (5):1500–1506, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sloss:2008:GIB

[SJFP08] Brian L. Sloss, Martin J. Jennings, Ryan Franckowiak, and Dennis M. Pratt. Genetic identity of brook trout in Lake Superior south shore streams: Potential for genetic monitoring of stocking and rehabilitation efforts. Transactions of the American Fisheries Society, 137(4):1244–1251, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sakaris:2005:BBI

[SJP05] Peter C. Sakaris, Roman V. Jesien, and Alfred E. Pinkney. Brown bullhead as an indicator species: Seasonal movement patterns and home ranges within the Anacostia River, Washington, D.C. Transactions of the American Fisheries Society, 134(5):1262– 1270, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sanchez-Jerez:2000:CFA

[SJRE00] Pablo Sánchez-Jerez and Alfonso Ramos-Esplá. Changes in fish assemblages associated with the deployment of an antitrawling reef in seagrass meadows. Transactions of the American Fisheries Society, 129(5):1150–1159, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sykes:2009:TFE

[SJS09] Gregory E. Sykes, Chris J. Johnson, and J. Mark Shrimpton. Temperature and flow effects on migration timing of Chinook salmon smolts. Transactions of the American Fisheries Society, 138(6):1252–1265, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2005:MHU

[SK05a] Kregg M. Smith and Donna K. King. Movement and habitat use of yearling and juvenile lake sturgeon in Black Lake, Michigan. Transactions of the American Fisheries Society, 134(5): 1159–1172, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2005:SFA

[SK05b] Tamara A. Smith and Clifford E. Kraft. Stream fish assemblages in relation to landscape position and local habitat variables. Transactions of the American Fisheries Society, 134(2): 430–440, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:2007:ELB

[SK07] Jay R. Stauffer, Jr. and Patrick M. Kocovsky. Exploring links between systematics and fisheries management. Transactions of the American Fisheries Society, 136(4):1122–1125, July 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stahl:2008:STV

[SK08] Jennifer P. Stahl and Gordon H. Kruse. Spatial and temporal variability in size at maturity of walleye pollock in the Eastern Bering Sea. Transactions of the American Fisheries Society, 137 (5):1543–1557, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schwindt:2009:RAT

[SKA^{+}09] Adam R. Schwindt, Michael L. Kent, Luke K. Ackerman, Staci L. Massey Simonich, Dixon H. Landers, Tamara Blett, and Carl B. Schreck. Reproductive abnormalities in trout from Western U.S. national parks. Transactions of the American Fisheries Society, 138(3):522–531, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Suski:2004:PSW

[SKC^{+}04] Cory D. Suski, Shaun S. Killen, Steven J. Cooke, James D. Kieffer, David P. Philipp, and Bruce L. Tufts. Physiological significance of the weigh-in during live-release angling tournaments for largemouth bass. Transactions of the American Fisheries Society, 133(6):1291–1303, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sheehan:2005:MGM

[SKC^{+}05] Timothy F. Sheehan, John F. Kocik, Steven X. Cadrin, Christopher M. Legault, Ernest Atkinson, and David Bengtson. Marine growth and morphometrics for three populations of Atlantic salmon from Eastern Maine, USA. Transactions of the American Fisheries Society, 134(3):775–788, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schooley:2008:DLR

[SKK^{+}08] Jason D. Schooley, Abraham P. Karam, Brian R. Kesner, Paul C. Marsh, Carol A. Pacey, and Darren J. Thornbrugh. Detection of larval remains after consumption by fishes. Transactions of the American Fisheries Society, 137(4):1044–1049, July 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spidle:2003:PSA

[SKL^{+}03] A. P. Spidle, S. T. Kalinowski, B. A. Lubinski, D. L. Perkins, K. F. Beland, J. F. Kocik, and T. L. King. Population structure of Atlantic salmon in Maine with reference to populations from Atlantic Canada. Transactions of the American Fisheries Society, 132(2):196–209, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schroder:2008:BSW

[SKP^{+}08] Steven L. Schroder, Curtis M. Knudsen, Todd N. Pearsons, Todd W. Kassler, Sewall F. Young, Craig A. Busack, and David E. Fast. Breeding success of wild and first-generation hatchery female spring Chinook salmon spawning in an artificial stream. Transactions of the American Fisheries Society, 137(5):1475– 1489, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stunz:2000:AGS

[SLC00] Gregory W. Stunz, Thomas L. Linton, and Robert L. Colura. Age and growth of southern flounder in Texas waters, with emphasis on Matagorda Bay. Transactions of the American Fisheries Society, 129(1):119–125, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sigourney:2006:IBA

[SLC06] Douglas B. Sigourney, Benjamin H. Letcher, and Richard A. Cunjak. Influence of beaver activity on summer growth and condition of age-2 Atlantic salmon parr. Transactions of the American Fisheries Society, 135(4):1068–1075, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Small:2005:TSG

[SLF^{+}05] Maureen P. Small, Janet L. Loxterman, Alice E. Frye, Jennifer F. Von Bargen, Cherril Bowman, and Sewall F. Young. Temporal and spatial genetic structure among some Pacific herring populations in Puget Sound and the Southern Strait of Georgia. Transactions of the American Fisheries Society, 134(5):1329– 1341, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Skalski:2008:GSC

[SLGH08] Garrick T. Skalski, Jacob B. Landis, Michael J. Grose, and Stephen P. Hudman. Genetic structure of creek chub, a headwater minnow, in an impounded river system. Transactions of the American Fisheries Society, 137(4):962–975, July 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schroeder:2001:OSH

[SLK01] R. Kirk Schroeder, Robert B. Lindsay, and Ken R. Kenaston. Origin and straying of hatchery winter steelhead in Oregon coastal rivers. Transactions of the American Fisheries Society, 130(3): 431–441, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simonin:2007:BEG

[SLM07] Paul W. Simonin, Karin E. Limburg, and Leonard S. Machut. Bridging the energy gap: Anadromous blueback herring feeding in the Hudson and Mohawk Rivers, New York. Transactions of the American Fisheries Society, 136(6):1614–1621, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stokesbury:2001:ISA

[SLP^{+}01] Michael J. W. Stokesbury, Gilles L. Lacroix, E. Lisa Price, Derek Knox, and Michael J. Dadswell. Identification by scale analysis of farmed Atlantic salmon juveniles in Southwestern New Brunswick rivers. Transactions of the American Fisheries Society, 130(5): 815–822, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Selbie:2007:LTP

[SLSF07] Daniel T. Selbie, Bert A. Lewis, John P. Smol, and Bruce P. Finney. Long-term population dynamics of the endangered Snake River sockeye salmon: Evidence of past influences on stock decline and impediments to recovery. Transactions of the American Fisheries Society, 136(3):800–821, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2007:PMR

[SM07] Stephen J. Smith and J. Ellen Marsden. Predictive morphometric relationships for estimating fecundity of sea lampreys from Lake Champlain and other landlocked populations. Transactions of the American Fisheries Society, 136(4):979–987, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schwartz:2008:GEI

[SM08a] Rachel S. Schwartz and Bernie May. Genetic evaluation of isolated populations for use in reintroductions reveals significant genetic bottlenecks in potential stocks of Sacramento perch. Transactions of the American Fisheries Society, 137(6):1764–1777, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scott:2008:SPF

[SM08b] Mandy K. Scott and Daniel D. Magoulick. Swimming performance of five warmwater stream fish species. Transactions of the American Fisheries Society, 137(1):209–215, January 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saillant:2003:GVR

[SMG03] Eric Saillant, Timothy A. Mousseau, and John R. Gold. Genetic variation and relatedness of juvenile red snapper sampled from shrimp trawls in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 132(6):1229–1235, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:2007:TPU

[SMK^{+}07] Jay Richard Stauffer, Jr., Henry Madsen, Adrianus Konings, Paul Bloch, Cecilia Paola Ferreri, Jeremy Likongwe, Kenneth R. McKaye, and Kristin E. Black. Taxonomy: a precursor to understanding ecological interactions among schistosomes, snail hosts, and snail-eating fishes. Transactions of the American Fisheries Society, 136(4):1136–1145, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Small:2007:FSP

[SML^{+}07] Maureen P. Small, Jason G. McLellan, Janet Loxterman, Jennifer Von Bargen, Alice Frye, and Cherril Bowman. Fine-scale population structure of rainbow trout in the Spokane River drainage in relation to hatchery stocking and barriers. Transactions of the American Fisheries Society, 136(2):301–317, March 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steinhart:2004:RGP

[SMS04] Geoffrey B. Steinhart, Elizabeth A. Marschall, and Roy A. Stein. Round goby predation on smallmouth bass offspring in nests during simulated catch-and-release angling. Transactions of the American Fisheries Society, 133(1):121–131, January 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stapanian:2006:ESL

[SMW06] Martin A. Stapanian, Charles P. Madenjian, and Larry D. Witzel. Evidence that sea lamprey control led to recovery of the burbot population in Lake Erie. Transactions of the American Fisheries Society, 135(4):1033–1043, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Selong:2001:ETG

[SMZB01] Jason H. Selong, Thomas E. McMahon, Alexander V. Zale, and Frederic T. Barrows. Effect of temperature on growth and survival of bull trout, with application of an improved method for determining thermal tolerance in fishes. Transactions of the American Fisheries Society, 130(6):1026–1037, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [Ano05].

Selong:2005:EHU

[SMZB05] Jason H. Selong, Thomas E. McMahon, Alexander V. Zale, and Frederic T. Barrows. Erratum: Habitat Use by Middle Mississippi River Pallid Sturgeon. Transactions of the American Fisheries Society, 134(2):548, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [HSH^{+}04].

Sutton:2001:SDM

[SN01] Trent M. Sutton and John J. Ney. Size-dependent mechanisms influencing first-year growth and winter survival of stocked striped bass in a Virginia mainstream reservoir. Transactions of the American Fisheries Society, 130(1):1–17, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savoy:2003:MIH

[SP03] Thomas Savoy and Deborah Pacileo. Movements and important habitats of subadult Atlantic sturgeon in Connecticut waters. Transactions of the American Fisheries Society, 132(1):1–8, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Suski:2004:FAV

[SP04] Cory D. Suski and David P. Philipp. Factors affecting the vulnerability to angling of nesting male largemouth and smallmouth bass. Transactions of the American Fisheries Society, 133(5):

1100–1106, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shea:2007:ERI

[SP07] Colin P. Shea and James T. Peterson. An evaluation of the relative influence of habitat complexity and habitat stability on fish assemblage structure in unregulated and regulated reaches of a large southeastern warmwater stream. Transactions of the American Fisheries Society, 136(4):943–958, July 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spina:2001:IDA

[Spi01] Anthony P. Spina. Incubation discharge and aspects of brown trout population dynamics. Transactions of the American Fisheries Society, 130(2):322–327, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scoppettone:2007:EPI

[SR07] G. Gary Scoppettone and Peter H. Rissler. Effects of population increase on cui-ui growth and maturation. Transactions of the American Fisheries Society, 136(2):331–340, March 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scoppettone:2000:RLF

[SRB00] G. Gary Scoppettone, Peter H. Rissler, and Mark E. Buettner. Reproductive longevity and fecundity associated with nonannual spawning in cui-ui. Transactions of the American Fisheries Society, 129(3):658–669, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swenson:2007:VED

[SRH07] Erica A. Swenson, Amanda E. Rosenberger, and Philip J. Howell. Validation of endoscopy for determination of maturity in small salmonids and sex of mature individuals. Transactions of the American Fisheries Society, 136(4):994–998, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schrank:2003:ELD

[SRJ03] Amy J. Schrank, Frank J. Rahel, and Helene C. Johnstone. Evaluating laboratory-derived thermal criteria in the field: an example involving Bonneville cutthroat trout. Transactions of the American Fisheries Society, 132(1):100–109, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarnecchia:2006:LTP

[SRL^{+}06] Dennis L. Scarnecchia, L. Fred Ryckman, Youngtaik Lim, Greg Power, Brad Schmitz, and Vic Riggs. A long-term program for validation and verification of dentaries for age estimation in the Yellowstone–Sakakawea paddlefish stock. Transactions of the American Fisheries Society, 135(4):1086–1094, July 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Szedlmayer:2005:LTR

[SS05] Stephen T. Szedlmayer and Rebecca L. Schroepfer. Long-term residence of red snapper on artificial reefs in the Northeastern Gulf of Mexico. Transactions of the American Fisheries Society, 134(2):315–325, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [HSH^{+}05].

Sheer:2006:LWB

[SS06] M. B. Sheer and E. A. Steel. Lost watersheds: Barriers, aquatic habitat connectivity, and salmon persistence in the willamette and Lower Columbia River Basins. Transactions of the American Fisheries Society, 135(6):1654–1669, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2008:NAP

[SS08a] Christian T. Smith and Lisa W. Seeb. Number of alleles as a predictor of the relative assignment accuracy of Short Tandem Repeat (STR) and Single-Nucleotide-Polymorphism (SNP) baselines for chum salmon. Transactions of the American Fisheries Society, 137(3):751–762, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stewart:2008:ERG

[SS08b] Christopher B. Stewart and Frederick S. Scharf. Estuarine recruitment, growth, and first-year survival of juvenile red drum in North Carolina. Transactions of the American Fisheries Society, 137(4):1089–1103, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sprowles:2006:FST

[SSCM06] A. E. Sprowles, M. R. Stephens, N. W. Clipperton, and B. P. May. Fishing for SNPs: a targeted locus approach for single nucleotide polymorphism discovery in rainbow trout. Transactions of the American Fisheries Society, 135(6):1698–1721, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreck:2006:MEJ

[SSD^{+}06] Carl B. Schreck, Thomas P. Stahl, Lawrence E. Davis, Daniel D. Roby, and Benjamin J. Clemens. Mortality estimates of juvenile spring–summer Chinook salmon in the Lower Columbia River and estuary, 1992–1998: Evidence for delayed mortality? Transactions of the American Fisheries Society, 135(2):457–475, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stafford:2002:CLT

[SSHB02] Craig P. Stafford, Jack A. Stanford, F. Richard Hauer, and Edward B. Brothers. Changes in lake trout growth associated with Mysis relicta establishment: a retrospective analysis using otoliths. Transactions of the American Fisheries Society, 131(5): 994–1003, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Suski:2003:ECR

[SSL^{+}03] C. D. Suski, J. H. Svec, J. B. Ludden, F. J. S. Phelan, and D. P. Philipp. The effect of catch-and-release angling on the parental care behavior of male smallmouth bass. Transactions of the American Fisheries Society, 132(2):210–218, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sturdevant:2009:SPJ

[SSO09] M. V. Sturdevant, M. F. Sigler, and J. A. Orsi. Sablefish predation on juvenile Pacific salmon in the coastal marine waters of Southeast Alaska in 1999. Transactions of the American Fisheries Society, 138(3):675–691, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2006:UOM

[SSR06] Nathan G. Smith, Patrick J. Sullivan, and Lars G. Rudstam. Using otolith microstructure to determine natal origin of Lake Ontario Chinook salmon. Transactions of the American Fisheries Society, 135(4):908–914, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spoelstra:2008:MRS

[SSRM08] Jeffrey A. Spoelstra, Roy A. Stein, J. Andrew Royle, and Elizabeth A. Marschall. Movement of reservoir-stocked riverine fish between tailwaters and rivers. Transactions of the American Fisheries Society, 137(5):1530–1542, September 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2005:UNR

[SSSS05] Christian T. Smith, James E. Seeb, Piper Schwenke, and Lisa W. Seeb. Use of the 5’-nuclease reaction for single nucleotide polymorphism genotyping in Chinook salmon. Transactions of the American Fisheries Society, 134(1):207–217, January 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sanderson:2009:NSF

[STC^{+}09] Beth L. Sanderson, Chau D. Tran, Holly J. Coe, Vija Pelekis, E. Ashley Steel, and William L. Reichert. Nonlethal sampling of fish caudal fins yields valuable stable isotope data for threatened and endangered fishes. Transactions of the American Fisheries Society, 138(5):1166–1177, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stokesbury:2002:ESS

[Sto02] Kevin D. E. Stokesbury. Estimation of sea scallop abundance in closed areas of Georges Bank, USA. Transactions of the American Fisheries Society, 131(6):1081–1092, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stetter:2007:IPC

[STR^{+}07] Sandra L. Parker Stetter, Jennifer L. Stritzel Thomson, Lars G. Rudstam, Donna L. Parrish, and Patrick J. Sullivan. Importance and predictability of cannibalism in rainbow smelt. Transactions of the American Fisheries Society, 136(1):227–237, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sutton:2001:BR

[Sut01] Trent M. Sutton. Book review. Transactions of the American Fisheries Society, 130(1):174–175, January 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sutherland:2007:EIS

[Sut07] Andrew B. Sutherland. Effects of increased suspended sediment on the reproductive success of an upland crevice-spawning minnow. Transactions of the American Fisheries Society, 136(2):416– 422, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaus:2002:BDD

[SVW02] Maynard H. Schaus, Michael J. Vanni, and Thomas E. Wissing. Biomass-dependent diet shifts in omnivorous gizzard shad: Implications for growth, food Web, and ecosystem effects. Transactions of the American Fisheries Society, 131(1):40–54, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Santucci:2003:EGP

[SW03a] Victor J. Santucci, Jr. and David H. Wahl. The effects of growth, predation, and first-winter mortality on recruitment of bluegill cohorts. Transactions of the American Fisheries Society, 132(2): 346–360, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steinhart:2003:WEK

[SW03b] Geoffrey B. Steinhart and Wayne A. Wurtsbaugh. Winter ecology of Kokanee: Implications for salmon management. Transactions of the American Fisheries Society, 132(6):1076–1088, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaffler:2008:TSV

[SW08a] Jason J. Schaffler and Dana L. Winkelman. Temporal and spatial variability in otolith trace-element signatures of juvenile striped bass from spawning locations in Lake Texoma, Oklahoma–Texas. Transactions of the American Fisheries Society, 137(3):818–829, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2008:ELB

[SW08b] Daniel E. Shoup and David H. Wahl. The effect of largemouth bass predation on overwinter survival of two size-classes of age-0 bluegills. Transactions of the American Fisheries Society, 137(4):

1063–1071, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2009:ETP

[SW09a] Daniel E. Shoup and David H. Wahl. The effects of turbidity on prey selection by piscivorous largemouth bass. Transactions of the American Fisheries Society, 138(5):1018–1027, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shoup:2009:FDA

[SW09b] Daniel E. Shoup and David H. Wahl. Fish diversity and abundance in relation to interannual and lake-specific variation in abiotic characteristics of floodplain lakes of the Lower Kaskaskia River, Illinois. Transactions of the American Fisheries Society, 138(5):1076–1092, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Slaughter:2004:EAB

[SWD04] Joe E. Slaughter IV, Russell A. Wright, and Dennis R. DeVries. The effects of age-0 body size on the predictive ability of a largemouth bass bioenergetics model. Transactions of the American Fisheries Society, 133(2):279–291, March 2004. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sogard:2009:SPA

[SWF09] Susan M. Sogard, Thomas H. Williams, and Heidi Fish. Seasonal patterns of abundance, growth, and site fidelity of juvenile steelhead in a small coastal California stream. Transactions of the American Fisheries Society, 138(3):549–563, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simons:2001:PSB

[SWH^{+}01] Andrew M. Simons, Robert M. Wood, Lucie S. Heath, Bernard R. Kuhajda, and Richard L. Mayden. Phylogenetics of Scaphirhynchus based on mitochondrial DNA sequences. Transactions of the American Fisheries Society, 130(3):359–366, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelton:2006:EAD

[SWHW06] Andrew O. Shelton, Douglas A. Woodby, Kyle Hebert, and Jon D. Witman. Evaluating age determination and spatial patterns of growth in red sea urchins in Southeast Alaska. Transactions of the American Fisheries Society, 135(6):1670–1680, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swanson:2004:STV

[SYC04] Christina Swanson, Paciencia S. Young, and Joseph J. Cech, Jr. Swimming in two-vector flows: Performance and behavior of juvenile Chinook salmon near a simulated screened water diversion. Transactions of the American Fisheries Society, 133(2):265–278, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swanson:2005:CEF

[SYC05] Christina Swanson, Paciencia S. Young, and Joseph J. Cech, Jr. Close encounters with a fish screen: Integrating physiological and behavioral results to protect endangered species in exploited ecosystems. Transactions of the American Fisheries Society, 134 (5):1111–1123, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shrimpton:2001:SRJ

[SZM01] J. Mark Shrimpton, Joseph D. Zydlewski, and Stephen D. Mc-Cormick. The stress response of juvenile American shad to handling and confinement is greater during migration in freshwater than in seawater. Transactions of the American Fisheries Society, 130(6):1203–1210, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steen:2008:CTM

[SZSS08] Paul J. Steen, Troy G. Zorn, Paul W. Seelbach, and Jeffrey S. Schaeffer. Classification tree models for predicting distributions of Michigan stream fish from landscape variables. Transactions of the American Fisheries Society, 137(4):976–996, July 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:2006:CDS

[TA06] David L. Taylor and Kenneth W. Able. Cohort dynamics of summer-spawned bluefish as determined by length-frequency and otolith microstructure analyses. Transactions of the American Fisheries Society, 135(4):955–969, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Timmerman:2000:DFL

[TAB00] Cindy M. Timmerman, Cynthia A. Annett, and Claudia F. Bailey. Determination of factors limiting prey size swallowed by larval and small juvenile largemouth bass. Transactions of the American Fisheries Society, 129(2):618–622, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tuten:2008:EBP

[TAC08] Travis Tuten, Micheal Allen, and Chuck Cichra. Effects of benthic prey composition and abundance on diet and growth of black crappies in three Florida lakes. Transactions of the American Fisheries Society, 137(6):1778–1790, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tymchuk:2005:CAM

[TAD05] W. E. Vandersteen Tymchuk, M. V. Abrahams, and R. H. Devlin. Competitive ability and mortality of growth-enhanced transgenic Coho salmon fry and parr when foraging for food. Transactions of the American Fisheries Society, 134(2):381–389, March 2005. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trippel:2007:STA

[TAM07] Nicholas A. Trippel, Micheal S. Allen, and Richard S. McBride. Seasonal trends in abundance and size of juvenile American shad, hickory shad, and blueback herring in the St. Johns River, Florida, and comparison with historical data. Transactions of the American Fisheries Society, 136(4):988–993, July 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyler:2008:IVB

[TB08] Jeffrey A. Tyler and Melanie B. Bolduc. Individual variation in bioenergetic rates of young-of-year rainbow trout. Transactions of the American Fisheries Society, 137(1):314–323, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tonn:2004:FAS

[TBA^{+}04] William M. Tonn, Shelly M. Boss, Peter K. M. Aku, Garry J. Scrimgeour, and Cynthia A. Paszkowski. Fish assemblages in subarctic lakes: Does fire affect fish–environment relations in Northern Alberta? Transactions of the American Fisheries Society, 133 (1):132–143, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teel:2009:GSC

[TBK^{+}09] David J. Teel, Cyndi Baker, David R. Kuligowski, Thomas A. Friesen, and Barbara Shields. Genetic stock composition of subyearling Chinook salmon in seasonal floodplain wetlands of the Lower Willamette River, Oregon. Transactions of the American Fisheries Society, 138(1):211–217, January 2009. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tymchuk:2006:GBC

[TBWD06] Wendy E. Tymchuk, Carlo Biagi, Ruth Withler, and Robert H. Devlin. Growth and behavioral consequences of introgression of a domesticated aquaculture genotype into a native strain of Coho salmon. Transactions of the American Fisheries Society, 135(2): 442–455, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tripp:2009:DRG

[TCG09] Sara J. Tripp, Robert E. Colombo, and James E. Garvey. Declining recruitment and growth of shovelnose sturgeon in the Middle Mississippi River: Implications for conservation. Transactions of the American Fisheries Society, 138(2):416–422, March 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

TeWinkel:2001:VMB

[TF01] Leslie M. TeWinkel and Guy W. Fleischer. Vertical migration of bloaters: Response to comment. Transactions of the American Fisheries Society, 130(1):167–169, January 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [EBC01].

Tallman:2007:OGC

[TF07] Jessica C. Tallman and Graham E. Forrester. Oyster grow-out cages function as artificial reefs for temperate fishes. Transactions of the American Fisheries Society, 136(3):790–799, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trudel:2009:DMJ

[TFO^{+}09] M. Trudel, J. Fisher, J. A. Orsi, J. F. T. Morris, M. E. Thiess, R. M. Sweeting, S. Hinton, E. A. Fergusson, and D. W. Welch.

Distribution and migration of juvenile Chinook salmon derived from coded wire tag recoveries along the Continental Shelf of Western North America. Transactions of the American Fisheries Society, 138(6):1369–1391, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trudel:2004:MOC

[TGW04] Marc Trudel, David R. Geist, and David W. Welch. Modeling the oxygen consumption rates in Pacific salmon and steelhead: an assessment of current models and practices. Transactions of the American Fisheries Society, 133(2):326–348, March 2004. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiemann:2004:ELD

[TGWE04] Jeremy S. Tiemann, David P. Gillette, Mark L. Wildhaber, and David R. Edds. Effects of lowhead dams on riffle-dwelling fishes and macroinvertebrates in a Midwestern river. Transactions of the American Fisheries Society, 133(3):705–717, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Timmons:2000:EMP

[TH00] Tom J. Timmons and Tyrone A. Hughbanks. Exploitation and mortality of paddlefish in the Lower Tennessee and Cumberland Rivers. Transactions of the American Fisheries Society, 129(5): 1171–1180, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:2003:HRA

[Tho03] William L. Thompson. Hankin and Reeves’ approach to estimating fish abundance in small streams: Limitations and alternatives. Transactions of the American Fisheries Society, 132(1): 69–75, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thedinga:2008:FAS

[TJNL08] John F. Thedinga, Scott W. Johnson, A. Darcie Neff, and Mandy R. Lindeberg. Fish assemblages in shallow, nearshore habitats of the Bering Sea. Transactions of the American Fisheries Society, 137(4):1157–1164, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyson:2001:RYP

[TK01] Jeff T. Tyson and Roger L. Knight. Response of yellow perch to changes in the benthic invertebrate community of Western Lake Erie. Transactions of the American Fisheries Society, 130(5):766– 782, September 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2009:WVT

[TKH^{+}09] Kenneth F. Tiffan, Tobias J. Kock, Craig A. Haskell, William P. Connor, and R. Kirk Steinhorst. Water velocity, turbulence, and migration rate of subyearling fall Chinook salmon in the freeflowing and impounded Snake River. Transactions of the American Fisheries Society, 138(2):373–384, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tolimieri:2006:ASE

[TL06] Nick Tolimieri and Phillip S. Levin. Assemblage structure of Eastern Pacific groundfishes on the U.S. continental slope in relation to physical and environmental variables. Transactions of the American Fisheries Society, 135(2):317–332, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tranah:2006:PII

[TM06] Gregory J. Tranah and Bernie May. Patterns of intra- and interspecies genetic diversity in Klamath River Basin suckers. Transactions of the American Fisheries Society, 135(2):306–316, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Terwilliger:2003:ABW

[TMK03] Mark R. Terwilliger, Douglas F. Markle, and Jacob Kann. Associations between water quality and daily growth of juvenile shortnose and Lost River suckers in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 132(4):691–708, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teel:2000:GPS

[TMWG00] David J. Teel, George B. Milner, Gary A. Winans, and W. Stewart Grant. Genetic population structure and origin of life history types in Chinook salmon in British Columbia, Canada. Transactions of the American Fisheries Society, 129(1):194–209, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomcko:2001:RBG

[TP01] Cynthia M. Tomcko and Rodney B. Pierce. The relationship of bluegill growth, lake morphometry, and water quality in Minnesota. Transactions of the American Fisheries Society, 130(2): 317–321, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tonn:2003:EFH

[TPS^{+}03] William M. Tonn, Cynthia A. Paszkowski, Garry J. Scrimgeour, Peter K. M. Aku, Marc Lange, Ellie E. Prepas, and Kim Westcott. Effects of forest harvesting and fire on fish assemblagesin boreal plains lakes: a reference condition approach. Transactions of the American Fisheries Society, 132(3):514–523, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyler:2007:RRE

[TR07] Jeffrey A. Tyler and Edward S. Rutherford. River restoration effects on steelhead populations in the Manistee River, Michigan: Analysis using an individual-based model. Transactions of the American Fisheries Society, 136(6):1654–1673, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tuomikoski:2008:EAS

[TRBH08] Jack E. Tuomikoski, Paul J. Rudershausen, Jeffrey A. Buckel, and Joseph E. Hightower. Effects of age-1 striped bass predation on juvenile fish in Western Albemarle Sound. Transactions of the American Fisheries Society, 137(1):324–339, January 2008. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2000:IJF

[TRGV00] Kenneth F. Tiffan, Dennis W. Rondorf, Rodney D. Garland, and Peter A. Verhey. Identification of juvenile fall versus spring Chinook salmon migrating through the Lower Snake River based on body morphology. Transactions of the American Fisheries Society, 129(6):1389–1395, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Troutman:2007:PHU

[TRK07] John P. Troutman, D. Allen Rutherford, and W. E. Kelso. Patterns of habitat use among vegetation-dwelling littoral fishes in the Atchafalaya River Basin, Louisiana. Transactions of the American Fisheries Society, 136(4):1063–1075, July 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffan:2005:DSB

[TRS05] Kenneth F. Tiffan, Dennis W. Rondorf, and Joseph J. Skalicky. Diel spawning behavior of chum salmon in the Columbia River. Transactions of the American Fisheries Society, 134(4):892–900, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tatara:2009:GSH

[TRS09] Christopher P. Tatara, Stephen C. Riley, and Julie A. Scheurer. Growth, survival, and habitat use of naturally reared and hatchery steelhead fry in streams: Effects of an enriched hatchery rearing environment. Transactions of the American Fisheries Society, 138(3):441–457, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:2009:SSS

[TTW^{+}09] S. Tucker, M. Trudel, D. W. Welch, J. R. Candy, J. F. T. Morris, M. E. Thiess, C. Wallace, D. J. Teel, W. Crawford, E. V. Farley, Jr., and T. D. Beacham. Seasonal stock-specific migrations of juvenile sockeye salmon along the West Coast of North America: Implications for growth. Transactions of the American Fisheries Society, 138(6):1458–1480, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:2005:DFE

[Tur05] Bruce J. Turner. Darwin’s fishes: an encyclopedia of ichthyology, ecology and evolution. Transactions of the American Fisheries Society, 134(5):1367–1368, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trudel:2005:MOC

[TW05] Marc Trudel and David W. Welch. Modeling the oxygen consumption rates in Pacific salmon and steelhead: Model development. Transactions of the American Fisheries Society, 134(6):

1542–1561, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tabor:2007:ODS

[TWF^{+}07] Roger A. Tabor, Eric J. Warner, Kurt L. Fresh, Brian A. Footen, and Jeffrey R. Chan. Ontogenetic diet shifts of prickly sculpin in the Lake Washington Basin, Washington. Transactions of the American Fisheries Society, 136(6):1801–1813, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Utz:2007:DRI

[URMH07] Ryan M. Utz, Brett C. Ratcliffe, Brett T. Moore, and Kyle J. Hartman. Disproportionate relative importance of a terrestrial beetle family (Coleoptera: Scarabaeidae) as a prey source for Central Appalachian brook trout. Transactions of the American Fisheries Society, 136(1):177–184, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDenHeuvel:2004:BIA

[Van04] Michael R. Van Den Heuvel. Biological indicators of aquatic ecosystem stress. Transactions of the American Fisheries Society, 133(2):492, March 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDeValk:2002:CAC

[VAR^{+}02] Anthony J. VanDeValk, Connie M. Adams, Lars G. Rudstam, John L. Forney, Thomas E. Brooking, Mindy A. Gerken, Brian P. Young, and Jacob T. Hooper. Comparison of angler and cormorant harvest of walleye and yellow perch in Oneida Lake, New York. Transactions of the American Fisheries Society, 131(1): 27–39, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Valley:2002:EMB

[VB02] Rahman D. Valley and Mary T. Bremigan. Effects of macrophyte bed architecture on largemouth bass foraging: Implications of exotic macrophyte invasions. Transactions of the American Fisheries Society, 131(2):234–244, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vigliano:2009:QPG

[VBM^{+}09] Pablo H. Vigliano, David A. Beauchamp, Daniela Milano, Patricio J. Macchi, Marcelo F. Alonso, Martín I. García Asorey, María A. Denegri, Javier E. Ciancio, Gustavo Lippolt, Magali Rechencq, and Juan P. Barriga. Quantifying predation on galaxiids and other native organisms by introduced rainbow trout in an ultraoligotrophic lake in Northern Patagonia, Argentina: a bioenergetics modeling approach. Transactions of the American Fisheries Society, 138(6):1405–1419, November 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vatland:2008:BAM

[VBT08] Shane Vatland, Phaedra Budy, and Gary P. Thiede. A bioenergetics approach to modeling striped bass and threadfin shad predator-prey dynamics in Lake Powell, Utah–Arizona. Transactions of the American Fisheries Society, 137(1):262–277, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Visser:2002:UAP

[VDG02] R. Visser, D. D. Dauble, and D. R. Geist. Use of aerial photography to monitor fall Chinook salmon spawning in the Columbia River. Transactions of the American Fisheries Society, 131(6): 1173–1179, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDoornik:2002:PTG

[VFT02] Donald M. Van Doornik, Michael J. Ford, and David J. Teel. Patterns of temporal genetic variation in Coho salmon: Estimates of the effective proportion of 2-year-olds in natural and hatchery populations. Transactions of the American Fisheries Society, 131 (6):1007–1019, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Volpe:2001:RAA

[VGA01] John P. Volpe, Barry W. Glickman, and Bradley R. Anholt. Reproduction of aquaculture Atlantic salmon in a controlled stream channel on Vancouver Island, British Columbia. Transactions of the American Fisheries Society, 130(3):489–494, May 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDusen:2005:AAS

[VHF05] Peter J. VanDusen, Casey J. F. Huckins, and David J. Flaspohler. Associations among selection logging history, brook trout, macroinvertebrates, and habitat in Northern Michigan headwater streams. Transactions of the American Fisheries Society, 134(3): 762–774, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanEenennaam:2006:RCK

[VLD^{+}06] Joel P. Van Eenennaam, Javier Linares, Serge I. Doroshov, David C. Hillemeier, Thomas E. Willson, and Arnold A. Nova. Reproductive conditions of the Klamath River green sturgeon. Transactions of the American Fisheries Society, 135(1):151–163, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanderKooi:2001:EEI

[VMS01] Scott P. VanderKooi, Alec G. Maule, and Carl B. Schreck. The effects of electroshock on immune function and disease progression in juvenile spring Chinook salmon. Transactions of the American Fisheries Society, 130(3):397–408, May 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vadas:2001:FHS

[VO01] Robert L. Vadas, Jr. and Donald J. Orth. Formulation of habitat suitability models for stream fish guilds: Do the standard methods work? Transactions of the American Fisheries Society, 130(2): 217–235, March 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Veinott:2009:UMP

[VPN09] Geoffrey Ian Veinott, T. Rex Porter, and Lutz Nasdala. Using Mg as a proxy for crystal structure and Sr as an indicator of marine growth in vaterite and aragonite otoliths of aquaculture rainbow trout. Transactions of the American Fisheries Society, 138(5): 1157–1165, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vollestad:2004:EFM

[VPQ04] Leif Asbjørn Vøllestad, Jeramie Peterson, and Thomas P. Quinn. Effects of freshwater and marine growth rates on early maturity in male Coho and Chinook salmon. Transactions of the American Fisheries Society, 133(3):495–503, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vokoun:2005:HRS

[VR05] Jason C. Vokoun and Charles F. Rabeni. Home range and space use patterns of flathead catfish during the summer–fall period in two Missouri streams. Transactions of the American Fisheries Society, 134(2):509–517, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanTamelen:2005:EHM

[vT05] Peter G. van Tamelen. Estimating handling mortality due to air exposure: Development and application of thermal models for the Bering Sea snow crab fishery. Transactions of the American Fisheries Society, 134(2):411–429, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Venturelli:2006:DGN

[VT06] Paul A. Venturelli and William M. Tonn. Diet and growth of northern pike in the absence of prey fishes: Initial consequences for persisting in disturbance-prone lakes. Transactions of the American Fisheries Society, 135(6):1512–1522, November 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Volkoff:2007:UOM

[VT07] Martha C. Volkoff and Robert G. Titus. Use of otolith microstructure to discriminate stocks of juvenile Central Valley, California, fall-run Chinook salmon. Transactions of the American Fisheries Society, 136(1):195–210, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanEenennaam:2001:ASL

[VWD^{+}01] Joel P. Van Eenennaam, Molly A. H. Webb, Xin Deng, Serge I. Doroshov, Ryan B. Mayfield, Joseph J. Cech, Jr., David C. Hillemeier, and Thomas E. Willson. Artificial spawning and larval rearing of Klamath River green sturgeon. Transactions of the American Fisheries Society, 130(1):159–165, January 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wheeler:2003:HDP

[WA03] A. P. Wheeler and Michael S. Allen. Habitat and diet partitioning between shoal bass and largemouth bass in the Chipola River, Florida. Transactions of the American Fisheries Society, 132(3): 438–449, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waddy:2005:IIB

[WA05] S. L. Waddy and D. E. Aiken. Impact of invalid biological assumptions and misapplication of maturity criteria on size-at-maturity estimates for American lobster. Transactions of the American Fisheries Society, 134(5):1075–1090, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wildhaber:2000:NAI

[WAS^{+}00] Mark L. Wildhaber, Ann L. Allert, Christopher J. Schmitt, Vernon M. Tabor, Daniel Mulhern, Kenneth L. Powell, and Scott P. Sowa. Natural and anthropogenic influences on the distribution of the threatened neosho madtom in a Midwestern warmwater stream. Transactions of the American Fisheries Society, 129 (1):243–261, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wenburg:2001:GBE

[WB01] John K. Wenburg and Paul Bentzen. Genetic and behavioral evidence for restricted gene flow among coastal cutthroat trout populations. Transactions of the American Fisheries Society, 130 (6):1049–1069, November 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watters:2008:VPM

[WB08] George M. Watters and Cindy Bessey. Variation in the probability that male Coho salmon will mature early: Inferences from hierarchical models. Transactions of the American Fisheries Society, 137(1):70–95, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wood:2009:REF

[WB09] Jeremiah Wood and Phaedra Budy. The role of environmental factors in determining early survival and invasion success of exotic brown trout. Transactions of the American Fisheries Society, 138(4):756–767, July 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wills:2004:VEH

[WBH04] Todd C. Wills, Mary T. Bremigan, and Daniel B. Hayes. Variable effects of habitat enhancement structures across species and habitats in Michigan reservoirs. Transactions of the American Fisheries Society, 133(2):399–411, March 2004. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2005:IEE

[WBH^{+}05] Lance R. Williams, Timothy H. Bonner, James D. Hudson III, Marsha G. Williams, Tracy R. Leavy, and Casey S. Williams. Interactive effects of environmental variability and military training on stream biota of three headwater drainages in Western Louisiana. Transactions of the American Fisheries Society, 134 (1):192–206, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitledge:2006:IBM

[WBH06] Gregory W. Whitledge, Przemyslaw G. Bajer, and Robert S. Hayward. Improvement of bioenergetics model predictions for fish undergoing compensatory growth. Transactions of the American Fisheries Society, 135(1):49–54, January 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walsh:2000:EIM

[WBI00] Maureen G. Walsh, Kimberly A. Bjorgo, and J. Jeffery Isely. Effects of implantation method and temperature on mortality and loss of simulated transmitters in hybrid striped bass. Transactions of the American Fisheries Society, 129(2):539–544, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waite:2000:AAF

[WC00] Ian R. Waite and Kurt D. Carpenter. Associations among fish assemblage structure and environmental variables in Willamette Basin streams, Oregon. Transactions of the American Fisheries Society, 129(3):754–770, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2003:MSB

[WC03] Madelaine A. Williams and Charles C. Coutant. Modification of schooling behavior in larval atherinid fish Atherina mochon by heat exposure of eggs and larvae. Transactions of the American Fisheries Society, 132(4):638–645, July 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welker:2005:OSM

[WC05] Thomas L. Welker and James L. Congleton. Oxidative stress in migrating spring Chinook salmon smolts of hatchery origin: Changes in vitamin E and lipid peroxidation. Transactions of the American Fisheries Society, 134(6):1499–1508, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Widmer:2006:UTT

[WCBS06] Ann M. Widmer, Corissa J. Carveth, Scott A. Bonar, and Jeffrey R. Simms. Upper temperature tolerance of loach minnow under acute, chronic, and fluctuating thermal regimes. Transactions of the American Fisheries Society, 135(3):755–762, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wells:2008:ETH

[WCC08] R. J. David Wells, Matthew M. Chumchal, and James H. Cowan, Jr. Effect of trawling and habitat on mercury concentration in juvenile red snapper from the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 137(6):1839–1850, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webber:2007:USP

[WCM^{+}07] Jason D. Webber, Stephanie N. Chun, Teresa R. MacColl, Leslie T. Mirise, Ayako Kawabata, Emily K. Anderson, Tae Sung Cheong, Lev Kavvas, Maureen Gee McRotondo, Karen L. Hochgraf, Roger Churchwell, and Joseph J. Cech, Jr. Upstream swimming performance of adult white sturgeon: Effects of partial baffles and a ramp. Transactions of the American Fisheries Society, 136(2):402–408, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilde:2008:DSR

[WD08a] Gene R. Wilde and Bart W. Durham. Daily survival rates for juveniles of six Great Plains cyprinid species. Transactions of the American Fisheries Society, 137(3):830–833, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilde:2008:LHM

[WD08b] Gene R. Wilde and Bart W. Durham. A life history model for peppered chub, a broadcast-spawning cyprinid. Transactions of the American Fisheries Society, 137(6):1657–1666, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:2009:MDA

[WDHW09] John Waldman, Robert Daniels, Michael Hickerson, and Isaac Wirgin. Mitochondrial DNA analysis indicates sea lampreys are indigenous to Lake Ontario: Response to comment. Transactions of the American Fisheries Society, 138(5):1190–1197, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Esh09].

Wertheimer:2005:DPS

[WE05] Robert H. Wertheimer and Allen F. Evans. Downstream passage of steelhead kelts through hydroelectric dams on the Lower Snake and Columbia Rivers. Transactions of the American Fisheries Society, 134(4):853–865, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [Ano06a].

Ward:2000:RAL

[WEB00] Camille Ward, Randy L. Eshenroder, and James R. Bence. Relative abundance of lake trout and burbot in the main basin of Lake Michigan in the early 1930s. Transactions of the American Fisheries Society, 129(1):282–295, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warren:2009:ISF

[WEB09] Dana R. Warren, Anne G. Ernst, and Barry P. Baldigo. Influence of spring floods on year-class strength of fall- and spring-spawning salmonids in Catskill Mountain streams. Transactions of the American Fisheries Society, 138(1):200–210, January 2009. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weitkamp:2008:BRH

[Wei08] Laurie A. Weitkamp. Buoyancy regulation by hatchery and wild Coho salmon during the transition from freshwater to marine environments. Transactions of the American Fisheries Society, 137(3): 860–868, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weber:2005:CBH

[WF05] Edward D. Weber and Kurt D. Fausch. Competition between hatchery-reared and wild juvenile Chinook salmon in enclosures in the Sacramento River, California. Transactions of the American Fisheries Society, 134(1):44–58, January 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webb:2002:PCS

[WFF^{+}02] Molly A. H. Webb, Grant W. Feist, Eugene P. Foster, Carl B. Schreck, and Martin S. Fitzpatrick. Potential classification of sex and stage of gonadal maturity of wild white sturgeon using blood plasma indicators. Transactions of the American Fisheries Society, 131(1):132–142, January 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wesp:2003:DER

[WG03] Heather M. Wesp and Alice C. Gibb. Do endangered razorback suckers have poor larval escape performance relative to introduced rainbow trout? Transactions of the American Fisheries Society, 132(6):1166–1178, November 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williamson:2005:GFD

[WG05] Christopher J. Williamson and James E. Garvey. Growth, fecundity, and diets of newly established silver carp in the Middle Mississippi River. Transactions of the American Fisheries Society, 134(6):1423–1430, November 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:2004:MDA

[WGRW04] John R. Waldman, Cheryl Grunwald, Nirmal K. Roy, and Isaac I. Wirgin. Mitochondrial DNA analysis indicates sea lampreys are indigenous to Lake Ontario. Transactions of the American Fisheries Society, 133(4):950–960, July 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:2006:ENS

[WGW06] John R. Waldman, Cheryl Grunwald, and Isaac Wirgin. Evaluation of the native status of sea lampreys in Lake Champlain based on mitochondrial DNA sequencing analysis. Transactions of the American Fisheries Society, 135(4):1076–1085, July 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2007:WFS

[WH07] Jason L. White and Bret C. Harvey. Winter feeding success of stream trout under different streamflow and turbidity conditions. Transactions of the American Fisheries Society, 136(5): 1187–1192, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Workman:2002:MSM

[WHC02] R. Douglas Workman, Daniel B. Hayes, and Thomas G. Coon. A model of steelhead movement in relation to water temperature in two Lake Michigan tributaries. Transactions of the American Fisheries Society, 131(4):463–475, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wipfli:2004:RPS

[WHC04] Mark S. Wipfli, John P. Hudson, and John P. Caouette. Restoring productivity of salmon-based food Webs: Contrasting effects of salmon carcass and salmon carcass analog additions on streamresident salmonids. Transactions of the American Fisheries Society, 133(6):1440–1454, November 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wipfli:2003:MSF

[WHCC03] Mark S. Wipfli, John P. Hudson, John P. Caouette, and Dominic T. Chaloner. Marine subsidies in freshwater ecosystems: Salmon carcasses increase the growth rates of stream-resident salmonids. Transactions of the American Fisheries Society, 132 (2):371–381, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2000:GVW

[Whi00] Matthew M. White. Genetic variation in white bass. Transactions of the American Fisheries Society, 129(3):879–885, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whittier:2007:FAT

[WHLP07] Thomas R. Whittier, Robert M. Hughes, Gregg A. Lomnicky, and David V. Peck. Fish and amphibian tolerance values and an assemblage tolerance index for streams and rivers in the Western USA. Transactions of the American Fisheries Society, 136 (1):254–271, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whittier:2001:CTI

[WHP01] Thomas R. Whittier, Robert M. Hughes, and David V. Peck. Comment: Test of an index of biotic integrity. Transactions of the American Fisheries Society, 130(1):169–172, January 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [SHC01].

Whittier:2007:SAD

[WHS^{+}07] Thomas R. Whittier, Robert M. Hughes, John L. Stoddard, Gregg A. Lomnicky, David V. Peck, and Alan T. Herlihy. A structured approach for developing indices of biotic integrity: Three examples from streams and rivers in the Western USA. Transactions of the American Fisheries Society, 136(3):718–735, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wertheimer:2000:SAP

[WHT^{+}00] Alex C. Wertheimer, Ron A. Heintz, John F. Thedinga, Jacek M. Maselko, and Stanley D. Rice. Straying of adult pink salmon from their natal stream following embryonic exposure to weathered Exxon Valdez crude oil. Transactions of the American Fisheries Society, 129(4):989–1004, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitledge:2003:DLE

[WHZR03] Gregory W. Whitledge, Robert S. Hayward, Richard D. Zweifel, and Charles F. Rabeni. Development and laboratory evaluation of a bioenergetics model for subadult and adult smallmouth bass. Transactions of the American Fisheries Society, 132(2):316–325, March 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiley:2007:SCT

[Wil07] E. O. Wiley. Species concepts and their importance in fisheries management and research. Transactions of the American Fisheries Society, 136(4):1126–1135, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:2006:CHA

[WJB06] Tyler Wagner, Aaron K. Jubar, and Mary T. Bremigan. Can habitat alteration and spring angling explain largemouth bass nest success? Transactions of the American Fisheries Society, 135(4): 843–852, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weyers:2003:EPH

[WJF03] Richard S. Weyers, Cecil A. Jennings, and Mary C. Freeman. Effects of pulsed, high-velocity water flow on larval robust redhorse and V-lip redhorse. Transactions of the American Fisheries Society, 132(1):84–91, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wuenschel:2006:EED

[WJH06] Mark J. Wuenschel, Amelia R. Jugovich, and Jonathan A. Hare. Estimating the energy density of fish: The importance of ontogeny. Transactions of the American Fisheries Society, 135(2): 379–385, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weidel:2007:LFC

[WJK07] Brian C. Weidel, Daniel C. Josephson, and Clifford E. Kraft. Littoral fish community response to smallmouth bass removal from an Adirondack lake. Transactions of the American Fisheries Society, 136(3):778–789, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitledge:2007:SNC

[WJMM07] Gregory W. Whitledge, Brett M. Johnson, Patrick J. Martinez, and Anita M. Martinez. Sources of nonnative centrarchids in the Upper Colorado River revealed by stable isotope and microchemical analyses of otoliths. Transactions of the American Fisheries Society, 136(5):1263–1275, September 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wetzel:2005:DBG

[WK05] James Edwin Wetzel and Christopher C. Kohler. Distinction between gastric digestion and evacuation in black bass fed piscine prey. Transactions of the American Fisheries Society, 134(2):533– 536, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:2008:IAY

[WKCC08] David M. Warner, Courtney S. Kiley, Randall M. Claramunt, and David F. Clapp. The influence of alewife year-class strength on prey selection and abundance of age-1 Chinook salmon in Lake Michigan. Transactions of the American Fisheries Society, 137 (6):1683–1700, November 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirgin:2007:SIA

[WKM^{+}07] Isaac Wirgin, Adrienne I. Kovach, Lorraine Maceda, Nirmal K. Roy, John Waldman, and David L. Berlinsky. Stock identification of Atlantic cod in U.S. waters using microsatellite and single nucleotide polymorphism DNA analyses. Transactions of the American Fisheries Society, 136(2):375–391, March 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2005:GAO

[WKPS05] Matthew M. White, Todd W. Kassler, David P. Philipp, and Scott A. Schell. A genetic assessment of Ohio River walleyes. Transactions of the American Fisheries Society, 134(3):661–675, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weinstein:2000:RTS

[WLB^{+}00] Michael P. Weinstein, Steven Y. Litvin, Keith L. Bosley, Charlotte M. Fuller, and Sam C. Wainright. The role of tidal salt Marsh as an energy source for marine transient and resident finfishes: a stable isotope approach. Transactions of the American Fisheries Society, 129(3):797–810, May 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:2009:UPA

[WLBBM09] Sarah Walters, Susan Lowerre-Barbieri, Joel Bickford, and David Mann. Using a passive acoustic survey to identify spotted seatrout spawning sites and associated habitat in Tampa Bay, Florida. Transactions of the American Fisheries Society, 138(1):88–98, January 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wang:2003:IUL

[WLK03] Lizhu Wang, John Lyons, and Paul Kanehl. Impacts of urban land cover on trout streams in Wisconsin and Minnesota. Transactions of the American Fisheries Society, 132(5):825–839, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2005:EMW

[WM05a] Lance R. Williams and Arthur E. L. Morris. The ecology and management of wood in world rivers. Transactions of the American Fisheries Society, 134(5):1366–1367, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williamson:2005:ISI

[WM05b] Kevin S. Williamson and Bernie May. Inheritance studies implicate a genetic mechanism for apparent sex reversal in Chinook salmon. Transactions of the American Fisheries Society, 134(5): 1253–1261, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williamson:2007:MDH

[WM07] Kevin S. Williamson and Bernie May. Mitochondrial DNA haplotype diversity in apparent XY female fall-run and spring-run Chinook salmon in California’s Central Valley. Transactions of the American Fisheries Society, 136(6):1480–1486, November 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:2002:ETF

[WMBM02] David L. Ward, O. Eugene Maughan, Scott A. Bonar, and William J. Matter. Effects of temperature, fish length, and exercise on swimming performance of age-0 flannelmouth sucker. Transactions of the American Fisheries Society, 131(4):492–497, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiley:2004:RBP

[WMH^{+}04] Derek J. Wiley, Raymond P. Morgan II, Robert H. Hilderbrand, Richard L. Raesly, and Durland L. Shumway. Relations between physical habitat and American eel abundance in five river basins in Maryland. Transactions of the American Fisheries Society, 133(3):515–526, May 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitman:2003:PTC

[WMO03] Matthew S. Whitman, Edward H. Moran, and Robert T. Ourso. Photographic techniques for characterizing streambed particle sizes. Transactions of the American Fisheries Society, 132(3): 605–610, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:2009:IPD

[WNF09] D. M. Ward, K. H. Nislow, and C. L. Folt. Increased population density and suppressed prey biomass: Relative impacts on juvenile Atlantic salmon growth. Transactions of the American Fisheries Society, 138(1):135–143, January 2009. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodland:2002:SDL

[WNG02] Jason E. Woodland, Andrew D. Noyes, and John M. Grizzle. A survey to detect largemouth bass virus among fish from hatcheries in the Southeastern USA. Transactions of the American Fisheries Society, 131(2):308–311, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:2005:FNM

[WNH05] D. Scott Waters, Richard L. Noble, and Joseph E. Hightower. Fishing and natural mortality of adult largemouth bass in a tropical reservoir. Transactions of the American Fisheries Society, 134(3):563–571, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wright:2003:CPC

[WNT03] P. J. Wright, D. B. Noltie, and D. E. Tillitt. Comparison of prehatch C-start responses in rainbow trout and lake trout embryos by means of a tactile stimulus test. Transactions of the American Fisheries Society, 132(5):988–996, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woody:2000:TVP

[WORB00] Carol Ann Woody, Jeff Olsen, Joel Reynolds, and Paul Bentzen. Temporal variation in phenotypic and genotypic traits in two sockeye salmon populations, Tustumena Lake, Alaska. Transactions of the American Fisheries Society, 129(4):1031–1043, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilde:2008:SMP

[WP08] Gene R. Wilde and Kevin L. Pope. A simple model for predicting survival of angler-caught and released largemouth bass. Transactions of the American Fisheries Society, 137(3):834–840, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:2008:BSM

[WPGH08] Raymond A. Webster, Kenneth H. Pollock, Sujit K. Ghosh, and David G. Hankin. Bayesian spatial modeling of data from unitcount surveys of fish in streams. Transactions of the American Fisheries Society, 137(2):438–453, February 2008. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weigel:2002:MUP

[WPS02] Dana E. Weigel, James T. Peterson, and Paul Spruell. A model using phenotypic characteristics to detect introgressive hybridization in wild westslope cutthroat trout and rainbow trout. Transactions of the American Fisheries Society, 131(4):389–403, May 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2008:CHB

[WR08] Seth M. White and Frank J. Rahel. Complementation of habitats for Bonneville cutthroat trout in watersheds influenced by beavers, livestock, and drought. Transactions of the American Fisheries Society, 137(3):881–894, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wells:2003:RBW

[WRC^{+}03] Brian K. Wells, Bruce E. Rieman, James L. Clayton, Dona L. Horan, and Cynthia M. Jones. Relationships between water, otolith, and scale chemistries of westslope cutthroat trout from the Coeur d’Alene River, Idaho: The potential application of hard-part chemistry to describe movements in freshwater. Transactions of the American Fisheries Society, 132(3):409–424, May 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wang:2007:MWL

[WRC^{+}07] Hui-Yu Wang, Edward S. Rutherford, H. Andrew Cook, Donald W. Einhouse, Robert C. Haas, Timothy B. Johnson, Roger Kenyon, Brian Locke, and Mark W. Turner. Movement of walleyes in Lakes Erie and St. Clair inferred from tag return and fisheries data. Transactions of the American Fisheries Society, 136(2): 539–551, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2007:GIN

[WRGN07] Ian Williams, Gordon H. Reeves, Sara L. Graziano, and Jennifer L. Nielsen. Genetic investigation of natural hybridization between rainbow and coastal cutthroat trout in the Copper River Delta, Alaska. Transactions of the American Fisheries Society, 136(4):926–942, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:2002:STS

[WRK02] David M. Warner, Lars G. Rudstam, and Robert A. Klumb. In situ target strength of alewives in freshwater. Transactions of the American Fisheries Society, 131(2):212–223, March 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wingate:2007:ITH

[WS07a] Rebecca L. Wingate and David H. Secor. Intercept telemetry of the Hudson River striped bass resident contingent: Migration and homing patterns. Transactions of the American Fisheries Society, 136(1):95–104, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodland:2007:YCS

[WS07b] Ryan J. Woodland and David H. Secor. Year-class strength and recovery of endangered shortnose sturgeon in the Hudson River, New York. Transactions of the American Fisheries Society, 136 (1):72–81, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wingate:2008:EWT

[WS08] Rebecca L. Wingate and David H. Secor. Effects of winter temperature and flow on a summer-fall nursery fish assemblage in the Chesapeake Bay, Maryland. Transactions of the American Fisheries Society, 137(4):1147–1156, July 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilde:2009:EHR

[WS09] Gene R. Wilde and William Sawynok. Effect of hook removal on recapture rates of 27 species of angler-caught fish in Australia. Transactions of the American Fisheries Society, 138(3):692–697, May 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wills:2000:RRC

[WSA00] Paul S. Wills, Robert J. Sheehan, and Standish K. Allen, Jr. Reduced reproductive capacity in diploid and triploid hybrid sunfish. Transactions of the American Fisheries Society, 129(1):30– 40, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:2000:EST

[WSB00] Glenn N. Wagner, E. Don Stevens, and Philip Byrne. Effects of suture type and patterns on surgical wound healing in rainbow trout. Transactions of the American Fisheries Society, 129(5): 1196–1205, September 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolfe:2009:ECC

[WSEW09] Matthew D. Wolfe, Victor J. Santucci, Jr., Lisa M. Einfalt, and David H. Wahl. Effects of common carp on reproduction, growth, and survival of largemouth bass and bluegills. Transactions of the American Fisheries Society, 138(5):975–983, September 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Werner:2005:HSP

[WSFJ05] Ingeborg Werner, Timothy B. Smith, Joaquin Feliciano, and Michael L. Johnson. Heat shock proteins in juvenile steelhead reflect thermal conditions in the Navarro River Watershed, California. Transactions of the American Fisheries Society, 134(2): 399–410, March 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warren:2005:AGD

[WSJ^{+}05] Dana R. Warren, Stephen D. Sebestyen, Daniel C. Josephson, Jesse M. Lepak, and Clifford E. Kraft. Acidic groundwater discharge and in situ egg survival in redds of lake-spawning brook trout. Transactions of the American Fisheries Society, 134(5):

1193–1201, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wessel:2006:VMA

[WSJ06] Maria Lang Wessel, William W. Smoker, and John E. Joyce. Variation of morphology among juvenile Chinook salmon of hatchery, hybrid, and wild origin. Transactions of the American Fisheries Society, 135(2):333–340, March 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wertheimer:2001:CRH

[WSJH01] A. C. Wertheimer, W. W. Smoker, T. L. Joyce, and W. R. Heard. Comment: a review of the hatchery programs for pink salmon in Prince William Sound and Kodiak Island, Alaska. Transactions of the American Fisheries Society, 130(4):712–720, July 2001. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [HE01].

Welsh:2007:TBE

[WSLT07] Stuart A. Welsh, David R. Smith, R. Wilson Laney, and Ronald C. Tipton. Tag-based estimates of annual fishing mortality of a mixed Atlantic coastal stock of striped bass. Transactions of the American Fisheries Society, 136(1):34–42, January 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walsh:2005:ELH

[WSP05] Harvey J. Walsh, Lawrence R. Settle, and David S. Peters. Early life history of blueback herring and alewife in the Lower Roanoke River, North Carolina. Transactions of the American Fisheries Society, 134(4):910–926, July 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whiteley:2006:FSG

[WSRA06] Andrew R. Whiteley, Paul Spruell, Bruce E. Rieman, and Fred W. Allendorf. Fine-scale genetic structure of bull trout at the southern limit of their distribution. Transactions of the American Fisheries Society, 135(5):1238–1253, September 2006. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weitkamp:2003:BRF

[WSSD03a] Don E. Weitkamp, Robert D. Sullivan, Tim Swant, and Joe Dos-Santos. Behavior of resident fish relative to total dissolved gas supersaturation in the Lower Clark Fork River. Transactions of the American Fisheries Society, 132(5):856–864, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weitkamp:2003:GBD

[WSSD03b] Don E. Weitkamp, Robert D. Sullivan, Tim Swant, and Joe Dos-Santos. Gas bubble disease in resident fish of the Lower Clark Fork River. Transactions of the American Fisheries Society, 132 (5):865–876, September 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2007:CEF

[WSY^{+}07] David K. White, Christina Swanson, Paciencia S. Young, Joseph J. Cech, Jr., Zhi Qiang Chen, and M. Levent Kavvas. Close encounters with a fish screen II: Delta smelt behavior before and during screen contact. Transactions of the American Fisheries Society, 136(2):528–538, March 2007. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welch:2006:LSM

[WTB06] D. W. Welch, S. Turo, and S. D. Batten. Large-scale marine and freshwater movements of white sturgeon. Transactions of the American Fisheries Society, 135(2):386–389, March 2006. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wells:2000:GVT

[WTJ00] Brian K. Wells, Simon R. Thorrold, and Cynthia M. Jones. Geographic variation in trace element composition of juvenile weakfish scales. Transactions of the American Fisheries Society, 129(4): 889–900, July 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walther:2008:GSO

[WTO08] Benjamin D. Walther, Simon R. Thorrold, and John E. Olney. Geochemical signatures in otoliths record natal origins of American shad. Transactions of the American Fisheries Society, 137(1): 57–69, January 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winemiller:2000:FAS

[WTSC00] Kirk O. Winemiller, Soner Tarim, David Shormann, and James B. Cotner. Fish assemblage structure in relation to environmental variation among Brazos River Oxbow Lakes. Transactions of the American Fisheries Society, 129(2):451–468, March 2000. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wildhaber:2000:IPR

[WTW^{+}00] Mark L. Wildhaber, Vernon M. Tabor, Jo Anne E. Whitaker, Ann L. Allert, Daniel W. Mulhern, Peter J. Lamberson, and Kenneth L. Powell. Ictalurid populations in relation to the presence of a main-stem reservoir in a Midwestern warmwater stream with emphasis on the threatened neosho madtom. Transactions of the American Fisheries Society, 129(6):1264–1280, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2003:IFP

[WTW03] Lance R. Williams, Christopher M. Taylor, and Melvin L. Warren, Jr. Influence of fish predation on assemblage structure of macroinvertebrates in an intermittent stream. Transactions of the American Fisheries Society, 132(1):120–130, January 2003. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wydoski:2002:AGR

[WW02] Richard G. Wydoski and Richard S. Wydoski. Age, growth, and reproduction of mountain suckers in Lost Creek Reservoir, Utah. Transactions of the American Fisheries Society, 131(2):320–328, March 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Witte:2007:SDB

[WWKM07] F. Witte, J. H. Wanink, and M. Kishe-Machumu. Species distinction and the biodiversity crisis in Lake Victoria. Transactions of the American Fisheries Society, 136(4):1146–1159, July 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wehrly:2007:FBE

[WWM07] Kevin E. Wehrly, Lizhu Wang, and Matthew Mitro. Field-based estimates of thermal tolerance limits for trout: Incorporating exposure time and temperature fluctuation. Transactions of the American Fisheries Society, 136(2):365–374, March 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirgin:2000:GSA

[WWR^{+}00] Isaac Wirgin, John R. Waldman, Jannine Rosko, Rachel Gross, Mark R. Collins, S. Gordon Rogers, and Joseph Stabile. Genetic structure of Atlantic sturgeon populations based on mitochondrial DNA control region sequences. Transactions of the American Fisheries Society, 129(2):476–486, March 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wehrly:2003:CRV

[WWS03] Kevin E. Wehrly, Michael J. Wiley, and Paul W. Seelbach. Classifying regional variation in thermal regime based on stream fish community patterns. Transactions of the American Fisheries Society, 132(1):18–38, January 2003. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wydoski:2001:LHF

[Wyd01] Richard S. Wydoski. Life history and fecundity of mountain whitefish from Utah streams. Transactions of the American Fisheries Society, 130(4):692–698, July 2001. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Xiao:2000:EIR

[XM00] Y. Xiao and P. McShane. Estimation of instantaneous rates of fishing and natural mortalities from mark–recapture data on the western king prawn Penaeus latisulcatus in the Gulf St. Vincent, Australia, by conditional likelihood. Transactions of the American Fisheries Society, 129(4):1005–1017, July 2000. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ye:2001:SSD

[YA01] Yimin Ye and Adel H. Alsaffar. Study of shrimp distributions by means of locally weighted regression. Transactions of the American Fisheries Society, 130(1):68–79, January 2001. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2004:ECS

[YBSL04] Bradley A. Young, Mara B. Bryan, Stacia A. Sower, and Weiming Li. The effect of chemosterilization on sex steroid production in male sea lampreys. Transactions of the American Fisheries Society, 133(5):1270–1276, September 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2002:SBA

[YI02] Shawn P. Young and J. Jeffery Isely. Striped bass annual site fidelity and habitat utilization in J. Strom Thurmond Reservoir, South Carolina–Georgia. Transactions of the American Fisheries Society, 131(5):828–837, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2007:DBA

[YI07] Shawn P. Young and J. Jeffery Isely. Diel behavior of adult striped bass using tailwater habitat as summer refuge. Transactions of the American Fisheries Society, 136(4):1104–1112, July 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yako:2000:ACA

[YMJ00] Lisa A. Yako, Martha E. Mather, and Francis Juanes. Assessing the contribution of anadromous herring to largemouth bass growth. Transactions of the American Fisheries Society, 129(1): 77–88, January 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yule:2008:HSA

[YSB^{+}08] Daniel L. Yule, Jason D. Stockwell, Jeff A. Black, Ken I. Cullis, Gary A. Cholwek, and Jared T. Myers. How systematic age underestimation can impede understanding of fish population dynamics: Lessons learned from a Lake Superior cisco stock. Transactions of the American Fisheries Society, 137(2):481–495, February 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yule:2006:EME

[YSC^{+}06] Daniel L. Yule, Jason D. Stockwell, Gary A. Cholwek, Lori M. Evrard, Steven Schram, Michael Seider, and Matthew Symbal. Evaluation of methods to estimate lake herring spawner abundance in Lake Superior. Transactions of the American Fisheries Society, 135(3):680–694, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yurk:2000:EAR

[YT00] H. Yurk and A. W. Trites. Experimental attempts to reduce predation by harbor seals on out-migrating juvenile salmonids. Transactions of the American Fisheries Society, 129(6):1360– 1366, November 2000. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:2007:SDS

[YW07] Daniel B. Young and Carol Ann Woody. Spawning distribution of sockeye salmon in a glacially influenced watershed: The importance of glacial habitats. Transactions of the American Fisheries Society, 136(2):452–459, March 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zale:2005:ESI

[ZBF05] Alexander V. Zale, Carrie Brooke, and William C. Fraser. Effects of surgically implanted transmitter weights on growth and swimming stamina of small adult westslope cutthroat trout. Transactions of the American Fisheries Society, 134(3):653–660, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zigler:2004:HHF

[ZDK^{+}04] Steven J. Zigler, Michael R. Dewey, Brent C. Knights, Ann L. Runstrom, and Mark T. Steingraeber. Hydrologic and hydraulic factors affecting passage of paddlefish through dams in the Upper Mississippi River. Transactions of the American Fisheries Society, 133(1):160–172, January 2004. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:2009:MOM

[ZEP09] Christian E. Zimmerman, George W. Edwards, and Kathleen Perry. Maternal origin and migratory history of steelhead and rainbow trout captured in rivers of the Central Valley, California. Transactions of the American Fisheries Society, 138(2):280–291, March 2009. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zamor:2007:TAF

[ZG07] Richard M. Zamor and Gary D. Grossman. Turbidity affects foraging success of drift-feeding rosyside dace. Transactions of the American Fisheries Society, 136(1):167–176, January 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhou:2002:EPD

[Zho02a] Shijie Zhou. Estimating parameters of derived random variables: Comparison of the delta and parametric bootstrap methods. Transactions of the American Fisheries Society, 131(4):667– 675, July 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhou:2002:SDR

[Zho02b] Shijie Zhou. Size-dependent recovery of Chinook salmon in carcass surveys. Transactions of the American Fisheries Society, 131(6): 1194–1202, November 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ziegeweid:2008:EST

[ZJPB08] Jeffrey R. Ziegeweid, Cecil A. Jennings, Douglas L. Peterson, and Marsha C. Black. Effects of salinity, temperature, and weight on the survival of young-of-year shortnose sturgeon. Transactions of the American Fisheries Society, 137(5):1490–1499, September 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:2006:PDL

[ZKE06] Mara S. Zimmerman, Charles C. Krueger, and Randy L. Eshenroder. Phenotypic diversity of lake trout in Great Slave Lake: Differences in morphology, buoyancy, and habitat depth. Transactions of the American Fisheries Society, 135(4):1056–1067, July 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zerrenner:2005:ILS

[ZM05] Adam Zerrenner and J. Ellen Marsden. Influence of larval sea lamprey density on transformer life history characteristics in Lewis Creek, Vermont. Transactions of the American Fisheries Society, 134(3):687–696, May 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zerrenner:2006:CLS

[ZM06a] Adam Zerrenner and J. Ellen Marsden. Comparison of larval sea lamprey life history characteristics in a lampricide-treated tributary and untreated tributary system of Lake Champlain. Transactions of the American Fisheries Society, 135(5):1301– 1311, September 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:2006:RAC

[ZM06b] Chang-Ik Zhang and Bernard A. Megrey. A revised alverson and carney model for estimating the instantaneous rate of natural mortality. Transactions of the American Fisheries Society, 135(3):620–633, May 2006. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zorn:2007:IBT

[ZN07a] Troy G. Zorn and Andrew J. Nuhfer. Influences on brown trout and brook trout population dynamics in a Michigan River. Transactions of the American Fisheries Society, 136(3):691–705, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zorn:2007:RSB

[ZN07b] Troy G. Zorn and Andrew J. Nuhfer. Regional synchrony of brown trout and brook trout population dynamics among Michigan Rivers. Transactions of the American Fisheries Society, 136(3):706–717, May 2007. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:2002:ISR

[ZR02] Christian E. Zimmerman and Gordon H. Reeves. Identification of steelhead and resident rainbow trout progeny in the Deschutes River, Oregon, revealed with otolith microchemistry. Transactions of the American Fisheries Society, 131(5):986–993, September 2002. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zorn:2002:DSF

[ZSW02] Troy G. Zorn, Paul W. Seelbach, and Michael J. Wiley. Distributions of stream fishes and their relationship to stream size and hydrology in Michigan’s Lower Peninsula. Transactions of the American Fisheries Society, 131(1):70–85, January 2002. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:2008:MBC

[ZTP^{+}08] Huizi Zhang, Teresa Thieling, Samantha C. Bates Prins, Eric P. Smith, and Mark Hudy. Model-based clustering in a brook trout classification study within the Eastern United States. Transactions of the American Fisheries Society, 137(3):841–851, May 2008. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:2007:IBS

[ZV07] Julie K. H. Zimmerman and Bruce Vondracek. Interactions between slimy sculpin and trout: Slimy sculpin growth and diet in relation to native and nonnative trout. Transactions of the American Fisheries Society, 136(6):1791–1800, November 2007. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zeug:2005:RBR

[ZWT05] Steven C. Zeug, Kirk O. Winemiller, and Soner Tarim. Response of Brazos River oxbow fish assemblages to patterns of hydrologic connectivity and environmental variability. Transactions of the American Fisheries Society, 134(5):1389–1399, September 2005. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).